### v1

Abstract
We develop an information-theoretic framework for bounding the expected supremum and tail probabilities of stochastic processes, offering a simpler and sharper alternative to classical chaining and slicing arguments for generalization bounds. The key idea is a lifting argument that produces information-theoretic analogues of empirical process bounds, such as Dudley’s entropy integral. The lifting introduces symmetry, yielding sharp bounds even when the classical Dudley integral is loose. As a by-product, we obtain a concise proof of the majorizing measure theorem, providing explicit constants. The information-theoretic approach provides a soft version of classical localized complexity bounds in generalization theory, but is more concise and does not require the slicing argument. We apply this approach to empirical risk minimization over Sobolev ellipsoids and weak $\ell_{q}$ balls, obtaining sharper convergence rates or extensions to settings not covered by classical methods.
1 Introduction
Sharp control of empirical processes is central to modern statistical learning theory, underpinning generalization bounds, minimax rates, and risk analysis for high-dimensional estimators. Classical tools such as Dudley’s entropy integral have played a central role in this analysis, with applications ranging from nonparametric regression (Van de Geer,, 1990; Wainwright,, 2019) to sparse high-dimensional models (Raskutti et al.,, 2011). Although certain special cases can be handled by simpler techniques, the localized version of Dudley’s integral (Van de Geer,, 1990) has remained a fundamental tool in many general settings. However, its application to statistical learning may still be limited since (i) Dudley’s integral, while widely used, can be suboptimal even for basic examples such as Sobolev ellipsoids, and (ii) the slicing (or peeling) technique for localization often requires delicate covering number estimates, sometimes invoking deep results from approximation theory or the local theory of Banach spaces, e.g. in sparse regression over the $\ell_{q}$ ball (Raskutti et al.,, 2011).
In this paper, we introduce a “lifting” method for deriving new information-theoretic generalization bounds, which overcomes the limitations discussed above. The key idea is to introduce replicated copies of a system and apply the method of types from information theory, so that traditional “hard” inequalities for the supremum of a stochastic process for an index set yield “soft” information-theoretic inequalities for a distribution on that set. The strength of the lifting approach lies in leveraging geometric insights involving covering and packing numbers, such as the chaining argument in Dudley’s integral, yielding new information-theoretic inequalities that are not easily derived from traditional information-theoretic methods. This method yields sharper localized generalization bounds and new results for empirical risk minimization over ellipsoids and sparse regression.
For example, consider a Gaussian process $(X_{t})_{t\in T}$ and a measure $\mu$ on $T$. With the lifting approach, both Dudley’s integral and the majorizing measure theorem yield the following equivalences for Gaussian processes (up to universal constant factors):
$\sup\mathbb{E}[X_{Z}]\asymp\int_{0}^{\infty}\sqrt{\inf_{P_{Z\hat{Z}}\in\Pi_{\sigma}(\mu)}I(Z;\hat{Z})}d\sigma\asymp\int_{0}^{\infty}\sqrt{R_{\mu}(\sigma^{2})}d\sigma$ (1)
where $\Pi_{\sigma}(\mu)$ denotes the set of couplings of $\mu$ and itself with mean square error $\sigma^{2}$, I(Z;Z^)I($\hat{Z}$) is the mutual information, and $R_{\mu}(\sigma^{2})$ denotes the rate-distortion function (Thomas and Joy,, 2006). The supremum in (1) is over all couplings of the index and the process satisfying the marginal condition $P_{Z}=\mu$. While the analogy between the rate-distortion function $R_{\mu}(\sigma^{2})$ and covering numbers is well-known (Thomas and Joy,, 2006; Csiszár and Körner,, 2011), we are not aware of any prior literature explicitly stating (1). Previously, similar lifting arguments were used for problems in selected problems from network information theory (Wu et al.,, 2018; Liu,, 2022, 2021, 2023), but the application to empirical processes appears to be new. More generally, “soft” versions of traditional measure concentration techniques are known to provide sharper convergence in some hypothesis testing and information theory problems (Liu,, 2018; Liu et al.,, 2020; Liu,, 2020; Liu and Özgür,, 2020).
We use information-theoretic inequalities such as (1) to control the generalization error in empirical risk minimization (ERM), and this approach offers several advantages over the traditional approach based on slicing and Dudley’s integral:
First, our approach yields sharper convergence rate in some basic estimation problems. The reason is that (1) is a sharp two-sided inequality for any $\mu$, which might be surprising since we derive (1) from the Dudley integral, and the latter is not sharp for some $T$. The reason for the sharpness of (1) is that the lifting argument transforms the process into a stationary one - i.e., invariant under a transitive group of permutations - even if the original process lacks this symmetry. Since Dudley’s integral is sharp for stationary processes (see Section 2.1 for background), (1) inherits the sharpness, even if the original process is nonstationary. The sharpness of (1) also has an interesting consequence: by essentially maximizing over $\mu$ and applying the date processing inequality, we obtain a concise new proof of the nontrivial lower bound part of the classical majorizing measure theorem. This new proof of the majorizing measure theorem is conceptually simple, as it closely mirrors the argument used for the Dudley integral upper bound, essentially reversing its key steps. Furthermore, many nuisance parameters in the traditional proof disappear due to taking limits in the lifting argument. Figure 1 illustrates the implication relations of these results.
Dudley’s IntegralMajorizing Measure TheoremRate-Distortion Integralstationary processliftingliftingdata processing
Figure 1: Arrows indicate implications of results, and the techniques/assumptions for the implications. For Gaussian processes, the majorizing measure theorem and the proposed rate-distortion integral are two-sided bounds. Dudley’s integral is only an upper bound, unless stationarity is assumed.
Second, our approach makes localization of generalization bounds more straightforward. Localization refers to the phenomenon where the index $Z$ of the process concentrates on a subset much smaller than the full index set $T$, so that taking the supremum over all of $T$ may significantly overestimate the typical value of the process at $Z$. For example, in empirical risk minimization (ERM), $Z$ denotes the estimator, which tends to concentrate near zero as the sample size increases. Traditionally, localization is achieved by the “slicing” (or “peeling”) argument, which yields sharp convergence rate for some empirical risk minimization problems Wainwright, (2019). Slicing typically involves controlling the modulus of continuity of the process through careful controls of the deviation magnitude for different slices. In contrast, using properties of the rate-distortion function, we can often directly upper bound the right side of (1) using the variance of $Z\sim\mu$, which is much smaller than the size of $T$. Similarly, tail probability bounds on $X_{Z}$, based on the moment generating function of $\|Z\|_{2}$, can also be derived using a lifting technique. This new method of localization does not require slicing.
To demonstrate our approach, we derive new and sharper statistical rates in two classical ERM problems:
  • For mean estimation using ERM over the Sobolev ellipsoid with smoothness parameter $\beta>1/2$, which plays a fundamental role in nonparametric statistics (Tsybakov,, 2009), the slicing method based on Dudley’s integral only gives an upper bound $O(\frac{1}{2\beta-1})n^{-\frac{2\beta}{2\beta+1}}$ for the mean square error. The factor $\frac{1}{2\beta-1}$ arises in Dudley’s integral for the supremum over the Sobolev ellipsoid, a classical example for the looseness of Dudley’s integral (Talagrand,, 2014). On the other hand, our method achieves the optimal $O(\frac{1}{\sqrt{2\beta-1}})n^{-\frac{2\beta}{2\beta+1}}$ bound.
  • For sparse regression over the $\ell_{q}$-ball, $q\in(0,1]$, Raskutti et al. (Raskutti et al.,, 2011) showed that the mean square prediction error is upper bounded by $(\frac{\ln d}{n})^{\frac{2-q}{2}}R$ up to a constant depending on $q$, where $d$ is the feature dimension, $n$ is the sample size, $R^{1/q}$ is the $\ell_{q}$ radius, the noise is standard normal, and the columns of the design matrix are normalized to have $\ell_{2}$ norm at most $\sqrt{n}$. The proof in Raskutti et al., (2011) was based on the slicing method and estimating covering numbers of the $q$-absolutely convex hull, using deep results from convex geometry (Guedon and Litvak,, 2000; Carl and Pajor,, 1988); see proof of (Raskutti et al.,, 2011, Lemma 3). We provide a new proof without citing the convex geometry literature, using only simple properties of the entropy, such as the chain rule, to extend elementary one-dimensional estimates to higher dimensions. In fact, we show that the same convergence rate holds for weak $\ell_{q}$ balls ($q\in(0,1)$), which is new and appears to be beyond the reach of prior techniques in the literature.
Related work: A closely related line of recent research, started by Russo and Zou, (2016) and further developed by other authors including Xu and Raginsky, (2017); Dziugaite and Roy, (2017); Asadi et al., (2018); Bu et al., (2020); Aminian et al., (2023), proved information-theoretic generalization bounds in terms of the mutual information between the random index and the process. A basic version (see for example Russo and Zou, (2016)) is
$\mathbb{E}[X_{Z}]\lesssim\sup_{t\in T}\|X_{t}\|_{\Psi_{2}}\sqrt{I(Y;Z)}$ (2)
where the sub-Gaussian norm $\|\cdot\|_{\Psi_{2}}$ will be defined in (6), and $Y:=(X_{t})_{t\in T}$ denotes the stochastic process (in the case of finite $T$, we have $Y\in\mathbb{R}^{|T|}$). While this can be interpreted as a strengthening of a maximal inequality (Jiao et al.,, 2018), further refinement using the chaining argument, strengthening the Dudley integral, is also possible (Asadi et al.,, 2018). This type of results are useful in statistics and machine learning, where $Z$ represents the output of an algorithm which has limited mutual information with the training data, potentially much smaller than the covering number of $T$. Though also concerning a random index framework, their results differ from and are not directly comparable with (1), since their strengthening is in terms of the mutual information between $Z$ and the data, rather than the rate-distortion function of $Z$. Their proofs typically involve information-theoretic arguments such as chain rules and nonnegativity of the relative entropy, but not the lifting technique in the present paper. Nevertheless, we will show in Theorem 2 a result that encompasses both (1) and (2), highlighting the versatility of the lifting technique.
In related work on empirical risk minimization (ERM), it has been shown that under favorable conditions, such as convex regression classes with Gaussian noise, the mean-squared error (MSE) can often be characterized by a fixed-point equation involving a localized Gaussian width (Chatterjee,, 2014). For ellipsoids, this localized Gaussian width can then be controlled directly via the Cauchy–Schwarz inequality (Wei et al.,, 2020), thereby avoiding the slicing and Dudley-integral machinery and mitigating the looseness of Dudley’s bound. However, this shortcut relies essentially on concentration of the error, which can fail when the regression class is nonconvex (e.g., sparse regression under the $\ell_{q}$ constraint with $q<1$) or when the noise is sub-Gaussian but does not satisfy a measure concentration inequality (such as log-Sobolev or transportation-cost inequality). Moreover, the Cauchy-Schwarz step exploits linear structure: for example, consider a regression class given by the image of a Sobolev ellipsoid under a Lipschitz map (e.g., induced by a nonlinear neural network). This generalizes the setting of Wei et al., (2020), but their argument no longer applies. These limitations are expected to be artifacts of the proof: tail probability bounds need not require concentration (as naturally follows from the Dudley integral; see Lemma 2), and Lipschitz maps do not increase the supremum of the relevant stochastic process (a consequence of the majorizing measure theorem; see (13)). Unfortunately, existing tools do not suffice to rigorously establish these intuitions, and our work introduces an alternative framework that overcomes these limitations.
Organization: The paper is organized as follows: Section 2 provides more background on upper bounds for stochastic processes from probability theory, the traditional slicing technique for ERM, and information-theoretic tools used in our proof. Section 3 presents our main results, including (1), its sharpness for Gaussian processes, a simple proof of the classical majorizing measure theorem based on (1), a general framework for error tail bounds in ERM including results on the ellipsoid and sparse regression examples mentioned above. The remainder consists of details and proofs supporting the main results.
2 Preliminaries
Notation: We use $\wedge$ and $\vee$ to denote the max and min of two numbers, and $x]_{+}:=x\vee 0$. Given probability measures $P$ and $Q$ on a metric space $(T,d)$, $W_{p}(P,Q):=\sup_{P_{XY}\in\Pi(P,Q)}\mathbb{E}^{\frac{1}{p}}[d(X,Y)^{p}]$ denotes the Wasserstein distance, where $\Pi(P,Q)$ denotes the set of couplings. For brevity, $\Pi(P):=\Pi(P,P)$. Product measures, such as $P\times Q$ or $P^{\otimes N}$, indicate independence. ${\rm diam}(T):=\sup_{x,y\in T}d(x,y)$ denotes the diameter of a set $T$. Standard Landau notations such as $o(\dots)$ are used. $\lesssim_{q}$ means less than or equal up to a constant depending on $q$, and $\asymp_{q}$ means $\lesssim_{q}$ and $\gtrsim_{q}$ are both true. For any $q>0$, the $\ell_{q}$ ball and the weak $\ell_{q}$ ball of radius $r$ centered at 0 are
$B_{q}(r)$ $:=\{x\in\mathbb{R}^{d}\colon\sum_{i=1}^{d}x_{i}^{q}\leq r^{q}\},$ (3)
$wB_{q}(r)$ $:=\{x\in\mathbb{R}^{d}\colon\sup_{t>0}t^{q}\sum_{i=1}^{d}1_{|x_{i}|>t}\leq r^{q}\}.$ (4)
More generally, $B_{q}(z,r):=z+B_{q}(r)$ if the center is $z$. For brevity, $B(z,r):=B_{2}(z,r)$. $\ell_{2}$ denotes the Hilbert space of square-summable infinite sequences. The empirical distribution of a sequence $x=(x(1),x(2),\dots,x(d))$ is denoted by $\widehat{P}_{x}$. Given a set $T$ in $\mathbb{R}^{d}$ and a matrix $A\in\mathbb{R}^{n\times d}$, define $T-T:=\{t-s\colon t,s\in T\}$ and $AT:=\{At\colon t\in T\}$.
2.1 Upper Bounds for Stochastic Processes
Classically, Dudley’s integral provides upper bounds on a sub-Gaussian process in terms of the covering numbers. Let $(T,d)$ be a metric space, and suppose that $(X_{t})_{t\in T}$ is a centered sub-Gaussian process, that is, $\mathbb{E}[X_{t}]=0$, and the natural metric
$d(t,t^{\prime}):=\|X_{t}-X_{t^{\prime}}\|_{\Psi_{2}},\quad\forall t,t^{\prime}\in T,$ (5)
where, following Vershynin, (2018), we defined the sub-Gaussian norm
$\|X\|_{\Psi_{2}}=\inf\{K>0\colon\mathbb{E}[e^{X^{2}/K^{2}}]\leq 2\}.$ (6)
Dudley’ integral states that
$\mathbb{E}[\sup_{t\in T}X_{t}]\leq K\int_{0}^{\Delta}\sqrt{\ln N(T,\lambda)}d\lambda$ (7)
where $K$ is a universal constant, $N(T,\lambda)$ denotes the $\lambda$-covering number of $T$, i.e. the minimal number of (open) balls of radius $\lambda$ required to cover $T$, and $\Delta$ denotes the diameter of $T$. While usually called Dudley’s integral, (7) is occasionally attributed to other authors as well; see Dudley, (2016); Maurey, (2024); Talagrand, (2014) for more about the history. In the case of Gaussian processes, the natural metric is defined as
$d(t,t^{\prime}):=\sqrt{\mathbb{E}[|X_{t}-X_{t^{\prime}}|^{2}]}.$ (8)
Note that (8) differs from (5) by a universal constant factor (Vershynin,, 2018). Throughout this paper, when we discuss Gaussian processes specifically, we define the natural metric by (8) rather than (5). In the Gaussian case, a commonly referenced explicit bound is $K<24$ (Ledoux and Talagrand,, 2013). As pointed out in (Dudley,, 2016, Section 5.3), better estimates $K$ exist in the literature, and the proof essentially relies on the estimate
$\mathbb{E}[\max_{1\leq j\leq N}X_{j}]\leq\sqrt{2\ln N}$ (9)
where $X_{1},\dots,X_{N}$ are centered, jointly Gaussian, with variances not exceeding 1. If we drop Gaussianity and instead assume that $\|X_{1}\|_{\Psi_{2}},\dots,\|X_{N}\|_{\Psi_{2}}\leq 1$, then $\mathbb{E}[\max_{1\leq j\leq N}X_{j}]\leq\sqrt{2\ln(2N)}$ for $N\geq 1$ (Vershynin,, 2018, Section 2.7.3), which is slightly worse than the Gaussian case (9). We will see, however, that the difference will vanish when we apply the lifting argument, which effectively takes $N\to\infty$.
For stationary Gaussian processes, it is known since Fernique, (1975) that Dudley’s integral is tight:
$\int_{0}^{\Delta}\sqrt{\ln N(T,\lambda)}d\lambda\leq K^{\prime}\mathbb{E}[\sup_{t\in T}X_{t}]$ (10)
for a universal constant $K^{\prime}>0$. Maurey, (2024) gives a bound of $K^{\prime}\leq 432$. In the stationary case, the smaller balls in the chaining argument are equivalent in each level, so the chaining process in the proof of the Dudley integral can be reversed using the so-called super-Sudakov minoration (see Lemma 10 ahead).
For general nonstationary processes, the same paper of Fernique (Fernique,, 1975) conjectured what was later known as the majorizing measure theorem, whose nontrivial lower bound part was proved more than a decade later by Talagrand (Talagrand,, 1987) using a more sophisticated generic chaining argument. Given a probability measure $\mu$ on $T$, define
$I_{\mu}(t):=\int_{0}^{\Delta}\sqrt{\ln\frac{1}{\mu(B(t,\lambda))}}d\lambda$ (11)
The majorizing measure theorem states that the left side of (7) equals, up to a universal constant factor,
$\gamma_{2}(T):=\inf_{\mu}\sup_{t\in T}I_{\mu}(t).$ (12)
The usual definition of $\gamma_{2}(T)$ in terms of admissible partitions (cf. Talagrand, (2014)) is equivalent to (12) up to a constant factor. Generally speaking, $\gamma_{2}(T)$ is more difficult to evaluate than Dudley’s integral: $\mu$ needs to be optimized for a quantity depending on $\mu(B(t,\lambda))$ over different scales of $\lambda$, whereas in Dudley’s integral, we can calculate the covering number by optimizing for each scale separately.
Despite this, the majorizing measure theorem is appealing due to its sharpness for the Gaussian process. A nontrivial application of $\gamma_{2}$ is sub-Gaussian comparison: suppose that $(X_{t})$ is a separable sub-Gaussian process, $(Y_{t})$ is a separable Gaussian process, and both processes are zero mean with the same natural metric on $T$. Then
$\mathbb{E}[\sup_{t\in T}X_{t}]\lesssim\mathbb{E}[\sup_{t\in T}Y_{t}].$ (13)
The special case where $(X_{t})$ is also Gaussian is called the Gaussian comparison inequality, usually proved by an interpolation argument which is highly Gaussian-specific (Ledoux and Talagrand,, 2013). The more general sub-Gaussian case can be immediately seen from the majorizing measure theorem. As noted in Van Handel, (2014), this version of sub-Gaussian comparison is a nontrivial consequence of the majorizing measure theorem, for which no other proof is known.
It is also known (see Talagrand, (2014)) that the following quantities are equivalent to $\gamma_{2}(T)$ up to constants:
${\rm Fer}(T)$ $:=\sup_{\mu}\int_{T}I_{\mu}(t)d\mu(t);$ (14)
$\delta_{2}(T)$ $:=\sup_{\mu}\inf_{t\in T}I_{\mu}(t).$ (15)
Once the expected value is understood, tail bounds can be established either by applying concentration of measure, or by reworking the chaining steps to control the tail probability instead of the expectation. For technical convenience, our discussion of tail bounds will be restricted to processes of the form $X_{t}=\left<Y,t\right$>, where $t\in T\subseteq\ell_{2}$. Following Vershynin, (2018), the sub-Gaussian norm of a random vector $Y\in\ell_{2}$ is defined by taking the supremum over one-dimensional projections:
$\|Y\|_{\Psi_{2}}:=\sup_{v\in\ell_{2}}\|\left<Y,v\right>\|_{\Psi_{2}}.$ (16)
A bound on $\|Y\|_{\Psi_{2}}$ can be guaranteed, for example, if $Y$ has independent coordinates with uniformly bounded sub-Gaussian norms (Vershynin,, 2018).
Concentration inequalities control the deviation probability from the mean. We say $P_{Y}$ satisfies the log-Sobolev inequality (LSI), if for any continuously differentiable function $f\colon\ell_{2}\to\mathbb{R}$,
${\rm Ent}(f^{2})\leq 2\mathbb{E}[\|\nabla f(Y)\|_{2}^{2}],$ (17)
where ${\rm Ent}(f):=\mathbb{E}[f(Y)\ln f(Y)]-\mathbb{E}[f(Y)]\ln\mathbb{E}[f(Y)]$. The log-Sobolev inequality is satisfied when the coordinates of $Y$ are independent and each satisfies the log-Sobolev inequality (tensorization property). The log-Sobolev inequality implies sub-Gaussian concentration. In fact, the following property is well-known:
Lemma 1.
Suppose that $X_{t}:=\left<t,Y\right$>, where $Y$ is a zero-mean random vector in $\ell_{2}$ satisfying the log-Sobolev inequality, and $T$ is totally bounded (i.e., covering numbers are all finite) in $\ell_{2}$. Let $\sigma^{2}:=\sup_{t\in T}\mathbb{E}[X_{t}^{2}]$, and let $\xi:=\sup_{t\in T}X_{t}$. Then for all $u>0$,
$\mathbb{P}[\xi-\mathbb{E}[\xi]>u]\leq e^{-\frac{u^{2}}{2\sigma^{2}}}.$ (18)
The proof follows by showing the Lipschitz condition $\|\nabla\max_{t\in T}\left<t,y\right>\|_{2}\leq\max_{t\in T}\|t\|_{2}$, and using the fact that the log-Sobolev inequality implies sub-Gaussian concentration; see (Boucheron et al.,, 2013, Theorem 5.8). In fact, from the proof we can see that the LSI assumption in Lemma 1 can be weakened to Talagrand’s optimal transport inequality, which also implies sub-Gaussian deviation of Lipschitz functions (Boucheron et al.,, 2013).
It is possible, however, to obtain tail bounds without assuming concentration:
Lemma 2.
Suppose that $Y$ is a zero-mean random vector in $\ell_{2}$ satisfying $\|Y\|_{\Psi_{2}}\leq 1$, and $\sup_{t\in T}\|t\|_{2}\leq\Delta$. Then there exists a universal constant $C>0$ such that for any $x\geq 0$,
$\mathbb{P}[\sup_{t\in T}\left<t,Y\right>\geq CJ(T)+x]\leq Ce^{-\frac{x^{2}}{C\Delta^{2}}},$ (19)
where $J(T):=\int_{0}^{\infty}\sqrt{\ln N(T,\epsilon)}\epsilon$.
A proof of Lemma 2 can be found in Van Handel, (2014) and the references therein. Statistical literature, such as Van de Geer, (1990), also used similar tail probability bounds for sub-Gaussian processes. As noted in Van Handel, (2014), although Lemma 2 requires weaker an assumption on the distribution than Lemma 1, it can give a weaker bound when $J(T)$ is larger than $\mathbb{E}[\xi]$.
2.2 Empirical Risk Minimization (ERM)
In a general mean estimation problem, $T$ is a closed subset of $\ell_{2}$, and $X_{1},\dots,X_{n}$ are i.i.d. random vectors in $\ell_{2}$ whose mean is an unknown $m\in T$. For any $t\in\ell_{2}$, define the empirical risk
$\hat{L}(t):=\frac{1}{n}\sum_{i=1}^{n}\|t-X_{i}\|_{2}^{2}.$ (20)
The (population) risk is $L(t):=\mathbb{E}[\hat{L}(t)]$. The empirical risk minimizer (ERM) is defined as
$\hat{m}:=\operatorname*{arg\,min}_{t\in T}\hat{L}(t).$ (21)
We have
$\|\hat{m}-m\|^{2}$ $=L(\hat{m})-L(m)$ (22)
$\leq-\hat{L}(\hat{m})+\hat{L}(m)+L(\hat{m})-L(m)$ (23)
$=\frac{1}{\sqrt{n}}\chi_{m-\hat{m}}$ (24)
where $\chi_{z}:=-2\sqrt{n}\left<z,\frac{1}{n}\sum_{i=1}^{n}X_{i}-m\right$>, and $(\chi_{z})_{z\in T-T}$ is a centered sub-Gaussian process, whose natural metric is the $\ell_{2}$ metric (up to a universal constant factor). An example of such a mean estimation problem, which plays a fundamental role in nonparametric statistics, will be discussed in Section 3.4, where $T$ is a Sobolev ellipsoid.
In a general regression problem, $T$ is a closed subset in $\mathbb{R}^{n}$, and $Y=\sqrt{n}m+w$, where $m\in T$ is an unknown vector to estimate, and $w$ is a zero-mean noise vector satisfying $\|w\|_{\Psi_{2}}\leq 1$. An example of sparse regression with fixed designs will be discussed in Section 3.5. The case of random designs, which is traditionally tackled by local Rademacher complexity techniques (Lugosi and Wegkamp,, 2004; Bartlett et al.,, 2005), is not considered here, although it would be interesting to consider such extensions for future work. Define
$\hat{L}(t):=\frac{1}{n}\|\sqrt{n}t-Y\|_{2}^{2}$ (25)
and define $L(t)$ and $\hat{m}$ the same way as in the mean estimation problem. Then (24) still holds, if we now define $\chi_{z}:=-2\left<z,w\right$>, and $(\chi_{z})_{z\in T-T}$ is a centered sub-Gaussian process, whose natural metric is the $\ell_{2}$ metric (up to a universal constant factor).
With a slight change of notation, it is possible to unify the formulations of the two problems: both are essentially about projecting a random vector onto a subset in $\ell_{2}$. We present them as estimation and regression problems separately simply to make the sample size $n$ emerge naturally and consistently with the nonparametric estimation and sparse regression literature. More generally, the framework also encompasses many other problems including functional estimation, matrix completion, shape-restricted regression, and isotonic regression; see Chatterjee, (2014); Chatterjee et al., (2015).
In both problems, the inequality (24) plays a key role in upper bounding the mean square error $\mathbb{E}[\|\hat{m}-m\|_{2}^{2}]$; see e.g. (Van de Geer,, 1990, eq (2)). Since $\mathbb{E}[\hat{L}(m)-L(m)]=0$, the expected generalization error equals $\mathbb{E}[\frac{1}{\sqrt{n}}\chi_{m-\hat{m}}$], which is also connected to (24). If we take the expectation of both sides of (24), and upper bound the right side by the Dudley integral using the covering number of $T$, we generally do not get sharp convergence rates of the mean square error. In fact, even for the simplest case of 1-dimensional parameter estimation, such a naive approach yields a $1/\sqrt{n}$ rate rather than the optimal $1/n$ rate.
The traditional approach for improving the bound is a “slicing” argument that yields local complexity bounds; see for example Van de Geer, (1990); Wainwright, (2019); Raskutti et al., (2011) or (Van Handel,, 2014, Theorem 5.32). Roughly speaking, if $\|\hat{m}-m\|^{2}=r$, then $\chi_{m-\hat{m}}\leq\sup_{z\in B(r)\cap(T-T)}\chi_{z}$, which is bounded by some function $\omega(r)$ with high probability. The function $\omega(r)$ typically grows slower than $r^{2}$, so $r^{2}\lesssim\frac{1}{\sqrt{n}}\omega(r)$ yields an upper bound on $r$. More precisely, a union bound argument is applied to control the generalization error at different scales of $r$ with high probability. While it may seem that the uniform control by the union bound could lose logarithmic factors, this can actually be avoided by carefully scheduling of the error tolerance at each scale.
In Section 8.1, we give a more detailed review of how the traditional slicing approach can yield sharp convergence for the one-dimensional toy problem. Local complexity bounds, which are based on Dudley’ integral for local sets, can still be loose for general $T$ where Dudley’ integral is loose. The majorizing measure theorem, though sharp, is hard to evaluate, and is rarely used in the analysis of ERM.
2.3 Information-Theoretic Tools
We recall standard information-theoretic measures in Thomas and Joy, (2006), using the natural base of the logarithm. The entropy of a discrete random variable $X\sim\mu$ is $H(X):=\sum_{x}\mu(x)\ln\frac{1}{\mu(x)}$. The relative entropy between $\mu$ and $\nu$ on the same measurable space is $D(\mu\|\nu):=\int\ln\frac{d\mu}{d\nu}d\mu$. Mutual information between $(X,Y)\sim P_{XY}$ is defined as $I(X;Y):=D(P_{XY}\|P_{X}\times P_{Y})$. Conditional versions of these quantities are defined by averaging over the random variable which is conditioned on. For example, $D(P_{\hat{Z}|Z}\|\mu|P_{Z}):=\int D(P_{\hat{Z}|Z=z}\|\mu)dP_{Z}(z)$. The lifting argument relies on the method of types, a standard method in information theory (Csiszár and Körner,, 2011) and large deviation analysis (Dembo,, 2009). The type of a sequence of length $N$ is simply its empirical distribution. Since not all distributions can be a type, we need the following:
Definition 1.
Suppose that $\mu$ is a distribution on a finite set $\mathcal{Z}$. We say $\mu$ is rational if there exists integer $N>0$ such that $N\mu$ is integer, i.e., $N\mu(z)$ is integer for any $z\in\mathcal{Z}$.
If $N\nu$ is integer, we denote by $C_{\nu}$ the $\nu$-type set, i.e., the set of sequences with type $\nu$. Then the following result is standard and can be found in (Csiszár and Körner,, 2011, Chapter 1):
Lemma 3.
Let $S_{N}$ be the set of all possible types of $\mathcal{Z}^{N}$, where $\mathcal{Z}$ is a set satisfying the cardinality bound $|\mathcal{Z}|\leq n$. Then $|S_{N}|={n+N-1\choose n-1}\leq(N+1)^{n}$. Furthermore, if $\mu$ is a distribution on $\mathcal{Z}$ and $N\mu$ is integer, then
$(N+1)^{-n}\exp(-ND(\nu\|\mu))\leq\mu^{\otimes N}(C_{\nu})\leq\exp(-ND(\nu\|\mu)).$ (26)
3 Main Results
3.1 A Rate-Distortion Integral
Our first main result concerns $\mathbb{E}[X_{Z}]$, where $Z$ is a random index that may depend on the realization of the process, which is a natural consequence of our information-theoretic technique. We first clarify some measure-theoretic aspects regarding the random index formulation. First, if $T=\{t_{0},t_{1},\dots\}$ is countable, we assume that for each $t\in T$, $X_{t}(\omega)$ is a measurable function of $\omega\in\Omega$ (where $\Omega$ denotes the sample space). Suppose that $Z(\omega)\in\{t_{0},t_{1},\dots\}$ is also a measurable function of $\omega\in\Omega$. Then $X_{Z(\omega)}(\omega)$ is a measurable function of $\omega$, because for any Borel set $B\subseteq\mathbb{R}$,
$\{\omega\colon X_{Z(\omega)}(\omega)\in B\}=\bigcup_{i=0}^{\infty}\{\omega\colon Z(\omega)=t_{i}\}\cap\{\omega\colon X_{t_{i}}(\omega)\in B\}$ (27)
which is measurable.
If $T$ is uncountable, measurability is not so immediate since the union in (27) is uncountable. The standard approach is to assume separability, i.e., there is a dense countable subset $T_{0}=\{t_{0},t_{1},\dots\}\subset T$, and $X_{t}$ is continuous in $t\in T$ almost surely, so that $X_{t}$ for $t\notin T_{0}$ can be defined by passing a limit. Continuity here is with respect to the topology compatible with the metric $d$. We note that separability and continuity ensure that $X_{Z(\omega)}(\omega)$ is measurable, as long as $Z(\omega)$ is measurable, which should suffice for most statistical applications. Indeed, we can construct random variable $Z_{k}:=t_{i}$ where
$i=\operatorname*{arg\,min}_{j\in\{0,1,\dots,k\}}d(Z(\omega),t_{j})$ (28)
with ties broken arbitrarily but respecting measurability. Assuming measurability of $Z(\omega)$, we have measurability of $d(Z(\omega),t_{j})$ for each $j$, so that $Z_{k}$, and hence $X_{Z_{k}}$, is measurable. Density of $T_{0}$ ensures $\lim_{k\to\infty}Z_{k}=Z$ almost surely, and continuity implies that $X_{Z}=\lim_{k\to\infty}X_{Z_{k}}$ is measurable, being a sequential limit of measurable functions.
We recall the rate-distortion function in information theory (Thomas and Joy,, 2006), which will be a counterpart of the covering number in Dudley’s integral:
Definition 2.
Given $\mu$ on a metric space $(T,d)$, the rate-distortion function $R_{\mu}(\sigma^{2}):=\inf_{P_{UZ}}I(U;Z)$, where the infimum is over $P_{UZ}$ satisfying $P_{Z}=\mu$ and $\mathbb{E}[d^{2}(U,Z)]\leq\sigma^{2}$.
Given a probability measure $\mu$ on $T$ and $\sigma>0$, define
$\Pi_{\sigma}(\mu):=\{P_{Z\hat{Z}}\colon P_{Z}=P_{\hat{Z}}=\mu;\,\mathbb{E}[d^{2}(Z,\hat{Z})]\leq\sigma^{2}\}.$ (29)
Then we have
$R_{\mu}(\sigma^{2})\leq\inf_{P_{Z\hat{Z}}\in\Pi_{\sigma}(P_{Z})}I(Z;\hat{Z})\leq R_{\mu}(\sigma^{2}/4).$ (30)
[…]
For $Z\sim\mu$, define
$\sigma_{\rm m}(\mu):=\inf_{z\in T}\sqrt{\mathbb{E}[d^{2}(Z,t)]},$ (31)
abbreviated as $\sigma_{\rm m}$ if no confusion. Clearly $R_{\mu}(\sigma^{2})=0$ whenever $\sigma>\sigma_{\rm m}$. Therefore $\sigma_{\rm m}$ can be thought of as a soft version of the diameter of $T$, since the in the upper limit of the traditional Dudley integral can be set as the diameter.
Definition 3.
Given a stochastic process $(X_{t}(\omega))_{t\in T}$, define the complexity
$\mathsf{w}(\mu):=\sup\mathbb{E}[X_{Z}]$ (32)
where the supremum is over all $Z=Z(\omega)$ satisfying $P_{Z}=\mu$.
Clearly $\mathsf{w}(\mu)$ is analogous to the notion of Gaussian complexity of a set.
Theorem 1.
If $(X_{t})_{t\in T}$ is a centered sub-Gaussian process on a countable set $T$, and $\mu$ is a distribution on $T$ for which $\sigma_{\rm m}<\infty$, then
$\mathsf{w}(\mu)\leq 48\int_{0}^{\sigma_{\rm m}}\sqrt{R_{\mu}(\sigma^{2})}d\sigma.$ (33)
The proof is given in Section 4. It will be seen in Lemma 6 that if the right side of (33) is finite, then $\mathbb{E}[|X_{Z}|]<\infty$, so no need to worry about absolute integrability on the left side of (33). The constant 48 in the theorem is from the commonly referenced estimate $K<24$ in the Dudley integral (7) for Gaussian processes (see comments after (7) for the extension to the sub-Gaussian case). Actually, according to Dudley, (2016), the best published estimate appears to be $K\leq 4\sqrt{2}$ given in Lifshits, (2012), which would imply that 48 in Theorem 1 can be improved to $8\sqrt{2}$.
Another type of information-theoretic improvements of traditional generalization bounds, discussed around (2), have recently been studied in the literature. While this type of results are not directly comparable to Theorem 2, we can use the lifting approach to obtain a strengthening of Theorem 2 that encompass both improvements. For simplicity, assume that $|T|<\infty$, and we have the following improvement of Theorem 1:
Theorem 2.
Suppose that $(X_{t})_{t\in T}$ is a centered sub-Gaussian process, and let $Y:=(X_{t})\in\mathbb{R}^{|T|}$. Let $Z$ be a random variable on $T$ arbitrarily correlated with $Y$, and set $\mu:=P_{Z}$. Then
$\mathbb{E}[X_{Z}]\leq 48\int_{0}^{\sigma_{\rm m}}\sqrt{R_{\mu}(\sigma^{2})\wedge I(Y;Z)}d\sigma.$ (34)
The proof of Theorem 2 is given in Section 4.1. Since $\sigma_{\rm m}\lesssim\sup_{t\in T}\|X_{t}\|_{\Psi_{2}}$, it is easy to see that Theorem 2 implies (2). Since the rest of the paper does not concern Theorem 2, we will not discuss its extension beyond finite $T$. While Theorem 1 does not cover the case of uncountable $T$, it is sufficient for our statistical applications due to the following observation:
Proposition 1.
If $(X_{t})$ is a separable, continuous process, then for $Z\sim\mu$,
$\mathbb{E}[X_{Z}]\leq 48\lim_{\epsilon\downarrow 0}\sup_{\mu^{\prime}\colon W_{\infty}(\mu,\mu^{\prime})\leq\epsilon}\int_{0}^{\sigma_{\rm m}}\sqrt{R_{\mu^{\prime}}(\sigma^{2})}d\sigma.$ (35)
The proof of Proposition 4.2 is in Section 4.2. Naturally, we would like to show that the right side of (35) reduces to the right side of (33) by dominated convergence, but this requires showing that $\sqrt{R_{\mu^{\prime}}(\sigma^{2})}$ can be dominated by an absolutely integrable function, which is challenging in the regime of vanishing $\sigma$ where the rate-distortion function may grow fast. While it is possible to impose on certain growth condition for dominated convergence to apply, we will not dive into those complications. Instead, in many statistical applications, we will see that $R_{\mu^{\prime}}(\sigma^{2})$ can controlled using the second moments of $\mu$, so the right side of (35) can be directly evaluated using the continuity of the bound in the second moment, without the need to establish the equivalence of the right sides of (35) and (33).
3.2 Sharpness for Gaussian Processes
While the traditional Dudley integral is not sharp, somewhat surprisingly, its information-theoretic counterpart, Theorem 1, is sharp:
Theorem 3.
In the setting of Theorem 1, suppose that $(X_{t})_{t\in T}$ is a Gaussian process (with metric defined in (8)). For any $\mu$, we have
$\mathbb{E}[X_{Z}]\geq c\int_{0}^{\sigma_{\rm m}}\sqrt{R_{\mu}(\sigma^{2})}d\sigma$ (36)
where the supremum is over all $Z(\omega)$ satisfying $P_{Z}=\mu$, and
$c:=\max_{\tau\in(0,4/5)}\left\{\frac{4-5\tau}{2\sqrt{2}(4-\tau)\ln\frac{4}{\tau}}\right\}\approx 0.0935918,$ (37)
numerically evaluated at $\tau=0.23457905$.
Theorem 3 suggests that Theorem 1 may yield sharper results in certain statistical applications where the traditional method based on Dudley’s integral is suboptimal. Three proofs of Theorem 3 are given in Section 5: First, the result can be proved using the lifting argument, by applying the majorizing measure theorem to the $\mu$-type set. Secondly, we can apply the lifting argument and use Fernique’s theorem that Dudley’s integral is sharp for stationary Gaussian processes, i.e., the case where the metric on $T$ is invariant under the action of a transitive group (see (Dudley,, 1999, Cor 2.7.5) or (Fernique,, 1975)), which preceded Talagrand’s striking proof of the majorizing measure theorem (Talagrand,, 1987). Even though $T$ in Theorem 3 may not be permutation-invariant, the lifting argument in our proof lower bounds the expected supremum over the typical set, which satisfies permutation invariance. The third proof derives an information-theoretic counterpart of the super-Sudakov inequality (Talagrand,, 1992; Van Handel,, 2014), and finishes off by purely analytic arguments (without chaining or packing constructions), which gives the constant stated in Theorem 3.
Classically, equivalent formulations of the majorizing measure theorem, (14)-(15), were established by Fernique convexity argument; see (Talagrand,, 2014, Chapter 3) for a proof using combinatorial arguments such as tree packing. Similarly, the rate-distortion integral can also be expressed in alternative forms by applying the minimax theorem to a certain convex-concave functional; see Remark 1 for discussions.
We can derive from Theorem 3 the classical the majorizing measure theorem, with an explicit estimate of constant. Recall $I_{\mu}$ defined in (11).
Corollary 1.
For any Gaussian process (not necessarily stationary),
$\mathbb{E}[\sup_{t\in T}X_{t}]\geq\bar{c}\inf_{\mu}\sup_{t\in T}I_{\mu}(t).$ (38)
Here $\bar{c}:=\sup_{a>1}\frac{\sqrt{1-a^{-2}}}{\frac{2a}{c}+\sqrt{2\pi h(a^{-2})}}\approx 0.0218988$ (evaluated at $a=1.4392$), where $h$ is the binary entropy function in nats and $c$ was defined in Theorem 3.
The proof of Corollary 1 is given in Section 6, which is essentially based on lower bounding the rate-distortion function by the probability of a set using the data processing inequality. We have not found a reference on the majoring measure theorem that estimated the constant to compare with Corollary 1.
Another application of Theorem 3 is sub-Gaussian comparison theorems, which, in statistical applications, sometimes allow us to reduce the problem to simpler “linear” cases where straightforward Cauchy-Schwarz or Hölder inequality arguments apply.
Corollary 2.
Suppose that $(X_{t}(\omega))_{t\in T}$ is a centered sub-Gaussian process on a countable $T$, with complexity function $\mathsf{w}_{1}$ (Definition 3) and natural metric $d_{1}$. Suppose that $(G_{t}(\omega^{\prime}))_{t\in T}$ is a Gaussian process with complexity function $\mathsf{w}_{2}$ and natural metric $d_{2}$, and $d_{1}$ is dominated by $d_{2}$. Then we have $\mathsf{w}_{1}(\mu)\lesssim\mathsf{w}_{2}(\mu)$ for any $\mu$ on $T$.
The proof immediately follows from Theorem 1-3, since the rate-distortion functions under the two metrics can be directly compared. Corollary 2 can be viewed as a “soft” version of (13) (in fact, it is easy to see that it implies (13)). The Euclidean special case of Corollary 2 is of particular interest in practice:
Corollary 3.
If $T$ is a countable subset of $\mathbb{R}^{N}$ and the covariance matrix $\operatorname*{cov}(Z)=\Sigma$, and the natural metric on $T$ for a centered sub-Gaussian process $(X_{t})_{t\in T}$ (see (5)) is dominated by the Euclidean metric, then for any $P_{Z}$ on $T$, we have $\mathbb{E}[X_{Z}]\lesssim\operatorname*{tr}(\sqrt{\Sigma})$.
Proof.
Consider a Gaussian process $(G_{t})_{t\in T}$, where $G_{t}=\left<G,t\right$> and $G\sim\mathcal{N}(0,I_{N})$. The natural metric for this Gaussian process coincides with the Euclidean metric. Then by Cauchy-Schwarz we have $\mathbb{E}[G_{Z}]=\mathbb{E}[\left<\Sigma^{1/4}G,\Sigma^{-1/4}Z\right>]\leq\operatorname*{tr}(\sqrt{\Sigma})$. The claim then follows from Corollary 2. ∎
Although Corollary 3 looks like a simple Cauchy-Schwarz inequality, it would not be true if sub-Gaussianity is replaced by a second-moment bound. More precisely, “the natural metric on $T$ for a centered sub-Gaussian process $(X_{t})_{t\in T}$ is dominated by the Euclidean metric” cannot be replaced by $(X_{t})$ is a stochastic process satisfying $\mathbb{E}[|X_{t}-X_{t^{\prime}}|^{2}]\lesssim\|t-t^{\prime}\|_{2}^{2}$”. Indeed, consider $T$ as a maximal $\epsilon$-packing of the unit ball in $\mathbb{R}^{N}$, where $N=3$, and suppose that $X_{t}$ is an independent zero mean random variable with $\epsilon^{2}$ variance for each $t\in T$. Then it is true that $\mathbb{E}[|X_{t}-X_{t^{\prime}}|^{2}]\lesssim\|t-t^{\prime}\|_{2}^{2}$. However, for fixed $N=3$ and vanishing $\epsilon$, it is possible that $\mathbb{E}[X_{Z}]=\mathbb{E}[\max_{t\in T}X_{t}]\asymp\epsilon\sqrt{|T|}\asymp\epsilon^{1-N/2}$ diverges, for certain heavy-tailed $X_{t}$. On the other hand, since $Z$ is in the unit ball, $\operatorname*{tr}(\sqrt{\Sigma})=O(1)$ as $\epsilon$ vanishes, confirming that this is a counterexample.
3.3 Tail Bounds for ERM
In this section we consider processes in the form of an inner product in $\ell_{2}$. In this case, the complexity of an index distribution $\mu$ (Definition 3) is written as
$\mathsf{w}(\mu):=\sup_{P_{YZ}\in\Pi(\mu,\gamma)}\mathbb{E}[\left<Y,Z\right$>], (39)
where $\gamma$ is a given distribution on $\ell_{2}$, which is assumed to have zero mean in this section. We may denote $\mathsf{w}(\mu)$ by $\mathsf{w}(\mu;\gamma)$ to make the dependence on $\gamma$ explicit. The following regularity condition will be used in this section:
Definition 4.
$T\subseteq\ell_{2}$ is totally bounded if the covering number $N(T,\epsilon)$ is finite for every $\epsilon>0$.
In empirical process theory, total boundedness is a rather mild condition which is almost always assumed; see for example Ledoux and Talagrand, (2013). It is sometimes also referred to as precompactness in topology. It is also equivalent to
$\lim_{n\to\infty}\sup_{t\in T}\sum_{i>n}t(i)^{2}=0.$ (40)
Indeed, if (40) is true, then for every $\epsilon>0$ there exists $n$ such that $\sup_{t\in T}\sum_{i>n}t(i)^{2}<(\epsilon/4)^{2}$. We can cover the projection of $T$ onto the first $n$ coordinates with finitely many balls of radius $\epsilon/2$, whose centers are the projections of $t_{1},\dots,t_{M}\in T$. Then $B(t_{1},\epsilon),\dots,B(t_{M},\epsilon)$ also cover $T$. Conversely, total boundedness implies that every sequence in $T$ has a convergent subsequence, which implies (40).
Definition 5.
Suppose that $T$ is a subset of $\ell_{2}$. We call $\mathsf{G}\colon[0,\infty)\to[0,\infty)$ a width-moment function of $T$ if for any random vector $Z\in T$ satisfying $\mathbb{E}[\|Z\|_{2}^{2}]\leq E$, we have
$\mathsf{w}(P_{Z})\leq\mathsf{G}(E).$ (41)
Consider the setting of the empirical risk minimization problem in Section 2.2 (either the mean estimation or the regression problem). Our main result, for $\gamma$ satisfying the log-Sobolev inequality, is the following:
Theorem 4.
Suppose that $m\in T$, where $T$ is a closed, totally bounded subset of $\ell_{2}$, and TTT-T has a width-moment function $\mathsf{G}$. Assume that $\gamma$ has zero mean and satisfies the log-Sobolev inequality (17). Let $\psi(\lambda):=\ln\mathbb{E}[\exp(\lambda\|m-\hat{m}\|_{2}^{2})]$. Then for any $\lambda\geq 0$,
$\psi(\lambda)\leq\sup_{\epsilon>0}\left\{-\epsilon+\frac{2\lambda}{\sqrt{n}}\mathsf{G}\left(\frac{\epsilon+\psi(\lambda)}{\lambda}\right)+\sqrt{\frac{8\epsilon\lambda(\epsilon+\psi(\lambda))}{n}}\right\}.$ (42)
The proof of Theorem 4 is given in Section 7. The log-Sobolev inequality assumption in Theorem 4 can in fact be weakened to sub-Gaussianity. In that case, define
$\mathsf{w}_{\rm g}(\mu):=\sup_{P_{YZ}\in\Pi(\gamma_{\rm g},\mu)}\mathbb{E}\left[\sum_{i=1}^{\infty}Y(i)Z(i)\right],$ (43)
where $\gamma_{\rm g}$ denotes the distribution of an infinite sequence of i.i.d. standard normal random variables, and the infinite sum is defined in terms of convergence in second moment. Similarly, define $\mathsf{G}_{\rm g}(E)$ by replacing the left side of (41) with $\mathsf{w}_{\rm g}(P_{Z})$. Note that if $\gamma$ satisfies the log-Sobolev inequality (hence also sub-Gaussian), then Corollary 2 implies that $\mathsf{w}_{\rm g}$ and $\mathsf{G}_{\rm g}$ dominate $\mathsf{w}$ and $\mathsf{G}$ respectively. We have the following result:
Theorem 5.
In the setting of Theorem 4, suppose that $\gamma$ is only assumed to have zero-mean and satisfy $\|Y\|_{\Psi_{2}}\leq 1$, $Y\sim\gamma$. Assume that $\mathsf{G}_{\rm g}(\cdot)$ is continuous. Then for any $\lambda\geq 0$,
$\psi(\lambda)\leq\sup_{\epsilon>0}\left\{-\frac{\epsilon}{C}+\frac{2K^{\prime}C\lambda}{\sqrt{n}}\mathsf{G}_{\rm g}\left(\frac{\epsilon+\psi(\lambda)}{\lambda}\right)+K^{\prime}C\sqrt{\frac{8\epsilon\lambda(\epsilon+\psi(\lambda))}{n}}\right\},$ (44)
where $C,K^{\prime}>0$ are universal constants in Lemma 2 and (10), respectively.
The proof of Theorem 5 is given in Section 7.2. Similar to the remark following Lemma 2, Theorem 5 assumes a weaker condition on $\gamma$ than (4), but may yield a weaker bound if $\mathsf{G}_{\rm g}$ is larger than $\mathsf{G}$ (for example, in the case where $Y$ has i.i.d. coordinates following the uniform distribution on [1,1][-1,1]).
3.4 ERM for the Sobolev Ellipsoid
We apply the general technique in Section 3.3 to Sobolev ellipsoids, which plays a prominent role in nonparametric statistics (Van de Geer,, 1990; Wainwright,, 2019; Tsybakov,, 2009). The precise problem of estimation over a Sobolev ellipsoid was previously studied in Wei et al., (2020). Generally, an ellipsoid in $\ell_{2}$ is defined as
$\mathcal{E}:=\left\{t\colon\sum_{i=1}^{\infty}\frac{t_{i}^{2}}{a_{i}^{2}}\leq 1\right\},$ (45)
where $a_{1},a_{2},\dots$ is a decreasing sequence of nonnegative numbers. This is recurring example in Talagrand, (2014), in particular an example showing that Dudley’s integral is loose. A Sobolev ellipsoid is an ellipsoid with $a_{i}=i^{-\beta}$ for some $\beta>0$. By considering $t(1),t(2),\dots$ as the Fourier coefficients of a regression function from the $\beta$-Sobolev space, we see that the Sobolev ellipsoid can be identified with a nonparametric function class (Tsybakov,, 2009). In Section 9.1, we will review the supremum of a Gaussian process indexed by $\mathcal{E}$, and show that the rate-distortion integral yields a sharp bound.
Consider ERM in the setting of Section 2.2, where $T=\phi(\mathcal{E})$ is the image of the Sobolev ellipsoid (for $\beta>1/2$, where $1/2$ is the critical exponent for ERM) under a 1-Lipschitz map $\phi\colon\ell_{2}\to\ell_{2}$. The purpose of introducing $\phi$ is that if $\phi$ is linear, an elementary proof exists, while the rate-distortion integral approach can handle more general settings. Furthermore, we assume that $Y$ is a random vector in $\ell_{2}$ satisfying $\|Y\|_{\Psi_{2}}\leq 1$. Under these assumptions, we have the following main result:
Theorem 6.
There exists a function $c(\beta)>0$ satisfying $c(\beta)=O(\frac{1}{\sqrt{2\beta-1}})$ as $\beta\downarrow\frac{1}{2}$, and a universal constant $b>0$ such that
$\mathbb{P}[\|m-\hat{m}\|_{2}>\sqrt{c(\beta)}n^{-\frac{\beta}{2\beta+1}}]\leq\exp(-bc(\beta)n^{\frac{1}{2\beta+1}}).$ (46)
Theorem 6 shows that ERM achieves square error $O(\frac{1}{\sqrt{2\beta-1}})n^{-\frac{2\beta}{2\beta+1}}$ with high probability, which is optimal not only in terms of the dependence on $n$, but also in terms of the behavior of the preconstant $O(\frac{1}{\sqrt{2\beta-1}})$ as $\beta\downarrow 1/2$. Furthermore, from the proof of Theorem 6 we can also see that the expected generalization error, $\mathbb{E}[\frac{1}{\sqrt{n}}\chi_{m-\hat{m}}$], also converges at the rate $O(\frac{1}{\sqrt{2\beta-1}})n^{-\frac{2\beta}{2\beta+1}}$.
We present two approaches towards Theorem 6 in Section 9.2. The first one does not use any new tools introduced in this paper, but only works for a special case of Theorem 6 where $\phi$ is the identity map and $Y$ satisfies the stronger log-Sobolev inequality: We use the Cauchy-Schwarz argument to bound the mean of the square error, and then use a concentration inequality to control the deviation from the mean. This approach is a slightly streamlined version of the method in Chatterjee, (2014); Wei et al., (2020). The second proof follows the general recipe in Section 3.3, and works for general sub-Gaussian $Y$ and nonlinear Lipschitz $\phi$, where the rate-distortion integral is used to bound $\mathsf{G}_{\rm g}$.
For comparison, the traditional techniques will only prove an upper bound at the rate $O(\frac{1}{2\beta-1})n^{-\frac{2\beta}{2\beta+1}}$, which is optimal in terms of the dependence on $n$, but not in terms of the dependence on $\beta$ in the preconstant as $\beta\downarrow 1/2$ (we analyze the preconstant using the traditional method in Remark 3). Our result shows that the optimal $O(\frac{1}{\sqrt{2\beta-1}})$ preconstant can be regained by using the rate-distortion integral instead.
3.5 ERM for Sparse Regression
This section provides new upper bounds for empirical risk minimization for the weak $\ell_{q}$ ball in $\mathbb{R}^{d}$ defined in (4), extending prior results for the $\ell_{q}$ ball (3). Consider a special case of the general regression model in Section 2.2, where $T$ is chosen as a linear transform of $wB_{q}(r)$. More precisely, suppose that the columns of $\frac{1}{\sqrt{n}}\mathbb{X}$ have unit $\ell_{2}$ norm, $\beta^{*}\in wB_{q}(r)$, $q\in(0,1)$, and $Y=\mathbb{X}\beta^{*}+w$, where $w$ is a zero mean random vector satisfying $\|w\|_{\Psi_{2}}\leq 1$. Let $\hat{\beta}:=\operatorname*{arg\,min}_{\beta\in wB_{q}(r)}\|Y-\mathbb{X}\beta\|_{2}$. Throughout, we set
$R:=r^{q},$ (47)
and we caution the reader that some prior work, including Raskutti et al., (2011), uses the notation $B_{q}(R)=B_{q}(r^{q})$ to refer to what we denote by $B_{q}(r)$. Applying the technique in Section 3.3, we obtain the following result:
Theorem 7.
There exists $C_{q}>0$ (depending only on $q\in(0,1)$) such that
$\mathbb{P}\left[\|m-\hat{m}\|_{2}^{2}>C_{q}(\frac{\ln d}{n})^{\frac{2-q}{2}}R\right]\leq\exp(-(\ln d)^{\frac{2-q}{2}}Rn^{\frac{q}{2}}).$ (48)
The proof of Section 10 is given in Theorem 7, which is based on the rate-distortion integral and the general technique for tail bounds in Section 3.3.
Previously, Raskutti et al., (2011) established the sharp convergence of ERM when $\beta\in B_{q}(r)$ where $q\in(0,1]$. The analysis for the case of $\beta\in B_{1}(r)$ follows by a simple Hölder inequality argument (see Raskutti et al., (2011)), whereas the $q<1$ case was proved by the slicing technique, based on the Dudley integral where the covering numbers were bounded using deep results from the convex geometry (Carl and Pajor,, 1988; Guedon and Litvak,, 2000). In contrast, we prove Theorem 7 using the rate-distortion integral, where the rate-distortion function (a priori smaller than the log covering number, hence easier to upper bound) is upper bounded using purely information-theoretic tools, without invoking the convex geometry literature. Furthermore, Theorem 7 shows that the minimax rate for $wB_{q}(r)$ is the same as $B_{q}(r)$ when $q\in(0,1)$, a nontrivial extension of the result of Raskutti et al., (2011) which appears to be beyond the previous techniques. The $q=1$ case is somewhat special: by considering $\mathbb{X}$ with identical rows, we can see that the rate for $wB_{1}(r)$ can be worse than that of $B_{1}(r)$ by a logarithmic factor. On the other hand, in the special case where $\frac{1}{\sqrt{n}}\mathbb{X}$ is the identity, the rate $(\frac{\ln d}{n})^{\frac{2-q}{2}}R$ is achievable for $wB_{q}(r)$ for all $q\in(0,2)$; see Rigollet and Hütter, (2023); Donoho and Johnstone, (1994); Zhang, (2010).
4 Proof of the Rate-Distortion Integral
In this section we prove Theorem 1, Theorem 2, and Proposition 1.
4.1 Proof in the Finite Case
Recall the following basic property of the sub-Gaussian distribution:
Proposition 2.
Suppose that $X_{1},X_{2}$ are independent random variables, and $\|X_{i}\|_{\Psi_{2}}\leq K_{i}$, $i=1,2$. Then $\|X_{1}+X_{2}\|_{\Psi_{2}}\leq\sqrt{K_{1}^{2}+K_{2}^{2}}$.
Proof.
By Jensen’s inequality,
$\mathbb{E}\left[\exp(\frac{X_{1}^{2}+X_{2}^{2}}{K_{1}^{2}+K_{2}^{2}})\right]$ $=\mathbb{E}\left[\exp(\frac{X_{1}^{2}}{K_{1}^{2}+K_{2}^{2}})\right]\mathbb{E}\left[\exp(\frac{X_{2}^{2}}{K_{1}^{2}+K_{2}^{2}})\right]$ (49)
$\leq\mathbb{E}^{\frac{K_{1}^{2}}{K_{1}^{2}+K_{2}^{2}}}\left[\exp(\frac{X_{1}^{2}}{K_{1}^{2}})\right]\mathbb{E}^{\frac{K_{2}^{2}}{K_{1}^{2}+K_{2}^{2}}}\left[\exp(\frac{X_{2}^{2}}{K_{2}^{2}})\right]$ (50)
$\leq 2.$ (51)
We also need the following estimate of the covering number:
Lemma 4.
For given $\sigma>0$ independent of $N$, we have
$NR_{\mu}(4\sigma^{2})+o(N)\leq\ln N(\mathcal{C},2\sqrt{N}\sigma)\leq NR_{\mu}(\sigma^{2})+o(N)$ (52)
where $o(N)$ hides constants depending on $P_{Y}$ and $\sigma$.
Proof.
For any $t^{N}\in\mathcal{C}$, a standard application of Sanov’s theorem (Csiszár and Körner,, 2011, Chapter 1) shows
$\ln\frac{1}{\mu^{\otimes N}(B(t^{N},\sqrt{N}\sigma))}=NR_{\mu}(\sigma^{2})+o(N)$ (53)
where the implicit constant in $o(N)$ may depend on $\mu$. Let $P(\mathcal{C},2\sqrt{N}\sigma)$ denote the packing number, i.e. the maximal number of points $\mathcal{C}$ separated by distance at least $2\sqrt{N}\sigma$. Then $N(\mathcal{C},2\sqrt{N}\sigma)\leq P(\mathcal{C},2\sqrt{N}\sigma)$ (see e.g. Van Handel, (2014)), and standard volume estimates yields $P(\mathcal{C},2\sqrt{N}\sigma)\leq\frac{1}{\min_{t^{N}\in\mathcal{C}}\mu^{\otimes N}(B(t^{N},\sqrt{N}\sigma))}$, proving the second inequality in (52). Moreover, by a volume argument we also have $N(\mathcal{C},2\sqrt{N}\sigma)\max_{t^{N}\in\mathcal{C}}\mu^{\otimes N}(B(t^{N},2\sqrt{N}\sigma))\geq\mu^{\otimes N}(\mathcal{C})=\exp(o(N))$. Then using (53) with $\sigma$ replaced by $2\sigma$, we obtain the first inequality in (52). ∎
Proof of Theorem 1 (for finite $T$).
For any integer $N>0$, let $\mathcal{C}$ be the $P_{Z}$-typical set (Thomas and Joy,, 2006):
$\mathcal{C}:=\mathcal{C}_{N}:=\{z^{N}\in T^{N}\colon\|\widehat{P}_{z^{N}}-\mu\|_{1}\leq N^{-1/3}\}$ (54)
where $\widehat{P}_{z^{N}}$ denotes the empirical distribution of $z^{N}$, and $\|\widehat{P}_{z^{N}}-\mu\|_{1}:=\sum_{z\in T}|\widehat{P}_{z^{N}}(z)-\mu(z)|$ can in fact be replaced by any norms, which are equivalent up to constants due to finiteness of $T$. Moreover, let $n:=|T|$. Without loss of generality, we can identify $T$ with the standard basis vectors $e_{1},e_{2},\dots,e_{n}$ of $\mathbb{R}^{n}$, and define $Y_{i}:=X_{e_{i}}$ for each $i\in\{1,\dots,n\}$. Then the process can be represented by inner product:
$X_{z}=\left<Y,z\right>,\quad\forall z\in T.$ (55)
It is not necessary to adopt the inner product representation for this proof, but it helps making the notations clear. Suppose that $P_{YZ}\in\Pi(P_{Y},\mu)$ is arbitrary, and let $(Y^{N},Z^{N})\sim P_{YZ}^{\otimes N}$. We have
$\mathbb{E}[\left<Y,Z\right>]=\lim_{N\to\infty}\mathbb{E}[\frac{1}{N}\left<Y^{N},Z^{N}\right$>]. (56)
By Proposition 2, $(\left<Y^{N},z^{N}\right>)_{z^{N}\in\mathcal{C}}$ is a sub-Gaussian process with respect to the metric $d_{N}(z^{N},\hat{z}^{N}):=\sqrt{\sum_{i=1}^{N}d^{2}(z_{i},\hat{z}_{i})}$. Let $1_{\mathcal{C}}$ be the indicator of the event that $Z^{N}\in\mathcal{C}$. Since $\left<Y^{N},Z^{N}\right>1_{\mathcal{C}}\leq\max_{z^{N}\in\mathcal{C}}\left<Y^{N},z^{N}\right>1_{\mathcal{C}}=\max_{z^{N}\in\mathcal{C}}\left<Y^{N},z^{N}\right>-\max_{z^{N}\in\mathcal{C}}\left<Y^{N},z^{N}\right>1_{\mathcal{C}^{c}}$, we have
$\mathbb{E}[\left<Y^{N},Z^{N}\right$>] $\leq\mathbb{E}[\left<Y^{N},Z^{N}\right>1_{\mathcal{C}^{c}}]+\mathbb{E}[\max_{z^{N}\in\mathcal{C}}\left<Y^{N},z^{N}\right>]-\mathbb{E}[\max_{z^{N}\in\mathcal{C}}\left<Y^{N},z^{N}\right>1_{\mathcal{C}^{c}}]$ (57)
$\leq\mathbb{E}[\max_{z^{N}\in\mathcal{C}}\left<Y^{N},z^{N}\right>]+\mathbb{E}[\max_{z^{N},\hat{z}^{N}\in\mathcal{Z}^{N}}\left<Y^{N},z^{N}-\hat{z}^{N}\right>1_{\mathcal{C}^{c}}].$ (58)
The last term above has vanishing contribution: Indeed, note that for each $z^{n},\hat{z}^{n}\in\mathcal{Z}^{N}$, we have $\|\left<Y^{N},z^{N}-\hat{z}^{N}\right>\|_{\Psi_{2}}\leq C\sqrt{N}$ for some $C$ independent of $N$ (may depend on $P_{YZ}$). Taking the max over $\exp(O(N))$ such sub-Gaussian random variables will increase the sub-Gaussian norm by a factor of no more than $\sqrt{N}$ (Vershynin,, 2018), so
$\|\max_{z^{N},\hat{z}^{N}\in\mathcal{Z}^{N}}\left<Y^{N},z^{N}-\hat{z}^{N}\right>\|_{\Psi_{2}}\leq CN.$ (59)
Moreover, by Hoeffding’s inequality we also have $\mathbb{E}[1_{\mathcal{C}^{c}}]=\exp(-\Theta(N^{1/3}))$, so we can use Cauchy-Schwarz (or (Talagrand,, 2014, Lemma 7.8.4)) to show the second expectation in (58) is $o(N)$. Therefore, we have verified that
$\mathbb{E}[\left<Y,Z\right>]\leq\liminf_{N\to\infty}\frac{1}{N}\mathbb{E}[\max_{z^{N}\in\mathcal{C}}\left<Y^{N},z^{N}\right$>]. (60)
Furthermore, applying Dudley’s integral (see (7) and the subsequent comments) to $\mathcal{C}$ yields
$\mathbb{E}[\max_{z^{N}\in\mathcal{C}}\left<Y^{N},z^{N}\right>]\leq 24\int_{0}^{\Delta_{\rm m}}\sqrt{\ln N(\mathcal{C},2\sqrt{N}\sigma)}2\sqrt{N}d\sigma$ (61)
where $\Delta_{\rm m}:=\sup_{N\geq 1}\frac{{\rm diam}(\mathcal{C}_{N})}{2\sqrt{N}}<\infty$. Now we multiply $1/N$ to (61), take $N\to\infty$, and apply Lemma 4. We can apply dominated convergence, since $\frac{1}{N}\sqrt{\ln N(\mathcal{C},2\sqrt{N}\sigma)}2\sqrt{N}\leq\frac{2}{\sqrt{N}}\sqrt{\ln(n^{N})}=2\sqrt{\ln n}$, which is independent of $N$ and absolutely integrable on $\sigma\in[0,\Delta_{\rm m}]$. We have thus proved Theorem 1 for finite $T$. ∎
The proof of Theorem 2 will be almost identical to the proof of Theorem 1, but we will use an additional technique about the projection to the set of sequences of a given type. More precisely:
Lemma 5.
Let $\mu$ be a rational distribution on $\mathbb{R}^{n}$. Fix $N\geq 1$ such that $N\mu$ is integer. For any $y^{N}\in\mathbb{R}^{nN}$, we have
$\frac{1}{\sqrt{N}}\min_{s^{N}\colon\widehat{P}_{s^{N}}=\mu}\|y^{N}-s^{N}\|_{2}=W_{2}(\widehat{P}_{y^{N}},\mu).$ (62)
Proof.
The $\geq$ part is obvious from the definition. For the $\leq$ part, fix an arbitrary $\bar{s}^{N}\in\mathbb{R}^{nN}$ satisfying $\widehat{P}_{\bar{s}^{N}}=\mu$. Denote the supports of $\widehat{P}_{y^{N}}$ and $\mu$ be $\mathcal{A}$ and $\mathcal{B}$, respectively. Consider the following function
$F(M)$ $=\frac{1}{N}\sum_{1\leq i,j\leq N}\|y_{i}-\bar{s}_{j}\|_{2}^{2}M_{i,j}$ (63)
$=\frac{1}{N}\sum_{y\in\mathcal{A},\,s\in\mathcal{B}}\|y-s\|_{2}^{2}\sum_{i\colon y_{i}=y}\sum_{j\colon\bar{s}_{j}=s}M_{i,j}$ (64)
on $\mathcal{S}$, which is defined as the set of doubly stochastic matrices $M\in\mathbb{R}^{N\times N}$. We can see from (64) that $\inf_{M\in\mathcal{S}}F(M)\leq W_{2}^{2}(\widehat{P}_{y^{N}},\mu)$. Since $F$ is linear, by the Birkhoff–von Neumann theorem, there exists a permutation matrix $M$ achieving $\inf_{M\in\mathcal{S}}F(M)$. This permutation sends $\bar{s}^{N}$ to an $s^{N}$, achieving $\inf_{M\in\mathcal{S}}F(M)=\frac{1}{N}\|y^{N}-s^{N}\|_{2}$. This proves the $\leq$ part. ∎
Proof of Theorem 2.
As in the proof of Theorem 1, we can identify $T=\{e_{1},\dots,e_{n}\}$ as the standard basis vectors, so that $X_{z}=\left<Y,z\right$>. The assumption of zero-mean process implies that $\mathbb{E}[Y]=0$. Then, shifting $Z\sim\mu$ by a deterministic vector does not change the value of $\mathbb{E}[\left<Y,Z\right$>], so we can assume without loss of generality that $\mathbb{E}[Z]=0$. Let us assume that $P_{YZ}$ is rational; the general case will then following by approximation. Define
$\mathcal{I}:=\{N\geq 1\colon\textrm{$NP_{YZ}$ is integer}\}.$ (65)
For each $N\in\mathcal{I}$, let $\mathcal{C}_{\mu}$ denote the $\mu$-type set, consisting of all sequences of type $\mu=P_{Z}$. Select
$L:=\exp(NI(Y;Z)+N^{1/3})$ (66)
sequences from $\mathcal{C}_{\mu}$ uniformly at random with replacement, and call this random set $\mathcal{A}$. Let $Y^{N}\sim P_{Y}^{\otimes N}$. Then we will show that
$\mathbb{E}[\left<Y,Z\right>]\leq\liminf_{N\to\infty}\mathbb{E}[\sup_{z^{N}\in\mathcal{A}}\frac{1}{N}\left<Y^{N},z^{N}\right$>], (67)
where the limit is for $N\in\mathcal{I}$. Indeed, set
$\mathcal{B}$ $:=\{(y^{N})\colon\widehat{P}_{y^{N}}=P_{Y}\};$ (68)
$\mathcal{D}$ $:=\{(y^{N},z^{N})\colon\widehat{P}_{y^{N}z^{N}}=P_{YZ}\},$ (69)
and for any $y^{N}\in\mathcal{B}$, define $\mathcal{D}(y^{N}):=\{z^{N}\colon(y^{N},z^{N})\in\mathcal{D}\}$. A standard result states that (see Lemma 3, or (Csiszár and Körner,, 2011, Chapter 1)):
$\frac{1}{N}\ln|\mathcal{D}(y^{N})|=H(Z|Y)+O(\frac{\ln N}{N}),\quad\forall y^{N}\in\mathcal{B}.$ (70)
Let $\pi$ be the projection from $\mathbb{R}^{nN}$ to $\mathcal{B}$ (i.e., map to the closest point in $\mathcal{B}$, with ties broken arbitrarily). Define $\mathcal{E}$ as the event that $\mathcal{D}(\pi(Y^{N}))\cap\mathcal{A}=\emptyset$. Then, (70) implies that
$\mathbb{P}[\mathcal{E}]=\left(1-\exp(-NI(Y;Z)+O(\ln N))\right)^{L}\leq\exp(-CN^{1/3})$ (71)
for some $C>0$ depending only on $P_{YZ}$, and for all $N\in\mathcal{I}$. From the definitions it is clear that
$\max_{z^{N}\in\mathcal{A}}\frac{1}{N}\left<\pi(Y^{N}),z^{N}\right>1_{\mathcal{E}^{c}}\geq\mathbb{E}[\left<Y,Z\right>]1_{\mathcal{E}^{c}}$ (72)
for all $N\in\mathcal{I}$ almost surely. Since $\mathbb{E}[Z]=0$ is assumed, we see that 0 belongs to the convex hull of $\mathcal{C}_{\mu}$, therefore
$\max_{z^{N}\in\mathcal{A}}\frac{1}{N}\left<\pi(Y^{N}),z^{N}\right>\geq\max_{z^{N}\in\mathcal{A}}\frac{1}{N}\left<\pi(Y^{N}),z^{N}\right>1_{\mathcal{E}^{c}}$ (73)
almost surely. Furthermore,
$\max_{z^{N}\in\mathcal{A}}\frac{1}{N}\left<Y^{N},z^{N}\right>\geq\max_{z^{N}\in\mathcal{A}}\frac{1}{N}\left<\pi(Y^{N}),z^{N}\right>-\max_{z^{N}\in\mathcal{A}}\frac{1}{N}\left<\pi(Y^{N})-Y^{N},z^{N}\right$> (74)
From Lemma 5, we have
$\max_{z^{N}\in\mathcal{A}}\frac{1}{N}\left<\pi(Y^{N})-Y^{N},z^{N}\right$> $\leq\frac{1}{\sqrt{N}}\|\pi(Y^{N})-Y^{N}\|_{2}\cdot\frac{1}{\sqrt{N}}\max_{z^{N}\in\mathcal{A}}\|z^{N}\|_{2}$ (75)
$\leq W_{2}(\mu,\widehat{P}_{Y^{N}})\cdot O(1)$ (76)
which converges to 0 in expectation as $N\in\mathcal{I}$ tends to infinity. Thus (67) is verified by collecting the bounds above.
It remains to apply the Dudley inequality to the right side of (67). Since $\mathcal{A}$ is contained in the set $\mathcal{C}$ in Lemma 4, the upper bound part of Lemma 4 also applies to covering numbers of $\mathcal{A}$. Furthermore, we have the trivial bound $|\mathcal{A}|\leq L$. Therefore,
$\limsup_{N\to\infty}\frac{1}{N}\ln N(\mathcal{A},2\sqrt{N}\sigma)\leq R_{\mu}(\sigma^{2})\wedge I(Y;Z).$ (77)
The claim then follows by applying dominated convergence (see proof of Theorem 1). ∎
4.2 Beyond the Finite Case
In this section we extend the proof of Theorem 1 for finite $T$ to the countable case by a limiting argument, and also prove Proposition 1. We first prove an auxiliary result about absolute integrability of $|X_{Z}|$, which will be used later to apply dominated convergence.
Lemma 6.
Suppose that $(X_{t})_{t\in T}$ is a zero-mean sub-Gaussian process on a countable set $T$. Let $Z\sim\mu$ be a random variable on $T$, for which $\sigma_{\rm m}:=\sigma_{\rm m}(\mu)<\infty$.
  • Consider $\bar{T}:=T\times\{-1,1\}$, and a process $(\bar{X}_{\bar{t}})_{\bar{t}\in\bar{T}}$, where $\bar{X}_{\bar{t}}:=sX_{t}$ for $\bar{t}=(t,s)$. Let $\bar{d}$ be the natural metric for this sub-Gaussian process on $\bar{T}$, and $\bar{Z}:=(Z,S)$ be a random variable on $\bar{T}$, satisfying $P_{Z}=\mu$. Define $\bar{\sigma}_{\rm m}^{2}:=\inf_{\bar{t}\in\bar{T}}\mathbb{E}[\bar{d}^{2}(\bar{Z},\bar{t})]$, and let $\bar{R}_{P_{\bar{Z}}}(\cdot)$ be the rate-distortion function for $\bar{Z}$ on $(\bar{T},\bar{d})$. Then
    $\int_{0}^{\bar{\sigma}_{\rm m}}\sqrt{\bar{R}_{P_{\bar{Z}}}(\sigma^{2})}d\sigma\leq\int_{0}^{\sigma_{\rm m}}\sqrt{R_{Z}(\sigma^{2})}d\sigma+\sqrt{2\ln 2}\sigma_{\rm m}.$ (78)
  • If $\int_{0}^{\sigma_{\rm m}}\sqrt{R_{\mu}(\sigma^{2})}d\sigma<\infty$, then $\mathbb{E}[|X_{Z}|]<\infty$.
It is interesting to compare Lemma 6 with the classical observation that for symmetric processes, $\mathbb{E}[\sup_{t\in T}|X_{t}|]<\infty$ is equivalent to $\mathbb{E}[\sup_{t\in T}X_{t}]<\infty$, both equivalent to boundedness of the process in the Gaussian case; see (Talagrand,, 1987, p2)(Talagrand,, 2014, Exercise 2.2.2.) and the references therein.
Proof.
We first prove (78). Clearly, the natural metric for the new process satisfies $\bar{d}((t,s),(t^{\prime},s))=d(t,t^{\prime})$, for any $s\in\{-1,+1\}$ and $t,t^{\prime}\in T$. Now suppose that $P_{\bar{Z}}=P_{Z,S}$ is an arbitrary distribution on $\bar{T}$. We can assume without loss of generality that there is $t_{0}\in T$ satisfying $X_{t_{0}}\equiv 0$ and $\sigma_{\rm m}=\sqrt{\mathbb{E}[d^{2}(t_{0},Z)]}$. By the triangle inequality we can show that $\bar{d}(\bar{Z},(t_{0},+1))\leq d(Z,t_{0})+\|2X_{t_{0}}\|_{\Psi_{2}}=d(Z,t_{0})$, implying that $\bar{\sigma}_{\rm m}^{2}\leq\mathbb{E}[\bar{d}^{2}(\bar{Z},(t_{0},+1))]<\sigma_{\rm m}^{2}$. In addition, suppose that $U$ is any random variable such that $R_{P_{Z}}(\sigma^{2})=I(U;Z)$ and $\mathbb{E}[d^{2}(U,Z)]\leq\sigma^{2}$ for some $\sigma^{2}$. If we set $\bar{U}=(U,S)$, we have $\mathbb{E}[\bar{d}^{2}(\bar{Z},\bar{U})]=\mathbb{E}[d^{2}(Z,U)]\leq\sigma^{2}$, and moreover, the rate-distortion function on $\bar{T}$ satisfies $\bar{R}_{P_{\bar{Z}}}(\sigma^{2})\leq I(\bar{U};\bar{Z})=I(U,S;Z,S)=I(U;Z)+I(S;Z|U)+H(S|Z)\leq R(\sigma^{2})+2\ln 2$. Then (78) follows by the sub-additivity of the square root function.
It remains to prove the second claim. Suppose that $T$ is enumerated as $t_{0},t_{1},\dots$, assuming without loss of generality $X_{t_{0}}\equiv 0$, $\sigma_{\rm m}=\sqrt{\mathbb{E}[d^{2}(t_{0},Z)]}$, and write $T_{k}:=\{t_{0},\dots,t_{k}\}$. Given $Z=Z(\omega)$, define $Z_{k}=Z$ if $Z\in T_{k}$ and $Z_{k}=t_{0}$ otherwise. We can show that
$\sqrt{R_{P_{Z_{k}}}(\sigma^{2})}\leq\sqrt{R_{P_{Z}}(\sigma^{2})+2\ln 2}\leq\sqrt{R_{P_{Z}}(\sigma^{2})}+\sqrt{2\ln 2}.$ (79)
Indeed, if $P_{ZU}$ is an optimal distribution in the definition of $R_{P_{Z}}(\sigma^{2})$, then setting $E=1\{Z\notin T_{k}\}$ and $U_{k}=U$ if $E=0$, $U_{k}=t_{0}$ otherwise, we have
$d(U_{k},Z_{k})\leq d(U,Z),\quad\textrm{a.s.}$ (80)
and
$I(U_{k};Z_{k})$ $\leq I(U_{k},E;Z_{k})$ (81)
$=I(E;Z_{k})+I(U_{k};Z_{k}|E)$ (82)
$\leq H(E)+I(U_{k};Z_{k}|E)$ (83)
$\leq H(E)+I(U;Z|E)$ (84)
$\leq H(E)+I(UE;Z)$ (85)
$\leq 2H(E)+I(U;Z),$ (86)
establishing (79). Now applying the result of Theorem 1 in the finite case, we have
$\mathbb{E}[|X_{Z_{k}}|]$ $=\mathbb{E}[{\rm sign}(X_{Z_{k}})X_{Z_{k}}]$ (87)
$\leq 48\int_{0}^{\bar{\sigma}_{\rm m}}\sqrt{\bar{R}_{P_{\bar{Z}_{k}}}(\sigma^{2})}d\sigma$ (88)
$\leq 48\int_{0}^{\sigma_{\rm m}}\sqrt{R_{P_{Z_{k}}}(\sigma^{2})}d\sigma+\sqrt{2\ln 2}\sigma_{\rm m},$ (89)
where we defined $\bar{Z}_{k}:=(Z_{k},{\rm sign}(X_{Z_{k}}))$, and (89) follows from (78) by replacing the arbitrary $P_{Z}$ with $P_{Z_{k}}$ and noting that $\sigma_{\rm m}(P_{Z_{k}})\leq\sigma_{\rm m}(P_{Z})=:\sigma_{\rm m}$. Thus (79) and (89) imply that $\lim_{k\to\infty}\mathbb{E}[|X_{Z_{k}}|]<\infty$. The proof is completed since $\lim_{k\to\infty}|X_{Z_{k}}|=|X_{Z}|$ almost surely, and Fatou’s lemma shows $\mathbb{E}[|X_{Z}|]\leq\lim_{k\to\infty}\mathbb{E}[|X_{Z_{k}}|]$. ∎
Proof of Theorem 1.
At this point, it only remains to extend the result of Theorem 1 from the finite case to the countable case. We can assume that the integral on the right side of (33) is finite, since otherwise there is nothing to prove. Again, assume without loss of generality that there is $t_{0}\in T$ satisfying $X_{t_{0}}\equiv 0$ and $\sigma_{\rm m}:=\sigma_{\rm m}(P_{Z})=\sqrt{\mathbb{E}[d^{2}(t_{0},Z)]}$. Construct $Z_{k}$ as in the proof of Lemma 6. Lemma 6 implies $\mathbb{E}[|X_{Z}|]<\infty$, and since $|X_{Z_{k}}|\leq|X_{Z}|\vee|X_{t_{0}}|$, dominated convergence implies
$\lim_{k\to\infty}\mathbb{E}[X_{Z_{k}}]=\mathbb{E}[\lim_{k\to\infty}X_{Z_{k}}]=\mathbb{E}[X_{Z}].$ (90)
From the definition of $Z_{k}$, we see that $\sigma_{\rm m}(P_{Z_{k}})\leq\sigma_{\rm m}(P_{Z})=:\sigma_{\rm m}$. Hence $R_{P_{Z_{k}}}(\sigma^{2})=0$ for all $\sigma\geq\sigma_{\rm m}$. Then from the estimate in (79), we see that $\sqrt{R_{P_{Z_{k}}}(\sigma^{2})}$ is dominated by the absolutely integrable function $(\sqrt{R_{P_{Z}}(\sigma^{2})}+\sqrt{2\ln 2})1_{\sigma\leq\sigma_{\rm m}}$. Furthermore, since $P(E)$ converges 0 as $k\to\infty$, (86) implies that $\lim_{k\to\infty}R_{P_{Z_{k}}}(\sigma^{2})=R_{P_{Z}}(\sigma^{2})=0$ for any $\sigma>0$. Therefore dominated convergence shows
$\lim_{k\to\infty}\int_{0}^{+\infty}\sqrt{R_{P_{Z_{k}}}(\sigma^{2})}d\sigma=\int_{0}^{\sigma_{\rm m}}\sqrt{R_{P_{Z}}(\sigma^{2})}d\sigma.$ (91)
Then the claim of the lemma follows by (90), (91) and applying the result of Theorem 1 in the finite setting (proved in Section 4.1).
Proof of Proposition 1.
Let $T_{0}$ be the countable subset in the definition of separability. For any $\epsilon>0$, let $Z_{\epsilon}\in T_{0}$ be a random variable such that $|X_{Z}-X_{Z_{\epsilon}}|\leq\epsilon$ and $d(Z,Z_{\epsilon})\leq\epsilon$, which is possible by the separability and continuity assumption. Then $\mathbb{E}[X_{Z}]\leq\mathbb{E}[X_{Z_{\epsilon}}]+\epsilon$, and $W_{\infty}(\mu,\mu^{\prime})\leq\epsilon$, if $\mu^{\prime}$ is the distribution of $Z_{\epsilon}$. Therefore the claim follows by the result for the countable case and taking $\epsilon\downarrow 0$. ∎
5 Sharpness of the Rate-Distortion Integral
We provide several proofs of Theorem 3 (up to constant factors) in this section. The proof in Section 5.3 gives the constant stated in Theorem 3.
5.1 Proof via the Majorizing Measure Theorem
We first give a proof of the lower bound direction assuming a version of the majorizing measure theorem, i.e. the $\delta_{2}$ functional defined in (15) lower bounds the supremum of a Gaussian process. It is relatively easy (compared to the proof of the majorizing measure theorem) to show the equivalence of $\delta_{2}$ and $\gamma_{2}$ up to universal constants, and the direction of $\gamma_{2}\gtrsim\delta_{2}$ can be found in (Talagrand,, 2014, Proposition 3.3.1 & Proposition 3.4.1).
We first consider the case where $n:=|T|<\infty$. Recall the representation of $X_{z}$ as an inner product in (55), which embeds $Y,Z$ in $\mathbb{R}^{n}$, where $Y$ is a normal vector, and recall the typical set $\mathcal{C}$ defined in (54). Define
$\mathcal{D}:=\{y^{N}\in\mathbb{R}^{nN}\colon W_{2}(P_{Y},\widehat{P}_{y^{N}})\leq N^{-\frac{1}{2n}}\}$ (92)
Then by the Wasserstein law of large numbers, $\mathbb{P}[Y^{N}\in\mathcal{D}]=1-o(1)$.
Given $P_{Z}=\mu$, let $P_{YZ}$ be the optimal coupling of $P_{Y}$ and $\mu$ maximizing $\mathbb{E}[\left<Y,Z\right$>] (equivalently, the optimal transport map). Then
$\mathbb{E}[X_{Z}]$ $\geq\lim_{N\to\infty}\frac{1}{N}\mathbb{E}[\sup_{z^{N}\in\mathcal{C}}\left<Y^{N},z^{N}\right>1_{Y^{N}\in\mathcal{D}}]$ (93)
$=\lim_{N\to\infty}\frac{1}{N}\mathbb{E}[\sup_{z^{N}\in\mathcal{C}}\left<Y^{N},z^{N}\right$>] (94)
$\gtrsim\lim_{N\to\infty}\frac{1}{N}\delta_{2}(\mathcal{C})$ (95)
$\geq\lim_{N\to\infty}\frac{1}{N}\inf_{t\in\mathcal{C}}I_{\mu^{\otimes N}}(t).$ (96)
The above steps are justified below: to show (93), note that for any $y^{N}\in\mathcal{D}$, we have
$\lim_{N\to\infty}\frac{1}{N}\sup_{z^{N}\in\mathcal{C}}\left<y^{N},z^{N}\right$> $\leq\lim_{\epsilon\to 0}\sup_{P_{Y^{\prime}Z^{\prime}}}\mathbb{E}[\left<Y^{\prime},Z^{\prime}\right$>] (97)
$=\lim_{\epsilon\to 0}\sup_{P_{Y^{\prime}Z^{\prime}}}\frac{\mathbb{E}\|Y^{\prime}\|_{2}^{2}+\mathbb{E}\|Z^{\prime}\|_{2}^{2}-W_{2}^{2}(P_{Y^{\prime}},P_{Z^{\prime}})}{2}$ (98)
$=\frac{\mathbb{E}\|Y\|_{2}^{2}+\mathbb{E}\|Z\|_{2}^{2}-W_{2}^{2}(P_{Y},P_{Z})}{2}$ (99)
$=\mathbb{E}[\left<Y,Z\right$>] (100)
where in (97) the sup is over $P_{Y^{\prime}Z^{\prime}}$ for which $W_{2}(P_{Y},P_{Y^{\prime}})\leq\epsilon$ and $W_{2}(P_{Z},P_{Z^{\prime}})\leq\epsilon$. This shows that the right side of (93) is upper bounded by $\mathbb{E}[\left<Y,Z\right>]\cdot\lim_{N\to\infty}\mathbb{E}[1\{Y^{N}\in\mathcal{D}\}]=\mathbb{E}[\left<Y,Z\right$>]. The step (94) follows from $\mathbb{E}[\sup_{z^{N}\in\mathcal{C}}\left<Y^{N},z^{N}\right>1_{Y^{N}\in\mathcal{D}^{c}}]=o(N)$, whose proof is the same as the proof that the second expectation in (58) is $o(N)$. (95) follows since $\delta_{2}(\mathcal{C})\asymp\gamma_{2}(\mathcal{C})$ lower bounds (94) according to the majorizing measure theorem (Talagrand,, 2014). We see (96) by specifying the supremum in (15) as $\mu^{\otimes N}$.
Using the estimate in Lemma 4, we obtain that for any $t\in\mathcal{C}$,
$\liminf_{N\to\infty}\frac{1}{N}I_{\mu^{\otimes N}}(t)\gtrsim\int_{0}^{\sigma_{\rm m}}\sqrt{R_{\mu}(\sigma^{2})}d\sigma$ (101)
by using Fatou’s lemma. This proves Theorem 3 for finite $T$ (except for estimating the constant factor).
We now use a limiting argument to extend the result to the case of $\mu$ supported on a countable $T$. Assume that $T=\{t_{0},t_{1},\dots\}$. Let $T_{k}:=\{t_{0},t_{1},\dots,t_{k}\}$, $\mu_{k}(t_{i}):=\frac{\mu(t_{i})}{\mu(T_{k})}1\{i\leq k\}$, and $\tilde{\mu}_{k}(t_{i}):=\frac{\mu(t_{i})}{1-\mu(T_{k})}1\{i>k\}$. If $Z_{k}(\omega)$ maximizes $\mathbb{E}[X_{Z_{k}}]$ subject to $P_{Z_{k}}=\mu_{k}$, we can construct a Bernoulli random variable $E\sim{\rm Bern}(\mu(T_{k}))$ independent of $\omega$, and let $Z=Z_{k}$ when $E=1$, and sample $Z\sim\tilde{\mu}_{k}$ independently of $\omega$ when $E=0$. Then $Z\sim\mu$, and
$\mathbb{E}[X_{Z}]=\mu(T_{k})\mathbb{E}[X_{Z_{k}}].$ (102)
Moreover, if $P_{U_{k}|Z_{k}}$ achieves the infimum in the definition of $R_{\mu_{k}}(\sigma^{2})$, we can construct $P_{U|Z}$ by setting $P_{U|Z}(|z)=P_{U_{k}|Z_{k}}(|z)$ for $z\in T_{k}$ and $U=t_{0}$ if $z\notin T_{k}$. Assume without loss of generality that $\mathbb{E}_{\mu}[d^{2}(t_{0},Z)]=\sigma_{\rm m}^{2}$. Then $\mathbb{E}[d^{2}(U,Z)]=\sigma^{2}\mu(T_{k})+\mathbb{E}[d^{2}(t_{0},Z)1_{Z\in T_{k}^{c}}]$. Since $\lim_{k\to\infty}\mathbb{P}[T_{k}^{c}]=0$ and $\mathbb{E}[d^{2}(t_{0},Z)]<\infty$, a basic result from measure theory (which can be proved by dominated convergence) states that $\lim_{k\to\infty}\mathbb{E}[d^{2}(t_{0},Z)1_{Z\in T_{k}^{c}}]=0$. In particular, for any given $\sigma>0$, there exists $k(\sigma)>0$ such that for all $k>k(\sigma)$, this construction of $P_{U|Z}$ ensures that $\mathbb{E}[d^{2}(U,Z)]<2\sigma^{2}$. Moreover, if $E_{k}:=1\{Z\in T_{k}\}$, we have $I(U;Z)\leq I(U;Z|E)+H(E)=R_{\mu_{k}}(\sigma^{2})+h(\mu(T_{k}))$, where $h(\cdot)$ is the binary entropy function. Taking $k\to\infty$ we have $\mu(T_{k})\to 1$ and obtain
$\lim_{k\to\infty}R_{\mu_{k}}(\sigma^{2})\geq R_{\mu}(2\sigma^{2})$ (103)
for any $\sigma>0$. Then using (102) we obtain
$\mathbb{E}[X_{Z}]\geq\lim_{k\to\infty}\mathbb{E}[X_{Z_{k}}]\gtrsim\lim_{k\to\infty}\int_{0}^{\infty}\sqrt{R_{\mu_{k}}(\sigma^{2})}d\sigma\gtrsim\int_{0}^{\infty}\sqrt{R_{\mu}(\epsilon^{2})}d\epsilon$ (104)
by Fatou’s lemma and change of variables.
5.2 Proof via Sharpness of Dudley’s Integral
The second proof does not use the majorizing measure theorem, but merely the sharpness of the Dudley’s Integral for stationary processes (10), which was due to Fernique (Fernique,, 1975) a decade before the majorizing measure theorem. Recall (94) that $\mathbb{E}[X_{Z}]\geq\lim_{N\to\infty}\frac{1}{N}\mathbb{E}[\sup_{z^{N}\in\mathcal{C}}\left<Y^{N},z^{N}\right$>], which is further lower bounded (up to a constant) by $\lim_{N\to\infty}\frac{1}{N}\int_{0}^{\infty}\sqrt{\ln N(\mathcal{C},\sqrt{N}\sigma)}\sqrt{N}d\sigma$. In turn, it is further lower bounded by $\int_{0}^{\infty}\sqrt{R_{\mu}(\epsilon^{2})}d\epsilon$, by invoking Lemma 4 and Fatou’s Lemma.
5.3 Proof via Super-Sudakov
In this section, we give another proof of Theorem 3 which will yield the constant stated therein. Traditionally, the majorizing measure theorem was proved by a sophisticated generic chaining argument (Talagrand,, 1987, 2014), and in the special case of stationary processes by a simpler reverse chaining argument based on a “super Sudakov” theorem (Van Handel,, 2014; Talagrand,, 1992). Here, we derive an information-theoretic counterpart of the super-Sudakov theorem, and then obtains Theorem 3 using purely analytic arguments, without any combinatorial constructions like packing.
Suppose that $\mu$, $\gamma$ are distributions on $\mathbb{R}^{n}$, and $\epsilon>0$. Define
$\phi(\epsilon,\mu,\gamma):=\sup_{P_{ZZ^{*}G}}\mathbb{E}[\left<G,Z\right$>] (105)
where the supremum is over $P_{ZZ^{*}G}$ satisfying $P_{GZ^{*}}=\gamma\times\mu$, $P_{Z}=\mu$, and $\mathbb{E}[\|Z-Z^{*}\|_{2}^{2}]\leq\epsilon^{2}$. The parameter $\epsilon$ can be thought of as an analogue of the scale in the reverse chaining argument. The geometric meaning of $\phi(\epsilon,\mu,\gamma)$ is shown by the following:
Lemma 7.
Consider $\phi$ defined in (105), where $\mu$ is rational, and $\gamma$ has finite second moment. For each $N$ such that $N\mu$ is integer, define
$T$ $:=\{z^{N}\in\mathbb{R}^{nN}\colon\widehat{P}_{z^{N}}=\mu\};$ (106)
$\mathcal{C}$ $:=\{z^{N}\in T\colon\|z^{N}-z^{*N}\|_{2}\leq\epsilon\sqrt{N}\}.$ (107)
Let $z^{*N}=(z_{1}^{*},\dots,z_{N}^{*})$ be an arbitrary deterministic sequence in $T$. Let $G^{N}\sim\gamma^{\otimes n}$. Then for any $\epsilon>0$, we have
$\lim_{N\to\infty}\frac{1}{N}\mathbb{E}[\sup_{z^{N}\in\mathcal{C}}\left<G^{N},z^{N}\right>]=\phi(\epsilon,\mu,\gamma)$ (108)
where the limit is along the subsequence for which $N\mu$ is integer.
The proof involves constructing coupling distributions and approximation by types, which uses the following standard result about the transportation polytope:
Lemma 8.
Let $A\in\mathbb{R}_{\geq 0}^{m\times n}$ have integer row sums and integer column sums. Then there exists $B\in\mathbb{Z}_{\geq 0}^{m\times n}$ with the same row and column sums as $A$ and
$\max_{i,j}\,|B_{ij}-A_{ij}|\;\leq\;1.$
Proof.
Write $F:=A-\lfloor A\rfloor$, so $F_{ij}\in[0,1)$ for all $i,j$. Let $r_{i}:=\sum_{j}F_{ij}$ and $c_{j}:=\sum_{i}F_{ij}$. Because the row and column sums of $A$ are integers and $\sum_{j}\lfloor A_{ij}\rfloor,\sum_{i}\lfloor A_{ij}\rfloor$ are integers, we have $r_{i},c_{j}\in\mathbb{Z}$.
Form the bipartite graph $G$ with left vertices $\{1,\dots,m\}$ (rows) and right vertices $\{1,\dots,n\}$ (columns), and put an edge $(i,j)$ iff $F_{ij}\in(0,1)$. If some connected component of $G$ were a tree, it would have a leaf $v$ incident to a unique edge $e$ of weight $F_{e}\in(0,1)$, forcing the corresponding $r_{i}$ or $c_{j}$ to equal $F_{e}\notin\mathbb{Z}$, a contradiction. Hence every component of $G$ contains a cycle.
Fix a simple cycle $C$ in $G$. Along the edges of $C$, alternately add $\varepsilon$ and $\varepsilon$ to the corresponding entries of $F$ (leaving all other entries unchanged), where
$\varepsilon\;:=\;\min\{\,F_{ij},\,1-F_{ij}:(i,j)\in C\,\}.$
This preserves all row and column sums, since at each vertex on $C$ the two incident edges change by $\varepsilon$ and $\varepsilon$. Moreover, all modified entries remain in $[0,1]$, and at least one becomes 0 or $1$. Remove edges for which $F_{ij}$ becomes 0 or 1. Iterating this operation, which terminates in finitely many steps, we obtain a modified matrix $B$ with the desired properties. ∎
Proof Lemma 7.
Define $\mathcal{I}:=\{N\in\mathbb{N}\colon N\mu\textrm{ is integer}\}$, where $\mathbb{N}$ denotes the set of natural numbers. Since the integers $\mathbb{Z}$ form a principal ideal domain, there exists a natural number $n_{0}$ such that $\mathcal{I}=n_{0}\mathbb{N}$. Since $\gamma$ has finite second moment, for each $N=n_{0}k$, $k\in\mathbb{N}$, we can find a distribution $\bar{\gamma}_{N}$ on $\mathbb{R}^{n}$ satisfying the following properties: 1) $k\bar{\gamma}_{N}$ is integer, 2) $W_{2}(\gamma,\bar{\gamma}_{N})=o(1)$, 3) ${\rm diam}(\operatorname*{supp}(\bar{\gamma}_{N}))\leq N^{a}$, where $a:=\frac{1}{4+4n}$, 4) Cardinality of the support $|\operatorname*{supp}(\bar{\gamma}_{N})|\leq N^{2an}$. Note that property 1) implies that $N\bar{\gamma}_{N}\times\mu$ is integer, The purpose of $\bar{\gamma}_{N}$ is to approximate $\gamma$ which might not be rational. Define
$\mathcal{D}:=\{g^{N}\colon\widehat{P}_{g^{N}z^{*N}}=\bar{\gamma}_{N}\times\mu\},$ (109)
which is nonempty. Let $\pi$ be the projection from $\mathbb{R}^{nN}$ to $\mathcal{D}$. More precisely, it sends any point to a closest point in $\mathcal{D}$ under the $\ell_{2}$ distance, and ties are broken so that $\pi$ is invariant under all $z^{*N}$-invariant permutations. This makes $\pi(G^{N})$ uniformly distributed on $\mathcal{D}$, and
[…]
$\leq\mathbb{E}\left[\|\pi(G^{N})-G^{N}\|_{2}\sup_{z^{N}\in\mathcal{C}}\|z^{N}\|_{2}\right]$ (111)
$=o(N),$ (112)
where the last step follows from the fact that $N^{-1/2}\|\pi(G^{N})-G^{N}\|_{2}\leq W_{2}(\bar{\gamma}_{N},\gamma)+W_{2}(\gamma,\widehat{P}_{G^{N}})$, due to Lemma 5, which vanishes in expected value as $N\to\infty$. Moreover, for any given $G^{N}$ and $z^{N}\in\mathcal{C}$, we have $\frac{1}{N}\left<\pi(G^{N}),z^{N}\right>]\leq\phi(\epsilon,\mu,\bar{\gamma}_{N})$ because $\widehat{P}_{z^{N}z^{*N}\pi(G^{N})}$ satisfies the constraints on $P_{ZZ^{*}G}$ in the definition of $\phi(\epsilon,\mu,\bar{\gamma}_{N})$ (cf. (105)). By property 2) of $\bar{\gamma}_{N}$, we also see that
$\lim_{N\to\infty}\phi(\epsilon,\mu,\bar{\gamma}_{N})=\phi(\epsilon,\mu,\gamma).$ (113)
Then the $\leq$ part of (108) follows.
To show that $\geq$ part, suppose that $Q^{N}_{ZZ^{*}G}$ is an optimizer for the supremum in the definition of $\phi(\epsilon,\mu,\bar{\gamma}_{N})$ for some $N\in\mathcal{I}$. Note that $NQ^{N}_{Z^{*}G}$ and $NQ^{N}_{Z}$ are integer by the construction of $\bar{\gamma}$. By Lemma 8, we can find a probability measure $P^{N}_{ZZ^{*}G}$ such that i) $NP^{N}_{ZZ^{*}G}$ is integer, ii) $P^{N}_{Z^{*}G}=Q^{N}_{Z^{*}G}$ and $P^{N}_{Z}=Q^{N}_{Z}$, iii) $\max_{z,z^{*},g}|P^{N}_{ZZ^{*}G}(z,z^{*},g)-Q^{N}_{ZZ^{*}G}(z,z^{*},g)|\leq\frac{1}{N}$. Then the total variation distance $\|P^{N}_{ZZ^{*}G}-Q^{N}_{ZZ^{*}G}\|_{1}\lesssim\frac{1}{N}|\operatorname*{supp}(Q^{N}_{ZZ^{*}G})|\lesssim N^{2an-1}$. Hence there exists $P_{ZZ^{*}G\bar{Z}\bar{Z}^{*}\bar{G}}\in\Pi(P^{N}_{ZZ^{*}G},Q^{N}_{ZZ^{*}G})$ under which $\mathbb{P}[(Z,Z^{*},G)\neq(\bar{Z},\bar{Z}^{*},\bar{G})]\lesssim N^{2an-1}$, and
$\mathbb{E}[\|(Z,Z^{*},G)-(\bar{Z},\bar{Z}^{*},\bar{G})\|_{2}^{2}]$ $\leq 2{\rm diam}^{2}(\operatorname*{supp}(Q^{N}_{ZZ^{*}G}))N^{2an-1}$ (114)
$\lesssim N^{2a+2an-1}$ (115)
$=o(1).$ (116)
In particular, we have
$W_{2}(P^{N}_{ZZ^{*}G},Q^{N}_{ZZ^{*}G})=o(1).$ (117)
This implies that $\mathbb{E}_{P^{N}_{ZG}}[\left<G,Z\right>]=\phi(\epsilon,\mu,\bar{\gamma}_{N})+o(1)$ and $\mathbb{E}_{P^{N}_{ZZ^{*}}}[\|Z-Z^{*}\|_{2}^{2}]\leq\epsilon_{N}^{2}$ for some $\epsilon_{N}=\epsilon+o(1)$. Let $\mathcal{C}(\epsilon_{N}):=\{z^{N}\in T\colon\|z^{N}-z^{*N}\|_{2}\leq\epsilon_{N}\sqrt{N}\}$. For any given $G^{N}$, we can find $z^{N}\in\mathcal{C}(\epsilon_{N})$ such that $\frac{1}{N}\left<\pi(G^{N}),z^{N}\right>=\mathbb{E}_{P^{N}_{ZG}}[\left<G,Z\right$>]. Thus,
$\liminf_{N\to\infty}\frac{1}{N}\mathbb{E}[\max_{z^{N}\in\mathcal{C}(\epsilon_{N})}\left<\pi(G^{N}),z^{N}\right$>] $\geq\liminf_{N\to\infty}\mathbb{E}_{P^{N}_{ZG}}[\left<G,Z\right$>] (118)
$\geq\liminf_{N\to\infty}\phi(\epsilon,\mu,\bar{\gamma}_{N})$ (119)
$=\phi(\epsilon,\mu,\gamma)$ (120)
where the limits are along $N\in\mathcal{I}$. Since any sequence in $\mathcal{C}(\epsilon_{N})$ can be approximated by a sequence in $\mathcal{C}$ within $o(\sqrt{N})$ error under the Euclidean distance, we see that $\liminf_{N\to\infty}\frac{1}{N}\mathbb{E}[\max_{z^{N}\in\mathcal{C}}\left<\pi(G^{N}),z^{N}\right>]\geq\phi(\epsilon,\mu,\gamma)$. Then the $\geq$ part of (108) follows from (112). ∎
In the remainder of the section we consider $\phi$ defined in (105), where $\mu$ is finitely supported and $\gamma=\mathcal{N}(0,I_{n})$ (the extension to non-finite case can be achieved by the argument in Section 5.1). We write $\phi(\epsilon)=\phi(\epsilon,\mu,\gamma)$ when no confusion. We also define the shorthand notation
[…]
where we recall that $\Pi_{\epsilon}(\mu,\mu)$ denotes the set of couplings satisfying $\mathbb{E}[\|Z-\hat{Z}\|_{2}^{2}\leq\epsilon^{2}$.
Lemma 9.
For any $0<\epsilon_{2}\leq\frac{4\epsilon_{1}}{5}$, we have
$\phi(\epsilon_{1})\geq\frac{\epsilon_{2}}{4}\sqrt{2(i(\epsilon_{2})-i(\epsilon_{1}-\epsilon_{2}/4))}+\phi(\epsilon_{2}/4).$ (122)
The super-Sudakov inequality, originally proved by Talagrand (Talagrand,, 1992), can be established for rather general settings by combining Sudakov minoration and concentration. The following version for Gaussian processes with explicit constant can be found in Marcus and Rosen, (2006); Van Handel, (2014).
Lemma 10 (Super-Sudakov).
Let $\{X_{t}\}_{t\in T}$ be a Gaussian process on a metric space $(T,d)$, and $\mathcal{N}$ be an $\epsilon$-packing of $(T,d)$. For every $s\in N$, let $T_{s}:=\{t\in T\colon d(t,s)\leq\frac{\epsilon}{4}\}$. Then
$\mathbb{E}[\sup_{t\in T}X_{t}]\geq\frac{\epsilon}{4}\sqrt{2\ln|\mathcal{N}|}+\min_{s\in\mathcal{N}}\mathbb{E}[\sup_{t\in T_{s}}X_{t}].$ (123)
Proof of Lemma 9.
It suffices to consider the case of rational $\mu$, as the general case will then follow by approximation by ration $\mu$ in the Wasserstein distribution. Suppose that $N>0$ is an integer for which $N\mu$ is integer. Define $T$ as in (106), and fix arbitrary $z^{*N}\in T$. For any $\epsilon>0$, define
$T_{\epsilon}:=T\cap B(z^{*N},\sqrt{N}\epsilon).$ (124)
Now let $\mathcal{N}$ be a maximal $\epsilon_{2}$-packing of $T_{r}$, where $r:=\epsilon_{1}-\epsilon_{2}/4$. Maximality implies that every point in $T_{r}$ is within $\epsilon_{2}$ distance to a point in $\mathcal{N}$, hence
$|\mathcal{N}|\mu^{\otimes N}(B(\sqrt{N}\epsilon_{2}))\geq\mu^{\otimes N}(B(\sqrt{N}r)).$ (125)
By a standard type counting argument (Csiszár and Körner,, 2011, Chapter 1), this implies
$\frac{1}{N}\ln|\mathcal{N}|\geq i(\epsilon_{2})-i(r)+o(1),$ (126)
and $i(\epsilon_{2})-i(r)\geq 0$ by monotonicity of $i(\cdot)$. Moreover, for every $z^{N}\in\mathcal{N}$, we have $T\cap B(z^{N},\sqrt{N}\epsilon_{2}/4)\subseteq T_{\epsilon_{1}}$ since $r+\epsilon_{2}/4\leq\epsilon_{1}$. Then Lemma 10 implies (122) by taking $N\to\infty$. ∎
Lemma 11.
Let $\tau\in(0,\frac{4}{5})$. We have
$\sup_{P_{ZG}\in\Pi(\mu,\gamma)}\mathbb{E}[\left<G,Z\right>]\geq\frac{\sqrt{2}}{4\ln\frac{4}{\tau}}\cdot\frac{4-5\tau}{4-\tau}\int_{0}^{\infty}\sqrt{i(\epsilon)}d\epsilon.$ (127)
Proof.
From Lemma 9 we have
$\phi(\epsilon)-\phi(\tfrac{\tau\epsilon}{4})$ $\geq\frac{\tau\epsilon}{4}\sqrt{2(i(\tau\epsilon)-i((1-\tfrac{\tau}{4})\epsilon))}$ (128)
$\geq\frac{\tau\epsilon}{4}\sqrt{2i(\tau\epsilon)}-\frac{\tau\epsilon}{4}\sqrt{2i((1-\tfrac{\tau}{4})\epsilon)}.$ (129)
Let $\epsilon_{0}:=\sqrt{2\mathbb{E}_{\mu}[\|Z-\mathbb{E}_{\mu}[Z]\|_{2}^{2}]}$. Taking $\int_{0}^{\frac{4}{\tau}\epsilon_{0}}\frac{d\epsilon}{\epsilon}$ on both sides above and changing variables, we obtain
$\int_{\epsilon_{0}}^{\frac{4}{\tau}\epsilon_{0}}\phi(\epsilon)\frac{d\epsilon}{\epsilon}\geq\int_{0}^{4\epsilon_{0}}\frac{d\epsilon}{4}\sqrt{2i(\epsilon)}-\int_{0}^{\frac{4(1-\tfrac{\tau}{4})}{\tau}\epsilon_{0}}\frac{\tau d\epsilon}{4(1-\tfrac{\tau}{4})}\sqrt{2i(\epsilon)}.$ (130)
Since $i(\epsilon)=0$ and $\phi(\epsilon)=\phi(\epsilon_{0})$ when $\epsilon\geq\epsilon_{0}$, we further have
$\phi(\epsilon_{0})\ln\frac{4}{\tau}\geq\frac{\sqrt{2}}{4}\left(1-\frac{\tau}{1-\frac{\tau}{4}}\right)\int_{0}^{\infty}\sqrt{i(\epsilon)}d\epsilon$ (131)
and the claim follows. ∎
Finally, since $i(\epsilon)\geq R_{\mu}(\epsilon^{2})$ by (30), and $\phi(\epsilon_{0})=\sup_{P_{Z}=\mu}\mathbb{E}[X_{Z}]$, we establish Theorem 3.
6 New Proof of the Majorizing Measure Theorem
We prove Corollary 3 from Theorem 3 in this section. First, observe the following auxiliary result, which allows us to rewrite integrals in a “penalized optimization” form:
Lemma 12.
Suppose that $y\in(0,\infty)$ is a decreasing function of $x\in(0,\infty)$. Then
$2\int_{0}^{\infty}ydx\leq\int_{0}^{\infty}\min_{x}\{\alpha^{-2}x^{2}+y^{2}\}d\alpha\leq 4\int_{0}^{\infty}ydx$ (132)
Proof.
For any $\alpha\in(0,\infty)$, there is unique $x(\alpha)$ satisfying $\alpha=\frac{x(\alpha)}{y(x(\alpha))}$. Then $x$ and $y$ can be parameterized by $\alpha\in(0,\infty)$, and
$\int_{0}^{\infty}(\alpha^{-2}x^{2}+y^{2})d\alpha$ $=\int_{\alpha=0}^{\infty}2y^{2}\cdot\frac{ydx-xdy}{y^{2}}$ (133)
$=4\int_{0}^{\infty}ydx.$ (134)
The claim then follows since for any $\alpha$, $\frac{\min_{x\in(0,\infty)}\{\alpha^{-2}x^{2}+y^{2}(x)\}}{\alpha^{-2}x^{2}(\alpha)+y^{2}(x(\alpha))}\in[1/2,1]$. ∎
A similar idea (though in the form of series summation rather than integral) was previously used in proving equivalent forms of the majorizing measure theorem (Talagrand,, 1994). Here, we find Lemma 12 very convenient in forming information-theoretic functionals that are exactly convex-concave (Remark 1), and also simplifying evaluations of the rate-distortion integral in applications to ERM.
To prove Corollary 1, it suffices to consider the case of finite $T$, since the general case (where the lattice supremum in (38) is defined as $\sup_{T_{0}\subseteq T,|T_{0}|<\infty}\sup_{t\in T_{0}}$; see Talagrand, (2014)) will then follows by taking the limit. Then $T$ can be identified with points in $\mathbb{R}^{|T|}$ equipped with the $\ell_{2}$ metric. For $i(\epsilon)$ defined in (121), Lemma 12 implies
$4\sup_{\mu}\int_{0}^{\infty}\sqrt{i(\epsilon)}d\epsilon\geq\sup_{P_{Z}}\int_{0}^{\infty}\min_{P_{Z\hat{Z}}\in\Pi(P_{Z},P_{Z})}\{\alpha^{-2}\mathbb{E}[\|Z-\hat{Z}\|_{2}^{2}]+I(Z;\hat{Z})\}d\alpha$ (135)
where both supremums are over distributions on $T$ and the minimum is over couplings of $P_{Z}$ and $P_{Z}$. The right side of (135) is lower bounded by
$\sup_{P_{Z}}\inf_{\mu}\int_{0}^{\infty}\min_{P_{U|Z}}\{\alpha^{-2}\mathbb{E}[\|Z-U\|_{2}^{2}]+D(P_{U|Z}\|\mu|P_{Z})\}d\alpha,$ (136)
since equality would be achieved if $\inf_{\mu}$ were restricted to $\mu=P_{Z}$ and $\min_{P_{U|Z}}$ were restricted to $P_{U|Z}$ satisfying $P_{U|Z}P_{Z}\in\Pi(P_{Z},P_{Z})$. Since the relative entropy is jointly convex (i.e., $D(\frac{\nu_{1}+\nu_{2}}{2}\|\frac{\mu_{1}+\mu_{2}}{2})\leq\frac{1}{2}D(\nu_{1}\|\mu_{1})+\frac{1}{2}D(\nu_{2}\|\mu_{2})$), we can show that the integral in (136) is convex in $\mu$ and linear in $P_{Z}$. Then Sion’s minimax theorem implies that (136) equals
$\quad\inf_{\mu}\sup_{P_{Z}}\int_{0}^{\infty}\min_{P_{U|Z}}\{\alpha^{-2}\mathbb{E}[\|Z-U\|_{2}^{2}]+D(P_{U|Z}\|\mu|P_{Z})\}d\alpha$
$=\inf_{\mu}\sup_{z}\int_{0}^{\infty}\min_{P_{U}}\{\alpha^{-2}\mathbb{E}[\|z-U\|_{2}^{2}]+D(P_{U}\|\mu)\}d\alpha.$ (137)
Now fix $\mu$, $z$, $\alpha$, and $P_{U}$. Define $r=a\sqrt{\mathbb{E}[\|z-U\|^{2}]}$, where $U\sim P_{U}$ and $a>1$ will be specified later. We obtain, from the data processing inequality,
$D(P_{U}\|\mu)\geq d(P_{U}(B(z,r))\|\mu(B(z,r)))$ (138)
where $d(a\|b):=a\ln\frac{a}{b}+(1-a)\ln\frac{1-a}{1-b}$ denotes the binary relative entropy function. By the Markov inequality we have $P_{U}(B(z,r))\geq 1-a^{-2}$. Then we have
$d(P_{U}(B(z,r))\|\mu(B(z,r)))\geq(1-a^{-2})\ln\frac{1-a^{-2}}{\mu(B(z,r))}-a^{-2}\ln a^{2}$ (139)
because if $\mu(B(z,r))>1-a^{-2}$ then the right side of(139) is negative and the inequality is obviously true; if $\mu(B(z,r))\leq 1-a^{-2}$ then by monotonicity of the binary relative entropy function we have $d(P_{\hat{Z}}(B(z,r))\|\mu(B(z,r)))\geq d(1-a^{-2}\|\mu(B(z,r)))=(1-a^{-2})\ln\frac{1-a^{-2}}{\mu(B(z,r))}+a^{-2}\ln\frac{a^{-2}}{1-\mu(B(z,r))}$ and (139) also holds. Now the integral in (137) is lower bounded by
$\int_{0}^{\infty}\inf_{r>0}\left\{\frac{r^{2}}{a^{2}\alpha^{2}}+\left[(1-a^{-2})\ln\frac{1-a^{-2}}{\mu(B(z,r))}-a^{-2}\ln a^{2}\right]_{+}\right\}d\alpha$ (140)
which, by Lemma 12, is lower bounded by twice of
$\int_{0}^{{\rm diam}(T)}\sqrt{\left[(1-a^{-2})\ln\frac{1-a^{-2}}{\mu(B(z,r))}-a^{-2}\ln a^{2}\right]_{+}}\frac{dr}{a}$
$\geq\int_{0}^{{\rm diam}(T)}\sqrt{(1-a^{-2})\ln\frac{1}{\mu(B(z,r))}}\frac{dr}{a}-\int_{0}^{{\rm diam}(T)}\sqrt{h(a^{-2})}\frac{dr}{a}$ (141)
$=\frac{\sqrt{1-a^{-2}}}{a}I_{\mu}(z)-\frac{{\rm diam}(T)}{a}\sqrt{h(a^{-2})}$ (142)
where we used the basic inequality $\sqrt{[x-y]_{+}}\geq\sqrt{x}-\sqrt{y}$ for $x,y\geq 0$. Thus
$\frac{2}{c}\mathbb{E}[\sup_{z\in T}X_{z}]\geq\inf_{\mu}\sup_{z}\frac{\sqrt{1-a^{-2}}}{a}I_{\mu}(z)-\frac{{\rm diam}(T)}{a}\sqrt{h(a^{-2})}.$ (143)
It is easy to see, by considering two points in $T$, that $\mathbb{E}[\sup_{z\in T}X_{z}]\geq{\rm diam}(T)\mathbb{E}[G\vee 0]=\sqrt{\frac{1}{2\pi}}{\rm diam}(T)$, where $G\sim\mathcal{N}(0,1)$. Therefore,
$\left(\frac{2}{c}+\frac{\sqrt{2\pi h(a^{-2})}}{a}\right)\mathbb{E}[\sup_{z\in T}X_{z}]\geq\frac{\sqrt{1-a^{-2}}}{a}\inf_{\mu}\sup_{z}I_{\mu}(z)$ (144)
as desired.
Remark 1.
The minimax inequality step (136)-(137) crucially leverages Lemma 12 and the joint convexity of the relative entropy. Specifically, the integral in (136) is convex in $\mu$ and linear in $P_{Z}$, and this property is not true for the functional $\int_{0}^{\infty}\min_{P_{U|Z}\colon\mathbb{E}[\|Z-U\|_{2}^{2}\leq\epsilon^{2}}\sqrt{D(P_{U|Z}\|\mu|P_{Z})}d\epsilon$ before we applied Lemma 12. Moreover, had we not used the relative entropy approach, then the integral in (136) might be replaced by the functional $\int_{0}^{\infty}\min_{r_{Z}}\{\alpha^{-2}\mathbb{E}[r_{Z}^{2}]+\ln\frac{1}{\mu(B(Z,r_{Z}))}\}d\alpha$ (see the $\theta_{i}(T)$ functional defined around Proposition 4.3 in Talagrand, (1994)), where $P_{U|Z}$ is replaced by the conditional distribution of $\mu$ on the ball $B(Z,r_{Z})$, $r_{Z}$ being a function of $Z$. Again, this functional is not exactly convex in $\mu$, so the application of the minimax theorem is less immediate. A classical approach that has a similar role as (136)-(137) is the Fernique convexity argument, which loses a constant factor (Talagrand,, 2014).
7 Tail Bounds for the ERM
In this section, we prove the results in Section 3.3.
7.1 Under the Log-Sobolev Inequality
The goal of this section is to prove Theorem 4. For any $\epsilon>0$, define
$s(\epsilon,\mu;\gamma)=\sup_{\nu\colon D(\nu\|\mu)\leq\epsilon}\{\mathsf{w}(\nu;\gamma)+\sqrt{2\epsilon\mathbb{E}_{\nu}[\|Z\|^{2}]}\}.$ (145)
We may abbreviate $\mathsf{w}(\mu;\gamma)$ as $\mathsf{w}(\mu)$, or $s(\epsilon,\mu;\gamma)$ as $s(\epsilon,\mu)$ or $s(\epsilon)$, when there is no confusion.
Our key tool is the following observation, which can be viewed as an information-theoretic counterpart of Lemma 1.
Lemma 13.
Suppose that $\mu$ is a probability measure supported on a totally bounded set $T\subseteq\ell_{2}$. Let $\gamma$ be a probability measure on $\ell_{2}$ which has zero-mean, satisfies the log-Sobolev inequality (17), and such that $\mathbb{E}_{\gamma}[\|Y\|_{2}]<\infty$. For any $\lambda\geq 0$ and any $P_{YZ}$ such that $P_{Z}=\mu$ and $P_{Y}=\gamma$, we have
$\ln\mathbb{E}[\exp(\lambda\left<Y,Z\right>)]\leq\sup_{\epsilon>0}\{\lambda s(\epsilon,\mu;\gamma)-\epsilon\}.$ (146)
Proof.
We first suppose that $\mu$ is a rational distribution supported on at most $n$ points. For $N>0$ such that $N\mu$ is integer, let $S_{N}$ be the set of types of $N$-sequences over the $n$ points. Then $|S_{N}|\leq{n+N-1\choose n-1}$ which grows polynomially in $N$. For each $\nu\in S_{N}$, let $C_{\nu}$ be the type class of $\nu$. Let $\xi_{\nu}:=\max_{z^{N}\in C_{\nu}}\left<Y^{N},z^{N}\right$> and $\sigma_{\nu}^{2}:=\max_{z^{N}\in C_{\nu}}\mathbb{E}[\left<Y^{N},z^{N}\right>^{2}]=\max_{z^{N}\in C_{\nu}}\|z^{N}\|_{2}^{2}=N\mathbb{E}_{\nu}[\|Z\|_{2}^{2}]$. Using Lemma 1, we see that for each $\nu$ and $\epsilon>0$,
$\mathbb{P}[\xi_{\nu}-\mathbb{E}[\xi_{\nu}]>\sigma_{\nu}\sqrt{2N\epsilon}]\leq\exp(-N\epsilon).$ (147)
Therefore, setting $S_{N}^{\prime}:=\{\nu\in S_{N}\colon D(\nu\|\mu)\leq\epsilon\}$,
$\mathbb{P}\left[\bigcup_{\nu\in S_{N}^{\prime}}\left\{\xi_{\nu}-\mathbb{E}[\xi_{\nu}]>\sigma_{\nu}\sqrt{2N\epsilon}\right\}\right]\leq|S_{N}|\exp(-N\epsilon)$ (148)
which implies that
$\mathbb{P}\left[\max_{\nu\in S_{N}^{\prime}}\xi_{\nu}>Ns(\epsilon)\right]\leq|S_{N}|\exp(-N\epsilon).$ (149)
Let $(Y^{N},Z^{N})\sim P_{YZ}^{\otimes N}$, and let $\mathcal{E}$ be the set of $Y^{N}$ in the event in (149). Let $\tilde{\mathcal{E}}$ be the event that $D(\widehat{P}_{Z^{N}}\|\mu)>\epsilon$. Note that under $\tilde{\mathcal{E}}^{c}$, we have $\left<Y^{N},Z^{N}\right>\leq\max_{\nu\in S_{N}^{\prime}}\xi_{\nu}$. Then
$\mathbb{P}\left[\left<Y^{N},Z^{N}\right>>Ns(\epsilon)\right]$ $\leq\mathbb{P}[\tilde{\mathcal{E}}]+\mathbb{P}\left[\left<Y^{N},Z^{N}\right>>Ns(\epsilon)\cap\tilde{\mathcal{E}}^{c}\right]$ (150)
[…]
$\leq\mathbb{P}[\tilde{\mathcal{E}}$$\mathbb{P}\left$$\mathcal{E}\right$] (152)
$\leq 2(N+1)^{n}\exp(-N\epsilon)$ (153)
where we used Lemma 3 to estimate $\mathbb{P}[\tilde{\mathcal{E}}$]. Then
[…]
$\leq\exp(N\lambda s(0))+\int_{\epsilon=0}^{+\infty}\mathbb{P}[\left<Y^{N},Z^{N}\right>>Ns(\epsilon)]d\exp(\lambda Ns(\epsilon))$ (155)
$\leq\exp(N\lambda s(0))+2(N+1)^{n}\int_{\epsilon=0}^{+\infty}\exp(-N\epsilon)d\exp(\lambda Ns(\epsilon)).$ (156)
Since the support of $\mu$ is finite, we see that $s(\epsilon)\lesssim_{\mu}\sqrt{\epsilon}+1$, which implies that the integral in (156) is finite, and
$\lim_{N\to\infty}\frac{1}{N}\ln\int_{\epsilon=0}^{+\infty}\exp(-N\epsilon)d\exp(\lambda Ns(\epsilon))\leq\sup_{\epsilon>0}\{\lambda s(\epsilon)-\epsilon\}.$ (157)
Therefore, (156) shows that
$\ln\mathbb{E}[\exp(\lambda\left<Y,Z\right>)]=\lim_{N\to\infty}\frac{1}{N}\ln\mathbb{E}[\exp(\lambda\left<Y^{N},Z^{N}\right>)]\leq\sup_{\epsilon>0}\{\lambda s(\epsilon)-\epsilon\}.$ (158)
[…]
Finally, we remove the assumption of finite dimensionality. Denote by $Y_{\leq n}$ the truncation of $Y$ to the first $n$ coordinates, and let $Y_{>n}$ be the remaining coordinates. Denote by $\mu_{n}$ the distribution of $Y_{\leq n}$. Since $Y,Z\in\ell_{2}$, we have $\lim_{n\to\infty}\|Y_{>n}\|_{2}=0$ and $\lim_{n\to\infty}\|Z_{>n}\|_{2}=0$ almost surely. By the Cauchy-Schwarz inequality, we can thus show that $\lim_{n\to\infty}\left<Y_{>n},Z_{>n}\right>=0$ almost surely. Then by Fatou’s lemma we have
$\ln\mathbb{E}[\exp(\lambda\left<Y,Z\right>)]\leq\lim_{n\to\infty}\ln\mathbb{E}[\exp(\lambda\left<Y_{\leq n},Z_{\leq n}\right>)]\leq\lim_{n\to\infty}\sup_{\epsilon>0}\{\lambda s(\epsilon,\mu_{n};\gamma)-\epsilon\}.$ (159)
It remains to prove that
$\lim_{n\to\infty}\sup_{\epsilon>0}\{\lambda s(\epsilon,\mu_{n};\gamma)-\epsilon\}\leq\sup_{\epsilon>0}\{\lambda s(\epsilon,\mu;\gamma)-\epsilon\}.$ (160)
For given $n>0$, suppose that $\nu_{n}$ is a distribution achieving the supremum in the definition of $s(\epsilon,\mu_{n};\gamma)$. Then we can extend $\nu_{n}$ to a $\nu$ such that $D(\nu\|\mu)=D(\nu_{n}\|\mu_{n})$, by using the same channel from $Z_{\leq n}$ to $Z_{>n}$ under $\mu$. Suppose that $P_{Y_{\leq n}Z}$ achieves the supremum in the definition of $\mathsf{w}(\nu_{n};\gamma)$. Then we can extend it to a $P_{YZ}\in\Pi(\nu,\gamma)$. For $(Y,Z)\sim P_{YZ}$, we have
$\mathsf{w}(\nu;\gamma)$ $\geq\mathbb{E}[\left<Y,Z\right$>] (161)
$=\mathbb{E}[\left<Y_{\leq n},Z\right>]+\mathbb{E}[\left<Y_{>n},Z\right$>] (162)
$\geq\mathsf{w}(\nu_{n};\gamma)-\mathbb{E}[\|Y_{>n}\|_{2}\|Z_{>n}\|_{2}]$ (163)
$\geq\mathsf{w}(\nu_{n};\gamma)-\sup_{t\in T}\|t_{>n}\|_{2}\mathbb{E}[\|Y_{>n}\|_{2}]$ (164)
$=\mathsf{w}(\nu_{n};\gamma)+o(1).$ (165)
Moreover, we clearly have $\mathbb{E}_{\nu_{n}}[\|Z\|_{2}^{2}]\leq\mathbb{E}_{\nu}[\|Z\|_{2}^{2}]$. Thus (160) is true, and we have extended the result to totally bounded $T$. ∎
Lemma 14.
Suppose that $Q$ is a probability measure on $\mathbb{R}$, and $\psi(\lambda):=\ln\mathbb{E}_{X\sim Q}[\exp(\lambda X^{2})]$. Then for any $P$ absolutely continuous with respect to $Q$, we have $\mathbb{E}_{X\sim P}[X^{2}]\leq\inf_{\lambda>0}\frac{D(P\|Q)+\psi(\lambda)}{\lambda}$.
Proof.
It suffices to prove the case of $P=P_{\lambda}$ defined according to $\frac{dP_{\lambda}}{dQ}(x)\propto\exp(\lambda x^{2})$. Note that $\psi$ is a nonnegative convex function, and $\psi(0)=0$. Then the claim follows since $\mathbb{E}_{X\sim P_{\lambda}}[X^{2}]=\psi^{\prime}(\lambda)$ and $D(P_{\lambda}\|Q)=\lambda\psi^{\prime}(\lambda)-\psi(\lambda)$. ∎
Lemma 15.
Let $\mu$ be a probability measure with finite support on $T$ for which $\mathsf{G}$ is a width-moment function, and $Z\sim\mu$. Let $\psi(\lambda):=\ln\mathbb{E}[e^{\lambda\|Z\|_{2}^{2}}]$. Then
$\ln\mathbb{E}[\exp(\eta\left<Y,Z\right>)]\leq\sup_{\epsilon>0}\left\{-\epsilon+\eta\mathsf{G}\left(\frac{\epsilon+\psi(\lambda)}{\lambda}\right)+\eta\sqrt{2\epsilon\cdot\frac{\epsilon+\psi(\lambda)}{\lambda}}\right\}$ (166)
for any $\lambda,\eta>0$ and any $P_{YZ}$ for which $P_{Y}$ is standard normal and $P_{Z}=\mu$.
Proof.
The result follows from Lemma 13, Lemma 14, and the data processing inequality for the KL divergence. ∎
Proof of Theorem 4.
Recall that
$\|\hat{m}-m\|_{2}^{2}\leq\frac{1}{\sqrt{n}}\chi_{m-\hat{m}}$ (167)
where $\chi_{z}:=-2\left<z,w\right$>. We then have
[…]
Using Lemma 15 we have that the logarithm of the right side of (168) is upper bounded by
$\sup_{\epsilon>0}\left\{-\epsilon+\frac{2\lambda}{\sqrt{n}}\mathsf{G}\left(\frac{\epsilon+\psi(\lambda)}{\lambda}\right)+\sqrt{\frac{8\epsilon\lambda(\epsilon+\psi(\lambda))}{n}}\right\}$ (169)
and the result is proved. ∎
7.2 Under Sub-Gaussianity
In this section, we prove Theorem 5, which is a counterpart of Theorem 4 under sub-Gaussian $\gamma$. Since the proof is mostly similar to the proof of Theorem 4, we only point out the parts that are changed. It suffices to first prove a counterpart of Lemma 13, namely
$\ln\mathbb{E}[\exp(\lambda\left<Y,Z\right>)]\leq\sup_{\epsilon>0}\{432C\lambda s(\epsilon)-\epsilon/C\}.$ (170)
We first suppose that $\mu$ is a rational distribution supported on at most $n$ points, which can be embedded in $\mathbb{R}^{n}$. Recall the notations associated with the type class in the proof of Lemma 13. Using Lemma 2, we now have that for each type $\nu$ and $\epsilon>0$,
$\mathbb{P}[\xi_{\nu}>CJ(C_{\nu})+\sigma_{\nu}\sqrt{N\epsilon}]\leq C\exp(-N\epsilon/C),$ (171)
where $J(C_{\nu})$ denotes the Dudley integral for $C_{\nu}$. Therefore, setting $S_{N}^{\prime}:=\{\nu\in S_{N}\colon D(\nu\|\mu)\leq\epsilon\}$,
$\mathbb{P}\left[\bigcup_{\nu\in S_{N}^{\prime}}\left\{\xi_{\nu}>CJ(C_{\nu})+\sigma_{\nu}\sqrt{N\epsilon}\right\}\right]\leq C|S_{N}|\exp(-N\epsilon/C).$ (172)
Since Dudley’s integral is sharp for stationary processes, with a constant bounded by 432 (see (10)), we see that
$J(C_{\nu})$ $\leq 432\mathbb{E}[\max_{z^{N}\in C_{\nu}}\left<G^{N},z^{N}\right$>] (173)
$\leq 432N\mathbb{E}_{P_{YZ}\in\Pi(\widehat{P}_{G^{N}},\nu)}[\left<Y,Z\right$>] (174)
where $G^{N}\sim\mathcal{N}(0,I_{nN})$. It is easy to see that $\rho\mapsto\mathbb{E}_{P_{YZ}\in\Pi(\rho,\nu)}[\left<Y,Z\right$>] is concave, so the last displayed item is upper bounded by $432N\mathbb{E}_{P_{GZ}}[\left<Y,Z\right$>]. This implies that
$\mathbb{P}\left[\max_{\nu\in S_{N}^{\prime}}\xi_{\nu}>432CNs_{\rm g}(\epsilon)\right]\leq C|S_{N}|\exp(-N\epsilon/C),$ (175)
where $s_{\rm g}$ is define analogously to (145), but with $\gamma$ replaced by the Gaussian measure $\gamma_{\rm g}$. The rest of the proof proceeds as in the proof of Lemma 13, which yields (170), and the extension to general totally bounded $T\in\ell_{2}$ follows by the same limiting argument.
8 ERM in One Dimension
In this section, we study the toy problem of empirical risk minimization in one dimension, merely for the purpose of illustrating the application of the rate-distortion integral in ERM. We first review the traditional slicing approach from generalization theory, adapting the key ideas from Van de Geer, (1990); Wainwright, (2019), and then give another proof using the new rate-distortion integral approach.
8.1 Review of the Traditional Slicing Approach
Suppose that $X_{1},\dots,X_{n}$ are i.i.d. whose mean $m\in\mathbb{R}$ is unknown, and $X_{i}-m$ is 1-sub-Gaussian. Recall the empirical loss, population loss, and empirical risk minimizer (ERM) defined in Section 2.2. For simplicity, let us assume that T=[1,1]T=[-1,1], and then the ERM becomes
$\hat{m}:=\operatorname*{arg\,min}_{a\in[-1,1]}\hat{L}(a).$ (176)
In this example, we can directly show that $\hat{m}$ is the projection of $\frac{1}{n}\sum_{i=1}^{n}X_{i}$ to the interval [1,1][-1,1], and obtain the estimate
$\mathbb{E}|\hat{m}-m|^{2}=O(\frac{1}{n}).$ (177)
However, in more sophisticated ERM problems, an explicit expression may be unavailable. In this section, we review how (177) can be derived from the slicing technique. We recall (24) and the process $(\chi_{z})_{z\in T-T}$ defined there, whose natural metric is equivalent to the usual distance on $\mathbb{R}$. The covering number is estimated as
$N([-1,1],\lambda)\asymp 1\vee\frac{1}{\lambda}.$ (178)
Applying Dudley’s integrals (7) to (24) yields
$\mathbb{E}[|\hat{m}-m|^{2}]\leq\frac{1}{\sqrt{n}}\int_{0}^{\infty}\sqrt{\ln^{+}\left(\frac{1}{\lambda}\right)}d\lambda\lesssim\frac{1}{\sqrt{n}}$ (179)
which is not sharp. The reason for the looseness is that the estimated parameter is close to the true parameter, but the generalization bounds is a worst-case estimate for any parameter in [1,1][-1,1]. Therefore, the bound is expected be improved by using a local version of the complexity. The classical literature improves the estimate using a “slicing” argument to control the modulus of continuity. For example, using (Van Handel,, 2014, Theorem 5.32) we obtain that with high probability,
$|\hat{m}-m|^{2}\lesssim\frac{1}{\sqrt{n}}\omega(C|\hat{m}-m|)$ (180)
for some constant $C$, where the modulus of continuity of the process $(\chi_{z})$ is bounded as
$\omega(\delta)$ $=\int_{0}^{\delta}\sqrt{\ln N([-1,1],\lambda)}d\lambda$ (181)
$\lesssim\int_{0}^{\delta}\sqrt{\ln^{+}\left(\frac{1}{\lambda}\right)}d\lambda$ (182)
$\lesssim\delta\sqrt{\ln\frac{1}{\delta}}$ (183)
for $\delta<0.1$. In particular, it follows that $|\hat{m}-m|\lesssim\sqrt{\frac{\ln n}{n}}$ with high probability, which is better than (180) but still has a logarithmic factor of suboptimality due to requiring uniform convergence over all slices. Such extra log factors are typical in the analysis of the modulus of continuity; see for example the local and global modulus of continuity of the Wiener process (Koralov and Sinai,, 2007). For the analysis of ERM, it is possible to remove the nuisance log factors by further noting that 1) we only need the inequality (180) to hold for all $\hat{m}$ but fixed $m$, and 2) we can design the slicing scheme so that $\omega(\delta)\asymp\delta$ for the “right scale” $\delta\asymp\frac{1}{\sqrt{n}}$ only, and slightly overestimate $\omega(\delta)$ for other scales, yielding a converging sum in the union bound, yet the inequality $\delta^{2}\lesssim\frac{1}{\sqrt{n}}\omega(C\delta)$ still yields $\delta\lesssim 1/\sqrt{n}$. Details about constructing the slicing scheme and justifying the regularity conditions can be found, for example, in Van de Geer, (1990); Wainwright, (2019).
8.2 Rate-Distortion Integral Approach
In this section, we give another proof of the optimal rate (177) using the rate-distortion integral approach. This new method controls the supremum directly using the second moment, without the need of slicing in the traditional method.
Recall (24), that $|\hat{m}-m|^{2}\leq\frac{1}{\sqrt{n}}\chi_{m-\hat{m}}$, where $\chi_{z}=-2\sqrt{n}z(\frac{1}{n}\sum_{i=1}^{n}X_{i}-m)$, $z\in T-T$ is a centered $O(1)$-sub-Gaussian process. Let $Z:=m-\hat{m}$ and $\sigma_{\rm m}:=\sqrt{\operatorname*{var}(Z)}$. Using Proposition 1, we obtain
$\mathbb{E}[\chi_{Z}]$ $\lesssim\lim_{\epsilon\downarrow 0}\sup_{\mu^{\prime}\colon W_{\infty}(P_{Z},\mu^{\prime})\leq\epsilon}\int_{0}^{+\infty}\sqrt{R_{\mu^{\prime}}(\sigma^{2})}d\sigma.$ (184)
Effectively, we can just think of the right side of (184) as $\int_{0}^{\sigma_{\rm m}}\sqrt{R_{P_{Z}}(\sigma^{2})}d\sigma$, since we will bound the rate-distortion integral by a continuous function of the second moment, and the limiting argument easily applies.
To move forward, one possible approach is to show that, under the constraint $\operatorname*{var}(Z)\leq\sigma_{\rm m}^{2}$, there is
$R_{P_{Z}}(\sigma^{2})$ $\leq R_{\mathcal{N}(0,\sigma_{\rm m}^{2})}(\sigma^{2})$ (185)
$=[\frac{1}{2}\ln\frac{\sigma_{\rm m}^{2}}{\sigma^{2}}]_{+},$ (186)
where (186) follows from the Gaussian rate-distribution function (Thomas and Joy,, 2006), then by change-of-variable we have $\int_{0}^{\sigma_{\rm m}}\sqrt{R_{P_{Z}}(\sigma^{2})}d\sigma\lesssim\sigma_{\rm m}$. We thus obtain $\mathbb{E}[|\hat{m}-m|^{2}]\lesssim\frac{1}{\sqrt{n}}\sqrt{\operatorname*{var}(\hat{m})}\lesssim\frac{1}{\sqrt{n}}\sqrt{\mathbb{E}[|\hat{m}-m|^{2}]}$ and hence $\mathbb{E}[|\hat{m}-m|^{2}]\lesssim\frac{1}{n}$, which is sharp. To justify the key step (185), one possible approach is to use Lieb’s rotation argument for proving Gaussian optimality; see Liu et al., (2018); Courtade and Liu, (2021) for the analysis of a similar problem called the reverse Brascamp-Lieb inequality. We will not provide the details of this approach here.
An alternative, simpler approach, is to leverage the penalized version of the rate-distortion integral: from Lemma 12, we have
$\int_{0}^{\sigma_{\rm m}}\sqrt{R_{P_{Z}}(\sigma^{2})}d\sigma\leq\frac{1}{2}\int_{0}^{\infty}\inf_{P_{U|Z}}\{\alpha^{-2}\mathbb{E}[|U-Z|^{2}]+I(U;Z)\}d\alpha.$ (187)
This penalized form is helpful, because $P_{U|Z}$ is optimized over a set independent of $P_{X}$, so that saddle point analysis applies. We have the following:
Theorem 8.
For any given $\alpha>0$, and $Z$ satisfying $\operatorname*{var}(Z)\leq\sigma_{\rm m}^{2}$, we have
$\inf_{P_{U|Z}}\{\alpha^{-2}\mathbb{E}[|U-Z|^{2}]+I(U;Z)\}$ $\leq(\frac{1}{2}+\frac{1}{2}\ln\frac{2\sigma_{\rm m}^{2}}{\alpha^{2}})\vee\alpha^{-2}\sigma_{\rm m}^{2}$ (188)
and hence $\int_{0}^{\infty}\inf_{P_{U|Z}}\{\alpha^{-2}\mathbb{E}[|U-Z|^{2}]+I(U;Z)\}d\alpha\lesssim\sigma_{\rm m}$.
Proof.
By translation invariance, it suffices to assume that $\mathbb{E}[|Z|^{2}]\leq\sigma_{\rm m}^{2}$. Define the functional $f(P_{Z},P_{U|Z}):=\alpha^{-2}\mathbb{E}[|U-Z|^{2}]+I(U;Z)$, which is convex in $P_{U|Z}$ and concave in $P_{Z}$, and we have the saddle point condition
$\sup_{P_{Z}\colon\mathbb{E}[|Z|^{2}]\leq\sigma_{\rm m}^{2}}f(P_{Z},P_{U|Z}^{*})\leq f(P_{Z}^{*},P_{U|Z}^{*})\leq\inf_{P_{U|Z}}f(P_{Z}^{*},P_{U|Z})$ (189)
where $P_{Z}^{*}=\mathcal{N}(0,\sigma_{\rm m}^{2})$ and $P_{U|Z}^{*}$ is an additive noise Gaussian channel that achieves the infimum; see Thomas and Joy, (2006). Then $\inf_{P_{U|Z}}f(P_{Z}^{*},P_{U|Z})\geq\sup_{P_{Z}\colon\mathbb{E}[|Z|^{2}]\leq\sigma_{\rm m}^{2}}f(P_{Z},P_{U|Z}^{*})\geq\sup_{P_{Z}\colon\mathbb{E}[|Z|^{2}]\leq\sigma_{\rm m}^{2}}\inf_{P_{U|Z}}f(P_{Z},P_{U|Z})$. This shows that under $\mathbb{E}[|Z|]\leq\sigma_{\rm m}^{2}$, the right side of (187) is maximized by $P_{Z}=\mathcal{N}(0,\sigma_{\rm m}^{2})$, in which case
$\inf_{P_{U|Z}}\{\alpha^{-2}\mathbb{E}[|U-Z|^{2}]+I(U;Z)\}$ $=\inf_{D>0}\left\{\alpha^{-2}D+\frac{1}{2}[\ln\frac{\sigma_{\rm m}^{2}}{D}]_{+}\right\}$ (190)
$=(\frac{1}{2}+\frac{1}{2}\ln\frac{2\sigma_{\rm m}^{2}}{\alpha^{2}})\vee\alpha^{-2}\sigma_{\rm m}^{2}.$ (191)
The last claim follows since $\int_{0}^{\sqrt{2}\sigma_{\rm m}}(\frac{1}{2}+\frac{1}{2}\ln\frac{2\sigma_{\rm m}^{2}}{\alpha^{2}})d\alpha+\int_{\sqrt{2}\sigma_{\rm m}}^{\infty}\alpha^{-2}\sigma_{\rm m}^{2}d\alpha\lesssim\sigma_{\rm m}$. ∎
To finish, if $\operatorname*{var}(Z)\leq\sigma_{\rm m}^{2}$ and $W_{\infty}(P_{Z},\mu^{\prime})\leq\epsilon$, then the variance under $\mu^{\prime}$ is bounded by $(\epsilon+\sigma_{\rm m})^{2}$, and we deduce from Theorem 8 and (187) that
$\int_{0}^{+\infty}\sqrt{R_{\mu^{\prime}}(\sigma^{2})}d\sigma\lesssim\sigma_{\rm m}+\epsilon.$ (192)
Taking $\epsilon\downarrow 0$ and applying (184) and (24), we obtain the sharp rate of $\mathbb{E}[\|\hat{m}-m\|_{2}^{2}]\lesssim\frac{1}{n}$.
9 ERM for the Sobolev Ellipsoid
This section provides details of the analysis of the Sobolev ellipsoid in Section 3.4, and prove Theorem 6.
9.1 Supremum over an Ellipsoid
Let $X_{z}:=\left<G,z\right$>, where $G=(G(1),G(2),\dots)$ has i.i.d. standard normal coordinates. The majorizing measure theorem shows that (Talagrand,, 2014, Section 2.13)
$\mathbb{E}[\sup_{z\in\mathcal{E}}X_{z}]\asymp\sqrt{\sum_{i=0}^{\infty}2^{i}a_{2^{i}}^{2}}\asymp\sqrt{\sum_{i=1}^{\infty}a_{i}^{2}}.$ (193)
On the other hand, a nontrivial calculation (Talagrand,, 2014) shows that the Dudley integral is $\asymp\sum_{i=0}^{\infty}2^{i/2}a_{2^{i}}$ hence may not be sharp.
In this section we show that the technique in Section 3.2 can recover the sharp upper bound in (193). More generally, the upper bound holds when the index set is an image of $\mathcal{E}$ under a Lipschitz map, in view of Corollary 3. Although this result is not new, it serves to illustrate the technique. Suppose that $Z$ is a random variable taking values in $\mathcal{E}$ and $\Sigma:=\operatorname*{cov}(Z)$. Henceforth, we will first assume that the support of $P_{Z}$ is countable to derive bounds on $\mathbb{E}[X_{Z}]$ in terms of $\Sigma$, and then use the approximation argument in Proposition 1 to extend the bound to the possibly uncountable case, noting that the covariance is continuous with respect to the Wasserstein-2 metric. Then by Corollary 3,
$\mathbb{E}[X_{Z}]$ $\lesssim\operatorname*{tr}(\sqrt{\Sigma})$ (194)
$\leq\sum_{i=1}^{\infty}\sqrt{\Sigma_{ii}}$ (195)
$\leq\|(a_{i})_{i=1}^{\infty}\|_{2}$ (196)
where (195) follows from the Schur concavity of the square root function and the fact that the eigenvalues majorizes the diagonal values, and (196) follows from the ellipsoid constraint $\sum_{i=1}^{\infty}\frac{\Sigma_{ii}}{a_{i}^{2}}\leq 1$ and Cauchy-Schwarz. This bound recovers the sharp upper bound in (193).
For the case of the Sobolev ellipsoid, see that for $\beta>1/2$, the optimal rate in (193) is $O(\sqrt{\frac{1}{2\beta-1}})$ for $\beta=O(1)$. In contrast, the Dudley integral is
$\sum_{i=0}^{\infty}2^{i/2}a_{2^{i}}=\sum_{i=0}^{\infty}2^{-\beta i+i/2}=O(\frac{1}{2\beta-1})$ (197)
for $\beta=O(1)$, which is worse; see (Wainwright,, 2019, Example 5.12) or (Talagrand,, 2014, Section 2.13).
As a further example, consider the near critical case $a_{i}=i^{-1/2}(\ln i)^{-1}$. The rate-distortion integral gives a finite bound, whereas the traditional Dudley gives a vacuous bound of $\infty$.
9.2 Error Bounds for ERM
In this section, we give two proofs of Theorem 6. The first one only applies to a special case of Theorem 6 where $T$ is the identity map and $Y$ satisfies the stronger log-Sobolev inequality. The second uses the tools developed in Section 3.3 and applies to the more general setting in Theorem 6.
9.2.1 A Simple Proof in a Special Case
We first provide a simple proof of Theorem 6 in a special setting, stated in the following theorem:
Theorem 9.
In Theorem 6, assume that $Y$ satisfies the log-Sobolev inequality (17), and $T=\mathcal{E}$. For any $C>1/2$, there exists $C^{\prime}>0$ such that the following holds: for any $\beta\in(1/2,C]$,
$\mathbb{E}[\|\hat{m}-m\|^{2}]\leq\frac{C^{\prime}}{\sqrt{2\beta-1}}n^{-\frac{2\beta}{2\beta+1}},$ (198)
and if $T$ is convex and $X_{i}$ satisfies the log-Sobolev inequality, we have
$\mathbb{P}\left[\|\hat{m}-m\|>\frac{C^{\prime}}{(2\beta-1)^{1/4}}n^{-\frac{\beta}{2\beta+1}}\right]\leq\exp\left(-\frac{1}{C^{\prime}\sqrt{2\beta-1}}n^{\frac{1}{2\beta+1}}\right).$ (199)
Remark 2.
For fixed $\beta>1/2$, this is the standard nonparametric rate (Tsybakov,, 2009). The assumption of $2\beta>1$ is standard for ensuring the finiteness of the Dudley integral, hence is natural for the analysis of ERM; see (Wainwright,, 2019, Example 5.12). Beyond ERM, there are other estimators which achieve the minimax rate $n^{-\frac{2\beta}{2\beta+1}}$ for $\beta\in(0,1/2)$. For example, we can project $\bar{X}:=\frac{1}{n}\sum_{i=1}^{n}X_{i}$ to a finite dimensional subspace rather than to $\mathcal{E}$ (which requires the knowledge of $\beta$ to determine that subspace); see the analysis of kernel estimators in (Tsybakov,, 2009).
The proof we provide here is essentially that of Chatterjee, (2014); Wei et al., (2020). We control the “soft” version of the Gaussian width in Definition 3, in lieu of the traditional Gaussian width in Chatterjee, (2014). This difference is not essential for convex $T$ due to concentration of the error. The bound from this approach is also sharp; see discussion on the converse in Chatterjee, (2014).
Proof of Theorem 6.
Recall (24), where $\chi_{z}=-2\sqrt{n}\left<z,\frac{1}{n}\sum_{i=1}^{n}X_{i}-m\right$>, $z\in 2\mathcal{E}$ is a centered sub-Gaussian process, whose metric is the $\ell_{2}$ metric (up to universal constant). Invoking Corollary 3, we obtain
$\mathbb{E}[\chi_{Z}]\lesssim\operatorname*{tr}(\sqrt{\Sigma})$ (200)
where we defined the covariance matrix $\Sigma:=\operatorname*{cov}(Z)$, and $Z:=m-\hat{m}$. Now define $\epsilon:=\operatorname*{tr}(\Sigma)$. From the definition of $\mathcal{E}$, we also have $\sum_{i=1}^{\infty}\Sigma_{ii}a_{i}^{-2}\leq 4$. Therefore for any $\lambda>0$, we have
$\sum_{i=1}^{\infty}(a_{i}^{-2}+\lambda)\Sigma_{ii}\leq\lambda\epsilon+4.$ (201)
Then choosing $\lambda=1/\epsilon$,
$(\operatorname*{tr}(\sqrt{\Sigma}))^{2}$ $\leq(\sum_{i=1}^{\infty}\sqrt{\Sigma_{ii}})^{2}$ (202)
$\leq\left(\sum_{i=1}^{\infty}\frac{a_{i}^{-2}+\lambda}{\lambda\epsilon+4}\Sigma_{ii}\right)\left(\sum_{i=1}^{\infty}\frac{\lambda\epsilon+4}{a_{i}^{-2}+\lambda}\right)$ (203)
$\leq 5\sum_{i=1}^{\infty}a_{i}^{2}\wedge\epsilon,$ (204)
where the first inequality follows from concavity of $\sqrt{\cdot}$ and the fact that the eigenvalues of $\Sigma$ majorizes its diagonal values. In the case of $a_{i}=i^{-\beta}$, we have
$\operatorname*{tr}(\sqrt{\Sigma})\lesssim\sqrt{\sum_{i=1}^{\infty}a_{i}^{2}\wedge\epsilon}\asymp\sqrt{\frac{1}{2\beta-1}\epsilon^{1-\frac{1}{2\beta}}}$ (205)
where we assumed $\beta=O(1)$. Combining this with (24) and (200), we obtain $\epsilon\lesssim\frac{1}{\sqrt{n}}\sqrt{\frac{1}{2\beta-1}\epsilon^{1-\frac{1}{2\beta}}}$, which yields $\epsilon\lesssim[(2\beta-1)n]^{-\frac{2\beta}{2\beta+1}}$, as desired.
To show the concentration inequality (199), we treat $\hat{m}=\hat{m}(\bar{X})$ as a function of $\bar{X}$, which is the projection onto the convex set $T$. For any two $x,x^{\prime}\in\ell_{2}$, we have $\left<\hat{m}(x)-\hat{m}(x^{\prime}),x-\hat{m}(x)\right>\geq 0$ and $\left<\hat{m}(x^{\prime})-\hat{m}(x),x^{\prime}-\hat{m}(x^{\prime})\right>\geq 0$, by the hyperplane separation. This implies that $\|\hat{m}(x)-\hat{m}(x^{\prime})\|_{2}\leq\|x-x^{\prime}\|_{2}$, hence $\|\hat{m}-m\|_{2}$ is a Lipschitz function. Therefore $\left\|\nabla\|\hat{m}-m\|_{2}^{2}\right\|_{2}\leq 2\|\hat{m}-m\|_{2}\leq 2$. Since the log-Sobolev inequality implies sub-Gaussian concentration (Boucheron et al.,, 2013), we have
$\mathbb{P}[\|\hat{m}-m\|>\mathbb{E}[\|\hat{m}-m\|]+t]\leq\exp(-cnt^{2})$ (206)
for any $t>0$, where $c>0$ is a universal constant. Then (199) follows by choosing $C^{\prime}$ large enough. ∎
The above simple proof leverages the linearity and concentration, and does not seem to extend to the case of general $\phi$ in Theorem 6. In that case, the entropy integral type argument seems to be essential.
9.2.2 Second Proof
We now give a full proof of Theorem 6 using the technique in Section 3.3. Recall in (24) we defined the $O(1)$-sub-Gaussian process $(\chi_{z})_{T-T}$, where $T$ is now taken to be a Lipschitz image of the ellipsoid, and recall definitions of $\mathsf{G}(E)$ and $\mathsf{G}_{\rm g}(E)$ in Section 3.3. Our proof is based on the following key estimate of the width-moment function:
Theorem 10.
For $(\chi_{z})_{z\in T-T}$, and any $E>0$, we have $\mathsf{G}(E)\lesssim\mathsf{G}_{\rm g}(E)\lesssim C_{\beta}E^{\frac{2\beta-1}{4\beta}}$.
Remark 3.
In (Van de Geer,, 1990, Lemma 5.1), the ERM error for a very similar Sobolev space was bounded by localizing the Dudley integral, where $\beta\geq 1$ is assumed an integer. Suppose that we extend the approach of Van de Geer, (1990) to the setting of Theorem 9. The covering number is bounded as
$\ln N(\delta,\mathcal{E})\lesssim\delta^{-\frac{1}{\beta}}$ (207)
when $\beta=O(1)$. (Van de Geer,, 1990, Example 2.1) estimated the covering number for a similar Sobolev space for fixed $\beta$. For the Sobolev ellipsoid $\mathcal{E}$, (Wainwright,, 2019, Example 5.12) provided a proof of (207), from which we can see that the best preconstant in (207) is the order of a constant independent of $\beta$ when $\beta=O(1)$. For local covering numbers, Wei et al., (2020) showed that the estimate in (207) still holds, where the preconstant is a universal constant. We can then calculate the Dudley integral up to the scale of the error $\sqrt{E}$, resulting in a bound of $\int_{0}^{\sqrt{E}}\delta^{-\frac{1}{2\beta}}d\delta\leq O(\frac{1}{2\beta-1})E^{\frac{1}{2}-\frac{1}{4\beta}}$ when $\beta=O(1)$. Using our proof instead, we get $\mathbb{E}[\chi_{Z}]\leq O(\frac{1}{\sqrt{2\beta-1}})E^{\frac{1}{2}-\frac{1}{4\beta}}$, which is almost the same, but with a better (optimal) preconstant. Thus Dudley’s integral results in $\mathbb{E}[\|\hat{m}-m\|^{2}]=O(\frac{1}{2\beta-1})n^{-\frac{2\beta}{2\beta+1}}$ for $\beta=O(1)$, where the preconstant $O(\frac{1}{2\beta-1})$ is not $O(\frac{1}{\sqrt{2\beta-1}})$ in Theorem 9. This suboptimality is fundamental and is due to the looseness of the Dudley integral: as we see in (197), Dudley’s integral gives an $O(\frac{1}{2\beta-1})$ bound on the supremum over $\mathcal{E}$, rather than the optimal $O(\sqrt{\frac{1}{2\beta-1}})$ bound.
How does the rate-distortion integral improve the preconstant? Naively, we can show a result that for a random variable $Z\in 2\mathcal{E}$, $R_{P_{Z}}(\delta^{2})\lesssim\delta^{-\frac{1}{\beta}}$, for $\beta=O(1)$, which is analogous to (207), and proceed by integrating. But this will still result in a suboptimal preconstant $O(\frac{1}{2\beta-1})$. The reason is that the $P_{Z}$ maximizing $R_{P_{Z}}(\delta^{2})$ changes with $\delta$. Instead, we leverage the penalized version Lemma 12 and evaluate the integrand therein for given $\operatorname*{cov}(Z)$, calculate the integral, and optimize $\operatorname*{cov}(Z)$ in the end. The proof is made simple by properties of information-theoretic measures such as chain rules of the entropy.
We first need a few preparations. Define
$f(t):=\left\{\begin{array}[]{cc}t/2&0\leq t\leq 1\\ \frac{1+\ln t}{2}&t>1\end{array}\right.$. (210)
We have
Lemma 16.
$t\int_{0}^{1/t}f(\frac{1}{u^{2}})du\leq 4f(t)$, for all $t>0$.
Proof.
For $0<u_{0}\leq 1$,
$\int_{0}^{u_{0}}f(\frac{1}{u^{2}})du=\int_{0}^{u_{0}}\frac{1+\ln\frac{1}{u^{2}}}{2}du=\frac{3}{2}u_{0}+u_{0}\ln\frac{1}{u_{0}}\leq\frac{3}{2}u_{0}(1+\ln\frac{1}{u_{0}}).$ (211)
For $u_{0}>1$, $\int_{1}^{u_{0}}f(\frac{1}{u^{2}})du=\int_{1}^{u_{0}}\frac{1}{2u^{2}}du=\frac{1}{2}(1-\frac{1}{u_{0}})$, so that
$\int_{0}^{u_{0}}f(\frac{1}{u^{2}})du=\frac{3}{2}+\frac{1}{2}(1-\frac{1}{u_{0}})=2-\frac{1}{2u_{0}}\leq 2.$ (212)
The result follows by taking $u_{0}=1/t$. ∎
Lemma 17.
Assume that $W\in\ell_{2}$ is a random variable and $\Sigma:=\operatorname*{cov}(W)$. Let $Z=\phi(W)$, where $\phi$ is a 1-Lipschitz function, and suppose that $\mathbb{E}[\|Z\|^{2}]\leq\epsilon$. Then, for any $\alpha\in(0,\infty)$,
$\inf_{P_{U|Z}}\{\alpha^{-2}\mathbb{E}[\|U-Z\|^{2}]+I(U;Z)\}\leq\alpha^{-2}\epsilon\wedge\sum_{i=1}^{\infty}f(2\alpha^{-2}\Sigma_{ii}).$ (213)
Proof.
By considering $U=0$, it is easy to see that $\inf_{P_{U|X}}\{\alpha^{-2}\mathbb{E}[\|U-Z\|^{2}]+I(U;Z)\}\leq\alpha^{-2}\epsilon$. Thus it remains to show that
$\inf_{P_{U|Z}}\{\alpha^{-2}\mathbb{E}[\|U-Z\|^{2}]+I(U;Z)\}\leq\sum_{i=1}^{\infty}f(2\alpha^{-2}\Sigma_{ii}).$ (214)
It suffices to prove (214) in the special case where the Lipschitz map $\phi$ is the identity, since the left side of (214) only decreases with Lipschitz $\phi$. Furthermore, the left side of (214) is upper bounded by $\sum_{i=1}^{\infty}\inf_{P_{U_{i}|Z_{i}}}\{\alpha^{-2}\mathbb{E}[|U_{i}-Z_{i}|^{2}]+I(U_{i};Z_{i})\}$, because given $P_{U_{i}|Z_{i}}$ we can define $P_{U|Z}=\prod_{i=1}^{\infty}P_{U_{i}|Z_{i}}$, and then $I(U;Z)\leq\sum_{i=1}^{\infty}I(U_{i};Z_{i})$. Then the result in (214) follows from the one-dimensional case, which we already proved in Theorem 8. ∎
Lemma 18.
Suppose that $\sum_{i=1}^{\infty}\Sigma_{ii}i^{\beta}\leq 4$, where $\beta>1/2$. For any $\epsilon>0$, we have
$\int_{0}^{\infty}\alpha^{-2}\epsilon\wedge\sum_{i=1}^{\infty}f(2\alpha^{-2}\Sigma_{ii})d\alpha\leq C_{\beta}\epsilon^{\frac{2\beta-1}{4\beta}},$ (215)
where $C_{\beta}$ is a constant defined around (236), which satisfies $C_{\beta}=O(\frac{1}{\sqrt{2\beta-1}})$ as $\beta\downarrow\frac{1}{2}$.
Proof.
Set
$\alpha_{0}:=\epsilon^{\frac{1+2\beta}{4\beta}}.$ (216)
We have
$\int_{\alpha_{0}}^{\infty}\alpha^{-2}\epsilon\wedge d\alpha$ $\leq\epsilon^{\frac{2\beta-1}{4\beta}}.$ (217)
Moreover, by Lemma 16
$\int_{0}^{\alpha_{0}}\sum_{i=1}^{\infty}f(2\alpha^{-2}\Sigma_{ii})d\alpha$ $=\sum_{i=1}^{\infty}\sqrt{2\Sigma_{ii}}\int_{0}^{\frac{\alpha_{0}}{\sqrt{2\Sigma_{ii}}}}f(\frac{1}{u^{2}})du$ (218)
$\leq\alpha_{0}\sum_{i=1}^{\infty}4f(\frac{\sqrt{2\Sigma_{ii}}}{\alpha_{0}}).$ (219)
It will become clear that $\alpha_{0}$ was chosen so that (217) and (219) are both upper bounded by $\epsilon^{\frac{2\beta-1}{4\beta}}$. Therefore, it remains to consider the optimization problem:
$M_{1}:=\sup_{(\Sigma_{i})}\sum_{i=1}^{\infty}f(\frac{\sqrt{2\Sigma_{i}}}{\alpha_{0}})$ (220)
under the constraint $\sum_{i=1}^{d}\Sigma_{i}a_{i}^{-2}\leq 4$. Also, consider another optimization problem:
$M_{2}:=\sup_{\Sigma(\cdot)}\int_{0}^{\infty}f(\frac{\sqrt{2\Sigma(t)}}{\alpha_{0}})dt$ (221)
under the constraint $\int_{0}^{\infty}\Sigma(t)t^{2\beta}dt\leq 1$, where $\Sigma(\cdot)$ is a nonnegative continuous function. Clearly $M_{2}\geq M_{1}$, since given $(\Sigma_{ii})$ we can define $\Sigma(t)=\Sigma_{\lceil t\rceil}$.
Since $f$ has continuous, strictly increasing first derivative, we can see that an optimizer $\Sigma(\cdot)$ exists, which is continuous. Since $f(\cdot)$ is increasing, $\Sigma(\cdot)$ must be decreasing. There exists unique $t_{0}>0$ such that $\frac{\sqrt{2\Sigma(t)}}{\alpha_{0}}\geq 1$, $t\in(0,t_{0}]$ and $\frac{\sqrt{2\Sigma(t)}}{\alpha_{0}}\leq 1$, $t\in[t_{0},\infty)$. Stationarity condition implies that there exists $\lambda>0$ such that
$\frac{1}{4\Sigma(t)}$ $=\lambda t^{2\beta},\quad t\in(0,t_{0}],$ (222)
$\frac{1}{2\alpha_{0}\sqrt{2\Sigma(t)}}$ $=\lambda t^{2\beta}\quad t\in[t_{0},\infty).$ (223)
Setting $t=t_{0}$ in both equations and canceling $\Sigma(t_{0})$, we obtain
$\lambda=\frac{1}{2\alpha_{0}^{2}t_{0}^{2\beta}}$ (224)
Also, we can express $\Sigma(t)$ as a function of $t$ from the above, and $\int_{0}^{\infty}\Sigma(t)t^{2\beta}dt=1$ must be achieved, implying
$4$ $=\int_{0}^{t_{0}}\frac{1}{4\lambda}dt+\int_{t_{0}}^{\infty}\frac{1}{8\alpha_{0}^{2}\lambda^{2}t^{2\beta}}dt$ (225)
$=\frac{t_{0}}{4\lambda}+\frac{1}{8(2\beta-1)\alpha_{0}^{2}\lambda^{2}t_{0}^{2\beta-1}}$ (226)
$=\frac{\beta}{2\beta-1}\alpha_{0}^{2}t_{0}^{2\beta+1}.$ (227)
Now that we are able to express $t_{0}$ and $\lambda$ as functions of $\alpha$, we can evaluate the optimal value:
$\int_{0}^{t_{0}}f(\frac{\sqrt{2\Sigma(t)}}{\alpha_{0}})dt$ $=\int_{0}^{t_{0}}\frac{1+\ln\frac{\sqrt{2\Sigma(t)}}{\alpha_{0}}}{2}dt$ (228)
$=\int_{0}^{t_{0}}\frac{1+\beta\ln\frac{t_{0}}{t}}{2}dt$ (229)
$=\frac{t_{0}(1+\beta)}{2},$ (230)
and
$\int_{t_{0}}^{\infty}f(\frac{\sqrt{2\Sigma(t)}}{\alpha_{0}})dt$ $=\int_{t_{0}}^{\infty}\frac{\sqrt{2\Sigma(t)}}{2\alpha_{0}}dt$ (231)
$=\int_{t_{0}}^{\infty}\frac{t_{0}^{2\beta}}{2t^{2\beta}}dt$ (232)
$=\frac{t_{0}}{2(2\beta-1)}.$ (233)
In summary, we have shown that
$M_{1}\leq M_{2}$ $\leq\left(\frac{1+\beta}{2}+\frac{1}{2(2\beta-1)}\right)t_{0}$ (234)
$\leq c_{\beta}\alpha_{0}^{-\frac{2}{1+2\beta}}$ (235)
where we defined
$c_{\beta}=\left(\frac{1+\beta}{2}+\frac{1}{2(2\beta-1)}\right)\left(\frac{\beta}{4(2\beta-1)}\right)^{-\frac{1}{1+2\beta}}.$ (236)
Hence
$\int_{0}^{\infty}\alpha^{-2}\epsilon\wedge\sum_{i=1}^{d}f(2\alpha^{-2}\Sigma_{ii})d\alpha\leq 4\alpha_{0}M_{1}+\epsilon^{\frac{2\beta-1}{4\beta}}=(4c_{\beta}+1)\epsilon^{\frac{2\beta-1}{4\beta}}.$ (237)
The proof is completed by setting $C_{\beta}=4c_{\beta}+1$. ∎
Proof of Theorem 10.
It is obvious from Corollary 2 that $\mathsf{G}(E)\lesssim\mathsf{G}_{\rm g}(E)$, hence we only need to prove the second inequality. Let $Z:=m-\hat{m}$, and suppose that $\mathbb{E}[\|Z\|_{2}^{2}]\leq E$, and $Z$ has countable support. Invoking Theorem 1, Lemma 12, Lemma 17 and Lemma 18, we obtain
$\mathbb{E}[\chi_{Z}]$ $\lesssim\int_{0}^{\infty}\inf_{P_{U|X}}\{\alpha^{-2}\mathbb{E}[\|U-Z\|^{2}]+I(U;Z)\}d\alpha$ (238)
$\lesssim C_{\beta}E^{\frac{2\beta-1}{4\beta}}.$ (239)
The result is then extended to general $P_{Z}$ by approximation by countable distributions in $W_{\infty}$ and noting the continuity of $E\mapsto E^{\frac{2\beta-1}{4\beta}}$. ∎
Proof of Theorem 6.
The proof is based on (24), and we recall that Theorem 10 shows
$\mathsf{G}(E)\lesssim\mathsf{G}_{\rm g}(E)\lesssim C_{\beta}E^{\frac{2\beta-1}{4\beta}}.$ (240)
Below we given the details of the proof when $Y$ is assumed to satisfy the log-Sobolev inequality, and use Theorem 4. The proof for general sub-Gaussian $Y$ is similar, using Theorem 5, but with uglier constants, hence we omit. We see that $\mathsf{G}$ is sub-additive, hence
$\psi(n/72)$ $\leq\sup_{\epsilon>0}\{-\frac{\epsilon}{3}+\frac{\sqrt{n}}{36}\mathsf{G}(\frac{72\epsilon}{n})\}+\frac{\sqrt{n}}{36}\mathsf{G}(\frac{72\psi(n/72)}{n})$
$\quad+\sup_{\epsilon>0}\{-\frac{\epsilon}{3}+\frac{\epsilon}{3}\}+\sup_{\epsilon>0}\{-\frac{\epsilon}{3}+\frac{\sqrt{\epsilon\psi(n/72)}}{3}\}.$ (241)
Let $I_{1},I_{2},I_{3},I_{4}$ denote the four terms in (241). Then $I_{3}=0$, and the other terms can be controlled using (240):
$I_{1}\lesssim\sup_{\epsilon>0}\{-\epsilon+\sqrt{n}C_{\beta}(\frac{\epsilon}{n})^{\frac{2\beta-1}{4\beta}}\}\lesssim n^{\frac{1}{\beta+1}}C_{\beta}^{\frac{4\beta}{2\beta+1}};$ (242)
$I_{2}\lesssim C_{\beta}n^{\frac{1}{4\beta}}\psi^{\frac{2\beta-1}{4\beta}}(\frac{n}{72});$ (243)
$I_{4}\leq\frac{1}{12}\psi(n/72).$ (244)
Thus we must have either $\frac{1}{3}\psi(n/72)\leq I_{1}$ or $\frac{1}{3}\psi(n/72)\leq I_{2}$, and either option shows that $\psi(n/72)\leq cC_{\beta}^{\frac{4\beta}{2\beta+1}}n^{\frac{1}{2\beta+1}}$ for some universal constant $c>0$.
Now by the Markov inequality (with $\lambda=n/72$) we have, for any $t>0$,
$\mathbb{P}[\|m-\hat{m}\|_{2}>t]\leq\exp(cC_{\beta}^{\frac{4\beta}{2\beta+1}}n^{\frac{1}{2\beta+1}}-nt^{2}/72).$ (245)
For $t_{0}:=10\sqrt{c}C_{\beta}^{\frac{2\beta}{2\beta+1}}n^{-\frac{\beta}{2\beta+1}}$, we have
$\mathbb{P}[\|m-\hat{m}\|_{2}>t_{0}]\leq\exp(-\frac{7}{18}cC_{\beta}^{\frac{4\beta}{2\beta+1}}n^{\frac{1}{2\beta+1}})$ (246)
which implies the claimed result. ∎
10 ERM for Sparse Regression
In this section, we give two proofs of Theorem 7. The first one, based on Hölder’s inequality, is simpler but only applies to a special case; the second uses the rate-distortion integral and proves the theorem in full generality.
10.1 Simple Proof in a Special Case
First, we give a simple and direct proof of the following special case of Theorem 7:
Theorem 11.
Suppose that $A:=\frac{1}{\sqrt{n}}\mathbb{X}$ is a square matrix, and its largest and smallest singular values satisfy $\sigma_{\rm max}(A)\lesssim 1$ and $\sigma_{\rm max}(A)\gtrsim 1$. Let $q\in(0,1)$. Then $\mathbb{E}[\|\beta^{*}-\hat{\beta}\|_{2}^{2}]\lesssim_{q}R\left(\frac{\ln d}{n}\right)^{\frac{2-q}{2}}$.
Before proving Theorem 11, let us recall some auxiliary results. Note that if $x\in wB_{q}(r)$, then for any $p>q$ we have
$\|x\|_{p}^{p}$ $=\int_{0}^{+\infty}\sum_{i=1}^{d}1_{|x_{i}|>t}d(t^{p})$ (247)
$\leq\int_{t\in(0,r)}(\frac{r}{t})^{q}d(t^{p})$ (248)
$=\frac{r^{p}}{1-q/p}$ (249)
implying the inclusion $wB_{q}(r)\subset B_{p}((1-q/p)^{-\frac{1}{p}}r)$. Therefore risk bounds for $\ell_{p}$ balls imply risk bounds for weak $\ell_{q}$ balls, but this naive argument will not give the sharp rate in Theorem 7, since $p$ must be chosen to satisfy the strict inequality $p>q$.
Nevertheless, we will see that in the setting of Theorem 11, which is a special case of Theorem 7, there is a simple proof based on the interpolation inequality, and it is well-known in functional analysis that interpolation can improve “weak $\ell_{p}$ estimates” at the boundary to an $\ell_{p}$ estimate” at the interpolant. More precisely, we will use the following result, which can be viewed as a special case of the Marcinkiewicz interpolation theorem:
Lemma 19.
Suppose that $X$ is a nonnegative random variable, $\sup_{t>0}t^{a}\mathbb{P}[X>At]\leq 1$, $\sup_{t>0}t^{a}\mathbb{P}[X>Bt]\leq 1$ for $A,B,a,b>0$ and $\alpha\in(0,1)$, where $a\neq b$ and $\frac{\alpha}{a}+\frac{1-\alpha}{b}=1$. Then $\mathbb{E}[X]\lesssim_{a,b}A^{\alpha}B^{1-\alpha}$.
Proof.
By a rescaling argument, we see that it suffices to consider the case of $A=B$. Then the proof is completed since
$\mathbb{E}[X]$ $=\int_{0}^{\infty}\mathbb{P}[X>s]ds$ (250)
$\leq\int_{0}^{\infty}1\wedge\frac{A^{a}}{s^{a}}\wedge\frac{A^{b}}{s^{b}}ds$ (251)
$=A\int_{0}^{\infty}1\wedge\frac{1}{t^{a}}\wedge\frac{1}{t^{b}}dt\lesssim_{a,b}A.$ (252)
Proof of Theorem 11.
Let $E:=\mathbb{E}[\|\hat{\beta}-\beta^{*}\|_{2}^{2}]$ and $Z:=\hat{\beta}-\beta^{*}$. Suppose that $\theta\in[0,\frac{q}{2-q})$ and
$\frac{1}{p}=\frac{\theta}{q}+\frac{1-\theta}{2}.$ (253)
We think of $p>1$ as a function of $\theta$, which is close to 1. Let $R_{j}:=\sup_{t}t\mathbb{P}^{1/q}[|Z_{j}|>t]$ and $E_{j}:=\mathbb{E}[|Z_{j}|^{2}]$ for each $j$. Then $\sum_{j=1}^{d}E_{j}=E$. Since $\beta^{*},\hat{\beta}\in wB_{q}(r)$, we have $\sum_{j=1}^{d}|Z_{j}|^{q}\leq 2r^{q}=2R$, hence $\sum_{j=1}^{d}R_{j}\leq 2R$. By the interpolation inequality (Lemma 19), we have
$\mathbb{E}[|Z_{j}|^{p}]$ $\lesssim_{q}R_{j}^{\frac{\theta p}{q}}E_{j}^{\frac{1-\theta}{2}p}.$ (254)
Hence for $\theta=0$ we have
$\mathbb{E}[\|Z\|_{p}^{p}]$ $\lesssim_{q}\sum_{j=1}^{d}R_{j}^{\frac{\theta p}{q}}E_{j}^{\frac{1-\theta}{2}p}$ (255)
$\leq\|(R_{j}^{\frac{\theta p}{q}})\|_{\frac{q}{\theta p}}\|(E_{j}^{\frac{1-\theta}{2}p})\|_{\frac{2}{(1-\theta)p}}$ (256)
$\leq(2R)^{\frac{\theta p}{q}}E^{\frac{(1-\theta)p}{2}}.$ (257)
Moreover, (257) is also trivially true when $\theta=0$ (in which case $p=2$).
Recall (324). We have, by Hölder’s inequality,
$E$ $\leq\frac{2}{\sqrt{n}}\mathbb{E}[\|Z\|_{p}\|A^{\top}w\|_{p^{\prime}}]$ (258)
$\leq\frac{2}{\sqrt{n}}\mathbb{E}^{\frac{1}{p}}[\|Z\|_{p}^{p}]\cdot\mathbb{E}^{\frac{1}{p^{\prime}}}[\|A^{\top}w\|_{p^{\prime}}^{p^{\prime}}]$ (259)
$\lesssim\frac{1}{\sqrt{n}}(2R)^{\frac{\theta}{q}}E^{\frac{1-\theta}{2}}d^{\frac{1}{p^{\prime}}}\sqrt{p^{\prime}},$ (260)
where we used the fact that for $G\sim\mathcal{N}(0,1)$ we have $\mathbb{E}^{1/p^{\prime}}[|G|^{p^{\prime}}]\leq 3\sqrt{p^{\prime}}$, and $p^{\prime}=(1-1/p)^{-1}$ is the Hölder conjugate. Rearranging, we obtain
$E^{\frac{1+\theta}{2}}\lesssim_{q}\sqrt{p^{\prime}}d^{\frac{1}{p^{\prime}}-\frac{1}{2}}R^{\frac{\theta}{q}}.$ (261)
Define
$\theta_{0}:=\frac{q}{2-q}.$ (262)
Comparing with (253), we see that $p$ is close to 1 when $\theta$ is close to $\theta_{0}$.
Case I: $(\frac{1}{q}-\frac{1}{2})\ln d+\frac{1}{2}\ln\frac{(1-q/2)E}{(2R)^{\frac{2}{q}}}\leq\frac{1}{\theta_{0}}$. The assumption of this case implies
$E\lesssim_{q}R^{\frac{2}{q}}d^{-\frac{2-q}{q}}.$ (263)
Take $\theta=0$, $p=p^{\prime}=2$, we get from (261) that
$\sqrt{E}\lesssim_{q}1.$ (264)
Raising (263) to the power of $q/2$ and (264) to the power of 2q2-q gives $E\lesssim_{q}Rn^{-\frac{2-q}{2}}$.
Case II: $(\frac{1}{q}-\frac{1}{2})\ln d+\frac{1}{2}\ln\frac{(1-q/2)E}{(2R)^{\frac{2}{q}}}>\frac{1}{\theta_{0}}$. From (253) and (262) we have
$\frac{1}{p^{\prime}}=(\frac{1}{q}-\frac{1}{2})(\theta_{0}-\theta).$ (265)
Choosing
$\theta=\theta_{0}-\frac{1}{(\frac{1}{q}-\frac{1}{2})\ln d+\frac{1}{2}\ln\frac{(1-q/2)E}{(2R)^{\frac{2}{q}}}}\in(0,\frac{q}{2-q})$ (266)
yields
$d^{\frac{1}{p^{\prime}}}\left(\frac{(1-q/2)E}{(2R)^{2/q}}\right)^{\frac{\theta_{0}-\theta}{2}}=\left(\frac{d^{\frac{1}{q}-\frac{1}{2}}\sqrt{1-q/2}E^{\frac{1}{2}}}{(2R)^{\frac{1}{q}}}\right)^{\theta_{0}-\theta}=e$ (267)
and
$\sqrt{p^{\prime}}=\sqrt{\ln d+(\frac{2}{q}-1)^{-1}\ln\frac{(1-q/2)E}{(2R)^{\frac{2}{q}}}}\leq\sqrt{\ln d}$ (268)
where we used $\ln\frac{(1-q/2)E}{(2R)^{\frac{2}{q}}}\leq 0$ due to (249). Then (261) shows that
$E^{\frac{1+\theta_{0}}{2}}$ $\lesssim_{q}d^{-1/2}R^{\frac{\theta_{0}}{q}}\cdot\sqrt{p^{\prime}}d^{\frac{1}{p^{\prime}}}\left(\frac{(1-q/2)E}{(2R)^{2/q}}\right)^{\frac{\theta_{0}-\theta}{2}}$ (269)
$\lesssim d^{-1/2}R^{\frac{\theta_{0}}{q}}\cdot\sqrt{\ln d}.$ (270)
Equivalently, we have
$E\lesssim_{q}R\left(\frac{\ln d}{n}\right)^{\frac{2-q}{2}}.$ (271)
Remark 4.
In (Raskutti et al.,, 2011, (11)), a condition $\frac{d}{Rn^{q/2}}\gtrsim d^{\delta}\gtrsim 1$ (for some $\delta\in(0,1)$). was globally imposed. This condition would imply the optimality of the bound in Theorem 11, but it is not necessary for the upper bound direction.
10.2 Proof Using the Rate-Distortion Integral
In this section, we give a proof of Theorem 7 in full generality, using the technique in Section 3.3. The key part is the following estimate of the width-moment function:
Lemma 20.
Suppose that $q\in(0,1)$, $Z\in wB_{q}(r)$ in $\mathbb{R}^{d}$, and $\mathbb{E}[\|AZ\|_{2}^{2}]\leq E$, where $A$ is an $n\times d$ matrix whose columns have $\ell_{2}$ norm at most 1. Then we have
$\mathsf{w}(P_{AZ})\lesssim_{q}\sqrt{\ln d}r^{\frac{q}{2-q}}E^{\frac{1-q}{2-q}}.$ (272)
In particular, for the set $A(wB_{q}(r))$, we have $\mathsf{G}(E)\lesssim_{q}\sqrt{\ln d}r^{\frac{q}{2-q}}E^{\frac{1-q}{2-q}}$.
To prove Lemma 20 using the rate-distortion integral, we need to establish upper bounds on the rate-distortion function. Let $g\colon(0,\infty)\mapsto(0,\infty)$ be the concave envelope of $x\ln\frac{1}{x}\vee\ln x$. It is easy to see that $g$ has continuous first derivative, and there exists $a_{1}\in(0,1)$ and $a_{2}\in(0,1)$ such that
$g(x)$ $=x\ln\frac{1}{x},\quad\forall x<a_{1};$ (273)
$g(x)$ $=\ln x,\quad\forall x>a_{2},$ (274)
and $g$ is linear for $x\in[a_{1},a_{2}]$. We show that this function controls the rate-distortion function of random variables in $\mathbb{R}$ satisfying a weak $\ell_{q}$ condition:
Lemma 21.
Suppose that $q\in(0,1)$, $r\in(0,\infty)$, and $Z\in\mathbb{R}$ is a random variable satisfying $\sup_{t>0}t^{q}\mathbb{P}[|Z|>t]\leq r^{q}$. Then there exists a random variable $U$ such that $|U|\leq|Z|$ almost surely, $\mathbb{E}[|U-Z|]\leq b$, and
$H(U)\lesssim_{q}g\left(\left(\frac{r}{b}\right)^{\frac{q}{1-q}}\right).$ (275)
Proof.
By scaling invariance of the entropy, it suffices to consider the case of $r=1$, which we shall assume henceforth. Let $t_{k}=k(\frac{1-q}{2-q}b)^{\frac{1}{1-q}}$, $k\in\mathbb{Z}$. Given $Z$, define $U=t_{k}$ where $k=\operatorname*{arg\,min}_{l\colon|t_{l}|\leq|Z|}|Z-t_{l}|$. Then
$\mathbb{E}[|U-Z|]$ $\leq\mathbb{E}[|Z|1_{|Z|\leq t_{1}}]+t_{1}P_{Z}(\mathbb{R}\setminus[t_{-1},t_{1}])$ (276)
$\leq\int_{0}^{t_{1}}\frac{dt}{t^{q}}+t_{1}^{1-q}$ (277)
$=\frac{2-q}{1-q}t_{1}^{1-q}=b.$ (278)
To bound the entropy, first consider the case of $t_{1}^{-q}\leq 2/3$. For $k\geq 1$, define $p_{k}:=p_{-k}:=\frac{1}{2}t_{1}^{-q}\left(\frac{1}{k^{q}}-\frac{1}{(k+1)^{q}}\right)$, and $p_{0}:=1-t_{1}^{-q}$. We see that $p_{k}$ is a symmetric sequence and decreasing for $k\geq 0$ by our assumption. Then $P_{\xi}(kt_{1})=p_{k}$, $k\in\mathbb{Z}$ is the solution to the following optimization problem:
$\max_{P_{\xi}}H(\xi),\textrm{ where $\xi\in t_{1}\mathbb{Z}$ satisfies $\mathbb{P}[|\xi|\geq kt_{1}]\leq(kt_{1})^{-q}$ for all $k\in\mathbb{Z}\setminus\{0\}$. }$ (279)
Indeed, this is a convex optimization problem, and using the Lagrange multiplier method, we see that the local optimality condition can be expressed as $P_{\xi}(kt_{1})=\sum_{l=|k|}^{\infty}\lambda_{l}$ for some nonnegative $\lambda_{0},\lambda,\dots$, where $\lambda_{k}=0$ if $\mathbb{P}[|\xi|\geq kt_{1}]<(kt_{1})^{-q}$ (complementary slackness). Since $P_{\xi}(kt_{1})=p_{k}$ satisfies local optimality, it is also globally optimal. Clearly $H(U)$ is upper bounded by the value of (279). Setting $l_{k}:=\frac{1}{k^{q}}-\frac{1}{(k+1)^{q}}$, we have $\sum_{k\geq 1}l_{k}=1$, and
$\sum_{k\geq 1}l_{k}\ln\frac{1}{l_{k}}$ $\leq\sum_{k\geq 1}\frac{q}{k^{q+1}}\ln\frac{(k+1)^{q+1}}{q}\lesssim_{q}1.$ (280)
Hence
$H(U)$ $\leq\sum_{k\in\mathbb{Z}}p_{k}\ln\frac{1}{p_{k}}$ (281)
$\leq p_{0}\ln\frac{1}{p_{0}}+\sum_{k\geq 1}2p_{k}\ln\frac{1}{p_{k}}$ (282)
$\leq t_{1}^{-q}+t_{1}^{-q}\left(\sum_{k\geq 1}l_{k}\ln\frac{2t_{1}^{q}}{l_{k}}\right)$ (283)
$\lesssim_{q}g(t_{1}^{-q}).$ (284)
Next, consider the case of $t_{1}^{-q}>2/3$. Since $p_{k}$ is no longer a decreasing sequence for $k\geq 0$, the optimizer for (279) cannot be expressed in the same way as before. However, it is possible to bound $H(U)$ for this case using the bound in the previous case. We choose $\bar{t}_{1}=mt_{1}$ where $m=\lceil 1.5^{1/q}t_{1}^{-1}\rceil\geq 2$ is the minimum integer such that $\bar{t}_{1}^{-q}\leq 2/3$. Define $\bar{t}_{k}=k\bar{t}_{1}$, $k\in\mathbb{Z}$. Given $Z$, define $\bar{U}=\bar{t}_{k}$ where $k=\operatorname*{arg\,min}_{l\colon|\bar{t}_{l}|\leq|Z|}|Z-\bar{t}_{l}|$. Then since $U$ can take at most 2m12m-1 possible values given $\bar{U}$, we have $H(U|\bar{U})\leq\ln(2m+1)$, hence
$H(U)$ $\leq H(\bar{U})+\ln(2m-1)$ (285)
$\lesssim g(\bar{t}_{1}^{-q})+\ln m$ (286)
where we used (284) to bound $H(\bar{U})$. Since $m\asymp_{q}t_{1}^{-1}$, we have $\bar{t}_{1}^{q}\asymp_{q}1$, and $H(U)\lesssim_{q}1+\ln\frac{1}{t_{1}}\lesssim_{q}g(t_{1}^{-q})$. In summary, in either case, we have established $H(U)\lesssim_{q}g(t_{1}^{-q})$ in the case of $r=1$, and the proof for general $r$ is completed by rescaling. ∎
By subadditivity of entropy, Lemma 21 implies a bound on the rate-distortion function of weak $\ell_{q}$ balls:
Lemma 22.
Suppose that $Z\in wB_{q}(r)$ is a random variable, and $b>0$ is arbitrary. Then we can find $U=(U_{1},\dots,U_{d})$ such that $U\in wB_{q}(r)$ almost surely, $\mathbb{E}[\|U-Z\|_{1}]\leq b$, and
$H(U)\lesssim_{q}dg\left(d^{-1}\left(\frac{r}{b}\right)^{\frac{q}{1-q}}\right).$ (287)
Proof.
Let $r_{j}=\sup_{t}t\mathbb{P}^{1/q}[|Z_{j}|>t]$, so that $\sum_{j}r_{j}^{q}\leq r^{q}$. Let $b_{j}:=\frac{r_{j}^{q}b}{r^{q}}$. By Lemma 21, we can find $U_{1},\dots,U_{d}$ such that $|U_{i}|\leq|Z_{i}|$, i=1i=1,…, and that
$\mathbb{E}[\|U-Z\|_{1}]=\sum_{j=1}^{d}\mathbb{E}[|U_{j}-Z_{j}|]\leq\sum_{j=1}^{d}b_{j},$ (288)
and that
$H(U)$ $\leq\sum_{j=1}^{d}H(U_{j})$ (289)
$\lesssim_{q}\sum_{j=1}^{d}g\left(\left(\frac{r_{j}}{b_{j}}\right)^{\frac{q}{1-q}}\right)$ (290)
$=\sum_{j=1}^{d}g\left(r_{j}^{q}\left(\frac{b}{r^{q}}\right)^{-\frac{q}{1-q}}\right)$ (291)
$\leq dg\left(d^{-1}\left(\frac{r}{b}\right)^{\frac{q}{1-q}}\right)$ (292)
where the last step used Jensen’s inequality to see that the maximum value is achieved when $r_{j}^{q}=r^{q}/d$. ∎
Next, we prove the following information-theoretic version of Sudakov minoration, which will be our main tool for upper bound on the $\ell_{2}$ rate-distortion function”.
Lemma 23.
Suppose that $Z$ is a random variable in $\mathbb{R}^{d}$, and $G\sim\mathcal{N}(0,I_{d})$. Then for any $\epsilon>0$,
$\inf_{P_{V|Z}\colon\mathbb{E}[\|V-Z\|_{2}^{2}]\leq\epsilon^{2}}I(V;Z)$ $\lesssim\frac{1}{\epsilon^{2}}\sup_{P_{ZG}}|\mathbb{E}[\left<Z,G\right>]|^{2}$ (293)
where the supremum is over couplings of $P_{Z}$ and $P_{G}$.
Proof.
The proof follows by applying the lifting argument to the standard Sudakov minoration (see for example (Boucheron et al.,, 2013, Theorem 13.4)), which should upper bound the left side of (293) by $\frac{2}{\epsilon^{2}}|\mathbb{E}[\left<Z,G\right>]|^{2}$. Alternatively, the result follows by taking $\epsilon_{1}\to\infty$ in Lemma 9 of the present paper, although this is somewhat overkill, as Lemma 9 itself is a consequence of the super-Sudakov inequality-a strengthening of the standard Sudakov inequality. ∎
Next, we prove auxiliary results showing that the inner product in (293) can be “essentially” controlled by $\mathbb{E}[\|Z\|_{\ell_{1}}]$.
Lemma 24.
Suppose that $Z\in\mathbb{R}$ satisfies $\mathbb{E}[|Z|]\leq a$, and $|Z|\leq b$ almost surely, where $a\geq 0$ and $b>0$. Let $G\sim\mathcal{N}(0,1)$. Then $\mathbb{E}[ZG]\leq 3ebf(a/b)$, where we defined the function:
$f(x)=\left\{\begin{array}[]{cc}x\sqrt{\ln\frac{1}{x}}&x\in(0,e^{-1})\\ e^{-1}&x\in[e^{-1},\infty)\end{array}\right.$. (296)
Proof.
First, consider the case of $a=1$ and $b>e$. For any $p>1$, we obtain $\mathbb{E}[|Z|^{p}]\leq b^{p-1}$ from the interpolation inequality for the norm. Therefore, if $p$ is such that $\frac{1}{p^{\prime}}=1-\frac{1}{p}=\frac{1}{\ln b}$, we have
$\mathbb{E}[ZG]$ $\leq\|G\|_{p^{\prime}}\|Z\|_{p}$ (297)
$\leq 3\sqrt{p^{\prime}}\left(b^{p-1}\right)^{1/p}$ (298)
$\leq 3e\sqrt{\ln b}$ (299)
where we defined $\|Z\|_{p}:=\mathbb{E}^{1/p}[|Z|^{p}]$. If $b\leq e$, set $p=\infty$ and $p^{\prime}=1$. We get $\mathbb{E}[ZG]\leq 3b$. This establishes the claim for $a=1$. For general $a\geq 1$, consider $Z^{\prime}=Z/a$ and we have $|Z^{\prime}|\leq ba^{-1}$ almost surely. The result follows by applying the $a=1$ case we just proved to $Z^{\prime}$. ∎
Lemma 25.
Suppose that $Z\in\mathbb{R}^{d}$ satisfies $\mathbb{E}[\sum_{i=1}^{d}|Z_{i}|]\leq a$, and $Z\in wB_{q}(r)$ almost surely for some $q>0$. Let $A$ be a matrix with $d$ columns and arbitrarily many rows, where the column vectors satisfy $\|A_{i}\|_{2}\leq 1$ for all $i=1,\dots,d$. Let $G$ be a standard normal vector whose dimension equals the number of rows of $A$. Then we have
$\mathbb{E}[\left<G,AZ\right>]\lesssim rdf(a/rd),$ (300)
and
$\inf_{P_{V|AZ}\colon\|V-AZ\|^{2}\leq\epsilon^{2}}I(V;AZ)$ $\lesssim\frac{1}{\epsilon^{2}}r^{2}d^{2}f^{2}(a/rd).$ (301)
Proof.
Let $g_{i}:=(A^{\top}G)_{i}$, and $a_{i}=\mathbb{E}[|Z_{i}|]$ $i=1,\dots,d$. Since $Z\in wB_{q}(r)$, we have $|Z_{i}|\leq r$, almost surely (this is the only place we use $Z\in wB_{q}(r)$). Then since $f$ is an increasing concave function, and $\operatorname*{var}(g_{i})\leq 1$, we apply Lemma 24 to obtain
$\mathbb{E}[\left<g,Z\right$>] $\lesssim\sum_{j=1}^{d}rf(a_{j}/r)$ (302)
$\leq rdf(a/rd).$ (303)
The claim then follows from Lemma 23. ∎
Lemma 26.
There exists $c_{q}\geq 2$ depending only $q\in(0,1)$, such that the following holds: if $d>c_{q}$, $Z\in wB_{q}(r)$, $\epsilon>0$, and the $\ell_{2}$ norms of the columns of $A$ are not exceeding 1, then we can find $P_{V|AZ}$ such that $\mathbb{E}[\|A(V-Z)\|_{2}^{2}]\leq\epsilon^{2}$, and
$I(V;AZ)\lesssim_{q}\left(\frac{r}{\epsilon}\right)^{\frac{2q}{2-q}}\ln d.$ (304)
Proof.
By Lemma 22, for any $b>0$, we can find $U$ such that $U\in wB_{q}(r)$ almost surely, $\mathbb{E}[\|U-Z\|_{1}]\leq b$, and
$H(U)$ $\lesssim_{q}dg\left(d^{-1}\left(\frac{r}{b}\right)^{\frac{q}{1-q}}\right).$ (305)
For any $u$, let $b_{u}:=\mathbb{E}[\|Z-U\|_{1}|U=u]$, and $\epsilon_{u}:=\sqrt{\frac{b_{u}}{b}}\epsilon$. Note that $Z-U\in wB(2^{1/q}r)$ almost surely. We can apply Lemma 25 to each $u$ to find $P_{V|AZ,U=u}$ such that $\mathbb{E}[\|AZ-V\|_{2}^{2}|U=u]\leq\epsilon_{u}^{2}$ and $I(V;AZ|U=u)\lesssim_{q}\frac{1}{\epsilon_{u}^{2}}r^{2}d^{2}f^{2}(b_{u}/rd)$. Hence
$I(V;AZ|U)$ $\lesssim_{q}\mathbb{E}\left[\frac{1}{\epsilon_{U}^{2}}r^{2}d^{2}f^{2}(b_{U}/rd)\right]$ (306)
$\leq\mathbb{E}\left[\frac{br^{2}d^{2}}{\epsilon^{2}b_{U}}f^{2}(\frac{b_{U}}{rd})\right]$ (307)
$\lesssim_{q}\mathbb{E}\left[\frac{br^{2}d^{2}}{\epsilon^{2}b_{U}}\cdot(\frac{b_{U}}{rd})^{2}\ln\frac{rd}{b_{U}}\right]$ (308)
$\leq\frac{b^{2}}{\epsilon^{2}}\ln\frac{rd}{b}$ (309)
where
  • (308) follows since by (249), we have $\frac{b_{U}}{r}\lesssim_{q}1$ almost surely. By choosing $c_{q}$ large enough, we can ensure that $\frac{b_{U}}{rd}<e^{-1}$.
  • (309) follows by applying Jensen’s inequality to the concave function $t\to t\ln\frac{1}{t}$.
Then (305) and (309) yield
$I(V;AZ)$ $\leq I(U;AZ)+I(V;AZ|U)$ (310)
$\lesssim_{q}dg\left(d^{-1}\left(\frac{r}{b}\right)^{\frac{q}{1-q}}\right)+\frac{b^{2}}{\epsilon^{2}}\ln\frac{rd}{b}.$ (311)
We now consider two cases:
Case I: $2d>\left(\frac{r}{\epsilon}\right)^{\frac{2q}{2-q}}$.
We take $b=r^{\frac{q}{2-q}}\epsilon^{\frac{2-2q}{2-q}}$, which is selected to essentially balance the two terms in (311). Then
$d^{-1}\left(\frac{r}{b}\right)^{\frac{q}{1-q}}=d^{-1}\left(\frac{r}{\epsilon}\right)^{\frac{2q}{2-q}}<2.$ (312)
Since $g(t)\lesssim t\ln\frac{4}{t}$ for $t\in(0,2)$, we have
[…]
$\lesssim_{q}\left(\frac{r}{\epsilon}\right)^{\frac{2q}{2-q}}\ln d.$ (314)
Case II: $2d\leq\left(\frac{r}{\epsilon}\right)^{\frac{2q}{2-q}}$.
We take $b=r^{\frac{q}{2-q}}\epsilon^{\frac{2-2q}{2-q}}(\ln\frac{r}{\epsilon})^{-1/2}$. Then
$\frac{b^{2}}{\epsilon^{2}}\ln\frac{rd}{b}$ $\lesssim_{q}\left(\frac{r}{\epsilon}\right)^{\frac{2q}{2-q}}\frac{\ln d+\ln\frac{r}{\epsilon}}{\ln\frac{r}{\epsilon}}$ (315)
$\lesssim\left(\frac{r}{\epsilon}\right)^{\frac{2q}{2-q}}\ln d,$ (316)
and
$dg\left(d^{-1}\left(\frac{r}{b}\right)^{\frac{q}{1-q}}\right)$ $=dg\left(d^{-1}\left(\frac{r}{\epsilon}\right)^{\frac{2q}{2-q}}(\ln\frac{r}{\epsilon})^{\frac{q}{2(1-q)}}\right)$ (317)
$\lesssim_{q}d\ln\left(d^{-1}\left(\frac{r}{\epsilon}\right)^{\frac{2q}{2-q}}(\ln\frac{r}{\epsilon})^{\frac{q}{2(1-q)}}\right).$ (318)
Let $M:=d^{-1}\left(\frac{r}{\epsilon}\right)^{\frac{2q}{2-q}}\geq 2$. Then
[…]
$\lesssim_{q}\left(\frac{r}{\epsilon}\right)^{\frac{2q}{2-q}}\ln d$ (320)
where the last step is justified as $\frac{\ln M}{M}\lesssim 1$ and $\frac{1}{M}\ln\ln\frac{r}{\epsilon}\lesssim\frac{1}{M}\ln\frac{r}{\epsilon}\lesssim_{q}\frac{\ln d+\ln M}{M}\lesssim\ln d$. Therefore in either case, we have established (304), as desired. ∎
We are now ready to prove the key estimate of the width-moment function:
Proof of Lemma 20.
From Lemma 26 and Theorem 1 we obtain
$\mathsf{w}(P_{AZ})$ $\lesssim\int_{0}^{\sqrt{E}}\sqrt{R_{P_{AZ}}(\sigma^{2})}d\sigma$ (321)
$\lesssim_{q}\int_{0}^{\sqrt{E}}\left(\frac{r}{\sigma}\right)^{\frac{q}{2-q}}\sqrt{\ln d}\,d\sigma$ (322)
$\lesssim_{q}\sqrt{\ln d}r^{\frac{q}{2-q}}E^{\frac{1-q}{2-q}}.$ (323)
A weaker version of Theorem 7, which controls the mean square error rather than providing a high-probability bound, can now be obtained by taking the expectation of both sides of (24):
Theorem 12.
$\frac{1}{n}\mathbb{E}[\|\mathbb{X}\hat{\beta}-\mathbb{X}\beta\|_{2}^{2}]\lesssim_{q}(\frac{\ln d}{n})^{\frac{2-q}{2}}r^{q}$.
Proof of Theorem 12.
Recall from (24) that
$\|\hat{m}-m\|^{2}\leq\frac{1}{\sqrt{n}}\chi_{m-\hat{m}}$ (324)
where $\chi_{z}:=-2\left<z,w\right$>, $z\in 2A(wB_{q}(r))$ is a centered sub-Gaussian process, whose metric is the $\ell_{2}$ metric (up to universal constant). Let $Z:=m-\hat{m}$ and $E:=\mathbb{E}[\|Z\|_{2}^{2}]$. From Lemma 20 we obtain
$\mathbb{E}[\chi_{m-\hat{m}}$] $\lesssim_{q}\sqrt{\ln d}r^{\frac{q}{2-q}}E^{\frac{1-q}{2-q}}.$ (325)
Then, using (324), we have
$\mathbb{E}[\|\hat{m}-m\|^{2}]\lesssim_{q}\sqrt{\frac{\ln d}{n}}r^{\frac{q}{2-q}}\mathbb{E}^{\frac{1-q}{2-q}}[\|\hat{m}-m\|^{2}],$ (326)
or equivalently,
$\mathbb{E}[\|\hat{m}-m\|^{2}]\lesssim_{q}(\frac{\ln d}{n})^{\frac{2-q}{2}}r^{q}.$ (327)
The proof of Theorem 7 is similar to that of Theorem 12, but we use the technique of Section 3.3 instead, which controls the moment generation function of both sides of (24):
Proof of Theorem 7.
We provide the details of the calculation based on Theorem 4, where $Y$ is assumed to satisfy the log-Sobolev inequality. The proof for $Y$ satisfying $\|Y\|_{\Psi_{2}}\leq 1$ is similar, but uses Theorem 5 instead. Using Lemma 20, and the fact that $wB_{q}(r)-wB_{q}(r)\subseteq 2^{1/q}wB_{q}(r)$, we have
$\frac{1}{\sqrt{n}}\mathsf{G}(E)\lesssim_{q}BE^{\frac{1-q}{2-q}}$ (328)
where
$B:=\sqrt{\frac{\ln d}{n}}R^{-\frac{1}{2-q}}.$ (329)
We see that $\mathsf{G}$ is a sub-additive function. Therefore
$\psi(n/72)$ $\leq\sup_{\epsilon>0}\{-\frac{\epsilon}{3}+\frac{\sqrt{n}}{36}\mathsf{G}(\frac{72\epsilon}{n})\}+\frac{\sqrt{n}}{36}\mathsf{G}(\frac{72\psi(n/72)}{n})$
$\quad+\sup_{\epsilon>0}\{-\frac{\epsilon}{3}+\frac{\epsilon}{3}\}+\sup_{\epsilon>0}\{-\frac{\epsilon}{3}+\frac{\sqrt{\epsilon\psi(n/72)}}{3}\}.$ (330)
Let $I_{1},I_{2},I_{3},I_{4}$ denote the four terms in (330). Then $I_{3}=0$, and the other terms can be controlled using (328):
$I_{1}\lesssim_{q}\sup_{\epsilon>0}\{-\epsilon+Bn(\epsilon/n)^{\frac{1-q}{2-q}}\}\lesssim_{q}B^{2-q}n;$ (331)
$I_{2}\lesssim_{q}Bn\left(\frac{\psi(n/72)}{n}\right)^{\frac{1-q}{2-q}};$ (332)
$I_{4}\leq\frac{1}{12}\psi(n/72).$ (333)
Thus we must have either $\frac{1}{3}\psi(n/72)\leq I_{1}$ or $\frac{1}{3}\psi(n/72)\leq I_{2}$, which shows that $\psi(n/72)\leq c_{q}B^{2-q}n$ for some $c_{q}>0$ that may depend on $q$.
Now by the Markov inequality (with $\lambda=n/72$) we have, for any $t>0$,
$\mathbb{P}[\|m-\hat{m}\|_{2}>t]\leq\exp(c_{q}nB^{2-q}-nt^{2}/72).$ (334)
The result then follows by taking $t=\sqrt{72}(\sqrt{c_{q}}+1)B^{\frac{2-q}{2}}$. ∎
11 Acknowledgment
The author would like to thank Professor Ramon van Handel for feedback on an early version of the manuscript, and for encouraging the pursuit of a new proof of the majorizing measure theorem based on the rate-distortion integral. The author also thanks Professor Garvesh Raskutti and Philippe Rigollet for discussion and feedback on sparse linear regression. This research was supported in part by NSF Grant DMS-2515510.

### v3

@@
 Abstract
-We develop an information-theoretic framework for bounding the expected supremum and tail probabilities of stochastic processes, offering a simpler and sharper alternative to classical chaining and slicing arguments for generalization bounds.
+We develop an information-theoretic framework for bounding the supremum of stochastic processes, offering a simpler and sharper alternative to classical chaining and slicing arguments for generalization bounds.
 The key idea is a lifting argument that produces information-theoretic analogues of empirical process bounds, such as Dudley’s entropy integral.
-The lifting introduces symmetry, yielding sharp bounds even when the classical Dudley integral is loose.
-As a by-product, we obtain a concise proof of the majorizing measure theorem, providing explicit constants.
-The information-theoretic approach provides a soft version of classical localized complexity bounds in generalization theory, but is more concise and does not require the slicing argument.
-We apply this approach to empirical risk minimization over Sobolev ellipsoids and weak $\ell_{q}$ balls, obtaining sharper convergence rates or extensions to settings not covered by classical methods.
+Lifting introduces permutation symmetry, yielding sharp bounds when the classical Dudley integral is loose.
+This gives a simple proof of the majorizing measure theorem via the sharpness of Dudley’s entropy integral for stationary processes, a result known well before the proof of the majorizing measure theorem.
+Furthermore, the information-theoretic formulation provides soft versions of classical localized complexity bounds in generalization theory, but is simpler and does not require the slicing argument.
+We apply this approach to empirical risk minimization over Sobolev ellipsoids, obtaining sharp convergence rates in settings where previous methods are suboptimal.
 1 Introduction
 Sharp control of empirical processes is central to modern statistical learning theory, underpinning generalization bounds, minimax rates, and risk analysis for high-dimensional estimators.
-Classical tools such as Dudley’s entropy integral have played a central role in this analysis, with applications ranging from nonparametric regression (Van de Geer,, 1990; Wainwright,, 2019) to sparse high-dimensional models (Raskutti et al.,, 2011).
-Although certain special cases can be handled by simpler techniques, the localized version of Dudley’s integral (Van de Geer,, 1990) has remained a fundamental tool in many general settings.
+Classical tools such as Dudley’s entropy integral have played a central role in this analysis, with applications ranging from nonparametric regression [41, 44] to sparse high-dimensional models [35].
+Although certain special cases can be handled by simpler techniques, the localized version of Dudley’s integral [41] has remained a fundamental tool in many general settings.
 However, its application to statistical learning may still be limited since (i) Dudley’s integral, while widely used, can be suboptimal even for basic examples such as Sobolev ellipsoids, and (ii) the slicing (or peeling) technique for localization often requires delicate covering number estimates, sometimes invoking deep results from approximation theory or the local theory of Banach spaces, e.g.
-in sparse regression over the $\ell_{q}$ ball (Raskutti et al.,, 2011).
+in sparse regression over the $\ell_{q}$ ball [35].
 In this paper, we introduce a “lifting” method for deriving new information-theoretic generalization bounds, which overcomes the limitations discussed above.
@@
 The strength of the lifting approach lies in leveraging geometric insights involving covering and packing numbers, such as the chaining argument in Dudley’s integral, yielding new information-theoretic inequalities that are not easily derived from traditional information-theoretic methods.
-This method yields sharper localized generalization bounds and new results for empirical risk minimization over ellipsoids and sparse regression.
+This method yields sharper localized generalization bounds and new results for empirical risk minimization in some settings.
 For example, consider a Gaussian process $(X_{t})_{t\in T}$ and a measure $\mu$ on $T$.
@@
 $\sup\mathbb{E}[X_{Z}]\asymp\int_{0}^{\infty}\sqrt{\inf_{P_{Z\hat{Z}}\in\Pi_{\sigma}(\mu)}I(Z;\hat{Z})}d\sigma\asymp\int_{0}^{\infty}\sqrt{R_{\mu}(\sigma^{2})}d\sigma$ (1)
-where $\Pi_{\sigma}(\mu)$ denotes the set of couplings of $\mu$ and itself with mean square error $\sigma^{2}$, I(Z;Z^)I($\hat{Z}$) is the mutual information, and $R_{\mu}(\sigma^{2})$ denotes the rate-distortion function (Thomas and Joy,, 2006).
-The supremum in (1) is over all couplings of the index and the process satisfying the marginal condition $P_{Z}=\mu$.
-While the analogy between the rate-distortion function $R_{\mu}(\sigma^{2})$ and covering numbers is well-known (Thomas and Joy,, 2006; Csiszár and Körner,, 2011), we are not aware of any prior literature explicitly stating (1).
-Previously, similar lifting arguments were used for problems in selected problems from network information theory (Wu et al.,, 2018; Liu,, 2022, 2021, 2023), but the application to empirical processes appears to be new.
-More generally, “soft” versions of traditional measure concentration techniques are known to provide sharper convergence in some hypothesis testing and information theory problems (Liu,, 2018; Liu et al.,, 2020; Liu,, 2020; Liu and Özgür,, 2020).
+where $\Pi_{\sigma}(\mu)$ denotes the set of couplings of $\mu$ and itself with mean squared error $\sigma^{2}$, I(Z;Z^)I($\hat{Z}$) is the mutual information, and $R_{\mu}(\sigma^{2})$ denotes the rate-distortion function (Definition 3).
+The supremum in (1) is over all conditional distributions of the index $Z$ given the process such that $P_{Z}=\mu$.
+While the analogy between the rate-distortion function $R_{\mu}(\sigma^{2})$ and covering numbers is well-known [12, 13], we are not aware of any prior literature explicitly stating (1).
+Previously, similar lifting arguments were used for problems in selected problems from network information theory [46, 26, 25, 27], but the application to empirical processes appears to be new.
+More generally, “soft” versions of traditional measure concentration techniques are known to provide sharper convergence in some hypothesis testing and information theory problems [23, 30, 24, 28].
 We use information-theoretic inequalities such as (1) to control the generalization error in empirical risk minimization (ERM), and this approach offers several advantages over the traditional approach based on slicing and Dudley’s integral:
 First, our approach yields sharper convergence rate in some basic estimation problems.
-The reason is that (1) is a sharp two-sided inequality for any $\mu$, which might be surprising since we derive (1) from the Dudley integral, and the latter is not sharp for some $T$.
-The reason for the sharpness of (1) is that the lifting argument transforms the process into a stationary one - i.e., invariant under a transitive group of permutations - even if the original process lacks this symmetry.
+The reason is that (1) is a sharp two-sided inequality for any $\mu$, which is perhaps surprising, since (1) is derived from the Dudley integral, and the latter is not sharp for some $T$.
+The reason for the sharpness of (1) is that the lifting argument transforms a general Gaussian process into a stationary one (i.e., invariant under the transitive group of permutations; see Definition 1), even if the original process lacks this symmetry.
 Since Dudley’s integral is sharp for stationary processes (see Section 2.1 for background), (1) inherits the sharpness, even if the original process is nonstationary.
@@
 This new proof of the majorizing measure theorem is conceptually simple, as it closely mirrors the argument used for the Dudley integral upper bound, essentially reversing its key steps.
-Furthermore, many nuisance parameters in the traditional proof disappear due to taking limits in the lifting argument.
+It may also be interpreted as saying that Fernique’s result about sharpness of Dudley’s inequality for stationary processes [17] already contains, in essence, Talagrand’s majorizing measure theorem [37], once combined with relatively standard information-theoretic techniques.
 Figure 1 illustrates the implication relations of these results.
@@
 Figure 1: Arrows indicate implications of results, and the techniques/assumptions for the implications.
-For Gaussian processes, the majorizing measure theorem and the proposed rate-distortion integral are two-sided bounds.
-Dudley’s integral is only an upper bound, unless stationarity is assumed.
+For Gaussian processes, the majorizing measure theorem and the proposed rate-distortion integral are two-sided bounds, whereas Dudley’s integral is only an upper bound unless stationarity is assumed.
 Second, our approach makes localization of generalization bounds more straightforward.
@@
 For example, in empirical risk minimization (ERM), $Z$ denotes the estimator, which tends to concentrate near zero as the sample size increases.
-Traditionally, localization is achieved by the “slicing” (or “peeling”) argument, which yields sharp convergence rate for some empirical risk minimization problems Wainwright, (2019).
+Traditionally, localization is achieved by the “slicing” (or “peeling”) argument, which yields sharp convergence rate for some empirical risk minimization problems [44].
 Slicing typically involves controlling the modulus of continuity of the process through careful controls of the deviation magnitude for different slices.
@@
 This new method of localization does not require slicing.
-To demonstrate our approach, we derive new and sharper statistical rates in two classical ERM problems:
-  • For mean estimation using ERM over the Sobolev ellipsoid with smoothness parameter $\beta>1/2$, which plays a fundamental role in nonparametric statistics (Tsybakov,, 2009), the slicing method based on Dudley’s integral only gives an upper bound $O(\frac{1}{2\beta-1})n^{-\frac{2\beta}{2\beta+1}}$ for the mean square error.
-The factor $\frac{1}{2\beta-1}$ arises in Dudley’s integral for the supremum over the Sobolev ellipsoid, a classical example for the looseness of Dudley’s integral (Talagrand,, 2014).
-On the other hand, our method achieves the optimal $O(\frac{1}{\sqrt{2\beta-1}})n^{-\frac{2\beta}{2\beta+1}}$ bound.
-  • For sparse regression over the $\ell_{q}$-ball, $q\in(0,1]$, Raskutti et al.
-(Raskutti et al.,, 2011) showed that the mean square prediction error is upper bounded by $(\frac{\ln d}{n})^{\frac{2-q}{2}}R$ up to a constant depending on $q$, where $d$ is the feature dimension, $n$ is the sample size, $R^{1/q}$ is the $\ell_{q}$ radius, the noise is standard normal, and the columns of the design matrix are normalized to have $\ell_{2}$ norm at most $\sqrt{n}$.
-The proof in Raskutti et al., (2011) was based on the slicing method and estimating covering numbers of the $q$-absolutely convex hull, using deep results from convex geometry (Guedon and Litvak,, 2000; Carl and Pajor,, 1988); see proof of (Raskutti et al.,, 2011, Lemma 3).
-We provide a new proof without citing the convex geometry literature, using only simple properties of the entropy, such as the chain rule, to extend elementary one-dimensional estimates to higher dimensions.
-In fact, we show that the same convergence rate holds for weak $\ell_{q}$ balls ($q\in(0,1)$), which is new and appears to be beyond the reach of prior techniques in the literature.
-Related work: A closely related line of recent research, started by Russo and Zou, (2016) and further developed by other authors including Xu and Raginsky, (2017); Dziugaite and Roy, (2017); Asadi et al., (2018); Bu et al., (2020); Aminian et al., (2023), proved information-theoretic generalization bounds in terms of the mutual information between the random index and the process.
-A basic version (see for example Russo and Zou, (2016)) is
-$\mathbb{E}[X_{Z}]\lesssim\sup_{t\in T}\|X_{t}\|_{\Psi_{2}}\sqrt{I(Y;Z)}$ (2)
-where the sub-Gaussian norm $\|\cdot\|_{\Psi_{2}}$ will be defined in (6), and $Y:=(X_{t})_{t\in T}$ denotes the stochastic process (in the case of finite $T$, we have $Y\in\mathbb{R}^{|T|}$).
-While this can be interpreted as a strengthening of a maximal inequality (Jiao et al.,, 2018), further refinement using the chaining argument, strengthening the Dudley integral, is also possible (Asadi et al.,, 2018).
+To demonstrate our approach, we obtain new results for mean estimation via ERM when the parameter set is the image of the Sobolev ellipsoid under a Lipschitz map.
+Sobolev ellipsoid play a fundamental role in nonparametric statistics [40], whereas the Lipschitz map may represent a nonlinear feature map and is a technically challenging ingredient that necessitates the majorizing measure theorem machinery (see sub-Gaussian comparison (10)).
+Assuming smoothness parameter $\beta>1/2$, the traditional slicing method based on Dudley’s integral [41] yields an upper bound of order $\tfrac{1}{2\beta-1}n^{-2\beta/(2\beta+1)}$ for the mean squared error.
+In contrast, our method achieves the sharper rate $\tfrac{1}{\sqrt{2\beta-1}}n^{-2\beta/(2\beta+1)}$, which is beyond previous techniques (when the Lipschitz map is involved).
+Note that the factor $\tfrac{1}{2\beta-1}$ also appears in Dudley’s integral for the supremum of a Gaussian process indexed by the Sobolev ellipsoid, a classical example of the looseness of Dudley’s bound [39], where the optimal dependence is order $\tfrac{1}{\sqrt{2\beta-1}}$.
+In [arXiv:2508.18682v2], we further present lifting techniques for tail probability bounds, and application to sparse regression, recovering results of Raskutti et al.
+[35] for $\ell_{q}$ balls, and further extending the rates to weak $\ell_{q}$-balls.
+The results for weak $\ell_{q}$-balls appears to be beyond prior techniques.
+Further related work: Connections between ERM and upper bounds of stochastic processes, especially characterizing the minimax rate as the solution to a fixed point equation involving the localized Gaussian complexity, can be traced back to [41].
+In the random design setting where the symmetrization argument applies, the localized Rademacher complexity was developed by [31, 4, 19].
+An alternative approach for bounding the ERM risk using anti-concentration (small ball probability) rather than concentration was developed in [21, 34], which can be used in cases beyond sub-Gaussian noise.
+On the other hand, for convex regression classes with Gaussian noise, concentration has been used to show that the mean-squared error (MSE) is tightly bounded by the solution to an equation involving a localized Gaussian width [7, 45].
+These latter methods crucially rely on convexity of the regression class, hence do not apply to some statistical examples we consider (see Remark 2).
+Talagrand’s majorizing measure theorem [37] provides a sharp “geometric” characterization of the supremum of a Gaussian process in term of the $\gamma_{2}$ functional (see definition in (9)).
+Talagrand also developed several equivalent characterizations of $\gamma_{2}$, including admissible nets, tree-packing and labeled nets; see [39, 42] and the references therein.
+In particular, the labeled net (called “index” of an atom in the original paper [38, Theorem 2.1]) is closely related to the length of a uniquely decodable code, hence has an information-theoretic interpretation.
+This connection was clarified and further explored through the Kraft-McMillan inequality by [32, 9].
+However, none of these prior works established (1).
+While general bounds for empirical processes and PAC-Bayes were derived based on $\gamma_{2}$ or its related forms [34, 3, 20], its application to computing minimax rates for specific classes appears to be limited, since $\gamma_{2}$ can often be more difficult to compute than directly evaluating the supremum of a process [38].
+A closely related line of research, started by [36] and further developed by other authors including [47, 16, 2, 5, 1], proved information-theoretic generalization bounds in terms of the mutual information between the random index and the process.
+A basic version (see for example [36]) is
+$\mathbb{E}[X_{Z}]\lesssim\sup_{t\in T}\|X_{t}\|_{G}\sqrt{I(Y;Z)}$ (2)
+where the exact sub-Gaussian norm $\|\cdot\|_{G}$ will be defined in (4), and $Y:=(X_{t})_{t\in T}$ denotes the stochastic process (in the case of finite $T$, we have $Y\in\mathbb{R}^{|T|}$).
+While this can be interpreted as a strengthening of a maximal inequality [18], further refinement using the chaining argument, strengthening the Dudley integral, is also possible [2].
 This type of results are useful in statistics and machine learning, where $Z$ represents the output of an algorithm which has limited mutual information with the training data, potentially much smaller than the covering number of $T$.
@@
 Nevertheless, we will show in Theorem 2 a result that encompasses both (1) and (2), highlighting the versatility of the lifting technique.
-In related work on empirical risk minimization (ERM), it has been shown that under favorable conditions, such as convex regression classes with Gaussian noise, the mean-squared error (MSE) can often be characterized by a fixed-point equation involving a localized Gaussian width (Chatterjee,, 2014).
-For ellipsoids, this localized Gaussian width can then be controlled directly via the Cauchy–Schwarz inequality (Wei et al.,, 2020), thereby avoiding the slicing and Dudley-integral machinery and mitigating the looseness of Dudley’s bound.
-However, this shortcut relies essentially on concentration of the error, which can fail when the regression class is nonconvex (e.g., sparse regression under the $\ell_{q}$ constraint with $q<1$) or when the noise is sub-Gaussian but does not satisfy a measure concentration inequality (such as log-Sobolev or transportation-cost inequality).
-Moreover, the Cauchy-Schwarz step exploits linear structure: for example, consider a regression class given by the image of a Sobolev ellipsoid under a Lipschitz map (e.g., induced by a nonlinear neural network).
-This generalizes the setting of Wei et al., (2020), but their argument no longer applies.
-These limitations are expected to be artifacts of the proof: tail probability bounds need not require concentration (as naturally follows from the Dudley integral; see Lemma 2), and Lipschitz maps do not increase the supremum of the relevant stochastic process (a consequence of the majorizing measure theorem; see (13)).
-Unfortunately, existing tools do not suffice to rigorously establish these intuitions, and our work introduces an alternative framework that overcomes these limitations.
-Organization: The paper is organized as follows: Section 2 provides more background on upper bounds for stochastic processes from probability theory, the traditional slicing technique for ERM, and information-theoretic tools used in our proof.
-Section 3 presents our main results, including (1), its sharpness for Gaussian processes, a simple proof of the classical majorizing measure theorem based on (1), a general framework for error tail bounds in ERM including results on the ellipsoid and sparse regression examples mentioned above.
-The remainder consists of details and proofs supporting the main results.
+Finally, another direction of extending the theory about supremum of processes, considered recently by [28, 27, 10], generalizes the supremum to a soft-max, also known as free energy, which recovers the supremum as the inverse temperature tends to infinity.
+It seems an interesting future work to extend the lifting approach to bound the soft-max, and we reported preliminary reports in [27].
+Organization: Section 2 recollects standard facts about upper bounds for stochastic processes, its applications to ERM, and relevant information-theoretic tools.
+Section 3 presents the main results: the rate-distortion integral, its sharpness, and an example of application to ERM over the Sobolev ellipsoid under a 1-Lipschitz map.
+Section 4-5 prove the upper and lower bounds for the rate-distortion integral.
+Section 6 presents an equivalent form of the rate-distortion integral which yields a new proof of the majorizing measure theorem, and Section 7 further applies it to the ellipsoid example.
 2 Preliminaries
-Notation: We use $\wedge$ and $\vee$ to denote the max and min of two numbers, and $x]_{+}:=x\vee 0$.
+Notation: $\wedge$ and $\vee$ to denote the max and min of two numbers, and $x]_{+}:=x\vee 0$.
 Given probability measures $P$ and $Q$ on a metric space $(T,d)$, $W_{p}(P,Q):=\sup_{P_{XY}\in\Pi(P,Q)}\mathbb{E}^{\frac{1}{p}}[d(X,Y)^{p}]$ denotes the Wasserstein distance, where $\Pi(P,Q)$ denotes the set of couplings.
 For brevity, $\Pi(P):=\Pi(P,P)$.
-Product measures, such as $P\times Q$ or $P^{\otimes N}$, indicate independence.
-${\rm diam}(T):=\sup_{x,y\in T}d(x,y)$ denotes the diameter of a set $T$.
-Standard Landau notations such as $o(\dots)$ are used.
-$\lesssim_{q}$ means less than or equal up to a constant depending on $q$, and $\asymp_{q}$ means $\lesssim_{q}$ and $\gtrsim_{q}$ are both true.
-For any $q>0$, the $\ell_{q}$ ball and the weak $\ell_{q}$ ball of radius $r$ centered at 0 are
-$B_{q}(r)$ $:=\{x\in\mathbb{R}^{d}\colon\sum_{i=1}^{d}x_{i}^{q}\leq r^{q}\},$ (3)
-$wB_{q}(r)$ $:=\{x\in\mathbb{R}^{d}\colon\sup_{t>0}t^{q}\sum_{i=1}^{d}1_{|x_{i}|>t}\leq r^{q}\}.$ (4)
-More generally, $B_{q}(z,r):=z+B_{q}(r)$ if the center is $z$.
-For brevity, $B(z,r):=B_{2}(z,r)$.
-$\ell_{2}$ denotes the Hilbert space of square-summable infinite sequences.
+Product measures are denoted by $P\times Q$ or $P^{\otimes N}$.
+The diameter of a set $T$ is ${\rm diam}(T):=\sup_{x,y\in T}d(x,y)$.
+$\lesssim$ means less than or equal to up to a universal constant, and $\asymp$ means both $\lesssim$ and $\gtrsim$.
+We denote by $\ell_{2}$ the Hilbert space of square-summable infinite sequences, equipped by the norm $\|\cdot\|_{2}$, abbreviated as $\|\cdot\|$.
 The empirical distribution of a sequence $x=(x(1),x(2),\dots,x(d))$ is denoted by $\widehat{P}_{x}$.
-Given a set $T$ in $\mathbb{R}^{d}$ and a matrix $A\in\mathbb{R}^{n\times d}$, define $T-T:=\{t-s\colon t,s\in T\}$ and $AT:=\{At\colon t\in T\}$.
+Given a set $T$ in $\mathbb{R}^{d}$ and a matrix $A\in\mathbb{R}^{n\times d}$, define $T-T:=\{t-s\colon t,s\in T\}$.
 2.1 Upper bounds for stochastic processes
 Classically, Dudley’s integral provides upper bounds on a sub-Gaussian process in terms of the covering numbers.
-Let $(T,d)$ be a metric space, and suppose that $(X_{t})_{t\in T}$ is a centered sub-Gaussian process, that is, $\mathbb{E}[X_{t}]=0$, and the natural metric
-$d(t,t^{\prime}):=\|X_{t}-X_{t^{\prime}}\|_{\Psi_{2}},\quad\forall t,t^{\prime}\in T,$ (5)
-where, following Vershynin, (2018), we defined the sub-Gaussian norm
-$\|X\|_{\Psi_{2}}=\inf\{K>0\colon\mathbb{E}[e^{X^{2}/K^{2}}]\leq 2\}.$ (6)
-Dudley’ integral states that
-$\mathbb{E}[\sup_{t\in T}X_{t}]\leq K\int_{0}^{\Delta}\sqrt{\ln N(T,\lambda)}d\lambda$ (7)
-where $K$ is a universal constant, $N(T,\lambda)$ denotes the $\lambda$-covering number of $T$, i.e.
+Following [43, Exercise 2.40], we define the sub-gaussian variance of a random variable $X$ as
+${\rm var}_{G}(X):=\inf\{\sigma^{2}\colon\mathbb{E}[e^{\lambda(X-\mathbb{E}X)}]\leq e^{\sigma^{2}\lambda^{2}/2}\textrm{ for all $\lambda\in\mathbb{R}$}\},$ (3)
+and define the exact sub-Gaussian norm by
+$\|X\|_{G}^{2}={\rm var}_{G}(X)+(\mathbb{E}[X])^{2}.$ (4)
+This is one of the several equivalent (up to universal constant) definitions of the sub-Gaussian norm; see [43, Proposition 2.6.1].
+Let $(X_{t})_{t\in T}$ be a centered (i.e., $\mathbb{E}[X_{t}]=0$) sub-Gaussian process, and define the natural metric
+$d(t,t^{\prime}):=\|X_{t}-X_{t^{\prime}}\|_{G},\quad\forall t,t^{\prime}\in T.$ (5)
+Then $(T,d)$ is a metric space. Dudley’ integral states that
+$\mathbb{E}[\sup_{t\in T}X_{t}]\leq K\int_{0}^{\Delta}\sqrt{\ln\mathsf{N}(T,\lambda)}d\lambda$ (6)
+where $K$ is a universal constant, $\mathsf{N}(T,\lambda)$ denotes the $\lambda$-covering number of $T$, i.e.
 the minimal number of (open) balls of radius $\lambda$ required to cover $T$, and $\Delta$ denotes the diameter of $T$.
-While usually called Dudley’s integral, (7) is occasionally attributed to other authors as well; see Dudley, (2016); Maurey, (2024); Talagrand, (2014) for more about the history.
-In the case of Gaussian processes, the natural metric is defined as
-$d(t,t^{\prime}):=\sqrt{\mathbb{E}[|X_{t}-X_{t^{\prime}}|^{2}]}.$ (8)
-Note that (8) differs from (5) by a universal constant factor (Vershynin,, 2018).
-Throughout this paper, when we discuss Gaussian processes specifically, we define the natural metric by (8) rather than (5).
-In the Gaussian case, a commonly referenced explicit bound is $K<24$ (Ledoux and Talagrand,, 2013).
-As pointed out in (Dudley,, 2016, Section 5.3), better estimates $K$ exist in the literature, and the proof essentially relies on the estimate
-$\mathbb{E}[\max_{1\leq j\leq N}X_{j}]\leq\sqrt{2\ln N}$ (9)
-where $X_{1},\dots,X_{N}$ are centered, jointly Gaussian, with variances not exceeding 1.
-If we drop Gaussianity and instead assume that $\|X_{1}\|_{\Psi_{2}},\dots,\|X_{N}\|_{\Psi_{2}}\leq 1$, then $\mathbb{E}[\max_{1\leq j\leq N}X_{j}]\leq\sqrt{2\ln(2N)}$ for $N\geq 1$ (Vershynin,, 2018, Section 2.7.3), which is slightly worse than the Gaussian case (9).
-We will see, however, that the difference will vanish when we apply the lifting argument, which effectively takes $N\to\infty$.
-For stationary Gaussian processes, it is known since Fernique, (1975) that Dudley’s integral is tight:
-$\int_{0}^{\Delta}\sqrt{\ln N(T,\lambda)}d\lambda\leq K^{\prime}\mathbb{E}[\sup_{t\in T}X_{t}]$ (10)
+While usually called Dudley’s integral, (6) is occasionally attributed to other authors as well; see [15, 33, 39] for more about the history.
+The Gaussian process is a special case of the sub-Gaussian process where $(X_{t_{1}},X_{t_{2}},\dots,X_{t_{n}})$ follows the Gaussian distribution for any $t_{1},\dots,t_{n}$.
+For Gaussian processes, a commonly referenced explicit bound is $K<24$ [22].
+Recall the following definition of stationary Gaussian processes (see e.g. [42, 39]):
+Definition 1.
+The Gaussian process $(X_{t})_{t\in T}$ is called stationary if there exists a group $G$ acting on $T$ such that
+  1. 1.
+    $d(g(t),g(s))=d(t,s)$ for all $t,s\in T$, $g\in G$ (translation invariance).
+  2. 2.
+    For every $t,s\in T$, there exists $g\in G$ such that $t=g(s)$ (transitivity).
+It is well-known that Dudley’s integral is not sharp (i.e., not a two-sided inequality up to universal constants).
+However, for stationary Gaussian processes, it is known since Fernique [17] that Dudley’s integral is sharp:
+$\int_{0}^{\Delta}\sqrt{\ln\mathsf{N}(T,\lambda)}d\lambda\leq K^{\prime}\mathbb{E}[\sup_{t\in T}X_{t}]$ (7)
 for a universal constant $K^{\prime}>0$.
-Maurey, (2024) gives a bound of $K^{\prime}\leq 432$.
-In the stationary case, the smaller balls in the chaining argument are equivalent in each level, so the chaining process in the proof of the Dudley integral can be reversed using the so-called super-Sudakov minoration (see Lemma 10 ahead).
-For general nonstationary processes, the same paper of Fernique (Fernique,, 1975) conjectured what was later known as the majorizing measure theorem, whose nontrivial lower bound part was proved more than a decade later by Talagrand (Talagrand,, 1987) using a more sophisticated generic chaining argument.
+[33] provides an exposition of the stationary process story and gives a bound of $K^{\prime}\leq 432$.
+The stationary case is particularly convenient, since the smaller balls in the chaining argument are equivalent in each level, so that the chaining process in the proof of the Dudley integral can be reversed (see [42] for discussions).
+For general nonstationary Gaussian processes, the same paper of Fernique [17] conjectured what was later known as the majorizing measure theorem, whose nontrivial lower bound part was proved more than a decade later by Talagrand [37] using a more sophisticated generic chaining argument.
 Given a probability measure $\mu$ on $T$, define
-$I_{\mu}(t):=\int_{0}^{\Delta}\sqrt{\ln\frac{1}{\mu(B(t,\lambda))}}d\lambda$ (11)
+$I_{\mu}(t):=\int_{0}^{\Delta}\sqrt{\ln\frac{1}{\mu(B(t,\lambda))}}d\lambda.$ (8)
 The majorizing measure theorem states that the left side of (6) equals, up to a universal constant factor,
 $\gamma_{2}(T):=\inf_{\mu}\sup_{t\in T}I_{\mu}(t).$ (9)
-The usual definition of $\gamma_{2}(T)$ in terms of admissible partitions (cf.
-Talagrand, (2014)) is equivalent to (12) up to a constant factor.
+The usual definition of $\gamma_{2}(T)$ in terms of admissible partitions (developed later by Talagrand in the 90s cf.
+[39]) is equivalent to (9) up to a constant factor.
 Generally speaking, $\gamma_{2}(T)$ is more difficult to evaluate than Dudley’s integral: $\mu$ needs to be optimized for a quantity depending on $\mu(B(t,\lambda))$ over different scales of $\lambda$, whereas in Dudley’s integral, we can calculate the covering number by optimizing for each scale separately.
@@
 $\mathbb{E}[\sup_{t\in T}X_{t}]\lesssim\mathbb{E}[\sup_{t\in T}Y_{t}].$ (10)
-The special case where $(X_{t})$ is also Gaussian is called the Gaussian comparison inequality, usually proved by an interpolation argument which is highly Gaussian-specific (Ledoux and Talagrand,, 2013).
+The special case where $(X_{t})$ is also Gaussian is called the Gaussian comparison inequality, usually proved by an interpolation argument which is highly Gaussian-specific [22].
 The more general sub-Gaussian case can be immediately seen from the majorizing measure theorem.
-As noted in Van Handel, (2014), this version of sub-Gaussian comparison is a nontrivial consequence of the majorizing measure theorem, for which no other proof is known.
-It is also known (see Talagrand, (2014)) that the following quantities are equivalent to $\gamma_{2}(T)$ up to constants:
+As noted in [42], this version of sub-Gaussian comparison is a nontrivial consequence of the majorizing measure theorem, for which no other proof is known.
+It is also known (see [39]) that the following quantities are equivalent to $\gamma_{2}(T)$ up to constants:
 ${\rm Fer}(T)$ $:=\sup_{\mu}\int_{T}I_{\mu}(t)d\mu(t);$ (11)
 $\delta_{2}(T)$ $:=\sup_{\mu}\inf_{t\in T}I_{\mu}(t).$ (12)
-Once the expected value is understood, tail bounds can be established either by applying concentration of measure, or by reworking the chaining steps to control the tail probability instead of the expectation.
-For technical convenience, our discussion of tail bounds will be restricted to processes of the form $X_{t}=\left<Y,t\right$>, where $t\in T\subseteq\ell_{2}$.
-Following Vershynin, (2018), the sub-Gaussian norm of a random vector $Y\in\ell_{2}$ is defined by taking the supremum over one-dimensional projections:
-$\|Y\|_{\Psi_{2}}:=\sup_{v\in\ell_{2}}\|\left<Y,v\right>\|_{\Psi_{2}}.$ (16)
-A bound on $\|Y\|_{\Psi_{2}}$ can be guaranteed, for example, if $Y$ has independent coordinates with uniformly bounded sub-Gaussian norms (Vershynin,, 2018).
-Concentration inequalities control the deviation probability from the mean.
-We say $P_{Y}$ satisfies the log-Sobolev inequality (LSI), if for any continuously differentiable function $f\colon\ell_{2}\to\mathbb{R}$,
-${\rm Ent}(f^{2})\leq 2\mathbb{E}[\|\nabla f(Y)\|_{2}^{2}],$ (17)
-where ${\rm Ent}(f):=\mathbb{E}[f(Y)\ln f(Y)]-\mathbb{E}[f(Y)]\ln\mathbb{E}[f(Y)]$.
-The log-Sobolev inequality is satisfied when the coordinates of $Y$ are independent and each satisfies the log-Sobolev inequality (tensorization property).
-The log-Sobolev inequality implies sub-Gaussian concentration.
-In fact, the following property is well-known:
-Lemma 1.
-Suppose that $X_{t}:=\left<t,Y\right$>, where $Y$ is a zero-mean random vector in $\ell_{2}$ satisfying the log-Sobolev inequality, and $T$ is totally bounded (i.e., covering numbers are all finite) in $\ell_{2}$.
-Let $\sigma^{2}:=\sup_{t\in T}\mathbb{E}[X_{t}^{2}]$, and let $\xi:=\sup_{t\in T}X_{t}$.
-Then for all $u>0$,
-$\mathbb{P}[\xi-\mathbb{E}[\xi]>u]\leq e^{-\frac{u^{2}}{2\sigma^{2}}}.$ (18)
-The proof follows by showing the Lipschitz condition $\|\nabla\max_{t\in T}\left<t,y\right>\|_{2}\leq\max_{t\in T}\|t\|_{2}$, and using the fact that the log-Sobolev inequality implies sub-Gaussian concentration; see (Boucheron et al.,, 2013, Theorem 5.8).
-In fact, from the proof we can see that the LSI assumption in Lemma 1 can be weakened to Talagrand’s optimal transport inequality, which also implies sub-Gaussian deviation of Lipschitz functions (Boucheron et al.,, 2013).
-It is possible, however, to obtain tail bounds without assuming concentration:
-Lemma 2.
-Suppose that $Y$ is a zero-mean random vector in $\ell_{2}$ satisfying $\|Y\|_{\Psi_{2}}\leq 1$, and $\sup_{t\in T}\|t\|_{2}\leq\Delta$.
-Then there exists a universal constant $C>0$ such that for any $x\geq 0$,
-$\mathbb{P}[\sup_{t\in T}\left<t,Y\right>\geq CJ(T)+x]\leq Ce^{-\frac{x^{2}}{C\Delta^{2}}},$ (19)
-where $J(T):=\int_{0}^{\infty}\sqrt{\ln N(T,\epsilon)}\epsilon$.
-A proof of Lemma 2 can be found in Van Handel, (2014) and the references therein.
-Statistical literature, such as Van de Geer, (1990), also used similar tail probability bounds for sub-Gaussian processes.
-As noted in Van Handel, (2014), although Lemma 2 requires weaker an assumption on the distribution than Lemma 1, it can give a weaker bound when $J(T)$ is larger than $\mathbb{E}[\xi]$.
 2.2 Empirical risk minimization (ERM)
-In a general mean estimation problem, $T$ is a closed subset of $\ell_{2}$, and $X_{1},\dots,X_{n}$ are i.i.d.
+Consider the mean estimation problem, where $T$ is a closed subset of $\ell_{2}$, and $X_{1},\dots,X_{n}$ are i.i.d.
 random vectors in $\ell_{2}$ whose mean is an unknown $m\in T$.
-For any $t\in\ell_{2}$, define the empirical risk
+We assume that $\|X_{1}-m\|_{G}\leq 1$, where following [43], the sub-Gaussian norm of a random vector $Y\in\ell_{2}$ is defined by taking the supremum over one-dimensional projections:
+$\|Y\|_{G}:=\sup_{v\colon\|v\|_{2}=1}\|\left<Y,v\right>\|_{G}.$ (13)
+A bound on $\|Y\|_{G}$ can be guaranteed, for example, if $Y$ has independent coordinates with uniformly bounded sub-Gaussian norms [43]. For any $t\in\ell_{2}$, define the empirical risk
 $\hat{L}(t):=\frac{1}{n}\sum_{i=1}^{n}\|t-X_{i}\|_{2}^{2}.$ (14)
@@
 $=\frac{1}{\sqrt{n}}\chi_{m-\hat{m}}$ (18)
-where $\chi_{z}:=-2\sqrt{n}\left<z,\frac{1}{n}\sum_{i=1}^{n}X_{i}-m\right$>, and $(\chi_{z})_{z\in T-T}$ is a centered sub-Gaussian process, whose natural metric is the $\ell_{2}$ metric (up to a universal constant factor).
-An example of such a mean estimation problem, which plays a fundamental role in nonparametric statistics, will be discussed in Section 3.4, where $T$ is a Sobolev ellipsoid.
-In a general regression problem, $T$ is a closed subset in $\mathbb{R}^{n}$, and $Y=\sqrt{n}m+w$, where $m\in T$ is an unknown vector to estimate, and $w$ is a zero-mean noise vector satisfying $\|w\|_{\Psi_{2}}\leq 1$.
-An example of sparse regression with fixed designs will be discussed in Section 3.5.
-The case of random designs, which is traditionally tackled by local Rademacher complexity techniques (Lugosi and Wegkamp,, 2004; Bartlett et al.,, 2005), is not considered here, although it would be interesting to consider such extensions for future work.
-Define
-$\hat{L}(t):=\frac{1}{n}\|\sqrt{n}t-Y\|_{2}^{2}$ (25)
-and define $L(t)$ and $\hat{m}$ the same way as in the mean estimation problem.
-Then (24) still holds, if we now define $\chi_{z}:=-2\left<z,w\right$>, and $(\chi_{z})_{z\in T-T}$ is a centered sub-Gaussian process, whose natural metric is the $\ell_{2}$ metric (up to a universal constant factor).
-With a slight change of notation, it is possible to unify the formulations of the two problems: both are essentially about projecting a random vector onto a subset in $\ell_{2}$.
-We present them as estimation and regression problems separately simply to make the sample size $n$ emerge naturally and consistently with the nonparametric estimation and sparse regression literature.
-More generally, the framework also encompasses many other problems including functional estimation, matrix completion, shape-restricted regression, and isotonic regression; see Chatterjee, (2014); Chatterjee et al., (2015).
-In both problems, the inequality (24) plays a key role in upper bounding the mean square error $\mathbb{E}[\|\hat{m}-m\|_{2}^{2}]$; see e.g.
-(Van de Geer,, 1990, eq (2)).
+where $\chi_{z}:=-2\sqrt{n}\left<z,\frac{1}{n}\sum_{i=1}^{n}X_{i}-m\right$>, and $(\frac{1}{2}\chi_{z})_{z\in T-T}$ is a centered sub-Gaussian process, whose natural metric is the $\ell_{2}$ metric.
+More generally, the ERM framework also applies to many other problems including regression, functional estimation, matrix completion, shape-restricted regression, and isotonic regression; see [7, 6].
+In any case, the inequality (18) plays a key role in upper bounding the mean squared error $\mathbb{E}[\|\hat{m}-m\|_{2}^{2}]$; see e.g.
+[41, eq (2)].
 Since $\mathbb{E}[\hat{L}(m)-L(m)]=0$, the expected generalization error equals $\mathbb{E}[\frac{1}{\sqrt{n}}\chi_{m-\hat{m}}$], which is also connected to (18).
-If we take the expectation of both sides of (24), and upper bound the right side by the Dudley integral using the covering number of $T$, we generally do not get sharp convergence rates of the mean square error.
+If we take the expectation of both sides of (18), and upper bound the right side by the Dudley integral using the covering number of $T$, we generally do not get sharp convergence rates of the mean squared error.
 In fact, even for the simplest case of 1-dimensional parameter estimation, such a naive approach yields a $1/\sqrt{n}$ rate rather than the optimal $1/n$ rate.
-The traditional approach for improving the bound is a “slicing” argument that yields local complexity bounds; see for example Van de Geer, (1990); Wainwright, (2019); Raskutti et al., (2011) or (Van Handel,, 2014, Theorem 5.32).
+The traditional approach for improving the bound is a “slicing” argument that yields local complexity bounds; see for example [41, 44, 35] or [42, Theorem 5.32].
 Roughly speaking, if $\|\hat{m}-m\|^{2}=r$, then $\chi_{m-\hat{m}}\leq\sup_{z\in B(r)\cap(T-T)}\chi_{z}$, which is bounded by some function $\omega(r)$ with high probability.
@@
 While it may seem that the uniform control by the union bound could lose logarithmic factors, this can actually be avoided by carefully scheduling of the error tolerance at each scale.
-In Section 8.1, we give a more detailed review of how the traditional slicing approach can yield sharp convergence for the one-dimensional toy problem.
-Local complexity bounds, which are based on Dudley’ integral for local sets, can still be loose for general $T$ where Dudley’ integral is loose.
+Local complexity bounds are usually based on Dudley’ integral, which can be loose for general $T$ when Dudley’ integral is loose.
 The majorizing measure theorem, though sharp, is hard to evaluate, and is rarely used in the analysis of ERM.
 2.3 Information-theoretic tools
-We recall standard information-theoretic measures in Thomas and Joy, (2006), using the natural base of the logarithm.
+We recall standard information-theoretic measures in [12], using the natural base of the logarithm.
 The entropy of a discrete random variable $X\sim\mu$ is $H(X):=\sum_{x}\mu(x)\ln\frac{1}{\mu(x)}$.
 The relative entropy between $\mu$ and $\nu$ on the same measurable space is $D(\mu\|\nu):=\int\ln\frac{d\mu}{d\nu}d\mu$.
-Mutual information between $(X,Y)\sim P_{XY}$ is defined as $I(X;Y):=D(P_{XY}\|P_{X}\times P_{Y})$.
+Mutual information for $(X,Y)\sim P_{XY}$ is defined as $I(X;Y):=D(P_{XY}\|P_{X}\times P_{Y})$.
 Conditional versions of these quantities are defined by averaging over the random variable which is conditioned on.
-For example, $D(P_{\hat{Z}|Z}\|\mu|P_{Z}):=\int D(P_{\hat{Z}|Z=z}\|\mu)dP_{Z}(z)$.
-The lifting argument relies on the method of types, a standard method in information theory (Csiszár and Körner,, 2011) and large deviation analysis (Dembo,, 2009).
-The type of a sequence of length $N$ is simply its empirical distribution.
-Since not all distributions can be a type, we need the following:
+For example, $D(P_{\hat{Z}|Z}\|\mu|P_{Z}):=\int D(P_{\hat{Z}|Z=z}\|\mu)dP_{Z}(z)$, and $H(Y|X):=\int H(Y|X=x)dP_{X}(x)$, where $H(Y|X=x):=\sum_{y}P_{Y|X=x}(y)\ln\frac{1}{P_{Y|X=x}(y)}$.
+The lifting argument relies on the method of types, a standard method in information theory [13] and large deviation analysis [14].
+The type of a sequence is defined as its empirical distribution.
+Since not all distributions can be a type for a given length of the sequence, we need the following:
 Definition 2.
 Suppose that $\mu$ is a distribution on a finite set $\mathcal{Z}$.
-We say $\mu$ is rational if there exists integer $N>0$ such that $N\mu$ is integer, i.e., $N\mu(z)$ is integer for any $z\in\mathcal{Z}$.
-If $N\nu$ is integer, we denote by $C_{\nu}$ the $\nu$-type set, i.e., the set of sequences with type $\nu$.
-Then the following result is standard and can be found in (Csiszár and Körner,, 2011, Chapter 1):
+We say $\mu$ is rational if there exists integer $N>0$ such that $N\mu$ is integer, i.e., $N\mu(z)\in\mathbb{Z}$ for any $z\in\mathcal{Z}$.
+If $N\nu$ is integer, the type class of $\nu$ is defined as the set of sequences with type $\nu$. The following result is standard and can be found in [13, Lemma 2.6]:
 Lemma 1.
-Let $S_{N}$ be the set of all possible types of $\mathcal{Z}^{N}$, where $\mathcal{Z}$ is a set satisfying the cardinality bound $|\mathcal{Z}|\leq n$.
-Then $|S_{N}|={n+N-1\choose n-1}\leq(N+1)^{n}$.
-Furthermore, if $\mu$ is a distribution on $\mathcal{Z}$ and $N\mu$ is integer, then
-$(N+1)^{-n}\exp(-ND(\nu\|\mu))\leq\mu^{\otimes N}(C_{\nu})\leq\exp(-ND(\nu\|\mu)).$ (26)
+Let $\mathcal{S}_{N}$ be the set of all possible types of $\mathcal{Z}^{N}$, where $\mathcal{Z}$ is a finite set with cardinality $n:=|\mathcal{Z}|$.
+Then $|\mathcal{S}_{N}|={n+N-1\choose n-1}\leq(N+1)^{n}$.
+Furthermore, if $\mu$ and $\nu$ are distributions on $\mathcal{Z}$, $N\nu$ is integer, and $\mathcal{C}$ denotes the type class of $\nu$, then
+$(N+1)^{-n}\exp(-ND(\nu\|\mu))\leq\mu^{\otimes N}(\mathcal{C})\leq\exp(-ND(\nu\|\mu)).$ (19)
+Lemma 1 implies the following result, which can be found in [13, Chapter 2, Problem 3]:
+Lemma 2.
+Suppose that $\mathcal{X}$ and $\mathcal{Y}$ are finite sets, $P_{XY}$ is a distribution on $\mathcal{X}\times\mathcal{Y}$, and $NP_{XY}$ is integer.
+Let $x^{N}$ be a sequence of type $P_{X}$, and let $\mathcal{C}\subseteq\mathcal{Y}^{N}$ be the type class of $P_{Y}$.
+Define
+$\mathcal{C}(x^{N}):=\{y^{N}\colon(x^{N},y^{N})\textrm{ has type $P_{XY}$}\}.$ (20)
+Then we have
+$(N+1)^{-|\mathcal{X}||\mathcal{Y}|}\exp(-NI(X;Y))\leq\frac{|\mathcal{C}(x^{N})|}{|\mathcal{C}|}\leq(N+1)^{|\mathcal{Y}|}\exp(-NI(X;Y)).$ (21)
+Note that sequences in $\mathcal{C}(x^{N})$ and $\mathcal{C}$ have the same probability under $P_{Y}^{\otimes N}$. Taking $\mu=\nu=P_{Y}$ in Lemma 1 yields
+$\quad(N+1)^{-2|\mathcal{X}||\mathcal{Y}|}\exp(-NI(X;Y))\leq P_{Y}^{\otimes N}(\mathcal{C}(x^{N}))\leq(N+1)^{|\mathcal{Y}|}\exp(-NI(X;Y)).$ (22)
+We recall the rate-distortion function in information theory [12], which will be a counterpart of the covering number in Dudley’s integral.
+Given a probability measure $\mu$ on $T$ and $\sigma>0$, define
+$\Pi_{\sigma}(\mu):=\{P_{Z\hat{Z}}\colon P_{Z}=P_{\hat{Z}}=\mu;\,\mathbb{E}[d^{2}(Z,\hat{Z})]\leq\sigma^{2}\}.$ (23)
+Definition 3.
+Given $\mu$ on a metric space $(T,d)$, define the rate-distortion function
+$R_{\mu}(\sigma^{2}):=\inf_{P_{UZ}}I(U;Z),$ (24)
+where the infimum is over probability measure $P_{UZ}$ on $T\times T$ satisfying $P_{Z}=\mu$ and $\mathbb{E}[d^{2}(U,Z)]\leq\sigma^{2}$. Define $r_{\mu}(\sigma):=R_{\mu}(\sigma^{2})$, and
+[…]
+Observe that
+$r_{\mu}(\sigma)\leq i_{\mu}(\sigma)\leq r_{\mu}(\sigma/2).$ (26)
+[…]
+For $Z\sim\mu$, define
+$\sigma_{\rm m}(\mu):=\inf_{z\in T}\sqrt{\mathbb{E}[d^{2}(Z,t)]},$ (27)
+abbreviated as $\sigma_{\rm m}$ if no confusion.
+Clearly $R_{\mu}(\sigma^{2})=0$ whenever $\sigma>\sigma_{\rm m}$.
+Therefore $\sigma_{\rm m}$ can be thought of as a soft version of the diameter of $T$, since the in the upper limit of the traditional Dudley integral can be set as the diameter.
 3 Main results
@@
 $i=\operatorname*{arg\,min}_{j\in\{0,1,\dots,k\}}d(Z(\omega),t_{j})$ (29)
-with ties broken arbitrarily but respecting measurability.
+with ties broken arbitrarily.
 Assuming measurability of $Z(\omega)$, we have measurability of $d(Z(\omega),t_{j})$ for each $j$, so that $Z_{k}$, and hence $X_{Z_{k}}$, is measurable.
 Density of $T_{0}$ ensures $\lim_{k\to\infty}Z_{k}=Z$ almost surely, and continuity implies that $X_{Z}=\lim_{k\to\infty}X_{Z_{k}}$ is measurable, being a sequential limit of measurable functions.
-We recall the rate-distortion function in information theory (Thomas and Joy,, 2006), which will be a counterpart of the covering number in Dudley’s integral:
 Definition 4.
-Given $\mu$ on a metric space $(T,d)$, the rate-distortion function $R_{\mu}(\sigma^{2}):=\inf_{P_{UZ}}I(U;Z)$, where the infimum is over $P_{UZ}$ satisfying $P_{Z}=\mu$ and $\mathbb{E}[d^{2}(U,Z)]\leq\sigma^{2}$.
-Given a probability measure $\mu$ on $T$ and $\sigma>0$, define
-$\Pi_{\sigma}(\mu):=\{P_{Z\hat{Z}}\colon P_{Z}=P_{\hat{Z}}=\mu;\,\mathbb{E}[d^{2}(Z,\hat{Z})]\leq\sigma^{2}\}.$ (29)
-Then we have
-$R_{\mu}(\sigma^{2})\leq\inf_{P_{Z\hat{Z}}\in\Pi_{\sigma}(P_{Z})}I(Z;\hat{Z})\leq R_{\mu}(\sigma^{2}/4).$ (30)
-[…]
-For $Z\sim\mu$, define
-$\sigma_{\rm m}(\mu):=\inf_{z\in T}\sqrt{\mathbb{E}[d^{2}(Z,t)]},$ (31)
-abbreviated as $\sigma_{\rm m}$ if no confusion.
-Clearly $R_{\mu}(\sigma^{2})=0$ whenever $\sigma>\sigma_{\rm m}$.
-Therefore $\sigma_{\rm m}$ can be thought of as a soft version of the diameter of $T$, since the in the upper limit of the traditional Dudley integral can be set as the diameter.
-Definition 3.
-Given a stochastic process $(X_{t}(\omega))_{t\in T}$, define the complexity
+Given a stochastic process $(X_{t}(\omega))_{t\in T}$ and distribution $\mu$ on $T$, define the complexity
 $\mathsf{w}(\mu):=\sup\mathbb{E}[X_{Z}]$ (30)
@@
 Theorem 1.
-If $(X_{t})_{t\in T}$ is a centered sub-Gaussian process on a countable set $T$, and $\mu$ is a distribution on $T$ for which $\sigma_{\rm m}<\infty$, then
-$\mathsf{w}(\mu)\leq 48\int_{0}^{\sigma_{\rm m}}\sqrt{R_{\mu}(\sigma^{2})}d\sigma.$ (33)
-The proof is given in Section 4.
-It will be seen in Lemma 6 that if the right side of (33) is finite, then $\mathbb{E}[|X_{Z}|]<\infty$, so no need to worry about absolute integrability on the left side of (33).
-The constant 48 in the theorem is from the commonly referenced estimate $K<24$ in the Dudley integral (7) for Gaussian processes (see comments after (7) for the extension to the sub-Gaussian case).
-Actually, according to Dudley, (2016), the best published estimate appears to be $K\leq 4\sqrt{2}$ given in Lifshits, (2012), which would imply that 48 in Theorem 1 can be improved to $8\sqrt{2}$.
+If $(X_{t})_{t\in T}$ is a centered sub-Gaussian process on a countable set $T$, and $\mu$ is a distribution on $T$ for which $\sigma_{\rm m}:=\sigma_{\rm m}(\mu)<\infty$, then
+$\mathsf{w}(\mu)\leq 2K\int_{0}^{\sigma_{\rm m}}\sqrt{i_{\mu}(\sigma)}d\sigma\leq 4K\int_{0}^{\sigma_{\rm m}}\sqrt{R_{\mu}(\sigma^{2})}d\sigma$ (31)
+where $K$ is the constant in (6).
+Theorem 1 is proved in Section 4 for finite $T$ and extended to countable $T$ in the appendix.
+It will be shown the appendix that if the right side of (31) is finite, then $\mathbb{E}[|X_{Z}|]<\infty$, so no need to worry about absolute integrability on the left side of (31).
 Another type of information-theoretic improvements of traditional generalization bounds, discussed around (2), have recently been studied in the literature.
@@
 Let $Z$ be a random variable on $T$ arbitrarily correlated with $Y$, and set $\mu:=P_{Z}$.
+Let $K$ be the constant in (6).
 Then
-$\mathbb{E}[X_{Z}]\leq 48\int_{0}^{\sigma_{\rm m}}\sqrt{R_{\mu}(\sigma^{2})\wedge I(Y;Z)}d\sigma.$ (34)
-The proof of Theorem 2 is given in Section 4.1.
-Since $\sigma_{\rm m}\lesssim\sup_{t\in T}\|X_{t}\|_{\Psi_{2}}$, it is easy to see that Theorem 2 implies (2).
-Since the rest of the paper does not concern Theorem 2, we will not discuss its extension beyond finite $T$.
+$\mathbb{E}[X_{Z}]\leq 2K\int_{0}^{\sigma_{\rm m}}\sqrt{i_{\mu}(\sigma)\wedge I(Y;Z)}d\sigma.$ (32)
+The proof of Theorem 2 is given in Section 4. Since $\sigma_{\rm m}\lesssim\sup_{t\in T}\|X_{t}\|_{G}$, it is easy to see that Theorem 2 implies (2).
 While Theorem 1 does not cover the case of uncountable $T$, it is sufficient for our statistical applications due to the following observation:
 Proposition 1.
-If $(X_{t})$ is a separable, continuous process, then for $Z\sim\mu$,
-$\mathbb{E}[X_{Z}]\leq 48\lim_{\epsilon\downarrow 0}\sup_{\mu^{\prime}\colon W_{\infty}(\mu,\mu^{\prime})\leq\epsilon}\int_{0}^{\sigma_{\rm m}}\sqrt{R_{\mu^{\prime}}(\sigma^{2})}d\sigma.$ (35)
-The proof of Proposition 4.2 is in Section 4.2.
+If $(X_{t})_{t\in T}$ is a separable, continuous process, then for $Z\sim\mu$,
+$\mathbb{E}[X_{Z}]\leq 2K\lim_{\epsilon\downarrow 0}\sup_{\mu^{\prime}\colon W_{\infty}(\mu,\mu^{\prime})\leq\epsilon}\int_{0}^{\sigma_{\rm m}}\sqrt{R_{\mu^{\prime}}(\sigma^{2})}d\sigma.$ (33)
+The proof of Proposition 1 is in the appendix.
 Naturally, we would like to show that the right side of (33) reduces to the right side of (31) by dominated convergence, but this requires showing that $\sqrt{R_{\mu^{\prime}}(\sigma^{2})}$ can be dominated by an absolutely integrable function, which is challenging in the regime of vanishing $\sigma$ where the rate-distortion function may grow fast.
 While it is possible to impose on certain growth condition for dominated convergence to apply, we will not dive into those complications.
-Instead, in many statistical applications, we will see that $R_{\mu^{\prime}}(\sigma^{2})$ can controlled using the second moments of $\mu$, so the right side of (35) can be directly evaluated using the continuity of the bound in the second moment, without the need to establish the equivalence of the right sides of (35) and (33).
+Instead, in many statistical applications, we will see that $R_{\mu^{\prime}}(\sigma^{2})$ can be controlled using the second moments of $\mu$, so the right side of (33) can be directly evaluated using the continuity of the bound in the second moment, without the need to establish the equivalence of the right sides of (33) and (31).
 3.2 Sharpness for Gaussian processes
@@
 Theorem 3.
-In the setting of Theorem 1, suppose that $(X_{t})_{t\in T}$ is a Gaussian process (with metric defined in (8)). For any $\mu$, we have
-$\mathbb{E}[X_{Z}]\geq c\int_{0}^{\sigma_{\rm m}}\sqrt{R_{\mu}(\sigma^{2})}d\sigma$ (36)
-where the supremum is over all $Z(\omega)$ satisfying $P_{Z}=\mu$, and
-$c:=\max_{\tau\in(0,4/5)}\left\{\frac{4-5\tau}{2\sqrt{2}(4-\tau)\ln\frac{4}{\tau}}\right\}\approx 0.0935918,$ (37)
-numerically evaluated at $\tau=0.23457905$.
+Suppose that $(X_{t})_{t\in T}$ is a centered Gaussian process, $T$ is a finite set, and $\mu$ is an arbitrary distribution on $T$. Then have
+$K^{\prime}\mathsf{w}(\mu)\geq\int_{0}^{\infty}\sqrt{i_{\mu}(\sigma)}d\sigma\geq\int_{0}^{\sigma_{\rm m}(\mu)}\sqrt{R_{\mu}(\sigma^{2})}d\sigma,$ (34)
+where $K^{\prime}$ is any universal constant such that (7) holds.
+We present a proof of Theorem 3 in Section 5, by using Fernique’s theorem that Dudley’s integral is sharp for stationary Gaussian processes (even though Theorem 3 does not assume stationary process).
+In the appendix we also extend the result to the case of countable $T$.
 Theorem 3 suggests that Theorem 1 may yield sharper results in certain statistical applications where the traditional method based on Dudley’s integral is suboptimal.
-Three proofs of Theorem 3 are given in Section 5: First, the result can be proved using the lifting argument, by applying the majorizing measure theorem to the $\mu$-type set.
-Secondly, we can apply the lifting argument and use Fernique’s theorem that Dudley’s integral is sharp for stationary Gaussian processes, i.e., the case where the metric on $T$ is invariant under the action of a transitive group (see (Dudley,, 1999, Cor 2.7.5) or (Fernique,, 1975)), which preceded Talagrand’s striking proof of the majorizing measure theorem (Talagrand,, 1987).
-Even though $T$ in Theorem 3 may not be permutation-invariant, the lifting argument in our proof lower bounds the expected supremum over the typical set, which satisfies permutation invariance.
-The third proof derives an information-theoretic counterpart of the super-Sudakov inequality (Talagrand,, 1992; Van Handel,, 2014), and finishes off by purely analytic arguments (without chaining or packing constructions), which gives the constant stated in Theorem 3.
-Classically, equivalent formulations of the majorizing measure theorem, (14)-(15), were established by Fernique convexity argument; see (Talagrand,, 2014, Chapter 3) for a proof using combinatorial arguments such as tree packing.
-Similarly, the rate-distortion integral can also be expressed in alternative forms by applying the minimax theorem to a certain convex-concave functional; see Remark 1 for discussions.
-We can derive from Theorem 3 the classical the majorizing measure theorem, with an explicit estimate of constant. Recall $I_{\mu}$ defined in (11).
+From Theorem 3 we can derive the classical majorizing measure theorem.
+Recall $I_{\mu}$ defined in (8).
 Corollary 1.
-For any Gaussian process (not necessarily stationary),
-$\mathbb{E}[\sup_{t\in T}X_{t}]\geq\bar{c}\inf_{\mu}\sup_{t\in T}I_{\mu}(t).$ (38)
-Here $\bar{c}:=\sup_{a>1}\frac{\sqrt{1-a^{-2}}}{\frac{2a}{c}+\sqrt{2\pi h(a^{-2})}}\approx 0.0218988$ (evaluated at $a=1.4392$), where $h$ is the binary entropy function in nats and $c$ was defined in Theorem 3.
+For any Gaussian process (not necessarily stationary), we have
+$\mathbb{E}[\sup_{t\in T}X_{t}]\geq\bar{c}\inf_{\mu}\sup_{t\in T}I_{\mu}(t),$ (35)
+where $\bar{c}:=\sup_{a>1}\frac{\sqrt{1-a^{-2}}}{2aK^{\prime}+\sqrt{2\pi h(a^{-2})}}$, $h(\cdot)$ is the binary entropy function, and $K^{\prime}$ is any universal constant such that (34) holds.
 The proof of Corollary 1 is given in Section 6, which is essentially based on lower bounding the rate-distortion function by the probability of a set using the data processing inequality.
-We have not found a reference on the majoring measure theorem that estimated the constant to compare with Corollary 1.
 Another application of Theorem 3 is sub-Gaussian comparison theorems, which, in statistical applications, sometimes allow us to reduce the problem to simpler “linear” cases where straightforward Cauchy-Schwarz or Hölder inequality arguments apply.
@@
 Corollary 2 can be viewed as a “soft” version of (10) (in fact, it is easy to see that it implies (10)).
-The Euclidean special case of Corollary 2 is of particular interest in practice:
+The Euclidean special case of Corollary 2 is of particular interest and will be used later in our statistical applications:
 Corollary 3.
@@
 However, for fixed $N=3$ and vanishing $\epsilon$, it is possible that $\mathbb{E}[X_{Z}]=\mathbb{E}[\max_{t\in T}X_{t}]\asymp\epsilon\sqrt{|T|}\asymp\epsilon^{1-N/2}$ diverges, for certain heavy-tailed $X_{t}$.
-On the other hand, since $Z$ is in the unit ball, $\operatorname*{tr}(\sqrt{\Sigma})=O(1)$ as $\epsilon$ vanishes, confirming that this is a counterexample.
-3.3 Tail Bounds for ERM
-In this section we consider processes in the form of an inner product in $\ell_{2}$. In this case, the complexity of an index distribution $\mu$ (Definition 3) is written as
-$\mathsf{w}(\mu):=\sup_{P_{YZ}\in\Pi(\mu,\gamma)}\mathbb{E}[\left<Y,Z\right$>], (39)
-where $\gamma$ is a given distribution on $\ell_{2}$, which is assumed to have zero mean in this section.
-We may denote $\mathsf{w}(\mu)$ by $\mathsf{w}(\mu;\gamma)$ to make the dependence on $\gamma$ explicit.
-The following regularity condition will be used in this section:
-Definition 4.
-$T\subseteq\ell_{2}$ is totally bounded if the covering number $N(T,\epsilon)$ is finite for every $\epsilon>0$.
-In empirical process theory, total boundedness is a rather mild condition which is almost always assumed; see for example Ledoux and Talagrand, (2013).
-It is sometimes also referred to as precompactness in topology.
-It is also equivalent to
-$\lim_{n\to\infty}\sup_{t\in T}\sum_{i>n}t(i)^{2}=0.$ (40)
-Indeed, if (40) is true, then for every $\epsilon>0$ there exists $n$ such that $\sup_{t\in T}\sum_{i>n}t(i)^{2}<(\epsilon/4)^{2}$.
-We can cover the projection of $T$ onto the first $n$ coordinates with finitely many balls of radius $\epsilon/2$, whose centers are the projections of $t_{1},\dots,t_{M}\in T$.
-Then $B(t_{1},\epsilon),\dots,B(t_{M},\epsilon)$ also cover $T$.
-Conversely, total boundedness implies that every sequence in $T$ has a convergent subsequence, which implies (40).
-Definition 5.
-Suppose that $T$ is a subset of $\ell_{2}$.
-We call $\mathsf{G}\colon[0,\infty)\to[0,\infty)$ a width-moment function of $T$ if for any random vector $Z\in T$ satisfying $\mathbb{E}[\|Z\|_{2}^{2}]\leq E$, we have
-$\mathsf{w}(P_{Z})\leq\mathsf{G}(E).$ (41)
-Consider the setting of the empirical risk minimization problem in Section 2.2 (either the mean estimation or the regression problem).
-Our main result, for $\gamma$ satisfying the log-Sobolev inequality, is the following:
-Theorem 4.
-Suppose that $m\in T$, where $T$ is a closed, totally bounded subset of $\ell_{2}$, and TTT-T has a width-moment function $\mathsf{G}$.
-Assume that $\gamma$ has zero mean and satisfies the log-Sobolev inequality (17).
-Let $\psi(\lambda):=\ln\mathbb{E}[\exp(\lambda\|m-\hat{m}\|_{2}^{2})]$.
-Then for any $\lambda\geq 0$,
-$\psi(\lambda)\leq\sup_{\epsilon>0}\left\{-\epsilon+\frac{2\lambda}{\sqrt{n}}\mathsf{G}\left(\frac{\epsilon+\psi(\lambda)}{\lambda}\right)+\sqrt{\frac{8\epsilon\lambda(\epsilon+\psi(\lambda))}{n}}\right\}.$ (42)
-The proof of Theorem 4 is given in Section 7. The log-Sobolev inequality assumption in Theorem 4 can in fact be weakened to sub-Gaussianity. In that case, define
-$\mathsf{w}_{\rm g}(\mu):=\sup_{P_{YZ}\in\Pi(\gamma_{\rm g},\mu)}\mathbb{E}\left[\sum_{i=1}^{\infty}Y(i)Z(i)\right],$ (43)
-where $\gamma_{\rm g}$ denotes the distribution of an infinite sequence of i.i.d.
-standard normal random variables, and the infinite sum is defined in terms of convergence in second moment.
-Similarly, define $\mathsf{G}_{\rm g}(E)$ by replacing the left side of (41) with $\mathsf{w}_{\rm g}(P_{Z})$.
-Note that if $\gamma$ satisfies the log-Sobolev inequality (hence also sub-Gaussian), then Corollary 2 implies that $\mathsf{w}_{\rm g}$ and $\mathsf{G}_{\rm g}$ dominate $\mathsf{w}$ and $\mathsf{G}$ respectively.
-We have the following result:
-Theorem 5.
-In the setting of Theorem 4, suppose that $\gamma$ is only assumed to have zero-mean and satisfy $\|Y\|_{\Psi_{2}}\leq 1$, $Y\sim\gamma$.
-Assume that $\mathsf{G}_{\rm g}(\cdot)$ is continuous.
-Then for any $\lambda\geq 0$,
-$\psi(\lambda)\leq\sup_{\epsilon>0}\left\{-\frac{\epsilon}{C}+\frac{2K^{\prime}C\lambda}{\sqrt{n}}\mathsf{G}_{\rm g}\left(\frac{\epsilon+\psi(\lambda)}{\lambda}\right)+K^{\prime}C\sqrt{\frac{8\epsilon\lambda(\epsilon+\psi(\lambda))}{n}}\right\},$ (44)
-where $C,K^{\prime}>0$ are universal constants in Lemma 2 and (10), respectively.
-The proof of Theorem 5 is given in Section 7.2.
-Similar to the remark following Lemma 2, Theorem 5 assumes a weaker condition on $\gamma$ than (4), but may yield a weaker bound if $\mathsf{G}_{\rm g}$ is larger than $\mathsf{G}$ (for example, in the case where $Y$ has i.i.d.
-coordinates following the uniform distribution on [1,1][-1,1]).
-3.4 ERM for the Sobolev Ellipsoid
-We apply the general technique in Section 3.3 to Sobolev ellipsoids, which plays a prominent role in nonparametric statistics (Van de Geer,, 1990; Wainwright,, 2019; Tsybakov,, 2009).
-The precise problem of estimation over a Sobolev ellipsoid was previously studied in Wei et al., (2020).
-Generally, an ellipsoid in $\ell_{2}$ is defined as
+On the other hand, since $Z$ is in the unit ball, $\operatorname*{tr}(\sqrt{\Sigma})$ remains bounded from above as $\epsilon$ vanishes, confirming that this is a counterexample.
+3.3 Application to ERM over the Sobolev ellipsoid
+We use the rate-distortion integral to upper bound the error in ERM over a Sobolev ellipsoid under a Lipschitz transformation, A general ellipsoid in $\ell_{2}$ is defined as
 $\mathcal{E}:=\left\{t\colon\sum_{i=1}^{\infty}\frac{t_{i}^{2}}{a_{i}^{2}}\leq 1\right\},$ (36)
 where $a_{1},a_{2},\dots$ is a decreasing sequence of nonnegative numbers.
-This is recurring example in Talagrand, (2014), in particular an example showing that Dudley’s integral is loose.
 A Sobolev ellipsoid is an ellipsoid with $a_{i}=i^{-\beta}$ for some $\beta>0$.
-By considering $t(1),t(2),\dots$ as the Fourier coefficients of a regression function from the $\beta$-Sobolev space, we see that the Sobolev ellipsoid can be identified with a nonparametric function class (Tsybakov,, 2009).
-In Section 9.1, we will review the supremum of a Gaussian process indexed by $\mathcal{E}$, and show that the rate-distortion integral yields a sharp bound.
+By considering $t(1),t(2),\dots$ as the Fourier coefficients of a regression function from the $\beta$-Sobolev space, we see that the Sobolev ellipsoid can be identified with a nonparametric function class [40].
+Sobolev ellipsoids therefore play a prominent role in nonparametric statistics [41, 44, 40].
+The precise problem of ERM over a Sobolev ellipsoid was previously studied in [45].
+The ellipsoid is recurring example in [39], in particular an example showing that Dudley’s integral is loose.
+In Section 7, we review the supremum of a Gaussian process indexed by $\mathcal{E}$, and show that the rate-distortion integral yields a sharp bound.
 Consider ERM in the setting of Section 2.2, where $T=\phi(\mathcal{E})$ is the image of the Sobolev ellipsoid (for $\beta>1/2$, where $1/2$ is the critical exponent for ERM) under a 1-Lipschitz map $\phi\colon\ell_{2}\to\ell_{2}$.
-The purpose of introducing $\phi$ is that if $\phi$ is linear, an elementary proof exists, while the rate-distortion integral approach can handle more general settings.
-Furthermore, we assume that $Y$ is a random vector in $\ell_{2}$ satisfying $\|Y\|_{\Psi_{2}}\leq 1$.
+In practice, $\phi$ may represent a nonlinear neural network or feature map.
+If $\phi$ is the identity map, the Cauchy-Schwarz inequality suffices for the error upper bound (Remark 2), while the rate-distortion integral machinery is essential for general $\phi$, analogous to the role of the majorizing measure theorem for sub-Gaussian comparison (10).
+Furthermore, we assume that $Y$ is a random vector in $\ell_{2}$ satisfying $\|Y\|_{G}\leq 1$.
 Under these assumptions, we have the following main result:
 Theorem 4.
-There exists a function $c(\beta)>0$ satisfying $c(\beta)=O(\frac{1}{\sqrt{2\beta-1}})$ as $\beta\downarrow\frac{1}{2}$, and a universal constant $b>0$ such that
-$\mathbb{P}[\|m-\hat{m}\|_{2}>\sqrt{c(\beta)}n^{-\frac{\beta}{2\beta+1}}]\leq\exp(-bc(\beta)n^{\frac{1}{2\beta+1}}).$ (46)
-Theorem 6 shows that ERM achieves square error $O(\frac{1}{\sqrt{2\beta-1}})n^{-\frac{2\beta}{2\beta+1}}$ with high probability, which is optimal not only in terms of the dependence on $n$, but also in terms of the behavior of the preconstant $O(\frac{1}{\sqrt{2\beta-1}})$ as $\beta\downarrow 1/2$.
-Furthermore, from the proof of Theorem 6 we can also see that the expected generalization error, $\mathbb{E}[\frac{1}{\sqrt{n}}\chi_{m-\hat{m}}$], also converges at the rate $O(\frac{1}{\sqrt{2\beta-1}})n^{-\frac{2\beta}{2\beta+1}}$.
-We present two approaches towards Theorem 6 in Section 9.2.
-The first one does not use any new tools introduced in this paper, but only works for a special case of Theorem 6 where $\phi$ is the identity map and $Y$ satisfies the stronger log-Sobolev inequality: We use the Cauchy-Schwarz argument to bound the mean of the square error, and then use a concentration inequality to control the deviation from the mean.
-This approach is a slightly streamlined version of the method in Chatterjee, (2014); Wei et al., (2020).
-The second proof follows the general recipe in Section 3.3, and works for general sub-Gaussian $Y$ and nonlinear Lipschitz $\phi$, where the rate-distortion integral is used to bound $\mathsf{G}_{\rm g}$.
-For comparison, the traditional techniques will only prove an upper bound at the rate $O(\frac{1}{2\beta-1})n^{-\frac{2\beta}{2\beta+1}}$, which is optimal in terms of the dependence on $n$, but not in terms of the dependence on $\beta$ in the preconstant as $\beta\downarrow 1/2$ (we analyze the preconstant using the traditional method in Remark 3).
-Our result shows that the optimal $O(\frac{1}{\sqrt{2\beta-1}})$ preconstant can be regained by using the rate-distortion integral instead.
-3.5 ERM for Sparse Regression
-This section provides new upper bounds for empirical risk minimization for the weak $\ell_{q}$ ball in $\mathbb{R}^{d}$ defined in (4), extending prior results for the $\ell_{q}$ ball (3).
-Consider a special case of the general regression model in Section 2.2, where $T$ is chosen as a linear transform of $wB_{q}(r)$.
-More precisely, suppose that the columns of $\frac{1}{\sqrt{n}}\mathbb{X}$ have unit $\ell_{2}$ norm, $\beta^{*}\in wB_{q}(r)$, $q\in(0,1)$, and $Y=\mathbb{X}\beta^{*}+w$, where $w$ is a zero mean random vector satisfying $\|w\|_{\Psi_{2}}\leq 1$.
-Let $\hat{\beta}:=\operatorname*{arg\,min}_{\beta\in wB_{q}(r)}\|Y-\mathbb{X}\beta\|_{2}$.
-Throughout, we set
-$R:=r^{q},$ (47)
-and we caution the reader that some prior work, including Raskutti et al., (2011), uses the notation $B_{q}(R)=B_{q}(r^{q})$ to refer to what we denote by $B_{q}(r)$.
-Applying the technique in Section 3.3, we obtain the following result:
-Theorem 7.
-There exists $C_{q}>0$ (depending only on $q\in(0,1)$) such that
-$\mathbb{P}\left[\|m-\hat{m}\|_{2}^{2}>C_{q}(\frac{\ln d}{n})^{\frac{2-q}{2}}R\right]\leq\exp(-(\ln d)^{\frac{2-q}{2}}Rn^{\frac{q}{2}}).$ (48)
-The proof of Section 10 is given in Theorem 7, which is based on the rate-distortion integral and the general technique for tail bounds in Section 3.3.
-Previously, Raskutti et al., (2011) established the sharp convergence of ERM when $\beta\in B_{q}(r)$ where $q\in(0,1]$.
-The analysis for the case of $\beta\in B_{1}(r)$ follows by a simple Hölder inequality argument (see Raskutti et al., (2011)), whereas the $q<1$ case was proved by the slicing technique, based on the Dudley integral where the covering numbers were bounded using deep results from the convex geometry (Carl and Pajor,, 1988; Guedon and Litvak,, 2000).
-In contrast, we prove Theorem 7 using the rate-distortion integral, where the rate-distortion function (a priori smaller than the log covering number, hence easier to upper bound) is upper bounded using purely information-theoretic tools, without invoking the convex geometry literature.
-Furthermore, Theorem 7 shows that the minimax rate for $wB_{q}(r)$ is the same as $B_{q}(r)$ when $q\in(0,1)$, a nontrivial extension of the result of Raskutti et al., (2011) which appears to be beyond the previous techniques.
-The $q=1$ case is somewhat special: by considering $\mathbb{X}$ with identical rows, we can see that the rate for $wB_{1}(r)$ can be worse than that of $B_{1}(r)$ by a logarithmic factor.
-On the other hand, in the special case where $\frac{1}{\sqrt{n}}\mathbb{X}$ is the identity, the rate $(\frac{\ln d}{n})^{\frac{2-q}{2}}R$ is achievable for $wB_{q}(r)$ for all $q\in(0,2)$; see Rigollet and Hütter, (2023); Donoho and Johnstone, (1994); Zhang, (2010).
-4 Proof of the Rate-Distortion Integral
-In this section we prove Theorem 1, Theorem 2, and Proposition 1.
-4.1 Proof in the Finite Case
-Recall the following basic property of the sub-Gaussian distribution:
+There exists a function $c(\beta)>0$ satisfying $\lim_{\beta\downarrow\frac{1}{2}}\sqrt{2\beta-1}c(\beta)<\infty$, such that $\mathbb{E}[\|m-\hat{m}\|_{2}^{2}]\leq c(\beta)n^{-\frac{2\beta}{2\beta+1}}$.
+We present a proof of Theorem 4 in Section 7.3.
+Remark 1.
+Theorem 4 shows that ERM achieves mean squared error of order $\frac{1}{\sqrt{2\beta-1}}n^{-\frac{2\beta}{2\beta+1}}$ with high probability.
+Furthermore, from the proof of Theorem 4 we can also see that the expected generalization error, $\mathbb{E}[\frac{1}{\sqrt{n}}\chi_{m-\hat{m}}$], also converges at the rate $\frac{1}{\sqrt{2\beta-1}}n^{-\frac{2\beta}{2\beta+1}}$.
+For comparison, the traditional technique of [41] based on Dudley’s integral will only prove an upper bound at the rate $\frac{1}{2\beta-1}n^{-\frac{2\beta}{2\beta+1}}$ (see Remark 4), which has the optimal exponent of $n$, but a worse preconstant.
+Remark 2.
+For convex regression classes with Gaussian noise, [7] showed that the mean-squared error (MSE) of ERM can be characterized by a fixed-point equation involving a localized Gaussian width.
+For ellipsoids, this localized Gaussian width can then be controlled directly via the Cauchy–Schwarz inequality [45], thereby avoiding the slicing and Dudley-integral machinery and mitigating the looseness of Dudley’s bound.
+However, the Cauchy-Schwarz step of [45] exploits linear structure, and cannot be applied to the setting of Theorem 4 when the map $\phi$ is nonlinear Lipschitz.
+This limitation is expected to be artifacts of the proof technique of [7, 45]: Lipschitz maps do not increase the supremum of the relevant stochastic process (a consequence of the majorizing measure theorem; see (10)).
+Our work thus introduces a new proof technique that overcomes this limitation.
+4 Proof of the rate-distortion integral: finite case
+In this section we prove Theorem 1 and Theorem 2 in the case of finite index set; extension beyond the finite case follows by a limiting argument, and is deferred to the appendix.
+Recall the following basic property of the sub-Gaussian distribution, whose proof can be found in [43, Exercise 2.40]:
 Proposition 2.
-Suppose that $X_{1},X_{2}$ are independent random variables, and $\|X_{i}\|_{\Psi_{2}}\leq K_{i}$, $i=1,2$. Then $\|X_{1}+X_{2}\|_{\Psi_{2}}\leq\sqrt{K_{1}^{2}+K_{2}^{2}}$.
-Proof.
-By Jensen’s inequality,
-$\mathbb{E}\left[\exp(\frac{X_{1}^{2}+X_{2}^{2}}{K_{1}^{2}+K_{2}^{2}})\right]$ $=\mathbb{E}\left[\exp(\frac{X_{1}^{2}}{K_{1}^{2}+K_{2}^{2}})\right]\mathbb{E}\left[\exp(\frac{X_{2}^{2}}{K_{1}^{2}+K_{2}^{2}})\right]$ (49)
-$\leq\mathbb{E}^{\frac{K_{1}^{2}}{K_{1}^{2}+K_{2}^{2}}}\left[\exp(\frac{X_{1}^{2}}{K_{1}^{2}})\right]\mathbb{E}^{\frac{K_{2}^{2}}{K_{1}^{2}+K_{2}^{2}}}\left[\exp(\frac{X_{2}^{2}}{K_{2}^{2}})\right]$ (50)
-$\leq 2.$ (51)
-We also need the following estimate of the covering number:
-Lemma 4.
-For given $\sigma>0$ independent of $N$, we have
-$NR_{\mu}(4\sigma^{2})+o(N)\leq\ln N(\mathcal{C},2\sqrt{N}\sigma)\leq NR_{\mu}(\sigma^{2})+o(N)$ (52)
-where $o(N)$ hides constants depending on $P_{Y}$ and $\sigma$.
-Proof.
-For any $t^{N}\in\mathcal{C}$, a standard application of Sanov’s theorem (Csiszár and Körner,, 2011, Chapter 1) shows
-$\ln\frac{1}{\mu^{\otimes N}(B(t^{N},\sqrt{N}\sigma))}=NR_{\mu}(\sigma^{2})+o(N)$ (53)
-where the implicit constant in $o(N)$ may depend on $\mu$.
-Let $P(\mathcal{C},2\sqrt{N}\sigma)$ denote the packing number, i.e.
-the maximal number of points $\mathcal{C}$ separated by distance at least $2\sqrt{N}\sigma$.
-Then $N(\mathcal{C},2\sqrt{N}\sigma)\leq P(\mathcal{C},2\sqrt{N}\sigma)$ (see e.g.
-Van Handel, (2014)), and standard volume estimates yields $P(\mathcal{C},2\sqrt{N}\sigma)\leq\frac{1}{\min_{t^{N}\in\mathcal{C}}\mu^{\otimes N}(B(t^{N},\sqrt{N}\sigma))}$, proving the second inequality in (52).
-Moreover, by a volume argument we also have $N(\mathcal{C},2\sqrt{N}\sigma)\max_{t^{N}\in\mathcal{C}}\mu^{\otimes N}(B(t^{N},2\sqrt{N}\sigma))\geq\mu^{\otimes N}(\mathcal{C})=\exp(o(N))$.
-Then using (53) with $\sigma$ replaced by $2\sigma$, we obtain the first inequality in (52).
-∎
-Proof of Theorem 1 (for finite $T$).
-For any integer $N>0$, let $\mathcal{C}$ be the $P_{Z}$-typical set (Thomas and Joy,, 2006):
-$\mathcal{C}:=\mathcal{C}_{N}:=\{z^{N}\in T^{N}\colon\|\widehat{P}_{z^{N}}-\mu\|_{1}\leq N^{-1/3}\}$ (54)
-where $\widehat{P}_{z^{N}}$ denotes the empirical distribution of $z^{N}$, and $\|\widehat{P}_{z^{N}}-\mu\|_{1}:=\sum_{z\in T}|\widehat{P}_{z^{N}}(z)-\mu(z)|$ can in fact be replaced by any norms, which are equivalent up to constants due to finiteness of $T$.
-Moreover, let $n:=|T|$.
+Suppose that $X_{1},X_{2},\dots,X_{n}$ are zero-mean, independent random variables, and $\|X_{i}\|_{G}\leq K_{i}$, $i=1,2,\dots,n$.
+Then $\|X_{1}+X_{2}+\dots+X_{n}\|_{G}\leq\sqrt{K_{1}^{2}+K_{2}^{2}+\dots+K_{n}^{2}}$.
+Let $n:=|T|$.
 Without loss of generality, we can identify $T$ with the standard basis vectors $e_{1},e_{2},\dots,e_{n}$ of $\mathbb{R}^{n}$, and define $Y_{i}:=X_{e_{i}}$ for each $i\in\{1,\dots,n\}$.
@@
 $X_{z}=\left<Y,z\right>,\quad\forall z\in T.$ (37)
-It is not necessary to adopt the inner product representation for this proof, but it helps making the notations clear.
-Suppose that $P_{YZ}\in\Pi(P_{Y},\mu)$ is arbitrary, and let $(Y^{N},Z^{N})\sim P_{YZ}^{\otimes N}$.
+It is not necessary to adopt the inner product representation for the proof of Theorem 1, but it helps making the notations clear.
+Now let $Y^{N}\sim P_{Y}^{\otimes N}$.
+Recall the definitions related to types in Section 2.3.
+For any $\mu$ on $T$, define
+$\mathcal{I}:=\{N\in\mathbb{Z}\colon\textrm{$N>0$, $N\mu$ is integer}\}$ (38)
+which is nonempty if $\mu$ is rational. For $N\in\mathcal{I}$, define the type class
+$\mathcal{C}_{\mu}$ $:=\{z^{N}\in T^{N}\colon\widehat{P}_{z^{N}}=\mu\}.$ (39)
+The following two lemmas show that a coupling can be approximated by a matching:
+Lemma 3.
+Let $\mu$ be a rational distribution on $\mathbb{R}^{n}$, and let $N\in\mathcal{I}$. For any $y^{N}\in\mathbb{R}^{nN}$, we have
+$\frac{1}{\sqrt{N}}\min_{s^{N}\in\mathcal{C}_{\mu}}\|y^{N}-s^{N}\|_{2}=W_{2}(\widehat{P}_{y^{N}},\mu).$ (40)
+Lemma 4.
+Suppose that $\mathcal{X}$ and $\mathcal{Y}$ are finite sets, $N>0$ is an integer, and $P_{XY}$ is a distribution with the property that both $NP_{X}$ and $NP_{Y}$ are integer.
+Then there exists $Q_{XY}$ satisfying $Q_{X}=P_{X}$, $Q_{Y}=P_{Y}$, $\max_{x,y}|Q_{XY}(x,y)-P_{XY}(x,y)|\leq\frac{1}{N}$, and that $NQ_{XY}$ is integer.
+Lemma 3 has no approximation error, whereas Lemma 4 shows approximation on the level of the distribution (not just the min value) but has a $1/N$ approximation error.
+For our purpose, it is in fact possible to use only Lemma 4, since the $1/N$ gap vanishes in the limit.
+The proofs of both lemmas can be found in the appendix.
+By Proposition 2, $(\left<Y^{N},z^{N}\right>)_{z^{N}\in\mathcal{C}}$ is a sub-Gaussian process with respect to the following metric on $T^{N}$:
+$d_{N}(z^{N},\hat{z}^{N}):=\sqrt{\sum_{i=1}^{N}d^{2}(z_{i},\hat{z}_{i})}.$ (41)
+The following observation is crucial:
+Lemma 5.
+For any rational $\mu$, we have
+$\lim_{\mathcal{I}\ni N\to\infty}\frac{1}{N}\mathbb{E}[\max_{z^{N}\in\mathcal{C}_{\mu}}\left<Y^{N},z^{N}\right>]=\mathsf{w}(\mu),$ (42)
+where $\lim_{\mathcal{I}\ni N\to\infty}$ denotes the limit along the subsequence in $\mathcal{I}$.
+Proof.
+By Lemma 3, we see that for any $N\in\mathcal{I}$,
+$\frac{1}{N}\max_{z^{N}\in\mathcal{C}_{\mu}}\left<y^{N},z^{N}\right$> $=\frac{\mathbb{E}_{\widehat{P}_{y^{N}}}[\|Y\|_{2}^{2}]+\mathbb{E}_{\mu}[\|Z\|_{2}^{2}]-W_{2}^{2}(\widehat{P}_{y^{N}},\mu)}{2}.$ (43)
+Define $M_{N}:=\sup_{P_{Y^{\prime}}\colon W_{2}(P_{Y},P_{Y^{\prime}})\leq N^{-\frac{1}{2n}}}\frac{\mathbb{E}_{P_{Y^{\prime}}}[\|Y\|_{2}^{2}]+\mathbb{E}_{\mu}[\|Z\|_{2}^{2}]-W_{2}^{2}(P_{Y^{\prime}},\mu)}{2}$.
+Since $W_{2}$ is a metric (satisfying the triangle inequality), we see that $\mathbb{E}_{P_{Y^{\prime}}}[\|Y\|_{2}^{2}]-W_{2}^{2}(P_{Y^{\prime}},\mu)$ is continuous in $P_{Y^{\prime}}$ with respect to $W_{2}$, and hence
+$\lim_{\mathcal{I}\ni N\to\infty}M_{N}$ $=\sup_{P_{YZ}\in\Pi(P_{Y},\mu)}\mathbb{E}[\left<Y,Z\right$>] (44)
+$=\mathsf{w}(\mu).$ (45)
+Define
+$\mathcal{D}$ $:=\{y^{N}\in\mathbb{R}^{nN}\colon W_{2}(P_{Y},\widehat{P}_{y^{N}})\leq N^{-\frac{1}{2n}}\}.$ (46)
+By the Wasserstein law of large numbers (cf. [8]), we have $\lim_{\mathcal{I}\ni N\to\infty}\mathbb{P}[Y^{N}\in\mathcal{D}]=1$. Therefore,
+$\limsup_{\mathcal{I}\ni N\to\infty}\frac{1}{N}\mathbb{E}[\max_{z^{N}\in\mathcal{C}}\left<Y^{N},z^{N}\right>1_{Y^{N}\in\mathcal{D}}]$ $=\limsup_{\mathcal{I}\ni N\to\infty}\mathbb{E}[\sup_{P_{YZ}\in\Pi(\widehat{P}_{Y^{N}},\mu)}\mathbb{E}[\left<Y,Z\right>]1_{Y^{N}\in\mathcal{D}}]$ (47)
+$\leq\lim_{\mathcal{I}\ni N\to\infty}M_{N}\mathbb{E}[1_{Y^{N}\in\mathcal{D}}]=\mathsf{w}(\mu),$ (48)
+where the first equality follows from Lemma 3.
+Similarly, if the $\sup$ in the definition of $M_{N}$ is replaced by $\inf$, then (45) still holds, but the inequality sign in (48) will be reversed, and the $\limsup$ in (47) will be replaced by $\liminf$.
+Hence we have
+$\lim_{\mathcal{I}\ni N\to\infty}\frac{1}{N}\mathbb{E}[\max_{z^{N}\in\mathcal{C}}\left<Y^{N},z^{N}\right>1_{Y^{N}\in\mathcal{D}}]=\mathsf{w}(\mu).$ (49)
+It remains to show that $\lim_{\mathcal{I}\ni N\to\infty}\frac{1}{N}\mathbb{E}[\max_{z^{N}\in\mathcal{C}_{\mu}}\left<Y^{N},z^{N}\right>1_{Y^{N}\in\mathcal{D}^{c}}]=0$, which, by the Cauchy-Schwarz inequality, would follow once we show
+$\limsup_{\mathcal{I}\ni N\to\infty}\mathbb{E}\left[\left(\frac{1}{N}\max_{z^{N}\in\mathcal{C}_{\mu}}\left<Y^{N},z^{N}\right>\right)^{2}\right]<\infty.$ (50)
+To show (50), first note that, due to $|T|<\infty$, we have $\|\frac{1}{N}\left<Y^{N},z^{N}\right>\|_{G}\leq\frac{C_{1}}{\sqrt{N}}$ for each $z^{N}\in\mathcal{C}$, where $C_{1},C_{2},C_{3}>0$ here are some constants independent of $N$.
+Since $\sqrt{\ln|\mathcal{C}_{\mu}|}\leq C_{2}\sqrt{N}$, by [43, Exercise 2.37] we have $\|\frac{1}{N}\max_{z^{N}\in\mathcal{C}_{\mu}}\left<Y^{N},z^{N}\right>\|_{G}\leq C_{3}\cdot\frac{C_{1}}{\sqrt{N}}\cdot C_{2}\sqrt{N}=C_{1}C_{2}C_{3}$.
+Then (50) follows since the sub-Gaussian norm dominates the square root of the second moment.
+∎
+We have the following estimate of the covering number $\mathsf{N}(\mathcal{C},\cdot)$ under the metric $d_{N}$:
+Lemma 6.
+Let $\mu$ be rational, and $\mathcal{C}:=\mathcal{C}_{\mu}$. For any $\sigma>0$, we have
+$\limsup_{\mathcal{I}\ni N\to\infty}\frac{1}{N}\ln\mathsf{N}(\mathcal{C},2\sqrt{N}\sigma)$ $\leq i_{\mu}(\sigma);$ (51)
+$\liminf_{\mathcal{I}\ni N\to\infty}\frac{1}{N}\ln\mathsf{N}(\mathcal{C},2\sqrt{N}\sigma)$ $\geq i_{\mu}(2\sigma).$ (52)
+Proof.
+For any $t^{N}\in\mathcal{C}$, we have
+$\mu^{\otimes N}(B(t^{N},\sqrt{N}\sigma))=\sum_{P_{Z\hat{Z}}\in\mathcal{S}_{N}}\mu^{\otimes N}(\mathcal{C}_{P_{Z\hat{Z}}}(t^{N}))$ (53)
+where $\mathcal{S}_{N}$ denotes set of all $P_{Z\hat{Z}}\in\Pi_{\sigma}(\mu)$ such that $NP_{Z\hat{Z}}$ is integer, and $\mathcal{C}_{P_{Z\hat{Z}}}(t^{N}):=\{u^{N}\in{T}^{N}\colon\textrm{$(t^{N},u^{N})$ has type $P_{Z\hat{Z}}$}\}$.
+Evidently,
+$\max_{P_{Z\hat{Z}}\in\mathcal{S}_{N}}\mu^{\otimes N}(\mathcal{C}_{P_{Z\hat{Z}}}(t^{N}))\leq\sum_{P_{Z\hat{Z}}\in\mathcal{S}_{N}}\mu^{\otimes N}(\mathcal{C}_{P_{Z\hat{Z}}}(t^{N}))\leq|\mathcal{S}_{N}|\max_{P_{Z\hat{Z}}\in\mathcal{S}_{N}}\mu^{\otimes N}(\mathcal{C}_{P_{Z\hat{Z}}}(t^{N}))$ (54)
+where $\lim_{\mathcal{I}\ni N\to\infty}\frac{1}{N}\ln|\mathcal{S}_{N}|=0$ by Lemma 1. By (22), we have
+$\left|\frac{1}{N}\ln\mu^{\otimes N}(\mathcal{C}_{P_{Z\hat{Z}}}(t^{N}))-I(Z;\hat{Z})\right|\leq c_{N}$ (55)
+where $c_{N}:=\frac{2|{T}|^{2}}{N}\ln(N+1)$. Next, we show that
+$\lim_{\mathcal{I}\ni N\to\infty}\max_{P_{Z\hat{Z}}\in\mathcal{S}_{N}}I(Z;\hat{Z})=i_{\mu}(\sigma).$ (56)
+Indeed, the $\leq$ part is obvious.
+For the $\geq$ part, let $P_{Z\hat{Z}}^{*}$ be the optimal distribution in the definition of $\Pi_{\sigma}(\mu,\mu)$.
+By Lemma 4, there exists $P^{N}_{Z\hat{Z}}\in\mathcal{S}_{N}$ such that $\max_{z,u}|P_{Z\hat{Z}}^{*}(z,u)-P^{N}_{Z\hat{Z}}(z,u)|\leq\frac{1}{N}$.
+Therefore $\lim_{\mathcal{I}\ni N\to\infty}I(Z;\hat{Z})=i_{\mu}(\sigma)$ and $\lim_{\mathcal{I}\ni N\to\infty}\mathbb{E}[d^{2}(Z,\hat{Z})]=\sigma^{2}$ where $(Z,\hat{Z})\sim P^{N}_{Z\hat{Z}}$.
+Since $i_{\mu}(\sigma)$ is continuous with respect to $\sigma$, this establishes the $\geq$ part of (56).
+Now combining (54), (55) and (56), we obtain
+$\left|\frac{1}{N}\ln\frac{1}{\mu^{\otimes N}(B(t^{N},\sqrt{N}\sigma))}-i_{\mu}(\sigma)\right|\leq c_{N}^{\prime}$ (57)
+where $c_{N}^{\prime}>0$ is a vanishing sequence possibly depending on $\sigma$ and $\mu$.
+Let ${\sf P}(\mathcal{C},2\sqrt{N}\sigma)$ denote the packing number, i.e.
+the maximal number of points $\mathcal{C}$ separated by distance at least $2\sqrt{N}\sigma$; then $\mathsf{N}(\mathcal{C},2\sqrt{N}\sigma)\leq{\sf P}(\mathcal{C},2\sqrt{N}\sigma)$ (see e.g.
+[42]), and standard volume estimates yields ${\sf P}(\mathcal{C},2\sqrt{N}\sigma)\leq\frac{1}{\min_{t^{N}\in\mathcal{C}}\mu^{\otimes N}(B(t^{N},\sqrt{N}\sigma))}$, proving (51).
+Moreover, by a volume argument we also have $\mathsf{N}(\mathcal{C},2\sqrt{N}\sigma)\cdot\max_{t^{N}\in\mathcal{C}}\mu^{\otimes N}(B(t^{N},2\sqrt{N}\sigma))\geq\mu^{\otimes N}(\mathcal{C})$.
+By Lemma 1, we have $\lim_{N\to\infty}\frac{1}{N}\ln\mu^{\otimes N}(\mathcal{C})=0$.
+Then using (57) with $\sigma$ replaced by $2\sigma$, we obtain (52).
+∎
+Proof of Theorem 1 (for finite $T$).
+When $|T|<\infty$, it suffices to consider the case of rational $\mu$, since the general case would then follow by a limiting argument.
+Applying Dudley’s integral (6) to $\mathcal{C}:=\mathcal{C}_{\mu}$ yields
+$\mathbb{E}[\max_{z^{N}\in\mathcal{C}}\left<Y^{N},z^{N}\right>]\leq K\int_{0}^{\Delta_{\rm m}}\sqrt{\ln\mathsf{N}(\mathcal{C},2\sqrt{N}\sigma)}2\sqrt{N}\,d\sigma$ (58)
+where $\Delta_{\rm m}:=\sup_{N\in\mathcal{I}}\frac{{\rm diam}(\mathcal{C})}{2\sqrt{N}}<\infty$.
+Now we multiply $1/N$ to (58), take $\mathcal{I}\ni N\to\infty$, and apply Lemma 6 and Lemma 5.
+We can apply the dominated convergence theorem, since $\frac{1}{N}\sqrt{\ln\mathsf{N}(\mathcal{C},2\sqrt{N}\sigma)}2\sqrt{N}\leq\frac{2}{\sqrt{N}}\sqrt{\ln(n^{N})}=2\sqrt{\ln n}$, which is independent of $N$ and absolutely integrable on $\sigma\in[0,\Delta_{\rm m}]$.
+Then Theorem 1 for finite $T$ follows from Lemma 6, Lemma 5, and (26).
+∎
+Proof of Theorem 2.
+As in the proof of Theorem 1, we can identify $T=\{e_{1},\dots,e_{n}\}$ as the standard basis vectors, so that $X_{z}=\left<Y,z\right$>.
+Furthermore, we assume that $|\mathcal{Y}|<\infty$ and $P_{YZ}$ is rational; the general case can then be proved by a limiting argument, since both sides of (2) are continuous with respect to the distribution when $|T|<\infty$.
+Define
+$\mathcal{I}:=\{N\in\mathbb{Z}\colon\textrm{$N>0$, $NP_{YZ}$ is integer}\}.$ (59)
+For each $N\in\mathcal{I}$, let $\mathcal{C}_{\mu}$ denote the type class of $\mu$, consisting of all sequences of type $\mu=P_{Z}$. Select
+$L:=\exp(NI(Y;Z)+N^{1/3})$ (60)
+sequences from $\mathcal{C}_{\mu}$ uniformly at random with replacement, and call this random set $\mathcal{A}$.
+Let $Y^{N}\sim P_{Y}^{\otimes N}$ (independent of this random set).
+Then we will show that
+$\mathbb{E}[\left<Y,Z\right>]\leq\liminf_{\mathcal{I}\ni N\to\infty}\mathbb{E}[\max_{z^{N}\in\mathcal{A}}\frac{1}{N}\left<Y^{N},z^{N}\right$>]. (61)
+Indeed, set
+$\mathcal{B}$ $:=\{(y^{N})\colon\widehat{P}_{y^{N}}=P_{Y}\};$ (62)
+$\mathcal{D}$ $:=\{(y^{N},z^{N})\colon\widehat{P}_{y^{N}z^{N}}=P_{YZ}\},$ (63)
+and for any $y^{N}\in\mathcal{B}$, define $\mathcal{D}(y^{N}):=\{z^{N}\colon(y^{N},z^{N})\in\mathcal{D}\}$. Standard results about type estimates (immediate from Lemma 2 and Lemma 1) state that:
+$\left|\frac{1}{N}\ln|\mathcal{D}(y^{N})|-H(Z|Y)\right|$ $\leq C_{1}\frac{\ln N}{N},\quad\forall y^{N}\in\mathcal{B},$ (64)
+$\left|\frac{1}{N}\ln|\mathcal{C}_{\mu}|-H(Z)\right|$ $\leq C_{1}\frac{\ln N}{N},$ (65)
+where $C_{1},C_{2}\dots$ in this proof are positive constants independent of $N$.
+Let $\pi$ be the projection from $\mathbb{R}^{nN}$ to $\mathcal{B}$ (i.e., map to the closest point in $\mathcal{B}$, with ties broken arbitrarily).
+Define $\mathcal{E}$ as the event that $\mathcal{D}(\pi(Y^{N}))\cap\mathcal{A}=\emptyset$.
+Then, (64) implies that
+$\mathbb{P}[\mathcal{E}]$ $=\left(1-|\mathcal{D}(y^{N})|/|\mathcal{C}_{\mu}|\right)^{L}$ (66)
+$\leq\left(1-\exp(-NI(Y;Z)-2C_{1}\ln N)\right)^{L}$ (67)
+$\leq\exp(-C_{2}N^{1/3}),$ (68)
+where $y^{N}$ is any sequence in $\mathcal{B}$. From the definitions it is clear that
+$\max_{z^{N}\in\mathcal{A}}\frac{1}{N}\left<\pi(Y^{N}),z^{N}\right>1_{\mathcal{E}^{c}}\geq\mathbb{E}[\left<Y,Z\right>]1_{\mathcal{E}^{c}}$ (69)
+for all $N\in\mathcal{I}$ almost surely.
+By our assumption of $|T|,|\mathcal{Y}|<\infty$, we have $\left|\max_{z^{N}\in\mathcal{A}}\frac{1}{N}\left<\pi(Y^{N}),z^{N}\right>\right|\leq\max_{y\in\mathcal{Y},z\in T}|\left<y,z\right>|\leq C_{3}$, therefore $\lim_{\mathcal{I}\ni N\to\infty}\mathbb{P}[\mathcal{E}]=0$ and the Cauchy-Schwarz inequality shows that
+$\limsup_{\mathcal{I}\ni N\to\infty}\mathbb{E}\left[\left|\max_{z^{N}\in\mathcal{A}}\frac{1}{N}\left<\pi(Y^{N}),z^{N}\right>\right|\cdot 1_{\mathcal{E}}\right]=0.$ (70)
+Then (69) implies
+$\liminf_{\mathcal{I}\ni N\to\infty}\left[\max_{z^{N}\in\mathcal{A}}\frac{1}{N}\left<\pi(Y^{N}),z^{N}\right>\right]\geq\mathbb{E}[\left<Y,Z\right$>]. (71)
+Furthermore,
+$\max_{z^{N}\in\mathcal{A}}\frac{1}{N}\left<Y^{N},z^{N}\right>\geq\max_{z^{N}\in\mathcal{A}}\frac{1}{N}\left<\pi(Y^{N}),z^{N}\right>-\max_{z^{N}\in\mathcal{A}}\frac{1}{N}\left<\pi(Y^{N})-Y^{N},z^{N}\right$> (72)
+From Lemma 3, we have
+$\max_{z^{N}\in\mathcal{A}}\frac{1}{N}\left<\pi(Y^{N})-Y^{N},z^{N}\right$> $\leq\frac{1}{\sqrt{N}}\|\pi(Y^{N})-Y^{N}\|_{2}\cdot\frac{1}{\sqrt{N}}\max_{z^{N}\in\mathcal{A}}\|z^{N}\|_{2}$ (73)
+$\leq C_{4}W_{2}(P_{Y},\widehat{P}_{Y^{N}})\cdot$ (74)
+which converges to 0 in expectation [8]. Thus (61) is verified by collecting the bounds above.
+It remains to apply the Dudley inequality to the right side of (61).
+Since $\mathcal{A}$ is contained in the set $\mathcal{C}$ in Lemma 6, the upper bound part of Lemma 6 also applies to covering numbers of $\mathcal{A}$.
+Furthermore, we have the trivial bound $|\mathcal{A}|\leq L$.
+Therefore,
+$\limsup_{\mathcal{I}\ni N\to\infty}\frac{1}{N}\ln\mathsf{N}(\mathcal{A},2\sqrt{N}\sigma)\leq i_{\mu}(\sigma)\wedge I(Y;Z).$ (75)
+The claim then follows by applying dominated convergence (see proof of Theorem 1). ∎
+5 Proof of sharpness of the rate-distortion integral
+In this section we present a proof of Theorem 3 using sharpness of Dudley’s integral for stationary Gaussian processes.
+We first consider the case where $n:=|T|<\infty$ and $\mu$ is rational.
+Recall the representation of $X_{z}$ as an inner product in (37), which embeds $Y,Z$ in $\mathbb{R}^{n}$ where $Y$ is a normal vector, and the definitions of $Y^{N}$, $\mathcal{I}$, and $\mathcal{C}:=\mathcal{C}_{\mu}$ as in (37), (38), (39).
+Then $(\left<Y^{N},z^{N}\right>)_{z^{N}\in\mathcal{C}}$ is a stationary Gaussian process with respect to the metric (41), according to Definition 1, due to invariance under the transitive group of coordinate permutations acting on $\mathcal{C}$.
+By Lemma 5 and the sharpness of Dudley’s integral for stationary Gaussian processes (7), we see that
+$K^{\prime}{\sf w}(\mu)$ $\geq\liminf_{\mathcal{I}\ni N\to\infty}\frac{1}{N}\int_{0}^{\infty}\sqrt{\ln\mathsf{N}(\mathcal{C},\sqrt{N}\sigma)}\sqrt{N}\,d\sigma$ (76)
+$\geq\int_{0}^{\infty}\sqrt{i_{\mu}(\epsilon)}d\epsilon$ (77)
+where $\mathcal{I}$ is defined in (38), and the last step follows from Lemma 6 and Fatou’s Lemma.
+The lower bound in terms of $R_{\mu}$ follows from (26).
+This establishes the theorem for finite $T$ and rational $\mu$.
+Next, we can drop the rationality assumption under $|T|<\infty$: It is easy to see that both $\mathsf{w}(\mu)$ and $\int_{0}^{\infty}\sqrt{i_{\mu}(\sigma)}d\sigma$ are continuous with respect to $\mu$, so that the rational case implies the general case by a limiting argument.
+For extension to countable case see the appendix.
+6 Penalization form and new proof of the majorizing measure theorem
+In this section, we prove Corollary 1.
+First, observe the following auxiliary result allowing us to write integrals in a “penalized optimization” form, which will be useful not only for proof of Corollary 3 but also in bounding the errors for empirical risk minimization.
+Lemma 7.
+Suppose that $y\in(0,\infty)$ is a decreasing function of $x\in(0,\infty)$. Then
+$2\int_{0}^{\infty}ydx\leq\int_{0}^{\infty}\min_{x}\{\alpha^{-2}x^{2}+y^{2}\}d\alpha\leq 4\int_{0}^{\infty}ydx$ (78)
+Proof.
+For any $\alpha\in(0,\infty)$, there is unique $x(\alpha)$ satisfying $\alpha=\frac{x(\alpha)}{y(x(\alpha))}$.
+Then $x$ and $y$ can be parameterized by $\alpha\in(0,\infty)$, and $\int_{0}^{\infty}(\alpha^{-2}x^{2}+y^{2})d\alpha=\int_{\alpha=0}^{\infty}2y^{2}\cdot\frac{ydx-xdy}{y^{2}}=4\int_{0}^{\infty}ydx$.
+The claim then follows since for any $\alpha$, $\frac{\min_{x\in(0,\infty)}\{\alpha^{-2}x^{2}+y^{2}(x)\}}{\alpha^{-2}x^{2}(\alpha)+y^{2}(x(\alpha))}\in[1/2,1]$.
+∎
+A similar idea (though in the form of series summation rather than integral) was previously used in proving equivalent forms of the majorizing measure theorem [38].
+Here, we find Lemma 7 very convenient in forming information-theoretic functionals that are exactly convex-concave (Remark 3), and also simplifying evaluations of the rate-distortion integral in applications to ERM.
+To prove Corollary 1, it suffices to consider the case of finite $T$, since the general case (where the lattice supremum in (35) is defined as $\sup_{T_{0}\subseteq T,|T_{0}|<\infty}\sup_{t\in T_{0}}$; see [39]) will then follows by taking the limit.
+Then $T$ can be identified with points in $\mathbb{R}^{|T|}$ equipped with the $\ell_{2}$ metric.
+For $i:=i_{\mu}$ in Definition 3, Lemma 7 implies
+$4\sup_{\mu}\int_{0}^{\infty}\sqrt{i(\epsilon)}d\epsilon\geq\sup_{P_{Z}}\int_{0}^{\infty}\min_{P_{Z\hat{Z}}\in\Pi(P_{Z},P_{Z})}\{\alpha^{-2}\mathbb{E}[\|Z-\hat{Z}\|_{2}^{2}]+I(Z;\hat{Z})\}d\alpha$ (79)
+where both supremums are over distributions on $T$ and the minimum is over couplings of $P_{Z}$ and $P_{Z}$. The right side of (79) is lower bounded by
+$\sup_{P_{Z}}\inf_{\mu}\int_{0}^{\infty}\min_{P_{U|Z}}\{\alpha^{-2}\mathbb{E}[\|Z-U\|_{2}^{2}]+D(P_{U|Z}\|\mu|P_{Z})\}d\alpha,$ (80)
+since equality would be achieved if $\inf_{\mu}$ were restricted to $\mu=P_{Z}$ and $\min_{P_{U|Z}}$ were restricted to $P_{U|Z}$ satisfying $P_{U|Z}P_{Z}\in\Pi(P_{Z},P_{Z})$.
+Now define, for each $z,\alpha$, $F_{\mu}(z,\alpha):=\min_{\nu}\{\alpha^{-2}\mathbb{E}_{U\sim\nu}[\|z-U\|_{2}^{2}+D(\nu\|\mu)]\}$.
+Since the relative entropy is jointly convex (i.e., $D(\frac{\nu_{1}+\nu_{2}}{2}\|\frac{\mu_{1}+\mu_{2}}{2})\leq\frac{1}{2}D(\nu_{1}\|\mu_{1})+\frac{1}{2}D(\nu_{2}\|\mu_{2})$), we see that $F_{\mu}(z,\alpha)$ is convex in $\mu$.
+The integral in (80) equals $\int_{0}^{\infty}\mathbb{E}[F_{\mu}(Z,\alpha)]d\alpha$, which is convex in $\mu$ and linear in $P_{Z}$.
+Then Sion’s minimax theorem implies that (80) equals
+$\quad\inf_{\mu}\sup_{P_{Z}}\int_{0}^{\infty}\mathbb{E}[F_{\mu}(Z,\alpha)]d\alpha=\inf_{\mu}\max_{z\in T}\int_{0}^{\infty}F_{\mu}(z,\alpha)d\alpha.$ (81)
+Now fix any $\mu$, $z$, $\alpha$, and let $P_{U}:=\nu$ achieve the minimum in the definition of $F_{\mu}(z,\alpha)$.
+Define $r=a\sqrt{\mathbb{E}[\|z-U\|^{2}]}$, where $U\sim P_{U}$ and $a>1$ can be optimized later.
+We obtain, from the data processing inequality [12],
+$D(P_{U}\|\mu)\geq d(P_{U}(B(z,r))\|\mu(B(z,r)))$ (82)
+where $d(a\|b):=a\ln\frac{a}{b}+(1-a)\ln\frac{1-a}{1-b}$ denotes the binary relative entropy function and $B(z,r)$ denotes the ball centered at $z$ with radius $r$.
+By the Markov inequality we have $P_{U}(B(z,r))\geq 1-a^{-2}$.
+Then we have
+$d(P_{U}(B(z,r))\|\mu(B(z,r)))\geq(1-a^{-2})\ln\frac{1-a^{-2}}{\mu(B(z,r))}-a^{-2}\ln a^{2}$ (83)
+because if $\mu(B(z,r))>1-a^{-2}$ then the right side of (83) is negative and the inequality is obviously true; if $\mu(B(z,r))\leq 1-a^{-2}$ then by monotonicity of the binary relative entropy function we have $d(P_{U}(B(z,r))\|\mu(B(z,r)))\geq d(1-a^{-2}\|\mu(B(z,r)))=(1-a^{-2})\ln\frac{1-a^{-2}}{\mu(B(z,r))}+a^{-2}\ln\frac{a^{-2}}{1-\mu(B(z,r))}$ and (83) also holds.
+Now the second integral in (81) is lower bounded by
+$\int_{0}^{\infty}\inf_{r>0}\left\{\frac{r^{2}}{a^{2}\alpha^{2}}+\left[(1-a^{-2})\ln\frac{1-a^{-2}}{\mu(B(z,r))}-a^{-2}\ln a^{2}\right]_{+}\right\}d\alpha$ (84)
+which, by Lemma 7, is lower bounded by twice of
+$\int_{0}^{{\rm diam}(T)}\sqrt{\left[(1-a^{-2})\ln\frac{1-a^{-2}}{\mu(B(z,r))}-a^{-2}\ln a^{2}\right]_{+}}\frac{dr}{a}$
+$\geq\int_{0}^{{\rm diam}(T)}\sqrt{(1-a^{-2})\ln\frac{1}{\mu(B(z,r))}}\frac{dr}{a}-\int_{0}^{{\rm diam}(T)}\sqrt{h(a^{-2})}\frac{dr}{a}$ (85)
+$=\frac{\sqrt{1-a^{-2}}}{a}I_{\mu}(z)-\frac{{\rm diam}(T)}{a}\sqrt{h(a^{-2})}$ (86)
+where we used the basic inequality $\sqrt{[x-y]_{+}}\geq\sqrt{x}-\sqrt{y}$ for $x,y\geq 0$. Thus (77) implies
+$2K^{\prime}\mathbb{E}[\sup_{z\in T}X_{z}]\geq\inf_{\mu}\sup_{z}\frac{\sqrt{1-a^{-2}}}{a}I_{\mu}(z)-\frac{{\rm diam}(T)}{a}\sqrt{h(a^{-2})}.$ (87)
+It is easy to see, by considering two points in $T$, that $\mathbb{E}[\sup_{z\in T}X_{z}]\geq{\rm diam}(T)\mathbb{E}[G\vee 0]=\sqrt{\frac{1}{2\pi}}{\rm diam}(T)$, where $G\sim\mathcal{N}(0,1)$.
+We finish the proof of Corollary 1 by
+$\left(2K^{\prime}+\frac{\sqrt{2\pi h(a^{-2})}}{a}\right)\mathbb{E}[\sup_{z\in T}X_{z}]\geq\frac{\sqrt{1-a^{-2}}}{a}\inf_{\mu}\sup_{z}I_{\mu}(z).$ (88)
+Remark 3.
+The minimax inequality step (80)-(81) crucially leverages Lemma 7 and the joint convexity of the relative entropy.
+Specifically, the integral in (80) is convex in $\mu$ and linear in $P_{Z}$, and this property is not true for the functional $\int_{0}^{\infty}\min_{P_{U|Z}\colon\mathbb{E}[\|Z-U\|_{2}^{2}\leq\epsilon^{2}}\sqrt{D(P_{U|Z}\|\mu|P_{Z})}d\epsilon$ before we applied Lemma 7.
+Moreover, had we not used the relative entropy approach, then the integral in (80) might be replaced by the functional $\int_{0}^{\infty}\min_{r_{Z}}\{\alpha^{-2}\mathbb{E}[r_{Z}^{2}]+\ln\frac{1}{\mu(B(Z,r_{Z}))}\}d\alpha$ (see the $\theta_{i}(T)$ functional defined around [38, Proposition 4.3]), where $P_{U|Z}$ is replaced by the conditional distribution of $\mu$ on the ball $B(Z,r_{Z})$, $r_{Z}$ being a function of $Z$.
+Again, this functional is not exactly convex in $\mu$, so the application of the minimax theorem is less immediate.
+A classical approach that has a similar role as (80)-(81) is the Fernique convexity argument, which establishes equivalences (11)-(12) but loses a constant factor [39].
+7 Proof of ERM over the Sobolev ellipsoid
+In this section we apply the rate-distortion integral to the ellipsoid, and in particular, prove Theorem 4.
+7.1 Bounding the rate-distortion integral by variance
+A key step in the analysis for the ellipsoid (and possibly many other examples) is to upper bound the rate-distortion integral using the covariance matrix.
+One possible approach is to show that, under the constraint $\operatorname*{cov}(Z)=\Sigma$, the Gaussian distribution maximizes the rate-distortion function, i.e.
+$R_{P_{Z}}(\sigma^{2})$ $\leq R_{\mathcal{N}(0,\Sigma)}(\sigma^{2})$ (89)
+for all $\sigma^{2}>0$.
+While it is possible to prove (89) using Lieb’s rotation argument for proving Gaussian optimality (see [29, 11] for the analysis of a similar problem called the reverse Brascamp-Lieb inequality), we will not provide the details here.
+Instead, we will use an alternative, simpler approach, leveraging the penalized version of the rate-distortion integral: from Lemma 7, we have
+$\int_{0}^{\infty}\sqrt{R_{P_{Z}}(\sigma^{2})}\,d\sigma\leq\frac{1}{2}\int_{0}^{\infty}\inf_{P_{U|Z}}\{\alpha^{-2}\mathbb{E}[\|U-Z\|^{2}]+I(U;Z)\}d\alpha.$ (90)
+This penalized form is more convenient, because $P_{U|Z}$ is optimized over a set independent of $P_{X}$, so that saddle point analysis applies, and because exact tensorization holds, allowing direct extension of 1-dimensional bounds to the vector case.
+Define
+$f(t):=\left\{\begin{array}[]{cc}t/2&0\leq t\leq 1\\ \frac{1+\ln t}{2}&t>1\end{array}\right.$. (93)
+We have the following:
+Lemma 8.
+For any distribution $P_{Z}$ on $\mathbb{R}$, we have
+$\inf_{P_{U|Z}}\{\alpha^{-2}\mathbb{E}[|U-Z|^{2}]+I(U;Z)\}$ $\leq f(2\alpha^{-2}\sigma_{\rm m}^{2})$ (94)
+for any given $\alpha>0$, where $\sigma_{\rm m}^{2}:=\operatorname*{var}(Z)$.
+Proof.
+By translation invariance, it suffices to assume that $\mathbb{E}[|Z|^{2}]\leq\sigma_{\rm m}^{2}$.
+Define the functional $f(P_{Z},P_{U|Z}):=\alpha^{-2}\mathbb{E}[|U-Z|^{2}]+I(U;Z)$, which is convex in $P_{U|Z}$ and concave in $P_{Z}$, and we have the saddle point condition
+$\sup_{P_{Z}\colon\mathbb{E}[|Z|^{2}]\leq\sigma_{\rm m}^{2}}f(P_{Z},P_{U|Z}^{*})\leq f(P_{Z}^{*},P_{U|Z}^{*})\leq\inf_{P_{U|Z}}f(P_{Z}^{*},P_{U|Z})$ (95)
+where $P_{Z}^{*}=\mathcal{N}(0,\sigma_{\rm m}^{2})$ and $P_{U|Z}^{*}$ is an additive noise Gaussian channel that achieves the infimum; see “Gaussian saddle point” in [12].
+Then we have $\inf_{P_{U|Z}}f(P_{Z}^{*},P_{U|Z})\geq\sup_{P_{Z}\colon\mathbb{E}[|Z|^{2}]\leq\sigma_{\rm m}^{2}}f(P_{Z},P_{U|Z}^{*})\geq\sup_{P_{Z}\colon\mathbb{E}[|Z|^{2}]\leq\sigma_{\rm m}^{2}}\inf_{P_{U|Z}}f(P_{Z},P_{U|Z})$.
+This shows that under $\mathbb{E}[|Z|]\leq\sigma_{\rm m}^{2}$, the right side of (90) is maximized by $P_{Z}=\mathcal{N}(0,\sigma_{\rm m}^{2})$, in which case
+$\inf_{P_{U|Z}}\{\alpha^{-2}\mathbb{E}[|U-Z|^{2}]+I(U;Z)\}$ $=\inf_{D>0}\left\{\alpha^{-2}D+\frac{1}{2}[\ln\frac{\sigma_{\rm m}^{2}}{D}]_{+}\right\}$ (96)
+$=(\frac{1}{2}+\frac{1}{2}\ln\frac{2\sigma_{\rm m}^{2}}{\alpha^{2}})1\{\alpha<\sqrt{2}\sigma_{\rm m}\}+\alpha^{-2}\sigma_{\rm m}^{2}1\{\alpha\geq\sqrt{2}\sigma_{\rm m}\}$ (97)
+as desired. ∎
+Next, we extend Lemma 8 to the case of vectors followed by a Lipschitz transformation:
+Lemma 9.
+Assume that $W\in\ell_{2}$ is a random variable and $\Sigma:=\operatorname*{cov}(W)$.
+Let $Z=\phi(W)$, where $\phi$ is a 1-Lipschitz function, and suppose that $\mathbb{E}[\|Z\|^{2}]\leq\epsilon$.
+Then, for any $\alpha\in(0,\infty)$,
+$\inf_{P_{U|Z}}\{\alpha^{-2}\mathbb{E}[\|U-Z\|^{2}]+I(U;Z)\}\leq\alpha^{-2}\epsilon\wedge\sum_{i=1}^{\infty}f(2\alpha^{-2}\Sigma_{ii}).$ (98)
+Proof.
+By considering $U=0$, it is easy to see that $\inf_{P_{U|X}}\{\alpha^{-2}\mathbb{E}[\|U-Z\|^{2}]+I(U;Z)\}\leq\alpha^{-2}\epsilon$. Thus it remains to show that
+$\inf_{P_{U|Z}}\{\alpha^{-2}\mathbb{E}[\|U-Z\|^{2}]+I(U;Z)\}\leq\sum_{i=1}^{\infty}f(2\alpha^{-2}\Sigma_{ii}).$ (99)
+It suffices to prove (99) in the special case where the 1-Lipschitz map $\phi$ is the identity, since the left side of (99) only decreases with the application of a 1-Lipschitz $\phi$.
+Furthermore, the left side of (99) is upper bounded by $\sum_{i=1}^{\infty}\inf_{P_{U_{i}|Z_{i}}}\{\alpha^{-2}\mathbb{E}[|U_{i}-Z_{i}|^{2}]+I(U_{i};Z_{i})\}$, because given $P_{U_{i}|Z_{i}}$ we can define $P_{U|Z}=\prod_{i=1}^{\infty}P_{U_{i}|Z_{i}}$, and then $I(U;Z)\leq\sum_{i=1}^{\infty}I(U_{i};Z_{i})$.
+Then the result in (99) follows from the one-dimensional case, which we already proved in Lemma 8.
+∎
+7.2 Warm-up: supremum over an ellipsoid
+Let $X_{z}:=\left<G,z\right$>, where $G=(G(1),G(2),\dots)$ has i.i.d. standard normal coordinates. The majorizing measure theorem shows that (see [39, Section 2.13]):
+$\mathbb{E}[\sup_{z\in\mathcal{E}}X_{z}]\asymp\sqrt{\sum_{i=0}^{\infty}2^{i}a_{2^{i}}^{2}}\asymp\sqrt{\sum_{i=1}^{\infty}a_{i}^{2}}.$ (100)
+On the other hand, a nontrivial calculation [39] shows that the Dudley integral is $\asymp\sum_{i=0}^{\infty}2^{i/2}a_{2^{i}}$ hence may not be sharp.
+Here we show that sharpness is regained using the rate-distortion integral.
+Suppose that $Z^{\prime}$ is a random variable taking values in $\mathcal{E}$ and $\Sigma:=\operatorname*{cov}(Z^{\prime})$.
 We have
-$\mathbb{E}[\left<Y,Z\right>]=\lim_{N\to\infty}\mathbb{E}[\frac{1}{N}\left<Y^{N},Z^{N}\right$>]. (56)
-By Proposition 2, $(\left<Y^{N},z^{N}\right>)_{z^{N}\in\mathcal{C}}$ is a sub-Gaussian process with respect to the metric $d_{N}(z^{N},\hat{z}^{N}):=\sqrt{\sum_{i=1}^{N}d^{2}(z_{i},\hat{z}_{i})}$.
-Let $1_{\mathcal{C}}$ be the indicator of the event that $Z^{N}\in\mathcal{C}$.
-Since $\left<Y^{N},Z^{N}\right>1_{\mathcal{C}}\leq\max_{z^{N}\in\mathcal{C}}\left<Y^{N},z^{N}\right>1_{\mathcal{C}}=\max_{z^{N}\in\mathcal{C}}\left<Y^{N},z^{N}\right>-\max_{z^{N}\in\mathcal{C}}\left<Y^{N},z^{N}\right>1_{\mathcal{C}^{c}}$, we have
-$\mathbb{E}[\left<Y^{N},Z^{N}\right$>] $\leq\mathbb{E}[\left<Y^{N},Z^{N}\right>1_{\mathcal{C}^{c}}]+\mathbb{E}[\max_{z^{N}\in\mathcal{C}}\left<Y^{N},z^{N}\right>]-\mathbb{E}[\max_{z^{N}\in\mathcal{C}}\left<Y^{N},z^{N}\right>1_{\mathcal{C}^{c}}]$ (57)
-$\leq\mathbb{E}[\max_{z^{N}\in\mathcal{C}}\left<Y^{N},z^{N}\right>]+\mathbb{E}[\max_{z^{N},\hat{z}^{N}\in\mathcal{Z}^{N}}\left<Y^{N},z^{N}-\hat{z}^{N}\right>1_{\mathcal{C}^{c}}].$ (58)
-The last term above has vanishing contribution: Indeed, note that for each $z^{n},\hat{z}^{n}\in\mathcal{Z}^{N}$, we have $\|\left<Y^{N},z^{N}-\hat{z}^{N}\right>\|_{\Psi_{2}}\leq C\sqrt{N}$ for some $C$ independent of $N$ (may depend on $P_{YZ}$).
-Taking the max over $\exp(O(N))$ such sub-Gaussian random variables will increase the sub-Gaussian norm by a factor of no more than $\sqrt{N}$ (Vershynin,, 2018), so
-$\|\max_{z^{N},\hat{z}^{N}\in\mathcal{Z}^{N}}\left<Y^{N},z^{N}-\hat{z}^{N}\right>\|_{\Psi_{2}}\leq CN.$ (59)
-Moreover, by Hoeffding’s inequality we also have $\mathbb{E}[1_{\mathcal{C}^{c}}]=\exp(-\Theta(N^{1/3}))$, so we can use Cauchy-Schwarz (or (Talagrand,, 2014, Lemma 7.8.4)) to show the second expectation in (58) is $o(N)$.
-Therefore, we have verified that
-$\mathbb{E}[\left<Y,Z\right>]\leq\liminf_{N\to\infty}\frac{1}{N}\mathbb{E}[\max_{z^{N}\in\mathcal{C}}\left<Y^{N},z^{N}\right$>]. (60)
-Furthermore, applying Dudley’s integral (see (7) and the subsequent comments) to $\mathcal{C}$ yields
-$\mathbb{E}[\max_{z^{N}\in\mathcal{C}}\left<Y^{N},z^{N}\right>]\leq 24\int_{0}^{\Delta_{\rm m}}\sqrt{\ln N(\mathcal{C},2\sqrt{N}\sigma)}2\sqrt{N}d\sigma$ (61)
-where $\Delta_{\rm m}:=\sup_{N\geq 1}\frac{{\rm diam}(\mathcal{C}_{N})}{2\sqrt{N}}<\infty$.
-Now we multiply $1/N$ to (61), take $N\to\infty$, and apply Lemma 4.
-We can apply dominated convergence, since $\frac{1}{N}\sqrt{\ln N(\mathcal{C},2\sqrt{N}\sigma)}2\sqrt{N}\leq\frac{2}{\sqrt{N}}\sqrt{\ln(n^{N})}=2\sqrt{\ln n}$, which is independent of $N$ and absolutely integrable on $\sigma\in[0,\Delta_{\rm m}]$.
-We have thus proved Theorem 1 for finite $T$.
-∎
-The proof of Theorem 2 will be almost identical to the proof of Theorem 1, but we will use an additional technique about the projection to the set of sequences of a given type. More precisely:
+$\int_{0}^{\infty}\sqrt{R_{P_{Z^{\prime}}}(\sigma^{2})}\,d\sigma$ $\lesssim\sum_{i=1}^{\infty}\sqrt{\Sigma_{ii}}$ (101)
+$\leq\|(a_{i})_{i=1}^{\infty}\|_{2}$ (102)
+where follows (101) from the (90) and Lemma 9, noting that $\int_{0}^{\infty}f(2\alpha^{-2}\Sigma_{ii})d\alpha\asymp\sqrt{\Sigma_{ii}}$, and (102) follows from the ellipsoid constraint $\sum_{i=1}^{\infty}\frac{\Sigma_{ii}}{a_{i}^{2}}\leq 1$ and the Cauchy-Schwarz inequality.
+Then for any random variable $Z\in\mathcal{E}$, we have $\mathbb{E}[X_{Z}]\lesssim\|(a_{i})_{i=1}^{\infty}\|_{2}$, which follows from Proposition 1, by taking $T$ and $\mu^{\prime}$ therein to be $\mathcal{E}$ and $P_{Z^{\prime}}$ in (102), respectively.
+Let us remark that the bound can also be obtained using Corollary 3, since $\operatorname*{tr}(\sqrt{\Sigma})\leq\sum_{i=1}^{\infty}\sqrt{\Sigma_{ii}}$ by the Schur concavity of the square root function and the fact that the eigenvalues majorizes the diagonal values.
+Furthermore, the argument here also shows that the same upper bound holds when $T$ is the image of the ellipsoid $\mathcal{E}$ under a 1-Lipschitz map.
+For the case of the Sobolev ellipsoid, we see the supremum in (100) is order $\sqrt{\frac{1}{2\beta-1}}$ as $\beta\downarrow 1/2$. In contrast, the Dudley integral is
+$\sum_{i=0}^{\infty}2^{i/2}a_{2^{i}}=\sum_{i=0}^{\infty}2^{-\beta i+i/2},$ (103)
+giving a worse upper bound of order $\frac{1}{2\beta-1}$ as $\beta\downarrow 1/2$; see [44, Example 5.12] or [39, Section 2.13].
+7.3 Error bound for ERM
+In this section we prove Theorem 4.
+Recall (18), where we defined the centered sub-Gaussian process $(\chi_{z})_{T-T}$, and $T$ is now taken to be a 1-Lipschitz image of the Sobolev ellipsoid.
+Our proof is based on the following key estimate:
+Theorem 5.
+If $\mathbb{E}[\|Z\|_{2}^{2}]\leq E$ for some $E>0$, then $\mathbb{E}[\chi_{Z}]\lesssim C_{\beta}E^{\frac{2\beta-1}{4\beta}}$ for some function $C_{\beta}>0$ (defined around (124)) satisfying $\lim_{\beta\downarrow\frac{1}{2}}\sqrt{2\beta-1}C_{\beta}<\infty$.
+Remark 4.
+In [41, Lemma 5.1], the ERM error for the Sobolev space of functions was bounded by localizing the Dudley integral, where $\beta\geq 1$ is assumed an integer.
+Suppose that we now extend the approach of [41] to the setting of Theorem 4: The covering number is bounded as
+$\ln N(\delta,\mathcal{E})\lesssim\delta^{-\frac{1}{\beta}}$ (104)
+where we assume throughout this Remark that $\beta$ is bounded above (i.e.
+we only focus on the behavior of the bound as $\beta\downarrow 1/2$).
+[41, Example 2.1] estimated the covering number for fixed $\beta$.
+For the Sobolev ellipsoid $\mathcal{E}$, [44, Example 5.12] provided a proof of (104), from which we can see that the best preconstant in (104) is the order of a constant independent of $\beta$.
+For local covering numbers, [45] showed that the estimate in (104) still holds, where the preconstant is a universal constant.
+We can then calculate the Dudley integral up to the scale of the error $\sqrt{E}$, resulting in a bound of $\int_{0}^{\sqrt{E}}\delta^{-\frac{1}{2\beta}}d\delta\lesssim\frac{1}{2\beta-1}E^{\frac{1}{2}-\frac{1}{4\beta}}$.
+Using our proof instead, we get $\mathbb{E}[\chi_{Z}]\lesssim\frac{1}{\sqrt{2\beta-1}}E^{\frac{1}{2}-\frac{1}{4\beta}}$, which is almost the same, but with a better preconstant.
+Thus Dudley’s integral results in $\mathbb{E}[\|\hat{m}-m\|^{2}]\lesssim\frac{1}{2\beta-1}n^{-\frac{2\beta}{2\beta+1}}$, where the preconstant $\frac{1}{2\beta-1}$ is worse than $\frac{1}{\sqrt{2\beta-1}}$ in Theorem 4.
+This gap is related to the looseness of the Dudley integral: as we see in (103), Dudley’s integral gives a bound of order $\frac{1}{2\beta-1}$ on the supremum over $\mathcal{E}$, rather than the optimal $\sqrt{\frac{1}{2\beta-1}}$.
+How does the rate-distortion integral improve the preconstant?
+A naive approach would be to first show that for a random variable $Z\in 2\mathcal{E}$, we have $R_{P_{Z}}(\delta^{2})\lesssim\delta^{-\frac{1}{\beta}}$, where we assume throughtout this paragraph that $\beta$ is bounded above (i.e.
+we only focus on the behavior of the bound as $\beta\downarrow 1/2$).
+This is analogous to (104), and we may proceed by plugging it into the rate-distortion integral.
+But this will still result in a preconstant $\frac{1}{2\beta-1}$ rather than $\frac{1}{\sqrt{2\beta-1}}$.
+The reason is that the $P_{Z}$ maximizing $R_{P_{Z}}(\delta^{2})$ changes with $\delta$.
+Instead, we leverage the penalized version (90) and evaluate the integrand therein for given $\operatorname*{cov}(Z)$, calculate the integral, and optimize $\operatorname*{cov}(Z)$ in the end.
+The proof is made simple by properties of information-theoretic measures such as chain rules of the entropy.
+We first need a few preparations. We have
 Lemma 10.
-Let $\mu$ be a rational distribution on $\mathbb{R}^{n}$. Fix $N\geq 1$ such that $N\mu$ is integer. For any $y^{N}\in\mathbb{R}^{nN}$, we have
-$\frac{1}{\sqrt{N}}\min_{s^{N}\colon\widehat{P}_{s^{N}}=\mu}\|y^{N}-s^{N}\|_{2}=W_{2}(\widehat{P}_{y^{N}},\mu).$ (62)
+$t\int_{0}^{1/t}f(\frac{1}{u^{2}})du\leq 4f(t)$, for all $t>0$.
 Proof.
+For $0<u_{0}\leq 1$,
+$\int_{0}^{u_{0}}f(\frac{1}{u^{2}})du=\int_{0}^{u_{0}}\frac{1+\ln\frac{1}{u^{2}}}{2}du=\frac{3}{2}u_{0}+u_{0}\ln\frac{1}{u_{0}}\leq\frac{3}{2}u_{0}(1+\ln\frac{1}{u_{0}}).$ (105)
+For $u_{0}>1$, $\int_{1}^{u_{0}}f(\frac{1}{u^{2}})du=\int_{1}^{u_{0}}\frac{1}{2u^{2}}du=\frac{1}{2}(1-\frac{1}{u_{0}})$, so that
+$\int_{0}^{u_{0}}f(\frac{1}{u^{2}})du=\frac{3}{2}+\frac{1}{2}(1-\frac{1}{u_{0}})=2-\frac{1}{2u_{0}}\leq 2.$ (106)
+The result follows by taking $u_{0}=1/t$. ∎
+Lemma 11.
+Suppose that $\sum_{i=1}^{\infty}\Sigma_{ii}i^{\beta}\leq 4$, where $\beta>1/2$. For any E>0{E}>0, we have
+$\int_{0}^{\infty}\alpha^{-2}{E}\wedge\sum_{i=1}^{\infty}f(2\alpha^{-2}\Sigma_{ii})d\alpha\leq C_{\beta}{E}^{\frac{2\beta-1}{4\beta}},$ (107)
+where $C_{\beta}$ is a constant defined around (124), which satisfies $\lim_{\beta\downarrow 1/2}\sqrt{2\beta-1}C_{\beta}<\infty$.
+Proof.
+Set $\alpha_{0}:={E}^{\frac{1+2\beta}{4\beta}}$. We have
+$\int_{\alpha_{0}}^{\infty}\alpha^{-2}{E}\wedge d\alpha$ $\leq{E}^{\frac{2\beta-1}{4\beta}}.$ (108)
+Moreover, by Lemma 10
+$\int_{0}^{\alpha_{0}}\sum_{i=1}^{\infty}f(2\alpha^{-2}\Sigma_{ii})d\alpha$ $=\sum_{i=1}^{\infty}\sqrt{2\Sigma_{ii}}\int_{0}^{\frac{\alpha_{0}}{\sqrt{2\Sigma_{ii}}}}f(\frac{1}{u^{2}})du$ (109)
+$\leq\alpha_{0}\sum_{i=1}^{\infty}4f(\frac{\sqrt{2\Sigma_{ii}}}{\alpha_{0}}).$ (110)
+It will become clear that $\alpha_{0}$ was chosen so that (108) and (110) are both upper bounded by ${E}^{\frac{2\beta-1}{4\beta}}$. Therefore, it remains to consider the optimization problem:
+$M_{1}:=\sup_{(\Sigma_{i})}\sum_{i=1}^{\infty}f(\frac{\sqrt{2\Sigma_{i}}}{\alpha_{0}})$ (111)
+under the constraint $\sum_{i=1}^{\infty}\Sigma_{i}i^{2\beta}\leq 4$. Also, consider another optimization problem:
+$M_{2}:=\sup_{\Sigma(\cdot)}\int_{0}^{\infty}f(\frac{\sqrt{2\Sigma(t)}}{\alpha_{0}})\,dt$ (112)
+under the constraint $\int_{0}^{\infty}\Sigma(t)t^{2\beta}\,dt\leq 4$, where $\Sigma(\cdot)$ is a nonnegative continuous function.
+Clearly $M_{2}\geq M_{1}$, since given $(\Sigma_{ii})$ we can define $\Sigma(t)=\Sigma_{\lceil t\rceil}$.
+Since $f$ has continuous, strictly increasing first derivative, we can see that an optimizer $\Sigma(\cdot)$ exists, which is continuous.
+Since $f(\cdot)$ is increasing, $\Sigma(\cdot)$ must be decreasing.
+There exists unique $t_{0}>0$ such that $\frac{\sqrt{2\Sigma(t)}}{\alpha_{0}}\geq 1$, $t\in(0,t_{0}]$ and $\frac{\sqrt{2\Sigma(t)}}{\alpha_{0}}\leq 1$, $t\in[t_{0},\infty)$.
+Stationarity condition implies that there exists $\lambda>0$ such that
+$\frac{1}{4\Sigma(t)}$ $=\lambda t^{2\beta},\quad t\in(0,t_{0}];$ (113)
+$\frac{1}{2\alpha_{0}\sqrt{2\Sigma(t)}}$ $=\lambda t^{2\beta},\quad t\in[t_{0},\infty).$ (114)
+Setting $t=t_{0}$ in both equations and canceling $\Sigma(t_{0})$, we obtain $\lambda=\frac{1}{2\alpha_{0}^{2}t_{0}^{2\beta}}$.
+Also, we can express $\Sigma(t)$ as a function of $t$ from the above, and $\int_{0}^{\infty}\Sigma(t)t^{2\beta}\,dt=4$ must be achieved, implying
+$4$ $=\int_{0}^{t_{0}}\frac{1}{4\lambda}\,dt+\int_{t_{0}}^{\infty}\frac{1}{8\alpha_{0}^{2}\lambda^{2}t^{2\beta}}\,dt$ (115)
+$=\frac{t_{0}}{4\lambda}+\frac{1}{8(2\beta-1)\alpha_{0}^{2}\lambda^{2}t_{0}^{2\beta-1}}$ (116)
+$=\frac{\beta}{2\beta-1}\alpha_{0}^{2}t_{0}^{2\beta+1}.$ (117)
+Now that we are able to express $t_{0}$ and $\lambda$ as functions of $\alpha$, we can evaluate the optimal value:
+$\int_{0}^{t_{0}}f(\frac{\sqrt{2\Sigma(t)}}{\alpha_{0}})\,dt$ $=\int_{0}^{t_{0}}\frac{1+\ln\frac{\sqrt{2\Sigma(t)}}{\alpha_{0}}}{2}\,dt$ (118)
+$=\int_{0}^{t_{0}}\frac{1+\beta\ln\frac{t_{0}}{t}}{2}\,dt$ (119)
+$=\frac{t_{0}(1+\beta)}{2};$ (120)
+$\int_{t_{0}}^{\infty}f(\frac{\sqrt{2\Sigma(t)}}{\alpha_{0}})\,dt$ $=\int_{t_{0}}^{\infty}\frac{\sqrt{2\Sigma(t)}}{2\alpha_{0}}\,dt$ (121)
+$=\int_{t_{0}}^{\infty}\frac{t_{0}^{2\beta}}{2t^{2\beta}}\,dt$ (122)
+$=\frac{t_{0}}{2(2\beta-1)}.$ (123)
+In summary, we have shown that $M_{1}\leq M_{2}\leq\left(\frac{1+\beta}{2}+\frac{1}{2(2\beta-1)}\right)t_{0}\leq c_{\beta}\alpha_{0}^{-\frac{2}{1+2\beta}}$, where we defined
+$c_{\beta}=\left(\frac{1+\beta}{2}+\frac{1}{2(2\beta-1)}\right)\left(\frac{\beta}{4(2\beta-1)}\right)^{-\frac{1}{1+2\beta}}.$ (124)
+Hence
+$\int_{0}^{\infty}\alpha^{-2}{E}\wedge\sum_{i=1}^{d}f(2\alpha^{-2}\Sigma_{ii})d\alpha\leq 4\alpha_{0}M_{1}+{E}^{\frac{2\beta-1}{4\beta}}=(4c_{\beta}+1){E}^{\frac{2\beta-1}{4\beta}}.$ (125)
+The proof is completed by setting $C_{\beta}=4c_{\beta}+1$. ∎
+Proof of Theorem 5.
+We will apply Proposition 1 with $\mu:=P_{Z}$, and let $\mu^{\prime}$ be an arbitrary distribution supported on TTT-T and satisfying $W_{\infty}(\mu,\mu^{\prime})\leq\epsilon$.
+Then there exists $Z^{\prime}\sim\mu^{\prime}$ such that $\|Z^{\prime}-Z\|_{2}\leq\epsilon$ almost surely, and hence $\mathbb{E}[\|Z^{\prime}\|_{2}^{2}]\leq E+\epsilon^{2}+2\sqrt{E}\epsilon$.
+From Lemma 9 and Lemma 11 we have
+$\int_{0}^{\infty}\inf_{P_{U|Z^{\prime}}}\{\alpha^{-2}\mathbb{E}[\|U-Z^{\prime}\|^{2}]+I(U;Z^{\prime})\}d\alpha\leq C_{\beta}(E+\epsilon^{2}+2\sqrt{E}\epsilon)^{\frac{2\beta-1}{4\beta}}.$ (126)
+Then Proposition 1 and (90) shows that
+$\mathbb{E}[\chi_{Z}]$ $\lesssim C_{\beta}\lim_{\epsilon\to 0}(E+\epsilon^{2}+2\sqrt{E}\epsilon)^{\frac{2\beta-1}{4\beta}}=C_{\beta}E^{\frac{2\beta-1}{4\beta}}.$ (127)
+Proof of Theorem 4.
+Taking the expectation of (18) and applying Theorem 5, we obtain
+$\mathbb{E}[\|\hat{m}-m\|_{2}^{2}]\lesssim\frac{1}{\sqrt{n}}C_{\beta}\mathbb{E}^{\frac{2\beta-1}{4\beta}}[\|\hat{m}-m\|_{2}^{2}]$ (128)
+which implies the claimed result. ∎
+8 Acknowledgment
+The author would like to thank professor Ramon van Handel for feedback on an early version of the manuscript, and for encouraging the pursuit of a new proof of the majorizing measure theorem based on the rate-distortion integral.
+The author also thanks professors Garvesh Raskutti, Philippe Rigollet and Yuting Wei for discussion and feedback on the regression examples.
+This research was supported in part by NSF Grant DMS-2515510.
+Appendix A From matching to coupling
+A.1 Proof of Lemma 3
 The $\geq$ part is obvious from the definition.
 For the $\leq$ part, fix an arbitrary $\bar{s}^{N}\in\mathbb{R}^{nN}$ satisfying $\widehat{P}_{\bar{s}^{N}}=\mu$.
-Denote the supports of $\widehat{P}_{y^{N}}$ and $\mu$ be $\mathcal{A}$ and $\mathcal{B}$, respectively.
+Denote the supports of $\widehat{P}_{y^{N}}$ and $\mu$ by $\mathcal{A}$ and $\mathcal{B}$, respectively.
 Consider the following function
@@
 We can see from (130) that $\inf_{M\in\mathcal{S}}F(M)\leq W_{2}^{2}(\widehat{P}_{y^{N}},\mu)$.
-Since $F$ is linear, by the Birkhoff–von Neumann theorem, there exists a permutation matrix $M$ achieving $\inf_{M\in\mathcal{S}}F(M)$.
+Since $F$ is linear, by the Birkhoff-von Neumann theorem, there exists a permutation matrix $M$ achieving $\inf_{M\in\mathcal{S}}F(M)$.
 This permutation sends $\bar{s}^{N}$ to an $s^{N}$, achieving $\inf_{M\in\mathcal{S}}F(M)=\frac{1}{N}\|y^{N}-s^{N}\|_{2}$.
 This proves the $\leq$ part.
+A.2 Proof of Lemma 4
+The proof follows immediately from the following observation, by taking $A:=[NP_{XY}(x,y)]_{x\in\mathcal{X},y\in\mathcal{Y}}$ and $B:=[NQ_{XY}(x,y)]_{x\in\mathcal{X},y\in\mathcal{Y}}$.
+Lemma 12.
+Let $A\in\mathbb{R}_{\geq 0}^{m\times n}$ have integer row sums and integer column sums. Then there exists $B\in\mathbb{Z}_{\geq 0}^{m\times n}$ with the same row and column sums as $A$ and
+$\max_{i,j}\,|B_{ij}-A_{ij}|\;\leq\;1.$
+Proof.
+Write $F:=A-\lfloor A\rfloor$, so $F_{ij}\in[0,1)$ for all $i,j$.
+Let $r_{i}:=\sum_{j}F_{ij}$ and $c_{j}:=\sum_{i}F_{ij}$.
+Because the row and column sums of $A$ are integers and $\sum_{j}\lfloor A_{ij}\rfloor,\sum_{i}\lfloor A_{ij}\rfloor$ are integers, we have $r_{i},c_{j}\in\mathbb{Z}$.
+Form the bipartite graph $G$ with left vertices $\{1,\dots,m\}$ (rows) and right vertices $\{1,\dots,n\}$ (columns), and put an edge $(i,j)$ iff $F_{ij}\in(0,1)$.
+If some connected component of $G$ were a tree, it would have a leaf $v$ incident to a unique edge $e$ of weight $F_{e}\in(0,1)$, forcing the corresponding $r_{i}$ or $c_{j}$ to equal $F_{e}\notin\mathbb{Z}$, a contradiction.
+Hence every component of $G$ contains a cycle.
+Fix a simple cycle $C$ in $G$. Along the edges of $C$, alternately add $\varepsilon$ and $\varepsilon$ to the corresponding entries of $F$ (leaving all other entries unchanged), where
+$\varepsilon\;:=\;\min\{\,F_{ij},\,1-F_{ij}:(i,j)\in C\,\}.$
+This preserves all row and column sums, since at each vertex on $C$ the two incident edges change by $\varepsilon$ and $\varepsilon$.
+Moreover, all modified entries remain in $[0,1]$, and at least one becomes 0 or $1$.
+Remove edges for which $F_{ij}$ becomes 0 or 1.
+Iterating this operation, which terminates in finitely many steps, we obtain a modified matrix $B$ with the desired properties.
 ∎
-Proof of Theorem 2.
-As in the proof of Theorem 1, we can identify $T=\{e_{1},\dots,e_{n}\}$ as the standard basis vectors, so that $X_{z}=\left<Y,z\right$>.
-The assumption of zero-mean process implies that $\mathbb{E}[Y]=0$.
-Then, shifting $Z\sim\mu$ by a deterministic vector does not change the value of $\mathbb{E}[\left<Y,Z\right$>], so we can assume without loss of generality that $\mathbb{E}[Z]=0$.
-Let us assume that $P_{YZ}$ is rational; the general case will then following by approximation.
-Define
-$\mathcal{I}:=\{N\geq 1\colon\textrm{$NP_{YZ}$ is integer}\}.$ (65)
-For each $N\in\mathcal{I}$, let $\mathcal{C}_{\mu}$ denote the $\mu$-type set, consisting of all sequences of type $\mu=P_{Z}$. Select
-$L:=\exp(NI(Y;Z)+N^{1/3})$ (66)
-sequences from $\mathcal{C}_{\mu}$ uniformly at random with replacement, and call this random set $\mathcal{A}$. Let $Y^{N}\sim P_{Y}^{\otimes N}$. Then we will show that
-$\mathbb{E}[\left<Y,Z\right>]\leq\liminf_{N\to\infty}\mathbb{E}[\sup_{z^{N}\in\mathcal{A}}\frac{1}{N}\left<Y^{N},z^{N}\right$>], (67)
-where the limit is for $N\in\mathcal{I}$. Indeed, set
-$\mathcal{B}$ $:=\{(y^{N})\colon\widehat{P}_{y^{N}}=P_{Y}\};$ (68)
-$\mathcal{D}$ $:=\{(y^{N},z^{N})\colon\widehat{P}_{y^{N}z^{N}}=P_{YZ}\},$ (69)
-and for any $y^{N}\in\mathcal{B}$, define $\mathcal{D}(y^{N}):=\{z^{N}\colon(y^{N},z^{N})\in\mathcal{D}\}$. A standard result states that (see Lemma 3, or (Csiszár and Körner,, 2011, Chapter 1)):
-$\frac{1}{N}\ln|\mathcal{D}(y^{N})|=H(Z|Y)+O(\frac{\ln N}{N}),\quad\forall y^{N}\in\mathcal{B}.$ (70)
-Let $\pi$ be the projection from $\mathbb{R}^{nN}$ to $\mathcal{B}$ (i.e., map to the closest point in $\mathcal{B}$, with ties broken arbitrarily).
-Define $\mathcal{E}$ as the event that $\mathcal{D}(\pi(Y^{N}))\cap\mathcal{A}=\emptyset$.
-Then, (70) implies that
-$\mathbb{P}[\mathcal{E}]=\left(1-\exp(-NI(Y;Z)+O(\ln N))\right)^{L}\leq\exp(-CN^{1/3})$ (71)
-for some $C>0$ depending only on $P_{YZ}$, and for all $N\in\mathcal{I}$. From the definitions it is clear that
-$\max_{z^{N}\in\mathcal{A}}\frac{1}{N}\left<\pi(Y^{N}),z^{N}\right>1_{\mathcal{E}^{c}}\geq\mathbb{E}[\left<Y,Z\right>]1_{\mathcal{E}^{c}}$ (72)
-for all $N\in\mathcal{I}$ almost surely. Since $\mathbb{E}[Z]=0$ is assumed, we see that 0 belongs to the convex hull of $\mathcal{C}_{\mu}$, therefore
-$\max_{z^{N}\in\mathcal{A}}\frac{1}{N}\left<\pi(Y^{N}),z^{N}\right>\geq\max_{z^{N}\in\mathcal{A}}\frac{1}{N}\left<\pi(Y^{N}),z^{N}\right>1_{\mathcal{E}^{c}}$ (73)
-almost surely. Furthermore,
-$\max_{z^{N}\in\mathcal{A}}\frac{1}{N}\left<Y^{N},z^{N}\right>\geq\max_{z^{N}\in\mathcal{A}}\frac{1}{N}\left<\pi(Y^{N}),z^{N}\right>-\max_{z^{N}\in\mathcal{A}}\frac{1}{N}\left<\pi(Y^{N})-Y^{N},z^{N}\right$> (74)
-From Lemma 5, we have
-$\max_{z^{N}\in\mathcal{A}}\frac{1}{N}\left<\pi(Y^{N})-Y^{N},z^{N}\right$> $\leq\frac{1}{\sqrt{N}}\|\pi(Y^{N})-Y^{N}\|_{2}\cdot\frac{1}{\sqrt{N}}\max_{z^{N}\in\mathcal{A}}\|z^{N}\|_{2}$ (75)
-$\leq W_{2}(\mu,\widehat{P}_{Y^{N}})\cdot O(1)$ (76)
-which converges to 0 in expectation as $N\in\mathcal{I}$ tends to infinity. Thus (67) is verified by collecting the bounds above.
-It remains to apply the Dudley inequality to the right side of (67).
-Since $\mathcal{A}$ is contained in the set $\mathcal{C}$ in Lemma 4, the upper bound part of Lemma 4 also applies to covering numbers of $\mathcal{A}$.
-Furthermore, we have the trivial bound $|\mathcal{A}|\leq L$.
-Therefore,
-$\limsup_{N\to\infty}\frac{1}{N}\ln N(\mathcal{A},2\sqrt{N}\sigma)\leq R_{\mu}(\sigma^{2})\wedge I(Y;Z).$ (77)
-The claim then follows by applying dominated convergence (see proof of Theorem 1). ∎
-4.2 Beyond the Finite Case
-In this section we extend the proof of Theorem 1 for finite $T$ to the countable case by a limiting argument, and also prove Proposition 1.
+Appendix B Extension of the rate-distortion integral beyond finite $T$
+In the main text, we proved the rate-distortion integral for finite $T$. In this section, we extend the proof beyond the finite case by a limiting argument. We also prove Proposition 1.
+B.1 Absolute integrability of $|X_{Z}|$
 We first prove an auxiliary result about absolute integrability of $|X_{Z}|$, which will be used later to apply dominated convergence.
@@
   • If $\int_{0}^{\sigma_{\rm m}}\sqrt{R_{\mu}(\sigma^{2})}\,d\sigma<\infty$, then $\mathbb{E}[|X_{Z}|]<\infty$.
-It is interesting to compare Lemma 6 with the classical observation that for symmetric processes, $\mathbb{E}[\sup_{t\in T}|X_{t}|]<\infty$ is equivalent to $\mathbb{E}[\sup_{t\in T}X_{t}]<\infty$, both equivalent to boundedness of the process in the Gaussian case; see (Talagrand,, 1987, p2)(Talagrand,, 2014, Exercise 2.2.2.) and the references therein.
+Lemma 13 may be compared with the classical observation that for symmetric processes, $\mathbb{E}[\sup_{t\in T}|X_{t}|]<\infty$ is equivalent to $\mathbb{E}[\sup_{t\in T}X_{t}]<\infty$, both equivalent to boundedness of the process in the Gaussian case; see [37, p2], [39, Exercise 2.2.2.] and the references therein.
 Proof.
 We first prove (131).
-Clearly, the natural metric for the new process satisfies $\bar{d}((t,s),(t^{\prime},s))=d(t,t^{\prime})$, for any $s\in\{-1,+1\}$ and $t,t^{\prime}\in T$.
+The natural metric for the new process satisfies $\bar{d}((t,s),(t^{\prime},s))=d(t,t^{\prime})$, for any $s\in\{-1,+1\}$ and $t,t^{\prime}\in T$.
 Now suppose that $P_{\bar{Z}}=P_{Z,S}$ is an arbitrary distribution on $\bar{T}$.
@@
 $\mathbb{E}[|X_{Z_{k}}|]$ $=\mathbb{E}[{\rm sign}(X_{Z_{k}})X_{Z_{k}}]$ (140)
-$\leq 48\int_{0}^{\bar{\sigma}_{\rm m}}\sqrt{\bar{R}_{P_{\bar{Z}_{k}}}(\sigma^{2})}d\sigma$ (88)
-$\leq 48\int_{0}^{\sigma_{\rm m}}\sqrt{R_{P_{Z_{k}}}(\sigma^{2})}d\sigma+\sqrt{2\ln 2}\sigma_{\rm m},$ (89)
+$\lesssim\int_{0}^{\bar{\sigma}_{\rm m}}\sqrt{\bar{R}_{P_{\bar{Z}_{k}}}(\sigma^{2})}\,d\sigma$ (141)
+$\lesssim\int_{0}^{\sigma_{\rm m}}\sqrt{R_{P_{Z_{k}}}(\sigma^{2})}\,d\sigma+\sqrt{2\ln 2}\sigma_{\rm m},$ (142)
 where we defined $\bar{Z}_{k}:=(Z_{k},{\rm sign}(X_{Z_{k}}))$, and (142) follows from (131) by replacing the arbitrary $P_{Z}$ with $P_{Z_{k}}$ and noting that $\sigma_{\rm m}(P_{Z_{k}})\leq\sigma_{\rm m}(P_{Z})=:\sigma_{\rm m}$.
@@
 ∎
-Proof of Theorem 1.
-At this point, it only remains to extend the result of Theorem 1 from the finite case to the countable case.
-We can assume that the integral on the right side of (33) is finite, since otherwise there is nothing to prove.
+B.2 Extension of Theorem 1 beyond finite $T$
+We extend the result of Theorem 1 from the finite case to the countable case.
+We can assume that $\int_{0}^{\sigma_{\rm m}}\sqrt{R_{\mu}(\sigma^{2})}\,d\sigma<\infty$, since otherwise there is nothing to prove.
 Again, assume without loss of generality that there is $t_{0}\in T$ satisfying $X_{t_{0}}\equiv 0$ and $\sigma_{\rm m}:=\sigma_{\rm m}(P_{Z})=\sqrt{\mathbb{E}[d^{2}(t_{0},Z)]}$.
-Construct $Z_{k}$ as in the proof of Lemma 6.
+Construct $Z_{k}$ and define $E$ as in the proof of Lemma 13.
 Lemma 13 implies $\mathbb{E}[|X_{Z}|]<\infty$, and since $|X_{Z_{k}}|\leq|X_{Z}|\vee|X_{t_{0}}|$, dominated convergence implies
@@
 $\lim_{k\to\infty}\int_{0}^{+\infty}\sqrt{R_{P_{Z_{k}}}(\sigma^{2})}\,d\sigma=\int_{0}^{\sigma_{\rm m}}\sqrt{R_{P_{Z}}(\sigma^{2})}\,d\sigma.$ (144)
-Then the claim of the lemma follows by (90), (91) and applying the result of Theorem 1 in the finite setting (proved in Section 4.1).
-Proof of Proposition 1.
+Then the claim of the lemma follows by (143), (144) and applying the result of Theorem 1 in the finite case (already proved in the main text).
+B.3 Proof of Proposition 1
 Let $T_{0}$ be the countable subset in the definition of separability.
@@
 Therefore the claim follows by the result for the countable case and taking $\epsilon\downarrow 0$.
-∎
-5 Sharpness of the Rate-Distortion Integral
-We provide several proofs of Theorem 3 (up to constant factors) in this section. The proof in Section 5.3 gives the constant stated in Theorem 3.
-5.1 Proof via the Majorizing Measure Theorem
-We first give a proof of the lower bound direction assuming a version of the majorizing measure theorem, i.e.
-the $\delta_{2}$ functional defined in (15) lower bounds the supremum of a Gaussian process.
-It is relatively easy (compared to the proof of the majorizing measure theorem) to show the equivalence of $\delta_{2}$ and $\gamma_{2}$ up to universal constants, and the direction of $\gamma_{2}\gtrsim\delta_{2}$ can be found in (Talagrand,, 2014, Proposition 3.3.1 & Proposition 3.4.1).
-We first consider the case where $n:=|T|<\infty$.
-Recall the representation of $X_{z}$ as an inner product in (55), which embeds $Y,Z$ in $\mathbb{R}^{n}$, where $Y$ is a normal vector, and recall the typical set $\mathcal{C}$ defined in (54).
-Define
-$\mathcal{D}:=\{y^{N}\in\mathbb{R}^{nN}\colon W_{2}(P_{Y},\widehat{P}_{y^{N}})\leq N^{-\frac{1}{2n}}\}$ (92)
-Then by the Wasserstein law of large numbers, $\mathbb{P}[Y^{N}\in\mathcal{D}]=1-o(1)$.
-Given $P_{Z}=\mu$, let $P_{YZ}$ be the optimal coupling of $P_{Y}$ and $\mu$ maximizing $\mathbb{E}[\left<Y,Z\right$>] (equivalently, the optimal transport map). Then
-$\mathbb{E}[X_{Z}]$ $\geq\lim_{N\to\infty}\frac{1}{N}\mathbb{E}[\sup_{z^{N}\in\mathcal{C}}\left<Y^{N},z^{N}\right>1_{Y^{N}\in\mathcal{D}}]$ (93)
-$=\lim_{N\to\infty}\frac{1}{N}\mathbb{E}[\sup_{z^{N}\in\mathcal{C}}\left<Y^{N},z^{N}\right$>] (94)
-$\gtrsim\lim_{N\to\infty}\frac{1}{N}\delta_{2}(\mathcal{C})$ (95)
-$\geq\lim_{N\to\infty}\frac{1}{N}\inf_{t\in\mathcal{C}}I_{\mu^{\otimes N}}(t).$ (96)
-The above steps are justified below: to show (93), note that for any $y^{N}\in\mathcal{D}$, we have
-$\lim_{N\to\infty}\frac{1}{N}\sup_{z^{N}\in\mathcal{C}}\left<y^{N},z^{N}\right$> $\leq\lim_{\epsilon\to 0}\sup_{P_{Y^{\prime}Z^{\prime}}}\mathbb{E}[\left<Y^{\prime},Z^{\prime}\right$>] (97)
-$=\lim_{\epsilon\to 0}\sup_{P_{Y^{\prime}Z^{\prime}}}\frac{\mathbb{E}\|Y^{\prime}\|_{2}^{2}+\mathbb{E}\|Z^{\prime}\|_{2}^{2}-W_{2}^{2}(P_{Y^{\prime}},P_{Z^{\prime}})}{2}$ (98)
-$=\frac{\mathbb{E}\|Y\|_{2}^{2}+\mathbb{E}\|Z\|_{2}^{2}-W_{2}^{2}(P_{Y},P_{Z})}{2}$ (99)
-$=\mathbb{E}[\left<Y,Z\right$>] (100)
-where in (97) the sup is over $P_{Y^{\prime}Z^{\prime}}$ for which $W_{2}(P_{Y},P_{Y^{\prime}})\leq\epsilon$ and $W_{2}(P_{Z},P_{Z^{\prime}})\leq\epsilon$.
-This shows that the right side of (93) is upper bounded by $\mathbb{E}[\left<Y,Z\right>]\cdot\lim_{N\to\infty}\mathbb{E}[1\{Y^{N}\in\mathcal{D}\}]=\mathbb{E}[\left<Y,Z\right$>].
-The step (94) follows from $\mathbb{E}[\sup_{z^{N}\in\mathcal{C}}\left<Y^{N},z^{N}\right>1_{Y^{N}\in\mathcal{D}^{c}}]=o(N)$, whose proof is the same as the proof that the second expectation in (58) is $o(N)$.
-(95) follows since $\delta_{2}(\mathcal{C})\asymp\gamma_{2}(\mathcal{C})$ lower bounds (94) according to the majorizing measure theorem (Talagrand,, 2014).
-We see (96) by specifying the supremum in (15) as $\mu^{\otimes N}$.
-Using the estimate in Lemma 4, we obtain that for any $t\in\mathcal{C}$,
-$\liminf_{N\to\infty}\frac{1}{N}I_{\mu^{\otimes N}}(t)\gtrsim\int_{0}^{\sigma_{\rm m}}\sqrt{R_{\mu}(\sigma^{2})}d\sigma$ (101)
-by using Fatou’s lemma. This proves Theorem 3 for finite $T$ (except for estimating the constant factor).
+B.4 Extension of Theorem 3 beyond finite $T$
 We now use a limiting argument to extend the result to the case of $\mu$ supported on a countable $T$.
@@
 Since $\lim_{k\to\infty}\mathbb{P}[T_{k}^{c}]=0$ and $\mathbb{E}[d^{2}(t_{0},Z)]<\infty$, a basic result from measure theory (which can be proved by dominated convergence) states that $\lim_{k\to\infty}\mathbb{E}[d^{2}(t_{0},Z)1_{Z\in T_{k}^{c}}]=0$.
-In particular, for any given $\sigma>0$, there exists $k(\sigma)>0$ such that for all $k>k(\sigma)$, this construction of $P_{U|Z}$ ensures that $\mathbb{E}[d^{2}(U,Z)]<2\sigma^{2}$.
+In particular, for any given $\sigma>0$, there exists $k(\sigma)>0$ such that for all $k>k(\sigma)$, this construction of $P_{U|Z}$ ensures that $\mathbb{E}[d^{2}(U,Z)]<(1+a)\sigma^{2}$, where $a>0$ is an arbitrary constant.
 Moreover, if $E_{k}:=1\{Z\in T_{k}\}$, we have $I(U;Z)\leq I(U;Z|E)+H(E)=R_{\mu_{k}}(\sigma^{2})+h(\mu(T_{k}))$, where $h(\cdot)$ is the binary entropy function.
 Taking $k\to\infty$ we have $\mu(T_{k})\to 1$ and obtain
-$\lim_{k\to\infty}R_{\mu_{k}}(\sigma^{2})\geq R_{\mu}(2\sigma^{2})$ (103)
+$\lim_{k\to\infty}R_{\mu_{k}}(\sigma^{2})\geq R_{\mu}((1+a)\sigma^{2})$ (146)
 for any $\sigma>0$. Then using (145) we obtain
-$\mathbb{E}[X_{Z}]\geq\lim_{k\to\infty}\mathbb{E}[X_{Z_{k}}]\gtrsim\lim_{k\to\infty}\int_{0}^{\infty}\sqrt{R_{\mu_{k}}(\sigma^{2})}d\sigma\gtrsim\int_{0}^{\infty}\sqrt{R_{\mu}(\epsilon^{2})}d\epsilon$ (104)
-by Fatou’s lemma and change of variables.
-5.2 Proof via Sharpness of Dudley’s Integral
-The second proof does not use the majorizing measure theorem, but merely the sharpness of the Dudley’s Integral for stationary processes (10), which was due to Fernique (Fernique,, 1975) a decade before the majorizing measure theorem.
-Recall (94) that $\mathbb{E}[X_{Z}]\geq\lim_{N\to\infty}\frac{1}{N}\mathbb{E}[\sup_{z^{N}\in\mathcal{C}}\left<Y^{N},z^{N}\right$>], which is further lower bounded (up to a constant) by $\lim_{N\to\infty}\frac{1}{N}\int_{0}^{\infty}\sqrt{\ln N(\mathcal{C},\sqrt{N}\sigma)}\sqrt{N}d\sigma$.
-In turn, it is further lower bounded by $\int_{0}^{\infty}\sqrt{R_{\mu}(\epsilon^{2})}d\epsilon$, by invoking Lemma 4 and Fatou’s Lemma.
-5.3 Proof via Super-Sudakov
-In this section, we give another proof of Theorem 3 which will yield the constant stated therein.
-Traditionally, the majorizing measure theorem was proved by a sophisticated generic chaining argument (Talagrand,, 1987, 2014), and in the special case of stationary processes by a simpler reverse chaining argument based on a “super Sudakov” theorem (Van Handel,, 2014; Talagrand,, 1992).
-Here, we derive an information-theoretic counterpart of the super-Sudakov theorem, and then obtains Theorem 3 using purely analytic arguments, without any combinatorial constructions like packing.
-Suppose that $\mu$, $\gamma$ are distributions on $\mathbb{R}^{n}$, and $\epsilon>0$. Define
-$\phi(\epsilon,\mu,\gamma):=\sup_{P_{ZZ^{*}G}}\mathbb{E}[\left<G,Z\right$>] (105)
-where the supremum is over $P_{ZZ^{*}G}$ satisfying $P_{GZ^{*}}=\gamma\times\mu$, $P_{Z}=\mu$, and $\mathbb{E}[\|Z-Z^{*}\|_{2}^{2}]\leq\epsilon^{2}$.
-The parameter $\epsilon$ can be thought of as an analogue of the scale in the reverse chaining argument.
-The geometric meaning of $\phi(\epsilon,\mu,\gamma)$ is shown by the following:
-Lemma 7.
-Consider $\phi$ defined in (105), where $\mu$ is rational, and $\gamma$ has finite second moment. For each $N$ such that $N\mu$ is integer, define
-$T$ $:=\{z^{N}\in\mathbb{R}^{nN}\colon\widehat{P}_{z^{N}}=\mu\};$ (106)
-$\mathcal{C}$ $:=\{z^{N}\in T\colon\|z^{N}-z^{*N}\|_{2}\leq\epsilon\sqrt{N}\}.$ (107)
-Let $z^{*N}=(z_{1}^{*},\dots,z_{N}^{*})$ be an arbitrary deterministic sequence in $T$. Let $G^{N}\sim\gamma^{\otimes n}$. Then for any $\epsilon>0$, we have
-$\lim_{N\to\infty}\frac{1}{N}\mathbb{E}[\sup_{z^{N}\in\mathcal{C}}\left<G^{N},z^{N}\right>]=\phi(\epsilon,\mu,\gamma)$ (108)
-where the limit is along the subsequence for which $N\mu$ is integer.
-The proof involves constructing coupling distributions and approximation by types, which uses the following standard result about the transportation polytope:
-Lemma 8.
-Let $A\in\mathbb{R}_{\geq 0}^{m\times n}$ have integer row sums and integer column sums. Then there exists $B\in\mathbb{Z}_{\geq 0}^{m\times n}$ with the same row and column sums as $A$ and
-$\max_{i,j}\,|B_{ij}-A_{ij}|\;\leq\;1.$
-Proof.
-Write $F:=A-\lfloor A\rfloor$, so $F_{ij}\in[0,1)$ for all $i,j$.
-Let $r_{i}:=\sum_{j}F_{ij}$ and $c_{j}:=\sum_{i}F_{ij}$.
-Because the row and column sums of $A$ are integers and $\sum_{j}\lfloor A_{ij}\rfloor,\sum_{i}\lfloor A_{ij}\rfloor$ are integers, we have $r_{i},c_{j}\in\mathbb{Z}$.
-Form the bipartite graph $G$ with left vertices $\{1,\dots,m\}$ (rows) and right vertices $\{1,\dots,n\}$ (columns), and put an edge $(i,j)$ iff $F_{ij}\in(0,1)$.
-If some connected component of $G$ were a tree, it would have a leaf $v$ incident to a unique edge $e$ of weight $F_{e}\in(0,1)$, forcing the corresponding $r_{i}$ or $c_{j}$ to equal $F_{e}\notin\mathbb{Z}$, a contradiction.
-Hence every component of $G$ contains a cycle.
-Fix a simple cycle $C$ in $G$. Along the edges of $C$, alternately add $\varepsilon$ and $\varepsilon$ to the corresponding entries of $F$ (leaving all other entries unchanged), where
-$\varepsilon\;:=\;\min\{\,F_{ij},\,1-F_{ij}:(i,j)\in C\,\}.$
-This preserves all row and column sums, since at each vertex on $C$ the two incident edges change by $\varepsilon$ and $\varepsilon$.
-Moreover, all modified entries remain in $[0,1]$, and at least one becomes 0 or $1$.
-Remove edges for which $F_{ij}$ becomes 0 or 1.
-Iterating this operation, which terminates in finitely many steps, we obtain a modified matrix $B$ with the desired properties.
-∎
-Proof Lemma 7.
-Define $\mathcal{I}:=\{N\in\mathbb{N}\colon N\mu\textrm{ is integer}\}$, where $\mathbb{N}$ denotes the set of natural numbers.
-Since the integers $\mathbb{Z}$ form a principal ideal domain, there exists a natural number $n_{0}$ such that $\mathcal{I}=n_{0}\mathbb{N}$.
-Since $\gamma$ has finite second moment, for each $N=n_{0}k$, $k\in\mathbb{N}$, we can find a distribution $\bar{\gamma}_{N}$ on $\mathbb{R}^{n}$ satisfying the following properties: 1) $k\bar{\gamma}_{N}$ is integer, 2) $W_{2}(\gamma,\bar{\gamma}_{N})=o(1)$, 3) ${\rm diam}(\operatorname*{supp}(\bar{\gamma}_{N}))\leq N^{a}$, where $a:=\frac{1}{4+4n}$, 4) Cardinality of the support $|\operatorname*{supp}(\bar{\gamma}_{N})|\leq N^{2an}$.
-Note that property 1) implies that $N\bar{\gamma}_{N}\times\mu$ is integer, The purpose of $\bar{\gamma}_{N}$ is to approximate $\gamma$ which might not be rational.
-Define
-$\mathcal{D}:=\{g^{N}\colon\widehat{P}_{g^{N}z^{*N}}=\bar{\gamma}_{N}\times\mu\},$ (109)
-which is nonempty.
-Let $\pi$ be the projection from $\mathbb{R}^{nN}$ to $\mathcal{D}$.
-More precisely, it sends any point to a closest point in $\mathcal{D}$ under the $\ell_{2}$ distance, and ties are broken so that $\pi$ is invariant under all $z^{*N}$-invariant permutations.
-This makes $\pi(G^{N})$ uniformly distributed on $\mathcal{D}$, and
-[…]
-$\leq\mathbb{E}\left[\|\pi(G^{N})-G^{N}\|_{2}\sup_{z^{N}\in\mathcal{C}}\|z^{N}\|_{2}\right]$ (111)
-$=o(N),$ (112)
-where the last step follows from the fact that $N^{-1/2}\|\pi(G^{N})-G^{N}\|_{2}\leq W_{2}(\bar{\gamma}_{N},\gamma)+W_{2}(\gamma,\widehat{P}_{G^{N}})$, due to Lemma 5, which vanishes in expected value as $N\to\infty$.
-Moreover, for any given $G^{N}$ and $z^{N}\in\mathcal{C}$, we have $\frac{1}{N}\left<\pi(G^{N}),z^{N}\right>]\leq\phi(\epsilon,\mu,\bar{\gamma}_{N})$ because $\widehat{P}_{z^{N}z^{*N}\pi(G^{N})}$ satisfies the constraints on $P_{ZZ^{*}G}$ in the definition of $\phi(\epsilon,\mu,\bar{\gamma}_{N})$ (cf.
-(105)).
-By property 2) of $\bar{\gamma}_{N}$, we also see that
-$\lim_{N\to\infty}\phi(\epsilon,\mu,\bar{\gamma}_{N})=\phi(\epsilon,\mu,\gamma).$ (113)
-Then the $\leq$ part of (108) follows.
-To show that $\geq$ part, suppose that $Q^{N}_{ZZ^{*}G}$ is an optimizer for the supremum in the definition of $\phi(\epsilon,\mu,\bar{\gamma}_{N})$ for some $N\in\mathcal{I}$.
-Note that $NQ^{N}_{Z^{*}G}$ and $NQ^{N}_{Z}$ are integer by the construction of $\bar{\gamma}$.
-By Lemma 8, we can find a probability measure $P^{N}_{ZZ^{*}G}$ such that i) $NP^{N}_{ZZ^{*}G}$ is integer, ii) $P^{N}_{Z^{*}G}=Q^{N}_{Z^{*}G}$ and $P^{N}_{Z}=Q^{N}_{Z}$, iii) $\max_{z,z^{*},g}|P^{N}_{ZZ^{*}G}(z,z^{*},g)-Q^{N}_{ZZ^{*}G}(z,z^{*},g)|\leq\frac{1}{N}$.
-Then the total variation distance $\|P^{N}_{ZZ^{*}G}-Q^{N}_{ZZ^{*}G}\|_{1}\lesssim\frac{1}{N}|\operatorname*{supp}(Q^{N}_{ZZ^{*}G})|\lesssim N^{2an-1}$.
-Hence there exists $P_{ZZ^{*}G\bar{Z}\bar{Z}^{*}\bar{G}}\in\Pi(P^{N}_{ZZ^{*}G},Q^{N}_{ZZ^{*}G})$ under which $\mathbb{P}[(Z,Z^{*},G)\neq(\bar{Z},\bar{Z}^{*},\bar{G})]\lesssim N^{2an-1}$, and
-$\mathbb{E}[\|(Z,Z^{*},G)-(\bar{Z},\bar{Z}^{*},\bar{G})\|_{2}^{2}]$ $\leq 2{\rm diam}^{2}(\operatorname*{supp}(Q^{N}_{ZZ^{*}G}))N^{2an-1}$ (114)
-$\lesssim N^{2a+2an-1}$ (115)
-$=o(1).$ (116)
-In particular, we have
-$W_{2}(P^{N}_{ZZ^{*}G},Q^{N}_{ZZ^{*}G})=o(1).$ (117)
-This implies that $\mathbb{E}_{P^{N}_{ZG}}[\left<G,Z\right>]=\phi(\epsilon,\mu,\bar{\gamma}_{N})+o(1)$ and $\mathbb{E}_{P^{N}_{ZZ^{*}}}[\|Z-Z^{*}\|_{2}^{2}]\leq\epsilon_{N}^{2}$ for some $\epsilon_{N}=\epsilon+o(1)$.
-Let $\mathcal{C}(\epsilon_{N}):=\{z^{N}\in T\colon\|z^{N}-z^{*N}\|_{2}\leq\epsilon_{N}\sqrt{N}\}$.
-For any given $G^{N}$, we can find $z^{N}\in\mathcal{C}(\epsilon_{N})$ such that $\frac{1}{N}\left<\pi(G^{N}),z^{N}\right>=\mathbb{E}_{P^{N}_{ZG}}[\left<G,Z\right$>].
-Thus,
-$\liminf_{N\to\infty}\frac{1}{N}\mathbb{E}[\max_{z^{N}\in\mathcal{C}(\epsilon_{N})}\left<\pi(G^{N}),z^{N}\right$>] $\geq\liminf_{N\to\infty}\mathbb{E}_{P^{N}_{ZG}}[\left<G,Z\right$>] (118)
-$\geq\liminf_{N\to\infty}\phi(\epsilon,\mu,\bar{\gamma}_{N})$ (119)
-$=\phi(\epsilon,\mu,\gamma)$ (120)
-where the limits are along $N\in\mathcal{I}$.
-Since any sequence in $\mathcal{C}(\epsilon_{N})$ can be approximated by a sequence in $\mathcal{C}$ within $o(\sqrt{N})$ error under the Euclidean distance, we see that $\liminf_{N\to\infty}\frac{1}{N}\mathbb{E}[\max_{z^{N}\in\mathcal{C}}\left<\pi(G^{N}),z^{N}\right>]\geq\phi(\epsilon,\mu,\gamma)$.
-Then the $\geq$ part of (108) follows from (112).
-∎
-In the remainder of the section we consider $\phi$ defined in (105), where $\mu$ is finitely supported and $\gamma=\mathcal{N}(0,I_{n})$ (the extension to non-finite case can be achieved by the argument in Section 5.1).
-We write $\phi(\epsilon)=\phi(\epsilon,\mu,\gamma)$ when no confusion.
-We also define the shorthand notation
-[…]
-where we recall that $\Pi_{\epsilon}(\mu,\mu)$ denotes the set of couplings satisfying $\mathbb{E}[\|Z-\hat{Z}\|_{2}^{2}\leq\epsilon^{2}$.
-Lemma 9.
-For any $0<\epsilon_{2}\leq\frac{4\epsilon_{1}}{5}$, we have
-$\phi(\epsilon_{1})\geq\frac{\epsilon_{2}}{4}\sqrt{2(i(\epsilon_{2})-i(\epsilon_{1}-\epsilon_{2}/4))}+\phi(\epsilon_{2}/4).$ (122)
-The super-Sudakov inequality, originally proved by Talagrand (Talagrand,, 1992), can be established for rather general settings by combining Sudakov minoration and concentration.
-The following version for Gaussian processes with explicit constant can be found in Marcus and Rosen, (2006); Van Handel, (2014).
-Lemma 10 (Super-Sudakov).
-Let $\{X_{t}\}_{t\in T}$ be a Gaussian process on a metric space $(T,d)$, and $\mathcal{N}$ be an $\epsilon$-packing of $(T,d)$.
-For every $s\in N$, let $T_{s}:=\{t\in T\colon d(t,s)\leq\frac{\epsilon}{4}\}$.
-Then
-$\mathbb{E}[\sup_{t\in T}X_{t}]\geq\frac{\epsilon}{4}\sqrt{2\ln|\mathcal{N}|}+\min_{s\in\mathcal{N}}\mathbb{E}[\sup_{t\in T_{s}}X_{t}].$ (123)
-Proof of Lemma 9.
-It suffices to consider the case of rational $\mu$, as the general case will then follow by approximation by ration $\mu$ in the Wasserstein distribution.
-Suppose that $N>0$ is an integer for which $N\mu$ is integer.
-Define $T$ as in (106), and fix arbitrary $z^{*N}\in T$.
-For any $\epsilon>0$, define
-$T_{\epsilon}:=T\cap B(z^{*N},\sqrt{N}\epsilon).$ (124)
-Now let $\mathcal{N}$ be a maximal $\epsilon_{2}$-packing of $T_{r}$, where $r:=\epsilon_{1}-\epsilon_{2}/4$.
-Maximality implies that every point in $T_{r}$ is within $\epsilon_{2}$ distance to a point in $\mathcal{N}$, hence
-$|\mathcal{N}|\mu^{\otimes N}(B(\sqrt{N}\epsilon_{2}))\geq\mu^{\otimes N}(B(\sqrt{N}r)).$ (125)
-By a standard type counting argument (Csiszár and Körner,, 2011, Chapter 1), this implies
-$\frac{1}{N}\ln|\mathcal{N}|\geq i(\epsilon_{2})-i(r)+o(1),$ (126)
-and $i(\epsilon_{2})-i(r)\geq 0$ by monotonicity of $i(\cdot)$.
-Moreover, for every $z^{N}\in\mathcal{N}$, we have $T\cap B(z^{N},\sqrt{N}\epsilon_{2}/4)\subseteq T_{\epsilon_{1}}$ since $r+\epsilon_{2}/4\leq\epsilon_{1}$.
-Then Lemma 10 implies (122) by taking $N\to\infty$.
-∎
-Lemma 11.
-Let $\tau\in(0,\frac{4}{5})$. We have
-$\sup_{P_{ZG}\in\Pi(\mu,\gamma)}\mathbb{E}[\left<G,Z\right>]\geq\frac{\sqrt{2}}{4\ln\frac{4}{\tau}}\cdot\frac{4-5\tau}{4-\tau}\int_{0}^{\infty}\sqrt{i(\epsilon)}d\epsilon.$ (127)
-Proof.
-From Lemma 9 we have
-$\phi(\epsilon)-\phi(\tfrac{\tau\epsilon}{4})$ $\geq\frac{\tau\epsilon}{4}\sqrt{2(i(\tau\epsilon)-i((1-\tfrac{\tau}{4})\epsilon))}$ (128)
-$\geq\frac{\tau\epsilon}{4}\sqrt{2i(\tau\epsilon)}-\frac{\tau\epsilon}{4}\sqrt{2i((1-\tfrac{\tau}{4})\epsilon)}.$ (129)
-Let $\epsilon_{0}:=\sqrt{2\mathbb{E}_{\mu}[\|Z-\mathbb{E}_{\mu}[Z]\|_{2}^{2}]}$.
-Taking $\int_{0}^{\frac{4}{\tau}\epsilon_{0}}\frac{d\epsilon}{\epsilon}$ on both sides above and changing variables, we obtain
-$\int_{\epsilon_{0}}^{\frac{4}{\tau}\epsilon_{0}}\phi(\epsilon)\frac{d\epsilon}{\epsilon}\geq\int_{0}^{4\epsilon_{0}}\frac{d\epsilon}{4}\sqrt{2i(\epsilon)}-\int_{0}^{\frac{4(1-\tfrac{\tau}{4})}{\tau}\epsilon_{0}}\frac{\tau d\epsilon}{4(1-\tfrac{\tau}{4})}\sqrt{2i(\epsilon)}.$ (130)
-Since $i(\epsilon)=0$ and $\phi(\epsilon)=\phi(\epsilon_{0})$ when $\epsilon\geq\epsilon_{0}$, we further have
-$\phi(\epsilon_{0})\ln\frac{4}{\tau}\geq\frac{\sqrt{2}}{4}\left(1-\frac{\tau}{1-\frac{\tau}{4}}\right)\int_{0}^{\infty}\sqrt{i(\epsilon)}d\epsilon$ (131)
-and the claim follows. ∎
-Finally, since $i(\epsilon)\geq R_{\mu}(\epsilon^{2})$ by (30), and $\phi(\epsilon_{0})=\sup_{P_{Z}=\mu}\mathbb{E}[X_{Z}]$, we establish Theorem 3.
-6 New Proof of the Majorizing Measure Theorem
-We prove Corollary 3 from Theorem 3 in this section. First, observe the following auxiliary result, which allows us to rewrite integrals in a “penalized optimization” form:
-Lemma 12.
-Suppose that $y\in(0,\infty)$ is a decreasing function of $x\in(0,\infty)$. Then
-$2\int_{0}^{\infty}ydx\leq\int_{0}^{\infty}\min_{x}\{\alpha^{-2}x^{2}+y^{2}\}d\alpha\leq 4\int_{0}^{\infty}ydx$ (132)
-Proof.
-For any $\alpha\in(0,\infty)$, there is unique $x(\alpha)$ satisfying $\alpha=\frac{x(\alpha)}{y(x(\alpha))}$. Then $x$ and $y$ can be parameterized by $\alpha\in(0,\infty)$, and
-$\int_{0}^{\infty}(\alpha^{-2}x^{2}+y^{2})d\alpha$ $=\int_{\alpha=0}^{\infty}2y^{2}\cdot\frac{ydx-xdy}{y^{2}}$ (133)
-$=4\int_{0}^{\infty}ydx.$ (134)
-The claim then follows since for any $\alpha$, $\frac{\min_{x\in(0,\infty)}\{\alpha^{-2}x^{2}+y^{2}(x)\}}{\alpha^{-2}x^{2}(\alpha)+y^{2}(x(\alpha))}\in[1/2,1]$. ∎
-A similar idea (though in the form of series summation rather than integral) was previously used in proving equivalent forms of the majorizing measure theorem (Talagrand,, 1994).
-Here, we find Lemma 12 very convenient in forming information-theoretic functionals that are exactly convex-concave (Remark 1), and also simplifying evaluations of the rate-distortion integral in applications to ERM.
-To prove Corollary 1, it suffices to consider the case of finite $T$, since the general case (where the lattice supremum in (38) is defined as $\sup_{T_{0}\subseteq T,|T_{0}|<\infty}\sup_{t\in T_{0}}$; see Talagrand, (2014)) will then follows by taking the limit.
-Then $T$ can be identified with points in $\mathbb{R}^{|T|}$ equipped with the $\ell_{2}$ metric.
-For $i(\epsilon)$ defined in (121), Lemma 12 implies
-$4\sup_{\mu}\int_{0}^{\infty}\sqrt{i(\epsilon)}d\epsilon\geq\sup_{P_{Z}}\int_{0}^{\infty}\min_{P_{Z\hat{Z}}\in\Pi(P_{Z},P_{Z})}\{\alpha^{-2}\mathbb{E}[\|Z-\hat{Z}\|_{2}^{2}]+I(Z;\hat{Z})\}d\alpha$ (135)
-where both supremums are over distributions on $T$ and the minimum is over couplings of $P_{Z}$ and $P_{Z}$. The right side of (135) is lower bounded by
-$\sup_{P_{Z}}\inf_{\mu}\int_{0}^{\infty}\min_{P_{U|Z}}\{\alpha^{-2}\mathbb{E}[\|Z-U\|_{2}^{2}]+D(P_{U|Z}\|\mu|P_{Z})\}d\alpha,$ (136)
-since equality would be achieved if $\inf_{\mu}$ were restricted to $\mu=P_{Z}$ and $\min_{P_{U|Z}}$ were restricted to $P_{U|Z}$ satisfying $P_{U|Z}P_{Z}\in\Pi(P_{Z},P_{Z})$.
-Since the relative entropy is jointly convex (i.e., $D(\frac{\nu_{1}+\nu_{2}}{2}\|\frac{\mu_{1}+\mu_{2}}{2})\leq\frac{1}{2}D(\nu_{1}\|\mu_{1})+\frac{1}{2}D(\nu_{2}\|\mu_{2})$), we can show that the integral in (136) is convex in $\mu$ and linear in $P_{Z}$.
-Then Sion’s minimax theorem implies that (136) equals
-$\quad\inf_{\mu}\sup_{P_{Z}}\int_{0}^{\infty}\min_{P_{U|Z}}\{\alpha^{-2}\mathbb{E}[\|Z-U\|_{2}^{2}]+D(P_{U|Z}\|\mu|P_{Z})\}d\alpha$
-$=\inf_{\mu}\sup_{z}\int_{0}^{\infty}\min_{P_{U}}\{\alpha^{-2}\mathbb{E}[\|z-U\|_{2}^{2}]+D(P_{U}\|\mu)\}d\alpha.$ (137)
-Now fix $\mu$, $z$, $\alpha$, and $P_{U}$. Define $r=a\sqrt{\mathbb{E}[\|z-U\|^{2}]}$, where $U\sim P_{U}$ and $a>1$ will be specified later. We obtain, from the data processing inequality,
-$D(P_{U}\|\mu)\geq d(P_{U}(B(z,r))\|\mu(B(z,r)))$ (138)
-where $d(a\|b):=a\ln\frac{a}{b}+(1-a)\ln\frac{1-a}{1-b}$ denotes the binary relative entropy function. By the Markov inequality we have $P_{U}(B(z,r))\geq 1-a^{-2}$. Then we have
-$d(P_{U}(B(z,r))\|\mu(B(z,r)))\geq(1-a^{-2})\ln\frac{1-a^{-2}}{\mu(B(z,r))}-a^{-2}\ln a^{2}$ (139)
-because if $\mu(B(z,r))>1-a^{-2}$ then the right side of(139) is negative and the inequality is obviously true; if $\mu(B(z,r))\leq 1-a^{-2}$ then by monotonicity of the binary relative entropy function we have $d(P_{\hat{Z}}(B(z,r))\|\mu(B(z,r)))\geq d(1-a^{-2}\|\mu(B(z,r)))=(1-a^{-2})\ln\frac{1-a^{-2}}{\mu(B(z,r))}+a^{-2}\ln\frac{a^{-2}}{1-\mu(B(z,r))}$ and (139) also holds.
-Now the integral in (137) is lower bounded by
-$\int_{0}^{\infty}\inf_{r>0}\left\{\frac{r^{2}}{a^{2}\alpha^{2}}+\left[(1-a^{-2})\ln\frac{1-a^{-2}}{\mu(B(z,r))}-a^{-2}\ln a^{2}\right]_{+}\right\}d\alpha$ (140)
-which, by Lemma 12, is lower bounded by twice of
-$\int_{0}^{{\rm diam}(T)}\sqrt{\left[(1-a^{-2})\ln\frac{1-a^{-2}}{\mu(B(z,r))}-a^{-2}\ln a^{2}\right]_{+}}\frac{dr}{a}$
-$\geq\int_{0}^{{\rm diam}(T)}\sqrt{(1-a^{-2})\ln\frac{1}{\mu(B(z,r))}}\frac{dr}{a}-\int_{0}^{{\rm diam}(T)}\sqrt{h(a^{-2})}\frac{dr}{a}$ (141)
-$=\frac{\sqrt{1-a^{-2}}}{a}I_{\mu}(z)-\frac{{\rm diam}(T)}{a}\sqrt{h(a^{-2})}$ (142)
-where we used the basic inequality $\sqrt{[x-y]_{+}}\geq\sqrt{x}-\sqrt{y}$ for $x,y\geq 0$. Thus
-$\frac{2}{c}\mathbb{E}[\sup_{z\in T}X_{z}]\geq\inf_{\mu}\sup_{z}\frac{\sqrt{1-a^{-2}}}{a}I_{\mu}(z)-\frac{{\rm diam}(T)}{a}\sqrt{h(a^{-2})}.$ (143)
-It is easy to see, by considering two points in $T$, that $\mathbb{E}[\sup_{z\in T}X_{z}]\geq{\rm diam}(T)\mathbb{E}[G\vee 0]=\sqrt{\frac{1}{2\pi}}{\rm diam}(T)$, where $G\sim\mathcal{N}(0,1)$.
-Therefore,
-$\left(\frac{2}{c}+\frac{\sqrt{2\pi h(a^{-2})}}{a}\right)\mathbb{E}[\sup_{z\in T}X_{z}]\geq\frac{\sqrt{1-a^{-2}}}{a}\inf_{\mu}\sup_{z}I_{\mu}(z)$ (144)
-as desired.
-Remark 1.
-The minimax inequality step (136)-(137) crucially leverages Lemma 12 and the joint convexity of the relative entropy.
-Specifically, the integral in (136) is convex in $\mu$ and linear in $P_{Z}$, and this property is not true for the functional $\int_{0}^{\infty}\min_{P_{U|Z}\colon\mathbb{E}[\|Z-U\|_{2}^{2}\leq\epsilon^{2}}\sqrt{D(P_{U|Z}\|\mu|P_{Z})}d\epsilon$ before we applied Lemma 12.
-Moreover, had we not used the relative entropy approach, then the integral in (136) might be replaced by the functional $\int_{0}^{\infty}\min_{r_{Z}}\{\alpha^{-2}\mathbb{E}[r_{Z}^{2}]+\ln\frac{1}{\mu(B(Z,r_{Z}))}\}d\alpha$ (see the $\theta_{i}(T)$ functional defined around Proposition 4.3 in Talagrand, (1994)), where $P_{U|Z}$ is replaced by the conditional distribution of $\mu$ on the ball $B(Z,r_{Z})$, $r_{Z}$ being a function of $Z$.
-Again, this functional is not exactly convex in $\mu$, so the application of the minimax theorem is less immediate.
-A classical approach that has a similar role as (136)-(137) is the Fernique convexity argument, which loses a constant factor (Talagrand,, 2014).
-7 Tail Bounds for the ERM
-In this section, we prove the results in Section 3.3.
-7.1 Under the Log-Sobolev Inequality
-The goal of this section is to prove Theorem 4. For any $\epsilon>0$, define
-$s(\epsilon,\mu;\gamma)=\sup_{\nu\colon D(\nu\|\mu)\leq\epsilon}\{\mathsf{w}(\nu;\gamma)+\sqrt{2\epsilon\mathbb{E}_{\nu}[\|Z\|^{2}]}\}.$ (145)
-We may abbreviate $\mathsf{w}(\mu;\gamma)$ as $\mathsf{w}(\mu)$, or $s(\epsilon,\mu;\gamma)$ as $s(\epsilon,\mu)$ or $s(\epsilon)$, when there is no confusion.
-Our key tool is the following observation, which can be viewed as an information-theoretic counterpart of Lemma 1.
-Lemma 13.
-Suppose that $\mu$ is a probability measure supported on a totally bounded set $T\subseteq\ell_{2}$.
-Let $\gamma$ be a probability measure on $\ell_{2}$ which has zero-mean, satisfies the log-Sobolev inequality (17), and such that $\mathbb{E}_{\gamma}[\|Y\|_{2}]<\infty$.
-For any $\lambda\geq 0$ and any $P_{YZ}$ such that $P_{Z}=\mu$ and $P_{Y}=\gamma$, we have
-$\ln\mathbb{E}[\exp(\lambda\left<Y,Z\right>)]\leq\sup_{\epsilon>0}\{\lambda s(\epsilon,\mu;\gamma)-\epsilon\}.$ (146)
-Proof.
-We first suppose that $\mu$ is a rational distribution supported on at most $n$ points.
-For $N>0$ such that $N\mu$ is integer, let $S_{N}$ be the set of types of $N$-sequences over the $n$ points.
-Then $|S_{N}|\leq{n+N-1\choose n-1}$ which grows polynomially in $N$.
-For each $\nu\in S_{N}$, let $C_{\nu}$ be the type class of $\nu$.
-Let $\xi_{\nu}:=\max_{z^{N}\in C_{\nu}}\left<Y^{N},z^{N}\right$> and $\sigma_{\nu}^{2}:=\max_{z^{N}\in C_{\nu}}\mathbb{E}[\left<Y^{N},z^{N}\right>^{2}]=\max_{z^{N}\in C_{\nu}}\|z^{N}\|_{2}^{2}=N\mathbb{E}_{\nu}[\|Z\|_{2}^{2}]$.
-Using Lemma 1, we see that for each $\nu$ and $\epsilon>0$,
-$\mathbb{P}[\xi_{\nu}-\mathbb{E}[\xi_{\nu}]>\sigma_{\nu}\sqrt{2N\epsilon}]\leq\exp(-N\epsilon).$ (147)
-Therefore, setting $S_{N}^{\prime}:=\{\nu\in S_{N}\colon D(\nu\|\mu)\leq\epsilon\}$,
-$\mathbb{P}\left[\bigcup_{\nu\in S_{N}^{\prime}}\left\{\xi_{\nu}-\mathbb{E}[\xi_{\nu}]>\sigma_{\nu}\sqrt{2N\epsilon}\right\}\right]\leq|S_{N}|\exp(-N\epsilon)$ (148)
-which implies that
-$\mathbb{P}\left[\max_{\nu\in S_{N}^{\prime}}\xi_{\nu}>Ns(\epsilon)\right]\leq|S_{N}|\exp(-N\epsilon).$ (149)
-Let $(Y^{N},Z^{N})\sim P_{YZ}^{\otimes N}$, and let $\mathcal{E}$ be the set of $Y^{N}$ in the event in (149).
-Let $\tilde{\mathcal{E}}$ be the event that $D(\widehat{P}_{Z^{N}}\|\mu)>\epsilon$.
-Note that under $\tilde{\mathcal{E}}^{c}$, we have $\left<Y^{N},Z^{N}\right>\leq\max_{\nu\in S_{N}^{\prime}}\xi_{\nu}$.
-Then
-$\mathbb{P}\left[\left<Y^{N},Z^{N}\right>>Ns(\epsilon)\right]$ $\leq\mathbb{P}[\tilde{\mathcal{E}}]+\mathbb{P}\left[\left<Y^{N},Z^{N}\right>>Ns(\epsilon)\cap\tilde{\mathcal{E}}^{c}\right]$ (150)
-[…]
-$\leq\mathbb{P}[\tilde{\mathcal{E}}$$\mathbb{P}\left$$\mathcal{E}\right$] (152)
-$\leq 2(N+1)^{n}\exp(-N\epsilon)$ (153)
-where we used Lemma 3 to estimate $\mathbb{P}[\tilde{\mathcal{E}}$]. Then
-[…]
-$\leq\exp(N\lambda s(0))+\int_{\epsilon=0}^{+\infty}\mathbb{P}[\left<Y^{N},Z^{N}\right>>Ns(\epsilon)]d\exp(\lambda Ns(\epsilon))$ (155)
-$\leq\exp(N\lambda s(0))+2(N+1)^{n}\int_{\epsilon=0}^{+\infty}\exp(-N\epsilon)d\exp(\lambda Ns(\epsilon)).$ (156)
-Since the support of $\mu$ is finite, we see that $s(\epsilon)\lesssim_{\mu}\sqrt{\epsilon}+1$, which implies that the integral in (156) is finite, and
-$\lim_{N\to\infty}\frac{1}{N}\ln\int_{\epsilon=0}^{+\infty}\exp(-N\epsilon)d\exp(\lambda Ns(\epsilon))\leq\sup_{\epsilon>0}\{\lambda s(\epsilon)-\epsilon\}.$ (157)
-Therefore, (156) shows that
-$\ln\mathbb{E}[\exp(\lambda\left<Y,Z\right>)]=\lim_{N\to\infty}\frac{1}{N}\ln\mathbb{E}[\exp(\lambda\left<Y^{N},Z^{N}\right>)]\leq\sup_{\epsilon>0}\{\lambda s(\epsilon)-\epsilon\}.$ (158)
-[…]
-Finally, we remove the assumption of finite dimensionality.
-Denote by $Y_{\leq n}$ the truncation of $Y$ to the first $n$ coordinates, and let $Y_{>n}$ be the remaining coordinates.
-Denote by $\mu_{n}$ the distribution of $Y_{\leq n}$.
-Since $Y,Z\in\ell_{2}$, we have $\lim_{n\to\infty}\|Y_{>n}\|_{2}=0$ and $\lim_{n\to\infty}\|Z_{>n}\|_{2}=0$ almost surely.
-By the Cauchy-Schwarz inequality, we can thus show that $\lim_{n\to\infty}\left<Y_{>n},Z_{>n}\right>=0$ almost surely.
-Then by Fatou’s lemma we have
-$\ln\mathbb{E}[\exp(\lambda\left<Y,Z\right>)]\leq\lim_{n\to\infty}\ln\mathbb{E}[\exp(\lambda\left<Y_{\leq n},Z_{\leq n}\right>)]\leq\lim_{n\to\infty}\sup_{\epsilon>0}\{\lambda s(\epsilon,\mu_{n};\gamma)-\epsilon\}.$ (159)
-It remains to prove that
-$\lim_{n\to\infty}\sup_{\epsilon>0}\{\lambda s(\epsilon,\mu_{n};\gamma)-\epsilon\}\leq\sup_{\epsilon>0}\{\lambda s(\epsilon,\mu;\gamma)-\epsilon\}.$ (160)
-For given $n>0$, suppose that $\nu_{n}$ is a distribution achieving the supremum in the definition of $s(\epsilon,\mu_{n};\gamma)$.
-Then we can extend $\nu_{n}$ to a $\nu$ such that $D(\nu\|\mu)=D(\nu_{n}\|\mu_{n})$, by using the same channel from $Z_{\leq n}$ to $Z_{>n}$ under $\mu$.
-Suppose that $P_{Y_{\leq n}Z}$ achieves the supremum in the definition of $\mathsf{w}(\nu_{n};\gamma)$.
-Then we can extend it to a $P_{YZ}\in\Pi(\nu,\gamma)$.
-For $(Y,Z)\sim P_{YZ}$, we have
-$\mathsf{w}(\nu;\gamma)$ $\geq\mathbb{E}[\left<Y,Z\right$>] (161)
-$=\mathbb{E}[\left<Y_{\leq n},Z\right>]+\mathbb{E}[\left<Y_{>n},Z\right$>] (162)
-$\geq\mathsf{w}(\nu_{n};\gamma)-\mathbb{E}[\|Y_{>n}\|_{2}\|Z_{>n}\|_{2}]$ (163)
-$\geq\mathsf{w}(\nu_{n};\gamma)-\sup_{t\in T}\|t_{>n}\|_{2}\mathbb{E}[\|Y_{>n}\|_{2}]$ (164)
-$=\mathsf{w}(\nu_{n};\gamma)+o(1).$ (165)
-Moreover, we clearly have $\mathbb{E}_{\nu_{n}}[\|Z\|_{2}^{2}]\leq\mathbb{E}_{\nu}[\|Z\|_{2}^{2}]$. Thus (160) is true, and we have extended the result to totally bounded $T$. ∎
-Lemma 14.
-Suppose that $Q$ is a probability measure on $\mathbb{R}$, and $\psi(\lambda):=\ln\mathbb{E}_{X\sim Q}[\exp(\lambda X^{2})]$.
-Then for any $P$ absolutely continuous with respect to $Q$, we have $\mathbb{E}_{X\sim P}[X^{2}]\leq\inf_{\lambda>0}\frac{D(P\|Q)+\psi(\lambda)}{\lambda}$.
-Proof.
-It suffices to prove the case of $P=P_{\lambda}$ defined according to $\frac{dP_{\lambda}}{dQ}(x)\propto\exp(\lambda x^{2})$.
-Note that $\psi$ is a nonnegative convex function, and $\psi(0)=0$.
-Then the claim follows since $\mathbb{E}_{X\sim P_{\lambda}}[X^{2}]=\psi^{\prime}(\lambda)$ and $D(P_{\lambda}\|Q)=\lambda\psi^{\prime}(\lambda)-\psi(\lambda)$.
-∎
-Lemma 15.
-Let $\mu$ be a probability measure with finite support on $T$ for which $\mathsf{G}$ is a width-moment function, and $Z\sim\mu$. Let $\psi(\lambda):=\ln\mathbb{E}[e^{\lambda\|Z\|_{2}^{2}}]$. Then
-$\ln\mathbb{E}[\exp(\eta\left<Y,Z\right>)]\leq\sup_{\epsilon>0}\left\{-\epsilon+\eta\mathsf{G}\left(\frac{\epsilon+\psi(\lambda)}{\lambda}\right)+\eta\sqrt{2\epsilon\cdot\frac{\epsilon+\psi(\lambda)}{\lambda}}\right\}$ (166)
-for any $\lambda,\eta>0$ and any $P_{YZ}$ for which $P_{Y}$ is standard normal and $P_{Z}=\mu$.
-Proof.
-The result follows from Lemma 13, Lemma 14, and the data processing inequality for the KL divergence. ∎
-Proof of Theorem 4.
-Recall that
-$\|\hat{m}-m\|_{2}^{2}\leq\frac{1}{\sqrt{n}}\chi_{m-\hat{m}}$ (167)
-where $\chi_{z}:=-2\left<z,w\right$>. We then have
-[…]
-Using Lemma 15 we have that the logarithm of the right side of (168) is upper bounded by
-$\sup_{\epsilon>0}\left\{-\epsilon+\frac{2\lambda}{\sqrt{n}}\mathsf{G}\left(\frac{\epsilon+\psi(\lambda)}{\lambda}\right)+\sqrt{\frac{8\epsilon\lambda(\epsilon+\psi(\lambda))}{n}}\right\}$ (169)
-and the result is proved. ∎
-7.2 Under Sub-Gaussianity
-In this section, we prove Theorem 5, which is a counterpart of Theorem 4 under sub-Gaussian $\gamma$.
-Since the proof is mostly similar to the proof of Theorem 4, we only point out the parts that are changed.
-It suffices to first prove a counterpart of Lemma 13, namely
-$\ln\mathbb{E}[\exp(\lambda\left<Y,Z\right>)]\leq\sup_{\epsilon>0}\{432C\lambda s(\epsilon)-\epsilon/C\}.$ (170)
-We first suppose that $\mu$ is a rational distribution supported on at most $n$ points, which can be embedded in $\mathbb{R}^{n}$.
-Recall the notations associated with the type class in the proof of Lemma 13.
-Using Lemma 2, we now have that for each type $\nu$ and $\epsilon>0$,
-$\mathbb{P}[\xi_{\nu}>CJ(C_{\nu})+\sigma_{\nu}\sqrt{N\epsilon}]\leq C\exp(-N\epsilon/C),$ (171)
-where $J(C_{\nu})$ denotes the Dudley integral for $C_{\nu}$. Therefore, setting $S_{N}^{\prime}:=\{\nu\in S_{N}\colon D(\nu\|\mu)\leq\epsilon\}$,
-$\mathbb{P}\left[\bigcup_{\nu\in S_{N}^{\prime}}\left\{\xi_{\nu}>CJ(C_{\nu})+\sigma_{\nu}\sqrt{N\epsilon}\right\}\right]\leq C|S_{N}|\exp(-N\epsilon/C).$ (172)
-Since Dudley’s integral is sharp for stationary processes, with a constant bounded by 432 (see (10)), we see that
-$J(C_{\nu})$ $\leq 432\mathbb{E}[\max_{z^{N}\in C_{\nu}}\left<G^{N},z^{N}\right$>] (173)
-$\leq 432N\mathbb{E}_{P_{YZ}\in\Pi(\widehat{P}_{G^{N}},\nu)}[\left<Y,Z\right$>] (174)
-where $G^{N}\sim\mathcal{N}(0,I_{nN})$.
-It is easy to see that $\rho\mapsto\mathbb{E}_{P_{YZ}\in\Pi(\rho,\nu)}[\left<Y,Z\right$>] is concave, so the last displayed item is upper bounded by $432N\mathbb{E}_{P_{GZ}}[\left<Y,Z\right$>].
-This implies that
-$\mathbb{P}\left[\max_{\nu\in S_{N}^{\prime}}\xi_{\nu}>432CNs_{\rm g}(\epsilon)\right]\leq C|S_{N}|\exp(-N\epsilon/C),$ (175)
-where $s_{\rm g}$ is define analogously to (145), but with $\gamma$ replaced by the Gaussian measure $\gamma_{\rm g}$.
-The rest of the proof proceeds as in the proof of Lemma 13, which yields (170), and the extension to general totally bounded $T\in\ell_{2}$ follows by the same limiting argument.
-8 ERM in One Dimension
-In this section, we study the toy problem of empirical risk minimization in one dimension, merely for the purpose of illustrating the application of the rate-distortion integral in ERM.
-We first review the traditional slicing approach from generalization theory, adapting the key ideas from Van de Geer, (1990); Wainwright, (2019), and then give another proof using the new rate-distortion integral approach.
-8.1 Review of the Traditional Slicing Approach
-Suppose that $X_{1},\dots,X_{n}$ are i.i.d.
-whose mean $m\in\mathbb{R}$ is unknown, and $X_{i}-m$ is 1-sub-Gaussian.
-Recall the empirical loss, population loss, and empirical risk minimizer (ERM) defined in Section 2.2.
-For simplicity, let us assume that T=[1,1]T=[-1,1], and then the ERM becomes
-$\hat{m}:=\operatorname*{arg\,min}_{a\in[-1,1]}\hat{L}(a).$ (176)
-In this example, we can directly show that $\hat{m}$ is the projection of $\frac{1}{n}\sum_{i=1}^{n}X_{i}$ to the interval [1,1][-1,1], and obtain the estimate
-$\mathbb{E}|\hat{m}-m|^{2}=O(\frac{1}{n}).$ (177)
-However, in more sophisticated ERM problems, an explicit expression may be unavailable.
-In this section, we review how (177) can be derived from the slicing technique.
-We recall (24) and the process $(\chi_{z})_{z\in T-T}$ defined there, whose natural metric is equivalent to the usual distance on $\mathbb{R}$.
-The covering number is estimated as
-$N([-1,1],\lambda)\asymp 1\vee\frac{1}{\lambda}.$ (178)
-Applying Dudley’s integrals (7) to (24) yields
-$\mathbb{E}[|\hat{m}-m|^{2}]\leq\frac{1}{\sqrt{n}}\int_{0}^{\infty}\sqrt{\ln^{+}\left(\frac{1}{\lambda}\right)}d\lambda\lesssim\frac{1}{\sqrt{n}}$ (179)
-which is not sharp.
-The reason for the looseness is that the estimated parameter is close to the true parameter, but the generalization bounds is a worst-case estimate for any parameter in [1,1][-1,1].
-Therefore, the bound is expected be improved by using a local version of the complexity.
-The classical literature improves the estimate using a “slicing” argument to control the modulus of continuity.
-For example, using (Van Handel,, 2014, Theorem 5.32) we obtain that with high probability,
-$|\hat{m}-m|^{2}\lesssim\frac{1}{\sqrt{n}}\omega(C|\hat{m}-m|)$ (180)
-for some constant $C$, where the modulus of continuity of the process $(\chi_{z})$ is bounded as
-$\omega(\delta)$ $=\int_{0}^{\delta}\sqrt{\ln N([-1,1],\lambda)}d\lambda$ (181)
-$\lesssim\int_{0}^{\delta}\sqrt{\ln^{+}\left(\frac{1}{\lambda}\right)}d\lambda$ (182)
-$\lesssim\delta\sqrt{\ln\frac{1}{\delta}}$ (183)
-for $\delta<0.1$.
-In particular, it follows that $|\hat{m}-m|\lesssim\sqrt{\frac{\ln n}{n}}$ with high probability, which is better than (180) but still has a logarithmic factor of suboptimality due to requiring uniform convergence over all slices.
-Such extra log factors are typical in the analysis of the modulus of continuity; see for example the local and global modulus of continuity of the Wiener process (Koralov and Sinai,, 2007).
-For the analysis of ERM, it is possible to remove the nuisance log factors by further noting that 1) we only need the inequality (180) to hold for all $\hat{m}$ but fixed $m$, and 2) we can design the slicing scheme so that $\omega(\delta)\asymp\delta$ for the “right scale” $\delta\asymp\frac{1}{\sqrt{n}}$ only, and slightly overestimate $\omega(\delta)$ for other scales, yielding a converging sum in the union bound, yet the inequality $\delta^{2}\lesssim\frac{1}{\sqrt{n}}\omega(C\delta)$ still yields $\delta\lesssim 1/\sqrt{n}$.
-Details about constructing the slicing scheme and justifying the regularity conditions can be found, for example, in Van de Geer, (1990); Wainwright, (2019).
-8.2 Rate-Distortion Integral Approach
-In this section, we give another proof of the optimal rate (177) using the rate-distortion integral approach.
-This new method controls the supremum directly using the second moment, without the need of slicing in the traditional method.
-Recall (24), that $|\hat{m}-m|^{2}\leq\frac{1}{\sqrt{n}}\chi_{m-\hat{m}}$, where $\chi_{z}=-2\sqrt{n}z(\frac{1}{n}\sum_{i=1}^{n}X_{i}-m)$, $z\in T-T$ is a centered $O(1)$-sub-Gaussian process.
-Let $Z:=m-\hat{m}$ and $\sigma_{\rm m}:=\sqrt{\operatorname*{var}(Z)}$.
-Using Proposition 1, we obtain
-$\mathbb{E}[\chi_{Z}]$ $\lesssim\lim_{\epsilon\downarrow 0}\sup_{\mu^{\prime}\colon W_{\infty}(P_{Z},\mu^{\prime})\leq\epsilon}\int_{0}^{+\infty}\sqrt{R_{\mu^{\prime}}(\sigma^{2})}d\sigma.$ (184)
-Effectively, we can just think of the right side of (184) as $\int_{0}^{\sigma_{\rm m}}\sqrt{R_{P_{Z}}(\sigma^{2})}d\sigma$, since we will bound the rate-distortion integral by a continuous function of the second moment, and the limiting argument easily applies.
-To move forward, one possible approach is to show that, under the constraint $\operatorname*{var}(Z)\leq\sigma_{\rm m}^{2}$, there is
-$R_{P_{Z}}(\sigma^{2})$ $\leq R_{\mathcal{N}(0,\sigma_{\rm m}^{2})}(\sigma^{2})$ (185)
-$=[\frac{1}{2}\ln\frac{\sigma_{\rm m}^{2}}{\sigma^{2}}]_{+},$ (186)
-where (186) follows from the Gaussian rate-distribution function (Thomas and Joy,, 2006), then by change-of-variable we have $\int_{0}^{\sigma_{\rm m}}\sqrt{R_{P_{Z}}(\sigma^{2})}d\sigma\lesssim\sigma_{\rm m}$.
-We thus obtain $\mathbb{E}[|\hat{m}-m|^{2}]\lesssim\frac{1}{\sqrt{n}}\sqrt{\operatorname*{var}(\hat{m})}\lesssim\frac{1}{\sqrt{n}}\sqrt{\mathbb{E}[|\hat{m}-m|^{2}]}$ and hence $\mathbb{E}[|\hat{m}-m|^{2}]\lesssim\frac{1}{n}$, which is sharp.
-To justify the key step (185), one possible approach is to use Lieb’s rotation argument for proving Gaussian optimality; see Liu et al., (2018); Courtade and Liu, (2021) for the analysis of a similar problem called the reverse Brascamp-Lieb inequality.
-We will not provide the details of this approach here.
-An alternative, simpler approach, is to leverage the penalized version of the rate-distortion integral: from Lemma 12, we have
-$\int_{0}^{\sigma_{\rm m}}\sqrt{R_{P_{Z}}(\sigma^{2})}d\sigma\leq\frac{1}{2}\int_{0}^{\infty}\inf_{P_{U|Z}}\{\alpha^{-2}\mathbb{E}[|U-Z|^{2}]+I(U;Z)\}d\alpha.$ (187)
-This penalized form is helpful, because $P_{U|Z}$ is optimized over a set independent of $P_{X}$, so that saddle point analysis applies. We have the following:
-Theorem 8.
-For any given $\alpha>0$, and $Z$ satisfying $\operatorname*{var}(Z)\leq\sigma_{\rm m}^{2}$, we have
-$\inf_{P_{U|Z}}\{\alpha^{-2}\mathbb{E}[|U-Z|^{2}]+I(U;Z)\}$ $\leq(\frac{1}{2}+\frac{1}{2}\ln\frac{2\sigma_{\rm m}^{2}}{\alpha^{2}})\vee\alpha^{-2}\sigma_{\rm m}^{2}$ (188)
-and hence $\int_{0}^{\infty}\inf_{P_{U|Z}}\{\alpha^{-2}\mathbb{E}[|U-Z|^{2}]+I(U;Z)\}d\alpha\lesssim\sigma_{\rm m}$.
-Proof.
-By translation invariance, it suffices to assume that $\mathbb{E}[|Z|^{2}]\leq\sigma_{\rm m}^{2}$.
-Define the functional $f(P_{Z},P_{U|Z}):=\alpha^{-2}\mathbb{E}[|U-Z|^{2}]+I(U;Z)$, which is convex in $P_{U|Z}$ and concave in $P_{Z}$, and we have the saddle point condition
-$\sup_{P_{Z}\colon\mathbb{E}[|Z|^{2}]\leq\sigma_{\rm m}^{2}}f(P_{Z},P_{U|Z}^{*})\leq f(P_{Z}^{*},P_{U|Z}^{*})\leq\inf_{P_{U|Z}}f(P_{Z}^{*},P_{U|Z})$ (189)
-where $P_{Z}^{*}=\mathcal{N}(0,\sigma_{\rm m}^{2})$ and $P_{U|Z}^{*}$ is an additive noise Gaussian channel that achieves the infimum; see Thomas and Joy, (2006).
-Then $\inf_{P_{U|Z}}f(P_{Z}^{*},P_{U|Z})\geq\sup_{P_{Z}\colon\mathbb{E}[|Z|^{2}]\leq\sigma_{\rm m}^{2}}f(P_{Z},P_{U|Z}^{*})\geq\sup_{P_{Z}\colon\mathbb{E}[|Z|^{2}]\leq\sigma_{\rm m}^{2}}\inf_{P_{U|Z}}f(P_{Z},P_{U|Z})$.
-This shows that under $\mathbb{E}[|Z|]\leq\sigma_{\rm m}^{2}$, the right side of (187) is maximized by $P_{Z}=\mathcal{N}(0,\sigma_{\rm m}^{2})$, in which case
-$\inf_{P_{U|Z}}\{\alpha^{-2}\mathbb{E}[|U-Z|^{2}]+I(U;Z)\}$ $=\inf_{D>0}\left\{\alpha^{-2}D+\frac{1}{2}[\ln\frac{\sigma_{\rm m}^{2}}{D}]_{+}\right\}$ (190)
-$=(\frac{1}{2}+\frac{1}{2}\ln\frac{2\sigma_{\rm m}^{2}}{\alpha^{2}})\vee\alpha^{-2}\sigma_{\rm m}^{2}.$ (191)
-The last claim follows since $\int_{0}^{\sqrt{2}\sigma_{\rm m}}(\frac{1}{2}+\frac{1}{2}\ln\frac{2\sigma_{\rm m}^{2}}{\alpha^{2}})d\alpha+\int_{\sqrt{2}\sigma_{\rm m}}^{\infty}\alpha^{-2}\sigma_{\rm m}^{2}d\alpha\lesssim\sigma_{\rm m}$.
-∎
-To finish, if $\operatorname*{var}(Z)\leq\sigma_{\rm m}^{2}$ and $W_{\infty}(P_{Z},\mu^{\prime})\leq\epsilon$, then the variance under $\mu^{\prime}$ is bounded by $(\epsilon+\sigma_{\rm m})^{2}$, and we deduce from Theorem 8 and (187) that
-$\int_{0}^{+\infty}\sqrt{R_{\mu^{\prime}}(\sigma^{2})}d\sigma\lesssim\sigma_{\rm m}+\epsilon.$ (192)
-Taking $\epsilon\downarrow 0$ and applying (184) and (24), we obtain the sharp rate of $\mathbb{E}[\|\hat{m}-m\|_{2}^{2}]\lesssim\frac{1}{n}$.
-9 ERM for the Sobolev Ellipsoid
-This section provides details of the analysis of the Sobolev ellipsoid in Section 3.4, and prove Theorem 6.
-9.1 Supremum over an Ellipsoid
-Let $X_{z}:=\left<G,z\right$>, where $G=(G(1),G(2),\dots)$ has i.i.d. standard normal coordinates. The majorizing measure theorem shows that (Talagrand,, 2014, Section 2.13)
-$\mathbb{E}[\sup_{z\in\mathcal{E}}X_{z}]\asymp\sqrt{\sum_{i=0}^{\infty}2^{i}a_{2^{i}}^{2}}\asymp\sqrt{\sum_{i=1}^{\infty}a_{i}^{2}}.$ (193)
-On the other hand, a nontrivial calculation (Talagrand,, 2014) shows that the Dudley integral is $\asymp\sum_{i=0}^{\infty}2^{i/2}a_{2^{i}}$ hence may not be sharp.
-In this section we show that the technique in Section 3.2 can recover the sharp upper bound in (193).
-More generally, the upper bound holds when the index set is an image of $\mathcal{E}$ under a Lipschitz map, in view of Corollary 3.
-Although this result is not new, it serves to illustrate the technique.
-Suppose that $Z$ is a random variable taking values in $\mathcal{E}$ and $\Sigma:=\operatorname*{cov}(Z)$.
-Henceforth, we will first assume that the support of $P_{Z}$ is countable to derive bounds on $\mathbb{E}[X_{Z}]$ in terms of $\Sigma$, and then use the approximation argument in Proposition 1 to extend the bound to the possibly uncountable case, noting that the covariance is continuous with respect to the Wasserstein-2 metric.
-Then by Corollary 3,
-$\mathbb{E}[X_{Z}]$ $\lesssim\operatorname*{tr}(\sqrt{\Sigma})$ (194)
-$\leq\sum_{i=1}^{\infty}\sqrt{\Sigma_{ii}}$ (195)
-$\leq\|(a_{i})_{i=1}^{\infty}\|_{2}$ (196)
-where (195) follows from the Schur concavity of the square root function and the fact that the eigenvalues majorizes the diagonal values, and (196) follows from the ellipsoid constraint $\sum_{i=1}^{\infty}\frac{\Sigma_{ii}}{a_{i}^{2}}\leq 1$ and Cauchy-Schwarz.
-This bound recovers the sharp upper bound in (193).
-For the case of the Sobolev ellipsoid, see that for $\beta>1/2$, the optimal rate in (193) is $O(\sqrt{\frac{1}{2\beta-1}})$ for $\beta=O(1)$. In contrast, the Dudley integral is
-$\sum_{i=0}^{\infty}2^{i/2}a_{2^{i}}=\sum_{i=0}^{\infty}2^{-\beta i+i/2}=O(\frac{1}{2\beta-1})$ (197)
-for $\beta=O(1)$, which is worse; see (Wainwright,, 2019, Example 5.12) or (Talagrand,, 2014, Section 2.13).
-As a further example, consider the near critical case $a_{i}=i^{-1/2}(\ln i)^{-1}$. The rate-distortion integral gives a finite bound, whereas the traditional Dudley gives a vacuous bound of $\infty$.
-9.2 Error Bounds for ERM
-In this section, we give two proofs of Theorem 6.
-The first one only applies to a special case of Theorem 6 where $T$ is the identity map and $Y$ satisfies the stronger log-Sobolev inequality.
-The second uses the tools developed in Section 3.3 and applies to the more general setting in Theorem 6.
-9.2.1 A Simple Proof in a Special Case
-We first provide a simple proof of Theorem 6 in a special setting, stated in the following theorem:
-Theorem 9.
-In Theorem 6, assume that $Y$ satisfies the log-Sobolev inequality (17), and $T=\mathcal{E}$. For any $C>1/2$, there exists $C^{\prime}>0$ such that the following holds: for any $\beta\in(1/2,C]$,
-$\mathbb{E}[\|\hat{m}-m\|^{2}]\leq\frac{C^{\prime}}{\sqrt{2\beta-1}}n^{-\frac{2\beta}{2\beta+1}},$ (198)
-and if $T$ is convex and $X_{i}$ satisfies the log-Sobolev inequality, we have
-$\mathbb{P}\left[\|\hat{m}-m\|>\frac{C^{\prime}}{(2\beta-1)^{1/4}}n^{-\frac{\beta}{2\beta+1}}\right]\leq\exp\left(-\frac{1}{C^{\prime}\sqrt{2\beta-1}}n^{\frac{1}{2\beta+1}}\right).$ (199)
-Remark 2.
-For fixed $\beta>1/2$, this is the standard nonparametric rate (Tsybakov,, 2009).
-The assumption of $2\beta>1$ is standard for ensuring the finiteness of the Dudley integral, hence is natural for the analysis of ERM; see (Wainwright,, 2019, Example 5.12).
-Beyond ERM, there are other estimators which achieve the minimax rate $n^{-\frac{2\beta}{2\beta+1}}$ for $\beta\in(0,1/2)$.
-For example, we can project $\bar{X}:=\frac{1}{n}\sum_{i=1}^{n}X_{i}$ to a finite dimensional subspace rather than to $\mathcal{E}$ (which requires the knowledge of $\beta$ to determine that subspace); see the analysis of kernel estimators in (Tsybakov,, 2009).
-The proof we provide here is essentially that of Chatterjee, (2014); Wei et al., (2020).
-We control the “soft” version of the Gaussian width in Definition 3, in lieu of the traditional Gaussian width in Chatterjee, (2014).
-This difference is not essential for convex $T$ due to concentration of the error.
-The bound from this approach is also sharp; see discussion on the converse in Chatterjee, (2014).
-Proof of Theorem 6.
-Recall (24), where $\chi_{z}=-2\sqrt{n}\left<z,\frac{1}{n}\sum_{i=1}^{n}X_{i}-m\right$>, $z\in 2\mathcal{E}$ is a centered sub-Gaussian process, whose metric is the $\ell_{2}$ metric (up to universal constant).
-Invoking Corollary 3, we obtain
-$\mathbb{E}[\chi_{Z}]\lesssim\operatorname*{tr}(\sqrt{\Sigma})$ (200)
-where we defined the covariance matrix $\Sigma:=\operatorname*{cov}(Z)$, and $Z:=m-\hat{m}$.
-Now define $\epsilon:=\operatorname*{tr}(\Sigma)$.
-From the definition of $\mathcal{E}$, we also have $\sum_{i=1}^{\infty}\Sigma_{ii}a_{i}^{-2}\leq 4$.
-Therefore for any $\lambda>0$, we have
-$\sum_{i=1}^{\infty}(a_{i}^{-2}+\lambda)\Sigma_{ii}\leq\lambda\epsilon+4.$ (201)
-Then choosing $\lambda=1/\epsilon$,
-$(\operatorname*{tr}(\sqrt{\Sigma}))^{2}$ $\leq(\sum_{i=1}^{\infty}\sqrt{\Sigma_{ii}})^{2}$ (202)
-$\leq\left(\sum_{i=1}^{\infty}\frac{a_{i}^{-2}+\lambda}{\lambda\epsilon+4}\Sigma_{ii}\right)\left(\sum_{i=1}^{\infty}\frac{\lambda\epsilon+4}{a_{i}^{-2}+\lambda}\right)$ (203)
-$\leq 5\sum_{i=1}^{\infty}a_{i}^{2}\wedge\epsilon,$ (204)
-where the first inequality follows from concavity of $\sqrt{\cdot}$ and the fact that the eigenvalues of $\Sigma$ majorizes its diagonal values. In the case of $a_{i}=i^{-\beta}$, we have
-$\operatorname*{tr}(\sqrt{\Sigma})\lesssim\sqrt{\sum_{i=1}^{\infty}a_{i}^{2}\wedge\epsilon}\asymp\sqrt{\frac{1}{2\beta-1}\epsilon^{1-\frac{1}{2\beta}}}$ (205)
-where we assumed $\beta=O(1)$.
-Combining this with (24) and (200), we obtain $\epsilon\lesssim\frac{1}{\sqrt{n}}\sqrt{\frac{1}{2\beta-1}\epsilon^{1-\frac{1}{2\beta}}}$, which yields $\epsilon\lesssim[(2\beta-1)n]^{-\frac{2\beta}{2\beta+1}}$, as desired.
-To show the concentration inequality (199), we treat $\hat{m}=\hat{m}(\bar{X})$ as a function of $\bar{X}$, which is the projection onto the convex set $T$.
-For any two $x,x^{\prime}\in\ell_{2}$, we have $\left<\hat{m}(x)-\hat{m}(x^{\prime}),x-\hat{m}(x)\right>\geq 0$ and $\left<\hat{m}(x^{\prime})-\hat{m}(x),x^{\prime}-\hat{m}(x^{\prime})\right>\geq 0$, by the hyperplane separation.
-This implies that $\|\hat{m}(x)-\hat{m}(x^{\prime})\|_{2}\leq\|x-x^{\prime}\|_{2}$, hence $\|\hat{m}-m\|_{2}$ is a Lipschitz function.
-Therefore $\left\|\nabla\|\hat{m}-m\|_{2}^{2}\right\|_{2}\leq 2\|\hat{m}-m\|_{2}\leq 2$.
-Since the log-Sobolev inequality implies sub-Gaussian concentration (Boucheron et al.,, 2013), we have
-$\mathbb{P}[\|\hat{m}-m\|>\mathbb{E}[\|\hat{m}-m\|]+t]\leq\exp(-cnt^{2})$ (206)
-for any $t>0$, where $c>0$ is a universal constant. Then (199) follows by choosing $C^{\prime}$ large enough. ∎
-The above simple proof leverages the linearity and concentration, and does not seem to extend to the case of general $\phi$ in Theorem 6.
-In that case, the entropy integral type argument seems to be essential.
-9.2.2 Second Proof
-We now give a full proof of Theorem 6 using the technique in Section 3.3.
-Recall in (24) we defined the $O(1)$-sub-Gaussian process $(\chi_{z})_{T-T}$, where $T$ is now taken to be a Lipschitz image of the ellipsoid, and recall definitions of $\mathsf{G}(E)$ and $\mathsf{G}_{\rm g}(E)$ in Section 3.3.
-Our proof is based on the following key estimate of the width-moment function:
-Theorem 10.
-For $(\chi_{z})_{z\in T-T}$, and any $E>0$, we have $\mathsf{G}(E)\lesssim\mathsf{G}_{\rm g}(E)\lesssim C_{\beta}E^{\frac{2\beta-1}{4\beta}}$.
-Remark 3.
-In (Van de Geer,, 1990, Lemma 5.1), the ERM error for a very similar Sobolev space was bounded by localizing the Dudley integral, where $\beta\geq 1$ is assumed an integer.
-Suppose that we extend the approach of Van de Geer, (1990) to the setting of Theorem 9.
-The covering number is bounded as
-$\ln N(\delta,\mathcal{E})\lesssim\delta^{-\frac{1}{\beta}}$ (207)
-when $\beta=O(1)$.
-(Van de Geer,, 1990, Example 2.1) estimated the covering number for a similar Sobolev space for fixed $\beta$.
-For the Sobolev ellipsoid $\mathcal{E}$, (Wainwright,, 2019, Example 5.12) provided a proof of (207), from which we can see that the best preconstant in (207) is the order of a constant independent of $\beta$ when $\beta=O(1)$.
-For local covering numbers, Wei et al., (2020) showed that the estimate in (207) still holds, where the preconstant is a universal constant.
-We can then calculate the Dudley integral up to the scale of the error $\sqrt{E}$, resulting in a bound of $\int_{0}^{\sqrt{E}}\delta^{-\frac{1}{2\beta}}d\delta\leq O(\frac{1}{2\beta-1})E^{\frac{1}{2}-\frac{1}{4\beta}}$ when $\beta=O(1)$.
-Using our proof instead, we get $\mathbb{E}[\chi_{Z}]\leq O(\frac{1}{\sqrt{2\beta-1}})E^{\frac{1}{2}-\frac{1}{4\beta}}$, which is almost the same, but with a better (optimal) preconstant.
-Thus Dudley’s integral results in $\mathbb{E}[\|\hat{m}-m\|^{2}]=O(\frac{1}{2\beta-1})n^{-\frac{2\beta}{2\beta+1}}$ for $\beta=O(1)$, where the preconstant $O(\frac{1}{2\beta-1})$ is not $O(\frac{1}{\sqrt{2\beta-1}})$ in Theorem 9.
-This suboptimality is fundamental and is due to the looseness of the Dudley integral: as we see in (197), Dudley’s integral gives an $O(\frac{1}{2\beta-1})$ bound on the supremum over $\mathcal{E}$, rather than the optimal $O(\sqrt{\frac{1}{2\beta-1}})$ bound.
-How does the rate-distortion integral improve the preconstant?
-Naively, we can show a result that for a random variable $Z\in 2\mathcal{E}$, $R_{P_{Z}}(\delta^{2})\lesssim\delta^{-\frac{1}{\beta}}$, for $\beta=O(1)$, which is analogous to (207), and proceed by integrating.
-But this will still result in a suboptimal preconstant $O(\frac{1}{2\beta-1})$.
-The reason is that the $P_{Z}$ maximizing $R_{P_{Z}}(\delta^{2})$ changes with $\delta$.
-Instead, we leverage the penalized version Lemma 12 and evaluate the integrand therein for given $\operatorname*{cov}(Z)$, calculate the integral, and optimize $\operatorname*{cov}(Z)$ in the end.
-The proof is made simple by properties of information-theoretic measures such as chain rules of the entropy.
-We first need a few preparations. Define
-$f(t):=\left\{\begin{array}[]{cc}t/2&0\leq t\leq 1\\ \frac{1+\ln t}{2}&t>1\end{array}\right.$. (210)
-We have
-Lemma 16.
-$t\int_{0}^{1/t}f(\frac{1}{u^{2}})du\leq 4f(t)$, for all $t>0$.
-Proof.
-For $0<u_{0}\leq 1$,
-$\int_{0}^{u_{0}}f(\frac{1}{u^{2}})du=\int_{0}^{u_{0}}\frac{1+\ln\frac{1}{u^{2}}}{2}du=\frac{3}{2}u_{0}+u_{0}\ln\frac{1}{u_{0}}\leq\frac{3}{2}u_{0}(1+\ln\frac{1}{u_{0}}).$ (211)
-For $u_{0}>1$, $\int_{1}^{u_{0}}f(\frac{1}{u^{2}})du=\int_{1}^{u_{0}}\frac{1}{2u^{2}}du=\frac{1}{2}(1-\frac{1}{u_{0}})$, so that
-$\int_{0}^{u_{0}}f(\frac{1}{u^{2}})du=\frac{3}{2}+\frac{1}{2}(1-\frac{1}{u_{0}})=2-\frac{1}{2u_{0}}\leq 2.$ (212)
-The result follows by taking $u_{0}=1/t$. ∎
-Lemma 17.
-Assume that $W\in\ell_{2}$ is a random variable and $\Sigma:=\operatorname*{cov}(W)$.
-Let $Z=\phi(W)$, where $\phi$ is a 1-Lipschitz function, and suppose that $\mathbb{E}[\|Z\|^{2}]\leq\epsilon$.
-Then, for any $\alpha\in(0,\infty)$,
-$\inf_{P_{U|Z}}\{\alpha^{-2}\mathbb{E}[\|U-Z\|^{2}]+I(U;Z)\}\leq\alpha^{-2}\epsilon\wedge\sum_{i=1}^{\infty}f(2\alpha^{-2}\Sigma_{ii}).$ (213)
-Proof.
-By considering $U=0$, it is easy to see that $\inf_{P_{U|X}}\{\alpha^{-2}\mathbb{E}[\|U-Z\|^{2}]+I(U;Z)\}\leq\alpha^{-2}\epsilon$. Thus it remains to show that
-$\inf_{P_{U|Z}}\{\alpha^{-2}\mathbb{E}[\|U-Z\|^{2}]+I(U;Z)\}\leq\sum_{i=1}^{\infty}f(2\alpha^{-2}\Sigma_{ii}).$ (214)
-It suffices to prove (214) in the special case where the Lipschitz map $\phi$ is the identity, since the left side of (214) only decreases with Lipschitz $\phi$.
-Furthermore, the left side of (214) is upper bounded by $\sum_{i=1}^{\infty}\inf_{P_{U_{i}|Z_{i}}}\{\alpha^{-2}\mathbb{E}[|U_{i}-Z_{i}|^{2}]+I(U_{i};Z_{i})\}$, because given $P_{U_{i}|Z_{i}}$ we can define $P_{U|Z}=\prod_{i=1}^{\infty}P_{U_{i}|Z_{i}}$, and then $I(U;Z)\leq\sum_{i=1}^{\infty}I(U_{i};Z_{i})$.
-Then the result in (214) follows from the one-dimensional case, which we already proved in Theorem 8.
-∎
-Lemma 18.
-Suppose that $\sum_{i=1}^{\infty}\Sigma_{ii}i^{\beta}\leq 4$, where $\beta>1/2$. For any $\epsilon>0$, we have
-$\int_{0}^{\infty}\alpha^{-2}\epsilon\wedge\sum_{i=1}^{\infty}f(2\alpha^{-2}\Sigma_{ii})d\alpha\leq C_{\beta}\epsilon^{\frac{2\beta-1}{4\beta}},$ (215)
-where $C_{\beta}$ is a constant defined around (236), which satisfies $C_{\beta}=O(\frac{1}{\sqrt{2\beta-1}})$ as $\beta\downarrow\frac{1}{2}$.
-Proof.
-Set
-$\alpha_{0}:=\epsilon^{\frac{1+2\beta}{4\beta}}.$ (216)
-We have
-$\int_{\alpha_{0}}^{\infty}\alpha^{-2}\epsilon\wedge d\alpha$ $\leq\epsilon^{\frac{2\beta-1}{4\beta}}.$ (217)
-Moreover, by Lemma 16
-$\int_{0}^{\alpha_{0}}\sum_{i=1}^{\infty}f(2\alpha^{-2}\Sigma_{ii})d\alpha$ $=\sum_{i=1}^{\infty}\sqrt{2\Sigma_{ii}}\int_{0}^{\frac{\alpha_{0}}{\sqrt{2\Sigma_{ii}}}}f(\frac{1}{u^{2}})du$ (218)
-$\leq\alpha_{0}\sum_{i=1}^{\infty}4f(\frac{\sqrt{2\Sigma_{ii}}}{\alpha_{0}}).$ (219)
-It will become clear that $\alpha_{0}$ was chosen so that (217) and (219) are both upper bounded by $\epsilon^{\frac{2\beta-1}{4\beta}}$. Therefore, it remains to consider the optimization problem:
-$M_{1}:=\sup_{(\Sigma_{i})}\sum_{i=1}^{\infty}f(\frac{\sqrt{2\Sigma_{i}}}{\alpha_{0}})$ (220)
-under the constraint $\sum_{i=1}^{d}\Sigma_{i}a_{i}^{-2}\leq 4$. Also, consider another optimization problem:
-$M_{2}:=\sup_{\Sigma(\cdot)}\int_{0}^{\infty}f(\frac{\sqrt{2\Sigma(t)}}{\alpha_{0}})dt$ (221)
-under the constraint $\int_{0}^{\infty}\Sigma(t)t^{2\beta}dt\leq 1$, where $\Sigma(\cdot)$ is a nonnegative continuous function.
-Clearly $M_{2}\geq M_{1}$, since given $(\Sigma_{ii})$ we can define $\Sigma(t)=\Sigma_{\lceil t\rceil}$.
-Since $f$ has continuous, strictly increasing first derivative, we can see that an optimizer $\Sigma(\cdot)$ exists, which is continuous.
-Since $f(\cdot)$ is increasing, $\Sigma(\cdot)$ must be decreasing.
-There exists unique $t_{0}>0$ such that $\frac{\sqrt{2\Sigma(t)}}{\alpha_{0}}\geq 1$, $t\in(0,t_{0}]$ and $\frac{\sqrt{2\Sigma(t)}}{\alpha_{0}}\leq 1$, $t\in[t_{0},\infty)$.
-Stationarity condition implies that there exists $\lambda>0$ such that
-$\frac{1}{4\Sigma(t)}$ $=\lambda t^{2\beta},\quad t\in(0,t_{0}],$ (222)
-$\frac{1}{2\alpha_{0}\sqrt{2\Sigma(t)}}$ $=\lambda t^{2\beta}\quad t\in[t_{0},\infty).$ (223)
-Setting $t=t_{0}$ in both equations and canceling $\Sigma(t_{0})$, we obtain
-$\lambda=\frac{1}{2\alpha_{0}^{2}t_{0}^{2\beta}}$ (224)
-Also, we can express $\Sigma(t)$ as a function of $t$ from the above, and $\int_{0}^{\infty}\Sigma(t)t^{2\beta}dt=1$ must be achieved, implying
-$4$ $=\int_{0}^{t_{0}}\frac{1}{4\lambda}dt+\int_{t_{0}}^{\infty}\frac{1}{8\alpha_{0}^{2}\lambda^{2}t^{2\beta}}dt$ (225)
-$=\frac{t_{0}}{4\lambda}+\frac{1}{8(2\beta-1)\alpha_{0}^{2}\lambda^{2}t_{0}^{2\beta-1}}$ (226)
-$=\frac{\beta}{2\beta-1}\alpha_{0}^{2}t_{0}^{2\beta+1}.$ (227)
-Now that we are able to express $t_{0}$ and $\lambda$ as functions of $\alpha$, we can evaluate the optimal value:
-$\int_{0}^{t_{0}}f(\frac{\sqrt{2\Sigma(t)}}{\alpha_{0}})dt$ $=\int_{0}^{t_{0}}\frac{1+\ln\frac{\sqrt{2\Sigma(t)}}{\alpha_{0}}}{2}dt$ (228)
-$=\int_{0}^{t_{0}}\frac{1+\beta\ln\frac{t_{0}}{t}}{2}dt$ (229)
-$=\frac{t_{0}(1+\beta)}{2},$ (230)
-and
-$\int_{t_{0}}^{\infty}f(\frac{\sqrt{2\Sigma(t)}}{\alpha_{0}})dt$ $=\int_{t_{0}}^{\infty}\frac{\sqrt{2\Sigma(t)}}{2\alpha_{0}}dt$ (231)
-$=\int_{t_{0}}^{\infty}\frac{t_{0}^{2\beta}}{2t^{2\beta}}dt$ (232)
-$=\frac{t_{0}}{2(2\beta-1)}.$ (233)
-In summary, we have shown that
-$M_{1}\leq M_{2}$ $\leq\left(\frac{1+\beta}{2}+\frac{1}{2(2\beta-1)}\right)t_{0}$ (234)
-$\leq c_{\beta}\alpha_{0}^{-\frac{2}{1+2\beta}}$ (235)
-where we defined
-$c_{\beta}=\left(\frac{1+\beta}{2}+\frac{1}{2(2\beta-1)}\right)\left(\frac{\beta}{4(2\beta-1)}\right)^{-\frac{1}{1+2\beta}}.$ (236)
-Hence
-$\int_{0}^{\infty}\alpha^{-2}\epsilon\wedge\sum_{i=1}^{d}f(2\alpha^{-2}\Sigma_{ii})d\alpha\leq 4\alpha_{0}M_{1}+\epsilon^{\frac{2\beta-1}{4\beta}}=(4c_{\beta}+1)\epsilon^{\frac{2\beta-1}{4\beta}}.$ (237)
-The proof is completed by setting $C_{\beta}=4c_{\beta}+1$. ∎
-Proof of Theorem 10.
-It is obvious from Corollary 2 that $\mathsf{G}(E)\lesssim\mathsf{G}_{\rm g}(E)$, hence we only need to prove the second inequality.
-Let $Z:=m-\hat{m}$, and suppose that $\mathbb{E}[\|Z\|_{2}^{2}]\leq E$, and $Z$ has countable support.
-Invoking Theorem 1, Lemma 12, Lemma 17 and Lemma 18, we obtain
-$\mathbb{E}[\chi_{Z}]$ $\lesssim\int_{0}^{\infty}\inf_{P_{U|X}}\{\alpha^{-2}\mathbb{E}[\|U-Z\|^{2}]+I(U;Z)\}d\alpha$ (238)
-$\lesssim C_{\beta}E^{\frac{2\beta-1}{4\beta}}.$ (239)
-The result is then extended to general $P_{Z}$ by approximation by countable distributions in $W_{\infty}$ and noting the continuity of $E\mapsto E^{\frac{2\beta-1}{4\beta}}$. ∎
-Proof of Theorem 6.
-The proof is based on (24), and we recall that Theorem 10 shows
-$\mathsf{G}(E)\lesssim\mathsf{G}_{\rm g}(E)\lesssim C_{\beta}E^{\frac{2\beta-1}{4\beta}}.$ (240)
-Below we given the details of the proof when $Y$ is assumed to satisfy the log-Sobolev inequality, and use Theorem 4.
-The proof for general sub-Gaussian $Y$ is similar, using Theorem 5, but with uglier constants, hence we omit.
-We see that $\mathsf{G}$ is sub-additive, hence
-$\psi(n/72)$ $\leq\sup_{\epsilon>0}\{-\frac{\epsilon}{3}+\frac{\sqrt{n}}{36}\mathsf{G}(\frac{72\epsilon}{n})\}+\frac{\sqrt{n}}{36}\mathsf{G}(\frac{72\psi(n/72)}{n})$
-$\quad+\sup_{\epsilon>0}\{-\frac{\epsilon}{3}+\frac{\epsilon}{3}\}+\sup_{\epsilon>0}\{-\frac{\epsilon}{3}+\frac{\sqrt{\epsilon\psi(n/72)}}{3}\}.$ (241)
-Let $I_{1},I_{2},I_{3},I_{4}$ denote the four terms in (241). Then $I_{3}=0$, and the other terms can be controlled using (240):
-$I_{1}\lesssim\sup_{\epsilon>0}\{-\epsilon+\sqrt{n}C_{\beta}(\frac{\epsilon}{n})^{\frac{2\beta-1}{4\beta}}\}\lesssim n^{\frac{1}{\beta+1}}C_{\beta}^{\frac{4\beta}{2\beta+1}};$ (242)
-$I_{2}\lesssim C_{\beta}n^{\frac{1}{4\beta}}\psi^{\frac{2\beta-1}{4\beta}}(\frac{n}{72});$ (243)
-$I_{4}\leq\frac{1}{12}\psi(n/72).$ (244)
-Thus we must have either $\frac{1}{3}\psi(n/72)\leq I_{1}$ or $\frac{1}{3}\psi(n/72)\leq I_{2}$, and either option shows that $\psi(n/72)\leq cC_{\beta}^{\frac{4\beta}{2\beta+1}}n^{\frac{1}{2\beta+1}}$ for some universal constant $c>0$.
-Now by the Markov inequality (with $\lambda=n/72$) we have, for any $t>0$,
-$\mathbb{P}[\|m-\hat{m}\|_{2}>t]\leq\exp(cC_{\beta}^{\frac{4\beta}{2\beta+1}}n^{\frac{1}{2\beta+1}}-nt^{2}/72).$ (245)
-For $t_{0}:=10\sqrt{c}C_{\beta}^{\frac{2\beta}{2\beta+1}}n^{-\frac{\beta}{2\beta+1}}$, we have
-$\mathbb{P}[\|m-\hat{m}\|_{2}>t_{0}]\leq\exp(-\frac{7}{18}cC_{\beta}^{\frac{4\beta}{2\beta+1}}n^{\frac{1}{2\beta+1}})$ (246)
-which implies the claimed result. ∎
-10 ERM for Sparse Regression
-In this section, we give two proofs of Theorem 7.
-The first one, based on Hölder’s inequality, is simpler but only applies to a special case; the second uses the rate-distortion integral and proves the theorem in full generality.
-10.1 Simple Proof in a Special Case
-First, we give a simple and direct proof of the following special case of Theorem 7:
-Theorem 11.
-Suppose that $A:=\frac{1}{\sqrt{n}}\mathbb{X}$ is a square matrix, and its largest and smallest singular values satisfy $\sigma_{\rm max}(A)\lesssim 1$ and $\sigma_{\rm max}(A)\gtrsim 1$.
-Let $q\in(0,1)$.
-Then $\mathbb{E}[\|\beta^{*}-\hat{\beta}\|_{2}^{2}]\lesssim_{q}R\left(\frac{\ln d}{n}\right)^{\frac{2-q}{2}}$.
-Before proving Theorem 11, let us recall some auxiliary results. Note that if $x\in wB_{q}(r)$, then for any $p>q$ we have
-$\|x\|_{p}^{p}$ $=\int_{0}^{+\infty}\sum_{i=1}^{d}1_{|x_{i}|>t}d(t^{p})$ (247)
-$\leq\int_{t\in(0,r)}(\frac{r}{t})^{q}d(t^{p})$ (248)
-$=\frac{r^{p}}{1-q/p}$ (249)
-implying the inclusion $wB_{q}(r)\subset B_{p}((1-q/p)^{-\frac{1}{p}}r)$.
-Therefore risk bounds for $\ell_{p}$ balls imply risk bounds for weak $\ell_{q}$ balls, but this naive argument will not give the sharp rate in Theorem 7, since $p$ must be chosen to satisfy the strict inequality $p>q$.
-Nevertheless, we will see that in the setting of Theorem 11, which is a special case of Theorem 7, there is a simple proof based on the interpolation inequality, and it is well-known in functional analysis that interpolation can improve “weak $\ell_{p}$ estimates” at the boundary to an $\ell_{p}$ estimate” at the interpolant.
-More precisely, we will use the following result, which can be viewed as a special case of the Marcinkiewicz interpolation theorem:
-Lemma 19.
-Suppose that $X$ is a nonnegative random variable, $\sup_{t>0}t^{a}\mathbb{P}[X>At]\leq 1$, $\sup_{t>0}t^{a}\mathbb{P}[X>Bt]\leq 1$ for $A,B,a,b>0$ and $\alpha\in(0,1)$, where $a\neq b$ and $\frac{\alpha}{a}+\frac{1-\alpha}{b}=1$.
-Then $\mathbb{E}[X]\lesssim_{a,b}A^{\alpha}B^{1-\alpha}$.
-Proof.
-By a rescaling argument, we see that it suffices to consider the case of $A=B$. Then the proof is completed since
-$\mathbb{E}[X]$ $=\int_{0}^{\infty}\mathbb{P}[X>s]ds$ (250)
-$\leq\int_{0}^{\infty}1\wedge\frac{A^{a}}{s^{a}}\wedge\frac{A^{b}}{s^{b}}ds$ (251)
-$=A\int_{0}^{\infty}1\wedge\frac{1}{t^{a}}\wedge\frac{1}{t^{b}}dt\lesssim_{a,b}A.$ (252)
-Proof of Theorem 11.
-Let $E:=\mathbb{E}[\|\hat{\beta}-\beta^{*}\|_{2}^{2}]$ and $Z:=\hat{\beta}-\beta^{*}$. Suppose that $\theta\in[0,\frac{q}{2-q})$ and
-$\frac{1}{p}=\frac{\theta}{q}+\frac{1-\theta}{2}.$ (253)
-We think of $p>1$ as a function of $\theta$, which is close to 1.
-Let $R_{j}:=\sup_{t}t\mathbb{P}^{1/q}[|Z_{j}|>t]$ and $E_{j}:=\mathbb{E}[|Z_{j}|^{2}]$ for each $j$.
-Then $\sum_{j=1}^{d}E_{j}=E$.
-Since $\beta^{*},\hat{\beta}\in wB_{q}(r)$, we have $\sum_{j=1}^{d}|Z_{j}|^{q}\leq 2r^{q}=2R$, hence $\sum_{j=1}^{d}R_{j}\leq 2R$.
-By the interpolation inequality (Lemma 19), we have
-$\mathbb{E}[|Z_{j}|^{p}]$ $\lesssim_{q}R_{j}^{\frac{\theta p}{q}}E_{j}^{\frac{1-\theta}{2}p}.$ (254)
-Hence for $\theta=0$ we have
-$\mathbb{E}[\|Z\|_{p}^{p}]$ $\lesssim_{q}\sum_{j=1}^{d}R_{j}^{\frac{\theta p}{q}}E_{j}^{\frac{1-\theta}{2}p}$ (255)
-$\leq\|(R_{j}^{\frac{\theta p}{q}})\|_{\frac{q}{\theta p}}\|(E_{j}^{\frac{1-\theta}{2}p})\|_{\frac{2}{(1-\theta)p}}$ (256)
-$\leq(2R)^{\frac{\theta p}{q}}E^{\frac{(1-\theta)p}{2}}.$ (257)
-Moreover, (257) is also trivially true when $\theta=0$ (in which case $p=2$).
-Recall (324). We have, by Hölder’s inequality,
-$E$ $\leq\frac{2}{\sqrt{n}}\mathbb{E}[\|Z\|_{p}\|A^{\top}w\|_{p^{\prime}}]$ (258)
-$\leq\frac{2}{\sqrt{n}}\mathbb{E}^{\frac{1}{p}}[\|Z\|_{p}^{p}]\cdot\mathbb{E}^{\frac{1}{p^{\prime}}}[\|A^{\top}w\|_{p^{\prime}}^{p^{\prime}}]$ (259)
-$\lesssim\frac{1}{\sqrt{n}}(2R)^{\frac{\theta}{q}}E^{\frac{1-\theta}{2}}d^{\frac{1}{p^{\prime}}}\sqrt{p^{\prime}},$ (260)
-where we used the fact that for $G\sim\mathcal{N}(0,1)$ we have $\mathbb{E}^{1/p^{\prime}}[|G|^{p^{\prime}}]\leq 3\sqrt{p^{\prime}}$, and $p^{\prime}=(1-1/p)^{-1}$ is the Hölder conjugate.
-Rearranging, we obtain
-$E^{\frac{1+\theta}{2}}\lesssim_{q}\sqrt{p^{\prime}}d^{\frac{1}{p^{\prime}}-\frac{1}{2}}R^{\frac{\theta}{q}}.$ (261)
-Define
-$\theta_{0}:=\frac{q}{2-q}.$ (262)
-Comparing with (253), we see that $p$ is close to 1 when $\theta$ is close to $\theta_{0}$.
-Case I: $(\frac{1}{q}-\frac{1}{2})\ln d+\frac{1}{2}\ln\frac{(1-q/2)E}{(2R)^{\frac{2}{q}}}\leq\frac{1}{\theta_{0}}$. The assumption of this case implies
-$E\lesssim_{q}R^{\frac{2}{q}}d^{-\frac{2-q}{q}}.$ (263)
-Take $\theta=0$, $p=p^{\prime}=2$, we get from (261) that
-$\sqrt{E}\lesssim_{q}1.$ (264)
-Raising (263) to the power of $q/2$ and (264) to the power of 2q2-q gives $E\lesssim_{q}Rn^{-\frac{2-q}{2}}$.
-Case II: $(\frac{1}{q}-\frac{1}{2})\ln d+\frac{1}{2}\ln\frac{(1-q/2)E}{(2R)^{\frac{2}{q}}}>\frac{1}{\theta_{0}}$. From (253) and (262) we have
-$\frac{1}{p^{\prime}}=(\frac{1}{q}-\frac{1}{2})(\theta_{0}-\theta).$ (265)
-Choosing
-$\theta=\theta_{0}-\frac{1}{(\frac{1}{q}-\frac{1}{2})\ln d+\frac{1}{2}\ln\frac{(1-q/2)E}{(2R)^{\frac{2}{q}}}}\in(0,\frac{q}{2-q})$ (266)
-yields
-$d^{\frac{1}{p^{\prime}}}\left(\frac{(1-q/2)E}{(2R)^{2/q}}\right)^{\frac{\theta_{0}-\theta}{2}}=\left(\frac{d^{\frac{1}{q}-\frac{1}{2}}\sqrt{1-q/2}E^{\frac{1}{2}}}{(2R)^{\frac{1}{q}}}\right)^{\theta_{0}-\theta}=e$ (267)
-and
-$\sqrt{p^{\prime}}=\sqrt{\ln d+(\frac{2}{q}-1)^{-1}\ln\frac{(1-q/2)E}{(2R)^{\frac{2}{q}}}}\leq\sqrt{\ln d}$ (268)
-where we used $\ln\frac{(1-q/2)E}{(2R)^{\frac{2}{q}}}\leq 0$ due to (249). Then (261) shows that
-$E^{\frac{1+\theta_{0}}{2}}$ $\lesssim_{q}d^{-1/2}R^{\frac{\theta_{0}}{q}}\cdot\sqrt{p^{\prime}}d^{\frac{1}{p^{\prime}}}\left(\frac{(1-q/2)E}{(2R)^{2/q}}\right)^{\frac{\theta_{0}-\theta}{2}}$ (269)
-$\lesssim d^{-1/2}R^{\frac{\theta_{0}}{q}}\cdot\sqrt{\ln d}.$ (270)
-Equivalently, we have
-$E\lesssim_{q}R\left(\frac{\ln d}{n}\right)^{\frac{2-q}{2}}.$ (271)
-Remark 4.
-In (Raskutti et al.,, 2011, (11)), a condition $\frac{d}{Rn^{q/2}}\gtrsim d^{\delta}\gtrsim 1$ (for some $\delta\in(0,1)$).
-was globally imposed.
-This condition would imply the optimality of the bound in Theorem 11, but it is not necessary for the upper bound direction.
-10.2 Proof Using the Rate-Distortion Integral
-In this section, we give a proof of Theorem 7 in full generality, using the technique in Section 3.3. The key part is the following estimate of the width-moment function:
-Lemma 20.
-Suppose that $q\in(0,1)$, $Z\in wB_{q}(r)$ in $\mathbb{R}^{d}$, and $\mathbb{E}[\|AZ\|_{2}^{2}]\leq E$, where $A$ is an $n\times d$ matrix whose columns have $\ell_{2}$ norm at most 1. Then we have
-$\mathsf{w}(P_{AZ})\lesssim_{q}\sqrt{\ln d}r^{\frac{q}{2-q}}E^{\frac{1-q}{2-q}}.$ (272)
-In particular, for the set $A(wB_{q}(r))$, we have $\mathsf{G}(E)\lesssim_{q}\sqrt{\ln d}r^{\frac{q}{2-q}}E^{\frac{1-q}{2-q}}$.
-To prove Lemma 20 using the rate-distortion integral, we need to establish upper bounds on the rate-distortion function.
-Let $g\colon(0,\infty)\mapsto(0,\infty)$ be the concave envelope of $x\ln\frac{1}{x}\vee\ln x$.
-It is easy to see that $g$ has continuous first derivative, and there exists $a_{1}\in(0,1)$ and $a_{2}\in(0,1)$ such that
-$g(x)$ $=x\ln\frac{1}{x},\quad\forall x<a_{1};$ (273)
-$g(x)$ $=\ln x,\quad\forall x>a_{2},$ (274)
-and $g$ is linear for $x\in[a_{1},a_{2}]$. We show that this function controls the rate-distortion function of random variables in $\mathbb{R}$ satisfying a weak $\ell_{q}$ condition:
-Lemma 21.
-Suppose that $q\in(0,1)$, $r\in(0,\infty)$, and $Z\in\mathbb{R}$ is a random variable satisfying $\sup_{t>0}t^{q}\mathbb{P}[|Z|>t]\leq r^{q}$.
-Then there exists a random variable $U$ such that $|U|\leq|Z|$ almost surely, $\mathbb{E}[|U-Z|]\leq b$, and
-$H(U)\lesssim_{q}g\left(\left(\frac{r}{b}\right)^{\frac{q}{1-q}}\right).$ (275)
-Proof.
-By scaling invariance of the entropy, it suffices to consider the case of $r=1$, which we shall assume henceforth.
-Let $t_{k}=k(\frac{1-q}{2-q}b)^{\frac{1}{1-q}}$, $k\in\mathbb{Z}$.
-Given $Z$, define $U=t_{k}$ where $k=\operatorname*{arg\,min}_{l\colon|t_{l}|\leq|Z|}|Z-t_{l}|$.
-Then
-$\mathbb{E}[|U-Z|]$ $\leq\mathbb{E}[|Z|1_{|Z|\leq t_{1}}]+t_{1}P_{Z}(\mathbb{R}\setminus[t_{-1},t_{1}])$ (276)
-$\leq\int_{0}^{t_{1}}\frac{dt}{t^{q}}+t_{1}^{1-q}$ (277)
-$=\frac{2-q}{1-q}t_{1}^{1-q}=b.$ (278)
-To bound the entropy, first consider the case of $t_{1}^{-q}\leq 2/3$.
-For $k\geq 1$, define $p_{k}:=p_{-k}:=\frac{1}{2}t_{1}^{-q}\left(\frac{1}{k^{q}}-\frac{1}{(k+1)^{q}}\right)$, and $p_{0}:=1-t_{1}^{-q}$.
-We see that $p_{k}$ is a symmetric sequence and decreasing for $k\geq 0$ by our assumption.
-Then $P_{\xi}(kt_{1})=p_{k}$, $k\in\mathbb{Z}$ is the solution to the following optimization problem:
-$\max_{P_{\xi}}H(\xi),\textrm{ where $\xi\in t_{1}\mathbb{Z}$ satisfies $\mathbb{P}[|\xi|\geq kt_{1}]\leq(kt_{1})^{-q}$ for all $k\in\mathbb{Z}\setminus\{0\}$. }$ (279)
-Indeed, this is a convex optimization problem, and using the Lagrange multiplier method, we see that the local optimality condition can be expressed as $P_{\xi}(kt_{1})=\sum_{l=|k|}^{\infty}\lambda_{l}$ for some nonnegative $\lambda_{0},\lambda,\dots$, where $\lambda_{k}=0$ if $\mathbb{P}[|\xi|\geq kt_{1}]<(kt_{1})^{-q}$ (complementary slackness).
-Since $P_{\xi}(kt_{1})=p_{k}$ satisfies local optimality, it is also globally optimal.
-Clearly $H(U)$ is upper bounded by the value of (279).
-Setting $l_{k}:=\frac{1}{k^{q}}-\frac{1}{(k+1)^{q}}$, we have $\sum_{k\geq 1}l_{k}=1$, and
-$\sum_{k\geq 1}l_{k}\ln\frac{1}{l_{k}}$ $\leq\sum_{k\geq 1}\frac{q}{k^{q+1}}\ln\frac{(k+1)^{q+1}}{q}\lesssim_{q}1.$ (280)
-Hence
-$H(U)$ $\leq\sum_{k\in\mathbb{Z}}p_{k}\ln\frac{1}{p_{k}}$ (281)
-$\leq p_{0}\ln\frac{1}{p_{0}}+\sum_{k\geq 1}2p_{k}\ln\frac{1}{p_{k}}$ (282)
-$\leq t_{1}^{-q}+t_{1}^{-q}\left(\sum_{k\geq 1}l_{k}\ln\frac{2t_{1}^{q}}{l_{k}}\right)$ (283)
-$\lesssim_{q}g(t_{1}^{-q}).$ (284)
-Next, consider the case of $t_{1}^{-q}>2/3$.
-Since $p_{k}$ is no longer a decreasing sequence for $k\geq 0$, the optimizer for (279) cannot be expressed in the same way as before.
-However, it is possible to bound $H(U)$ for this case using the bound in the previous case.
-We choose $\bar{t}_{1}=mt_{1}$ where $m=\lceil 1.5^{1/q}t_{1}^{-1}\rceil\geq 2$ is the minimum integer such that $\bar{t}_{1}^{-q}\leq 2/3$.
-Define $\bar{t}_{k}=k\bar{t}_{1}$, $k\in\mathbb{Z}$.
-Given $Z$, define $\bar{U}=\bar{t}_{k}$ where $k=\operatorname*{arg\,min}_{l\colon|\bar{t}_{l}|\leq|Z|}|Z-\bar{t}_{l}|$.
-Then since $U$ can take at most 2m12m-1 possible values given $\bar{U}$, we have $H(U|\bar{U})\leq\ln(2m+1)$, hence
-$H(U)$ $\leq H(\bar{U})+\ln(2m-1)$ (285)
-$\lesssim g(\bar{t}_{1}^{-q})+\ln m$ (286)
-where we used (284) to bound $H(\bar{U})$.
-Since $m\asymp_{q}t_{1}^{-1}$, we have $\bar{t}_{1}^{q}\asymp_{q}1$, and $H(U)\lesssim_{q}1+\ln\frac{1}{t_{1}}\lesssim_{q}g(t_{1}^{-q})$.
-In summary, in either case, we have established $H(U)\lesssim_{q}g(t_{1}^{-q})$ in the case of $r=1$, and the proof for general $r$ is completed by rescaling.
-∎
-By subadditivity of entropy, Lemma 21 implies a bound on the rate-distortion function of weak $\ell_{q}$ balls:
-Lemma 22.
-Suppose that $Z\in wB_{q}(r)$ is a random variable, and $b>0$ is arbitrary. Then we can find $U=(U_{1},\dots,U_{d})$ such that $U\in wB_{q}(r)$ almost surely, $\mathbb{E}[\|U-Z\|_{1}]\leq b$, and
-$H(U)\lesssim_{q}dg\left(d^{-1}\left(\frac{r}{b}\right)^{\frac{q}{1-q}}\right).$ (287)
-Proof.
-Let $r_{j}=\sup_{t}t\mathbb{P}^{1/q}[|Z_{j}|>t]$, so that $\sum_{j}r_{j}^{q}\leq r^{q}$.
-Let $b_{j}:=\frac{r_{j}^{q}b}{r^{q}}$.
-By Lemma 21, we can find $U_{1},\dots,U_{d}$ such that $|U_{i}|\leq|Z_{i}|$, i=1i=1,…, and that
-$\mathbb{E}[\|U-Z\|_{1}]=\sum_{j=1}^{d}\mathbb{E}[|U_{j}-Z_{j}|]\leq\sum_{j=1}^{d}b_{j},$ (288)
-and that
-$H(U)$ $\leq\sum_{j=1}^{d}H(U_{j})$ (289)
-$\lesssim_{q}\sum_{j=1}^{d}g\left(\left(\frac{r_{j}}{b_{j}}\right)^{\frac{q}{1-q}}\right)$ (290)
-$=\sum_{j=1}^{d}g\left(r_{j}^{q}\left(\frac{b}{r^{q}}\right)^{-\frac{q}{1-q}}\right)$ (291)
-$\leq dg\left(d^{-1}\left(\frac{r}{b}\right)^{\frac{q}{1-q}}\right)$ (292)
-where the last step used Jensen’s inequality to see that the maximum value is achieved when $r_{j}^{q}=r^{q}/d$. ∎
-Next, we prove the following information-theoretic version of Sudakov minoration, which will be our main tool for upper bound on the $\ell_{2}$ rate-distortion function”.
-Lemma 23.
-Suppose that $Z$ is a random variable in $\mathbb{R}^{d}$, and $G\sim\mathcal{N}(0,I_{d})$. Then for any $\epsilon>0$,
-$\inf_{P_{V|Z}\colon\mathbb{E}[\|V-Z\|_{2}^{2}]\leq\epsilon^{2}}I(V;Z)$ $\lesssim\frac{1}{\epsilon^{2}}\sup_{P_{ZG}}|\mathbb{E}[\left<Z,G\right>]|^{2}$ (293)
-where the supremum is over couplings of $P_{Z}$ and $P_{G}$.
-Proof.
-The proof follows by applying the lifting argument to the standard Sudakov minoration (see for example (Boucheron et al.,, 2013, Theorem 13.4)), which should upper bound the left side of (293) by $\frac{2}{\epsilon^{2}}|\mathbb{E}[\left<Z,G\right>]|^{2}$.
-Alternatively, the result follows by taking $\epsilon_{1}\to\infty$ in Lemma 9 of the present paper, although this is somewhat overkill, as Lemma 9 itself is a consequence of the super-Sudakov inequality-a strengthening of the standard Sudakov inequality.
-∎
-Next, we prove auxiliary results showing that the inner product in (293) can be “essentially” controlled by $\mathbb{E}[\|Z\|_{\ell_{1}}]$.
-Lemma 24.
-Suppose that $Z\in\mathbb{R}$ satisfies $\mathbb{E}[|Z|]\leq a$, and $|Z|\leq b$ almost surely, where $a\geq 0$ and $b>0$.
-Let $G\sim\mathcal{N}(0,1)$.
-Then $\mathbb{E}[ZG]\leq 3ebf(a/b)$, where we defined the function:
-$f(x)=\left\{\begin{array}[]{cc}x\sqrt{\ln\frac{1}{x}}&x\in(0,e^{-1})\\ e^{-1}&x\in[e^{-1},\infty)\end{array}\right.$. (296)
-Proof.
-First, consider the case of $a=1$ and $b>e$.
-For any $p>1$, we obtain $\mathbb{E}[|Z|^{p}]\leq b^{p-1}$ from the interpolation inequality for the norm.
-Therefore, if $p$ is such that $\frac{1}{p^{\prime}}=1-\frac{1}{p}=\frac{1}{\ln b}$, we have
-$\mathbb{E}[ZG]$ $\leq\|G\|_{p^{\prime}}\|Z\|_{p}$ (297)
-$\leq 3\sqrt{p^{\prime}}\left(b^{p-1}\right)^{1/p}$ (298)
-$\leq 3e\sqrt{\ln b}$ (299)
-where we defined $\|Z\|_{p}:=\mathbb{E}^{1/p}[|Z|^{p}]$.
-If $b\leq e$, set $p=\infty$ and $p^{\prime}=1$.
-We get $\mathbb{E}[ZG]\leq 3b$.
-This establishes the claim for $a=1$.
-For general $a\geq 1$, consider $Z^{\prime}=Z/a$ and we have $|Z^{\prime}|\leq ba^{-1}$ almost surely.
-The result follows by applying the $a=1$ case we just proved to $Z^{\prime}$.
-∎
-Lemma 25.
-Suppose that $Z\in\mathbb{R}^{d}$ satisfies $\mathbb{E}[\sum_{i=1}^{d}|Z_{i}|]\leq a$, and $Z\in wB_{q}(r)$ almost surely for some $q>0$.
-Let $A$ be a matrix with $d$ columns and arbitrarily many rows, where the column vectors satisfy $\|A_{i}\|_{2}\leq 1$ for all $i=1,\dots,d$.
-Let $G$ be a standard normal vector whose dimension equals the number of rows of $A$.
-Then we have
-$\mathbb{E}[\left<G,AZ\right>]\lesssim rdf(a/rd),$ (300)
-and
-$\inf_{P_{V|AZ}\colon\|V-AZ\|^{2}\leq\epsilon^{2}}I(V;AZ)$ $\lesssim\frac{1}{\epsilon^{2}}r^{2}d^{2}f^{2}(a/rd).$ (301)
-Proof.
-Let $g_{i}:=(A^{\top}G)_{i}$, and $a_{i}=\mathbb{E}[|Z_{i}|]$ $i=1,\dots,d$.
-Since $Z\in wB_{q}(r)$, we have $|Z_{i}|\leq r$, almost surely (this is the only place we use $Z\in wB_{q}(r)$).
-Then since $f$ is an increasing concave function, and $\operatorname*{var}(g_{i})\leq 1$, we apply Lemma 24 to obtain
-$\mathbb{E}[\left<g,Z\right$>] $\lesssim\sum_{j=1}^{d}rf(a_{j}/r)$ (302)
-$\leq rdf(a/rd).$ (303)
-The claim then follows from Lemma 23. ∎
-Lemma 26.
-There exists $c_{q}\geq 2$ depending only $q\in(0,1)$, such that the following holds: if $d>c_{q}$, $Z\in wB_{q}(r)$, $\epsilon>0$, and the $\ell_{2}$ norms of the columns of $A$ are not exceeding 1, then we can find $P_{V|AZ}$ such that $\mathbb{E}[\|A(V-Z)\|_{2}^{2}]\leq\epsilon^{2}$, and
-$I(V;AZ)\lesssim_{q}\left(\frac{r}{\epsilon}\right)^{\frac{2q}{2-q}}\ln d.$ (304)
-Proof.
-By Lemma 22, for any $b>0$, we can find $U$ such that $U\in wB_{q}(r)$ almost surely, $\mathbb{E}[\|U-Z\|_{1}]\leq b$, and
-$H(U)$ $\lesssim_{q}dg\left(d^{-1}\left(\frac{r}{b}\right)^{\frac{q}{1-q}}\right).$ (305)
-For any $u$, let $b_{u}:=\mathbb{E}[\|Z-U\|_{1}|U=u]$, and $\epsilon_{u}:=\sqrt{\frac{b_{u}}{b}}\epsilon$.
-Note that $Z-U\in wB(2^{1/q}r)$ almost surely.
-We can apply Lemma 25 to each $u$ to find $P_{V|AZ,U=u}$ such that $\mathbb{E}[\|AZ-V\|_{2}^{2}|U=u]\leq\epsilon_{u}^{2}$ and $I(V;AZ|U=u)\lesssim_{q}\frac{1}{\epsilon_{u}^{2}}r^{2}d^{2}f^{2}(b_{u}/rd)$.
-Hence
-$I(V;AZ|U)$ $\lesssim_{q}\mathbb{E}\left[\frac{1}{\epsilon_{U}^{2}}r^{2}d^{2}f^{2}(b_{U}/rd)\right]$ (306)
-$\leq\mathbb{E}\left[\frac{br^{2}d^{2}}{\epsilon^{2}b_{U}}f^{2}(\frac{b_{U}}{rd})\right]$ (307)
-$\lesssim_{q}\mathbb{E}\left[\frac{br^{2}d^{2}}{\epsilon^{2}b_{U}}\cdot(\frac{b_{U}}{rd})^{2}\ln\frac{rd}{b_{U}}\right]$ (308)
-$\leq\frac{b^{2}}{\epsilon^{2}}\ln\frac{rd}{b}$ (309)
-where
-  • (308) follows since by (249), we have $\frac{b_{U}}{r}\lesssim_{q}1$ almost surely. By choosing $c_{q}$ large enough, we can ensure that $\frac{b_{U}}{rd}<e^{-1}$.
-  • (309) follows by applying Jensen’s inequality to the concave function $t\to t\ln\frac{1}{t}$.
-Then (305) and (309) yield
-$I(V;AZ)$ $\leq I(U;AZ)+I(V;AZ|U)$ (310)
-$\lesssim_{q}dg\left(d^{-1}\left(\frac{r}{b}\right)^{\frac{q}{1-q}}\right)+\frac{b^{2}}{\epsilon^{2}}\ln\frac{rd}{b}.$ (311)
-We now consider two cases:
-Case I: $2d>\left(\frac{r}{\epsilon}\right)^{\frac{2q}{2-q}}$.
-We take $b=r^{\frac{q}{2-q}}\epsilon^{\frac{2-2q}{2-q}}$, which is selected to essentially balance the two terms in (311). Then
-$d^{-1}\left(\frac{r}{b}\right)^{\frac{q}{1-q}}=d^{-1}\left(\frac{r}{\epsilon}\right)^{\frac{2q}{2-q}}<2.$ (312)
-Since $g(t)\lesssim t\ln\frac{4}{t}$ for $t\in(0,2)$, we have
-[…]
-$\lesssim_{q}\left(\frac{r}{\epsilon}\right)^{\frac{2q}{2-q}}\ln d.$ (314)
-Case II: $2d\leq\left(\frac{r}{\epsilon}\right)^{\frac{2q}{2-q}}$.
-We take $b=r^{\frac{q}{2-q}}\epsilon^{\frac{2-2q}{2-q}}(\ln\frac{r}{\epsilon})^{-1/2}$. Then
-$\frac{b^{2}}{\epsilon^{2}}\ln\frac{rd}{b}$ $\lesssim_{q}\left(\frac{r}{\epsilon}\right)^{\frac{2q}{2-q}}\frac{\ln d+\ln\frac{r}{\epsilon}}{\ln\frac{r}{\epsilon}}$ (315)
-$\lesssim\left(\frac{r}{\epsilon}\right)^{\frac{2q}{2-q}}\ln d,$ (316)
-and
-$dg\left(d^{-1}\left(\frac{r}{b}\right)^{\frac{q}{1-q}}\right)$ $=dg\left(d^{-1}\left(\frac{r}{\epsilon}\right)^{\frac{2q}{2-q}}(\ln\frac{r}{\epsilon})^{\frac{q}{2(1-q)}}\right)$ (317)
-$\lesssim_{q}d\ln\left(d^{-1}\left(\frac{r}{\epsilon}\right)^{\frac{2q}{2-q}}(\ln\frac{r}{\epsilon})^{\frac{q}{2(1-q)}}\right).$ (318)
-Let $M:=d^{-1}\left(\frac{r}{\epsilon}\right)^{\frac{2q}{2-q}}\geq 2$. Then
-[…]
-$\lesssim_{q}\left(\frac{r}{\epsilon}\right)^{\frac{2q}{2-q}}\ln d$ (320)
-where the last step is justified as $\frac{\ln M}{M}\lesssim 1$ and $\frac{1}{M}\ln\ln\frac{r}{\epsilon}\lesssim\frac{1}{M}\ln\frac{r}{\epsilon}\lesssim_{q}\frac{\ln d+\ln M}{M}\lesssim\ln d$.
-Therefore in either case, we have established (304), as desired.
-∎
-We are now ready to prove the key estimate of the width-moment function:
-Proof of Lemma 20.
-From Lemma 26 and Theorem 1 we obtain
-$\mathsf{w}(P_{AZ})$ $\lesssim\int_{0}^{\sqrt{E}}\sqrt{R_{P_{AZ}}(\sigma^{2})}d\sigma$ (321)
-$\lesssim_{q}\int_{0}^{\sqrt{E}}\left(\frac{r}{\sigma}\right)^{\frac{q}{2-q}}\sqrt{\ln d}\,d\sigma$ (322)
-$\lesssim_{q}\sqrt{\ln d}r^{\frac{q}{2-q}}E^{\frac{1-q}{2-q}}.$ (323)
-A weaker version of Theorem 7, which controls the mean square error rather than providing a high-probability bound, can now be obtained by taking the expectation of both sides of (24):
-Theorem 12.
-$\frac{1}{n}\mathbb{E}[\|\mathbb{X}\hat{\beta}-\mathbb{X}\beta\|_{2}^{2}]\lesssim_{q}(\frac{\ln d}{n})^{\frac{2-q}{2}}r^{q}$.
-Proof of Theorem 12.
-Recall from (24) that
-$\|\hat{m}-m\|^{2}\leq\frac{1}{\sqrt{n}}\chi_{m-\hat{m}}$ (324)
-where $\chi_{z}:=-2\left<z,w\right$>, $z\in 2A(wB_{q}(r))$ is a centered sub-Gaussian process, whose metric is the $\ell_{2}$ metric (up to universal constant).
-Let $Z:=m-\hat{m}$ and $E:=\mathbb{E}[\|Z\|_{2}^{2}]$.
-From Lemma 20 we obtain
-$\mathbb{E}[\chi_{m-\hat{m}}$] $\lesssim_{q}\sqrt{\ln d}r^{\frac{q}{2-q}}E^{\frac{1-q}{2-q}}.$ (325)
-Then, using (324), we have
-$\mathbb{E}[\|\hat{m}-m\|^{2}]\lesssim_{q}\sqrt{\frac{\ln d}{n}}r^{\frac{q}{2-q}}\mathbb{E}^{\frac{1-q}{2-q}}[\|\hat{m}-m\|^{2}],$ (326)
-or equivalently,
-$\mathbb{E}[\|\hat{m}-m\|^{2}]\lesssim_{q}(\frac{\ln d}{n})^{\frac{2-q}{2}}r^{q}.$ (327)
-The proof of Theorem 7 is similar to that of Theorem 12, but we use the technique of Section 3.3 instead, which controls the moment generation function of both sides of (24):
-Proof of Theorem 7.
-We provide the details of the calculation based on Theorem 4, where $Y$ is assumed to satisfy the log-Sobolev inequality.
-The proof for $Y$ satisfying $\|Y\|_{\Psi_{2}}\leq 1$ is similar, but uses Theorem 5 instead.
-Using Lemma 20, and the fact that $wB_{q}(r)-wB_{q}(r)\subseteq 2^{1/q}wB_{q}(r)$, we have
-$\frac{1}{\sqrt{n}}\mathsf{G}(E)\lesssim_{q}BE^{\frac{1-q}{2-q}}$ (328)
-where
-$B:=\sqrt{\frac{\ln d}{n}}R^{-\frac{1}{2-q}}.$ (329)
-We see that $\mathsf{G}$ is a sub-additive function. Therefore
-$\psi(n/72)$ $\leq\sup_{\epsilon>0}\{-\frac{\epsilon}{3}+\frac{\sqrt{n}}{36}\mathsf{G}(\frac{72\epsilon}{n})\}+\frac{\sqrt{n}}{36}\mathsf{G}(\frac{72\psi(n/72)}{n})$
-$\quad+\sup_{\epsilon>0}\{-\frac{\epsilon}{3}+\frac{\epsilon}{3}\}+\sup_{\epsilon>0}\{-\frac{\epsilon}{3}+\frac{\sqrt{\epsilon\psi(n/72)}}{3}\}.$ (330)
-Let $I_{1},I_{2},I_{3},I_{4}$ denote the four terms in (330). Then $I_{3}=0$, and the other terms can be controlled using (328):
-$I_{1}\lesssim_{q}\sup_{\epsilon>0}\{-\epsilon+Bn(\epsilon/n)^{\frac{1-q}{2-q}}\}\lesssim_{q}B^{2-q}n;$ (331)
-$I_{2}\lesssim_{q}Bn\left(\frac{\psi(n/72)}{n}\right)^{\frac{1-q}{2-q}};$ (332)
-$I_{4}\leq\frac{1}{12}\psi(n/72).$ (333)
-Thus we must have either $\frac{1}{3}\psi(n/72)\leq I_{1}$ or $\frac{1}{3}\psi(n/72)\leq I_{2}$, which shows that $\psi(n/72)\leq c_{q}B^{2-q}n$ for some $c_{q}>0$ that may depend on $q$.
-Now by the Markov inequality (with $\lambda=n/72$) we have, for any $t>0$,
-$\mathbb{P}[\|m-\hat{m}\|_{2}>t]\leq\exp(c_{q}nB^{2-q}-nt^{2}/72).$ (334)
-The result then follows by taking $t=\sqrt{72}(\sqrt{c_{q}}+1)B^{\frac{2-q}{2}}$. ∎
-11 Acknowledgment
-The author would like to thank Professor Ramon van Handel for feedback on an early version of the manuscript, and for encouraging the pursuit of a new proof of the majorizing measure theorem based on the rate-distortion integral.
-The author also thanks Professor Garvesh Raskutti and Philippe Rigollet for discussion and feedback on sparse linear regression.
-This research was supported in part by NSF Grant DMS-2515510.
+$K^{\prime}\mathbb{E}[X_{Z}]\geq K^{\prime}\lim_{k\to\infty}\mathbb{E}[X_{Z_{k}}]\geq\lim_{k\to\infty}\int_{0}^{\infty}\sqrt{R_{\mu_{k}}(\sigma^{2})}\,d\sigma\geq\int_{0}^{\infty}\sqrt{R_{\mu}((1+a)\sigma^{2})}d\sigma$ (147)
+by Fatou’s lemma. Finally, taking $a\downarrow 0$ shows that the countable case holds with the same constant as the finite case.